A Complete Bibliography of Publications in Social Studies of Science

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

23 August 2025 Version 1.21

Title word cross-reference

$10.95 [Hol81b]. $100.00 [Ash07]. $12.00 [Nev80]. $13.50 [O’R83]. $13.95/£ [Kim94]. $135.00 [Hyy09]. $14.95 [Wri80]. $15.00 [Wri80]. $15.95 [Alb95, Mos79]. $16.00 [Doi11]. $16.50 [Wri80]. $17.50 [Blo77]. $17.95 [Bro95, Del05, Pra09]. $18.50 [Nev80]. $19.00 [Sis10]. $19.95 [Del05]. $20.00 [Gin10]. $22.50 [GM78]. $23.95 [Del02]. $24.95 [Col07, Gre10, Sto08]. $24.99 [Del05]. $25.00 [Sis10]. $25.95 [Mae07]. $27.50 [Col07]. $27.95 [Del02, Fis10, Ros10]. $28.00 [Moo10]. $29.95 [Del02]. $30.00 [GM78, Rob99]. $32.00 [Kim94]. $32.50 [Doi11]. $32.99 [Rei06]. $34.30 [Hol81b]. $39.00 [Ash07]. $39.50 [Wes88]. $41.00 [Del05]. $42.00 [Goo10]. $5.75 [Bru77]. $54.95 [Alb95]. $55.00 [Gin10]. $56.00 [Moo10]. $70.00 [Del05, Mae07]. $75.00 [Rei06]. $8.75 [Wri80]. $80.00 [Del05]. $85.00 [Col07, Gre10]. $9.50 [O’R83]. $9.95 [Nev80]. $90.00 [Del05, Pra09]. $93.83 [Col07].

-06097-7 [Kim94].

/$79.95 [Eps09].

0 [AB80, Blo77, Bro95, Hol81b, Mir07, Mos79, Rei06, Rob99, Sto08, Wes88, Wri80]. 0-00-262761-2 [Hol81b]. 0-03-021821-7 [Hol81b]. 0-07-010070-5 [Blo77]. 0-07-010112-4 [Blo77]. 0-19-515107-0 [Col07]. 0-19-516941-7 [Col07]. 0-19-928718-X [Sto08]. 0-202-30284-9 [GM78]. 0-223-96814-5 [Mos79]. 0-226-06096-9 [Kim94]. 0-226-06747-5 [Gin10]. 0-226-06751-3 [Gin10]. 0-226-11377-9 [Ash07]. 0-226-11378-7 [Ash07]. 0-226-21309-9 [Sis10]. 0-226-21310-2 [Sis10]. 0-262-02589-2 [Pra09]. 0-262-52489-9 [Pra09]. 0-300-07417-4 [Rob99]. 0-309-02329-7 [Bru77]. 0-309-08619-1 [Del05]. 0-333-23758-7 [Hol81b]. 0-333-23759-5 [Hol81b]. 0-335-00041-X [Mos79]. 0-375 [Mir07]. 0-375-50846-5 [Mir07]. 0-393-06435-2 [Hol81b]. 0-394-46631-4 [Mos79]. 0-415-094755 [Bro95]. 0-415-21358-4 [Del02]. 0-415-92607-6 [Del02]. 0-415-92696-3 [Del02]. 0-435-54060-2 [Mos79]. 0-444-99037-2 [Wri80]. 0-471-23273-4 [GM78]. 0-521-25109-5 [Alb95]. 0-521-27205-X [Alb95]. 0-521-60265-3 [Rei06]. 0-521-62299-9 [Del05]. 0-521-62672-2 [Del05]. 0-521-84177-1 [Rei06]. 0-674-00456-6 [Del05]. 0-674-01108-2 [Del05]. 0-674-02499-0 [Goo10]. 0-691-05265-4 [Hol81b]. 0-691-10063-2 [Hol81b]. 0-7100-8307-6 [Blo77]. 0-7129-0891-9 [Hol81b]. 0-7450-0416-4 [Rip91]. 0-7450-0787-2 [Rip91]. 0-801-8910-0 [Sis10]. 0-8018-8477-2 [Sis10]. 0-8039-0958-6 [Nev80]. 0-8039-0959-4 [Nev80]. 0-8047-3971-4 [Mae07]. 0-8047-3972-2 [Mae07]. 0-8047-4463-7 [Col07]. 0-8047-4464-5 [Col07]. 0-8078-2844-0 [Mir07]. 0-8223-3659-6 [Eps09]. 0-8223-3671-5 [Eps09]. 0-8419-0837-0 [Wes88]. 0-85345-438-8 [Wri80]. 0-85520-248-3 [Wri80]. 0-85952-013-7 [Out76]. 0-85952-014-5 [Out76]. 0-88202-032-3 [AB80]. 0-88202-046-3 [Mos79]. 0-88202-116-8 [Wri80]. 0-88202-161-3 [AB80]. 0-88936-174-6 [AB80]. 0-89158-450-1 [Hol81b]. 0-904425-06-1 [Nev80]. 0-906450-02-0 [O’R83]. 094755 [Bro95].

1 [Zie88]. 1-57607-230-4 [Del05]. 1-57607-742-X [Del05]. 1000-page [Ste12]. 115 [Har83a]. 15.50/18.22/$24.95 [Cri10]. 15.99/$22.95 [Eps09]. 1600 [Mos79]. 17th [GHR00, Muk07b]. 17th-Century [GHR00, Muk07b]. 18 [Nur03]. 1844 [Mos79]. 1890 [Hom90]. 19 [And21, CLCMW23, Did25, EMSL23, Sis20a]. 1914’ [Ano77d]. 1920 [AB80]. 1920s [Waz13]. 1930 [Kuk94]. 1930s [AA82, Bak17]. 1933 [Wei94b]. 1935 [Hol81b]. 1939’ [Ano77k]. 1940 [Hol81b]. 1941 [Hol81b]. 1950s [Ger01, GT00, McC88]. 1960s [Bak17, GT00, Sol01a]. 1970/71 [Roc74]. 1970s [FOC^{+}71, dSP71]. 1978 [Ano79, Opp79, Rip79]. 1980s [SM93a]. 1990s [LPKL12]. 1999 [Jan18]. 19th [Bay83].

2 [Hor25b, Hsu13]. 2.10 [Pri71]. 2000 [Lyn01b, Mil04a]. 2000s [LPKL12]. 20th [Avi15, Met06, Nur03, Som06a]. 20th-century [Avi15, Met06, Nur03, Som06a]. 28 [Pfe71].

3 [AB80, Rei06]. 3.75 [Jah72]. 33kop [Rab76]. 376pp [Del02].

4 [Bug93, Mos79, Rob99]. 41483-5 [Mir07]. 46.50/54.68/$75.00 [Cri10]. 4NR [O’R83]. 4S [Edg03].

6 [Blo77].

7 [Bar79].

8 [Wri80].

9 [Hol81b]. 9-11 [Edg03]. 90-279-7743-7 [GM78]. 978 [Hsu13]. 978-0-203-89194-0 [Hyy09]. 978-0-226-11380-7 [Doi11]. 978-0-226-67298-4 [Doi11]. 978-0-262-01324-6 [Moo10]. 978-0-262-51304-3 [Moo10]. 978-0-415-40397-9 [Hyy09]. 978-0-674-03266-8 [Fis10, Ros10]. 978-0-691-12755-2 [Hsu13]. 978-0-691-13741-4 [Gre10]. 978-0-691-13940-1 [Gre10]. 978-0-8166-5045-3 [Cri10]. 978-0-8166-5046-0 [Cri10]. 99056-9 [Wri80].

Abandoning [Mer98a]. Abatement [Bij01]. ABC [Del05]. Ability [Abr91]. able [Fuj87, MS14a]. ablowin [Hin14]. Aboriginal [Ver02]. Aborted [RH90]. Abortion [Sol92]. abortions [BJ12]. absence [Kna13]. Absolute [Tur91]. Absolutely [Mil04b]. Abstracting [Har02a]. Abstracts [Ano77a, Ano77b]. Absurdum} [Mey74]. Abundance [Ala09b, PH14]. Abyssal [Sch15]. academia [Foc16]. Academic [Bla76, Bru77, Fis10, GS02, Han86, HR87, Lam10a, Mir95, dSP71, Rek14, Ros10, Rou17, Sis19, TS24, Del21, FFB11, Ham19, MN23, SO17, VI17]. Academics [Don86, Rif72, Sok23]. Academy [Blo77, Bru77, CG89, Lyn95b, Ma19, Sim95, Wol95, Can75]. Accelerators [Hod83, IM85]. Acceptance [Sch90]. Accepting [Jeo19]. Access [Pic19, VFWC15, WH21]. accident [Ial21]. accidental [Sis20b]. Accidents [Wil88, MB22]. Accommodation [McC92]. According [Day08]. Account [Lat88, Dav24, Kan21, UB25]. accountabilities [Goo18]. Accountability [BS98, AB19, AABF10, Hoe14, HBP19, Ken15, LR20, Pic19]. accountable [Coo19, Hog19]. Accounting [De 24, Den87, Hog19, LC13, Por92b, Por96, RS89a, Ken15]. Accounts [Hen90, Woo76, BL13, May19, Sah12]. accrued [KHM10]. accuser [PBEW24]. Achievement [Ste78b]. Achieving [May19, TB97]. acquire [US20]. Across [BPBG06, CK05, Shi05, DS20]. Act [Jan18, Mul84, Ber08]. Acting [LTC18]. Action [Blo91a, Bou05, Cam85, Mar97, MMB09, Sis12b, TB97, Bar20, Gin10, KW15, RCB18, RIA^{+}11]. Active [GP09, Rit16, Swa24]. Activism [DH08, Eps97, WHBM02, Pop21, Tir18]. Activist [Rob96]. Activity [Gor90, Lyn11a]. Actor [Akr23, CE14, Gar97, Lag12, Say14, SM93b]. Actor-Network [Say14, Lag12]. Actor-Networks [Gar97, SM93b]. Actors [GR08, Law77]. Actual [Woo06]. actuando [Lyo18]. actuarial [Sad24]. actuaries [vdH20]. Ad [Lyn11a, Mey74, Sis16b]. Adam [Rei06]. Adaptation [GB09]. Addendum [KC93]. Addiction [Hsu13]. Address [Har01, Har02a]. Addressing [Lau25]. Adequate [Lev22]. adhaerens} [Dav20]. adherence [Dav20]. adhesion [BM14]. Adjudicating [DU15]. Adjunct [CM75]. Adjustment [SL91]. Administrative [SS17b]. Administrator [O’R83]. Admixture [Ful08]. Adolescent [dG75]. adoptive [CGP25]. Advance [Win76]. Advancement [Ora71]. Advances [Ake07, KC82a]. Advancing [Smi14]. advantage [SF19]. Adverse [Cra10]. adversity [Tol19]. Advice [Bri09, CD84, vEG91, Dem24, Eva22, MB25a]. Advisers [GKM75, LSS20]. Advisory [BR79, Mul72]. advocacy [Hor25b, Naf24, Pan11]. Aerodynamics [VB03]. Aesthetic [CS17]. Afar [Col94b]. Affair [Dol75, FLE97, Wil73]. Affect [Oik17, Hug24]. Affective [Lor08, HJ23, Mic20, Oik17]. Affects [SH19]. affirmative [KW15]. afraid [SAC^{+}20]. Africa [AB91, HHW^{+}19, Hec02, Pri12, RELHR79]. African [Cha22, Cra10, Egl97, Nel08, Pol14]. After [Bud78, Edg03, Jas00, Mil84, Sim99a, Sim07a, vW23, Bie18, Doe03, Lyn11d, SG73, SH12, Tre13, Wil91]. Afterlife [Goo10]. Aftermath [Hil07, Sch15]. Afterword [Bar20]. Against [Sto08, KW15, Max17]. Age [Hec02, Kim07, Mar10, Mes88, Oro77b, Ste78b, Ste90, Bol18, Gum20, JS17, Ken13, Mes17, RG16, Bij01, Rie15, SAC^{+}20, Wil91]. agencies [Hoo17, Oud15]. Agency [Cus96, LR01, PKS88, Pre06, Bel20, Ell11, Say14, Sor23, Tak24]. Agenda [Hag99]. Agendas [WHBM02]. agent [Car21a, Swa24]. agential [UB25]. aggressive [HJ13]. agility [Var20]. agnotology [Pin15]. Agonistes [Mir99]. agree [MN23]. Agreement [Div90, Mal98, vW23]. Agri [WG06]. Agri-biotechnologies [WG06]. Agricultural [Ano77d, Bar12, Kim83, Ros77, She06, Shr00, Tod92, VK84, Vel90, Pih25]. Agriculture [Fit86, Sch72a, Sah25]. agrochemicals [Dur20]. AI [CWvdVdB23, Col92a, Dah24, HB23, Jat23, JS23]. AI-Vey [Col92a]. AI/human [Dah24]. Aid [Rib07, Vin91]. AIDS [Eps97, Mar96b, Mar10, RP15a]. air [Cal18, Eri72, Hin14, Law20a, Naf24]. Airmail [LT92]. Airport [Kom22]. airs [Cal18]. airspace [Law20a]. Akademgorodok [For99]. Al [Mye91]. Alarm [ML93]. Alaska [PT18]. Alaskan [Lan20]. Aldermaston [Spi97]. Aldine [GM78]. Aldrin [GEJ79]. Aldrin/Dieldrin [GEJ79]. Alfred [Mos79]. Algebraic [Coh86]. Algorithm [Hel98, AB19, Ens12]. Algorithmic [Sta18, MB22, Sor23]. algorithms [Jat17]. Alien [Hed06, PH14]. alienate [WC18]. Aliens [Col99c]. Aligning [Eng20, VI17]. Alignment [Fuj87, Nur03]. Allan [Ano77c]. Alleged [Woo86]. Allies [Ano77e, Bea77]. Alliteration [Ros96a]. Allocations [Orl73]. Alphabetical [Rud77]. AlphaGo [Sor23]. already [Wyl15]. Alter [Pap11, RC16]. Alter-ontologies [Pap11]. alter-standardization [RC16]. Alternated [Col04b]. Alternates [MS00]. Alternative [Ada02, CM75, Kim07, Red75, Ver02, Mos16]. altitude [Heg16]. Always [Ash96, OL18]. Alzheimer [Hed06]. Amani [GK16a]. Amateur [Bru15]. Amateurs [Lan81, Rot81, SG89]. Amazon [MR17]. Amazonian [She20]. Ambien [ZDI^{+}23]. Ambiguities [Rap01]. Ambiguity [PBEW24, RB22]. Ambivalence [FLK^{+}14, Mul95b, SM93b, WC99]. America [Bri14, Cue90, Dun91, Ink91, Kra89, Mis85, Mos79, Ras04, Rem88, Rie15, Ves87, WLBRS15]. American [Ano77d, Ano77l, Blo77, Ler18, SS91, AS99, Bij07, Bij01, BS01, Bud78, CFS01, Den87, Ell75, Ezr71, Fit86, FG73, Ger01, Gre04, Kim83, Kop79, LT92, Mil04a, Nel77, Ros77, Ros81, Rot81, ST16, Tho99, Tor83, Whi75b]. Americans [Sah12]. Amerikanische [Pfe71]. Amnesia [Bar85a]. Among [Cam85, Rif72, Tür73, AB91, Fox83, Fox05, FFB11, Gra07, Gra98, Lau13, Law13, Mar20, MRL01, PD19, Sok23, Son95]. Amorphous [Pra07]. Amsterdam [Wri80, Rip79]. Amsterdamska [Blu09]. Anaemia [ML94]. Analogies [Pic80a]. analogue [AH11]. Analyses [SWB77a, DSJ19, Mat13]. Analysis [AQL94, BH97, BR79, Bri72, CKM04, CM75, Coz85, DC76, FG86, FO86, GHR00, Hal88, Hic87, HP91, Kra72, Len79, Lew80, Lyn85, Mar97, MC79, MG82a, MG82b, Nad80, PN87, Pin85, Por77, Pot87, Pra05, Ros96a, Ruf79, Sco92, Smi09, SR77, Vin91, Whi89, Woo89, Wri80, Yea99, Bir23, JL18, Man21, Oik17, Sha20, VFWC15, Hic88, Gin10, Ano77n]. Analysis/ [Hic88]. analyst [CE14]. analytic [dLDV21]. Analyzing [CLP91, CS20, Mar88]. Anatomy [Bro82, Mül18, Pra07, Pre05, Sør93, Tom87, Fuk16]. Ancestry [Ham82, Nel08, FR11, PD19, SMW15, Tsa24]. Anchoring [vdSvESW98, Rou17]. Ancien [Ald98]. Ancient [Wat14]. Andre [Mos79]. Android [Ala09a]. Angell [Mir07, Mir07]. Angels [Mia99]. Angiography [Mor00]. angst [Dam25, WZH^{+}25]. Animal [Asd08, Cri96, FC12, Lyn88, MB94, Bra24, DA24, Nel13]. animals [Suz21]. Ann [Del05]. Annihilation [Gus06]. Announcement [Ano76]. Anomalies [Blo85, Wes77, Wes78, Ras20]. Answer [Sha04]. Answers [MM84b]. ANT [Lag12, Tre13]. Antarctic [OOO12, DNTS08, Heg16]. Antebellum [Ell75]. Antecedents [Arm19]. Anthropo [Lor17]. Anthropo-scene [Lor17]. Anthropocene [Dam25, WZH^{+}25]. Anthropologists [Bro95]. Anthropology [For94, Bro95]. Anthropophagy [da 22]. Anti [EM06, lL99, RS11, MD20]. Anti-bureaucratic [RS11]. anti-influenza [MD20]. Anti-Malarial [lL99]. Anti-social [EM06]. Antibody [CKM04]. Anticipation [Uru22]. anticipatory [Gus14, Sad24]. Antidepressant [AD09, Lak04]. antiprograms [KH18]. Antisepsis [Lyn94a]. Anxieties [Lak04]. Anxiety [BV95, Lee16, SH19]. anxious [Lee16]. Any [Mar04]. Anyone [Mod01, FW17]. Anything [Rus83]. Anyway [Kai06]. Apart [MG82a, Sim07b, Woo06]. Apartheid [Cha22]. Apologetic [AA84]. Appealing [Car05, Joy05]. appearance [SMW15]. Applbaum [NK14b]. Applicable [Bri72]. Applicants [Jab95]. Application [AdSPG^{+}76, MCBDT00, MK84, MPA87, Ste86]. Applied [Ano77e, Bea77, Bre97, Don86, LC99, MK84, WG74, Hof15]. Appraising [BHC16]. Appreciating [dLDV21]. Appreciation [Mac84]. Apprehensions [LM75]. Approach [Ano77o, Blo75, CT89, Cri96, EES96, Har76, Har77, Lon71, MG92, SS94, Twe93, Kru13, Blo74]. Approaches [Gor89b, Rad92, SBHY15]. Appropriation [MG92, Has12]. April [Lyn11c, PCL98]. Arbeiterbewegung [Bay83]. Arbitrary [Ale08]. Archaeology [Jah72]. Archeology [Kim14]. Architectural [Yan05]. architecture [Skr16, Yan22]. Architectures [Par05]. Archives [Muk07a, CS20, Del16]. Archiving [Gal24]. Arctic [Lan20, Mar08b, Var20]. Area [Bro89]. Areas [DYS^{+}05, HMH80]. Arena [Sun07, Gre16a]. Arenas [KC82c]. Argentina [Lak04]. Argos [Ben12]. Argument [GHR00, Her07, Sto08]. Argumentation [Yea85]. Aristotle [dV07]. Arithmetic [Mac93]. Arms [Bar03b, Hol81a]. arousal [WE15]. ARPA [Gui98]. arranged [PW19]. arrangements [AMT11]. Array [Sch72b]. Arriving [Fis10]. Arrowsmith [Bru77]. Art [Hor04, Law97, MM84a, PW95, Waj00, Woo82, Ano16b, Bor17, CS17, Cho13, Hor25a, Pel21, Sok23]. Artefacts [Elz86, Joe99a, Pet00, PB84, Rus86, Str11, WC99, CB24]. Arthur [Mir07]. Arti [Sim99a]. Article [Gor90, Car21a, KBvLA22, SS17b]. Articles [Baz84, HSP02, Mye85, Fro12]. Articulating [Fuj87, Sil11]. Articulation [Pre05, Gra16]. articulations [Tal13]. Artifacts [Ros96a]. Artificial [Bra89, CL89, Dre92, For93, Gor89a, Sle92, vHVdW07, Dav24, Ens12]. artisanal [PH14]. artist [Wyl15]. Arts [Hud73, Ell11]. arXiv [Rit16, Sis16b]. Ashby [Pic02]. Asia [AB80, Sne12]. Aspect [Lan81]. Aspects [Ano77k, Blo77, Bow87, Lew72, Mos77, RB90b, Wil73, Tra81]. Assault [Jac78]. assemblage [Lak17]. assemble [US20]. Assembling [Law20a, dlB11]. assembly [Sab14]. Assess [Lev79, TY85]. Assessing [IM83b, Mor86a]. Assessment [Abr93, AS99, Ano79, GEJ79, Ham19, Joe94, Mor86a, dMdF98, Gar13, LOBB23, Lak17]. Assessorship [Chu75]. assetization [VT21]. assets [MB25b]. assignment [Sud22]. Assimilation [Gas90, Win76]. Assisted [CKM04]. Associate [Nel22]. Association [Bro95, Ora71, Kim83, Roc81, Whi75a]. Assumption [Koe99, Mac99c]. assurance [KH19]. Astbury [Olb84]. Asteroid [Mel07]. Asteroids [Cle86]. Astrology [Wri75]. astronomical [Hoe14]. Astronomy [GM78, McC00, Rot81, Tat86, GM78]. Astrophysics [Lan81, Sun11]. Asymmetric [MS00]. asymmetries [VT21]. at/on [Old90b]. athlete [Joh13]. Atlantis [For99, God98]. Atmospheres [Dav21]. Atmospheric [Buc92]. Atom [Sch06, RA22]. Atomic [Hol81a, Tho99, Zie88]. Atomism [Jac78]. Atoms [GP09]. attachments [Sat21, TS21, dLDV21]. Attack [Mar96a]. Attention [Gre04]. Attitude [Col82b, McD76]. Attitudes [LM75, Tor83, Tua71, Zin71]. attributes [Del23]. Attribution [Smi85, SL91]. Attributional [Bra89]. attuned [Cal18]. Audiences [Col99c, Ros11]. audio [Mar20]. Audiophilia [Per04]. audits [Mat14]. August [Bre02, HDS^{+}98, Jan18, Rip79]. austerity [Rei21]. Australia [Dun85, Ron78]. Australian [Hil74, Hin14, Ink85, Lan01a, MK84]. Austrian [FFMW08, Foc16]. Authentic [Dam25, WZH^{+}25]. Author [Rud77, STGG95]. Authored [LMA80]. Authoritative [Joy05, Kan21]. Authority [Cas09, Kuk94, Mah80, Per04, Tho02, Tou91, Kan21, Tak24]. Authors [Gor84, SR72]. Authorship [Hef79, Lin80, Bia22, Del16]. Autism [DU15, Kir12, Lap16, Met25]. Autogiro [Mod00]. Automated [Mac17]. Automatic [Vos04]. Automating [Har96, Hyy06, RG16, Suc23]. Automation [BS08, Jon17a, LB19]. Automotive [Bra92a, Leo10]. autonomou [TS21]. Autonomous [LC99]. Autonomy [Kri78, Sut84, Win02a, Dah24]. average [Jon23]. Aversion [JLM73]. Aviation [LT92, Mil76, Mod00]. Avocation [Mos79]. avoided [Tut11]. Avoiding [Pyc93]. Avorn [Mir07, Mir07]. Awards [LPKL12]. Axes [Blo98a].

Babies [Tim03]. Back [CS20, Ful89a, Mul84, GS11, HH22, HLRK14]. Backcasting [Blo85]. Background [Ora71]. backup [Cur22]. Baconian [Cur94]. Bacteriology [Ams87]. Bad [Mas10, SS99a, Sto08, IJJ13, SG21]. bag [WL13]. Bailes [Hol81b]. Balance [Ale08]. Balint [Osb93]. ballet [Mül18]. Ballistic [MS88a, MS88b]. Ballot [Lyn01b]. Bank [Gar73b, SAG11]. banking [Cur22, MS14b]. Banneker [Egl97]. Banu [Del02]. Barad [Pin11, HFGP17]. Barbara [Del05]. Barber [RD08]. Barbercheck [Del02]. Barcelona [CGPP24]. Bare [FS08]. Barkla [Wyn76]. Barnes [Ano77q, Luk75, Nev80, Mer98a, Mer98b, RH85]. Barred [Dow07b]. Barrel [Clo01]. Barrels [Gra98]. Barrett [Pin90]. Barrier [Rib07]. Barry [Ano77q, Luk75, Nev80]. Bars [PK09]. Bartels [AA84]. Bartsch [Del02]. based [Ben12, Bor17, CGP25, CC12, Hog19, Kna13, LSW24, Lau13]. Basement [Mir07]. Basic [IM83b, IM85, LC99, Hof15]. basis [SBHY15]. batteries [WN22]. battle [Joh18]. battles [Kir12]. Baümler [Wes88]. Bay [PT18]. Bayesian [Kru13]. Bayh [Ber08]. Bazerman [Gor90]. BBC [Kan21]. BBMRI [Aar23]. BBMRI-ERIC [Aar23]. Be [Ash96, BP80, Dis04, Joe99b, SK13, Bow93, Ful94b, Mar13, Mia12, Tut11]. Beach [CE03]. Beaches [Bed91]. beagles [Bol18]. Beam [Edg97a]. Beams [Del16]. Beardsley [Ano77e]. Beautiful [Fau92]. beauty [Lee16]. Became [LP11, Mis85, Shr05, Pic13]. Become [LR01, Son09, Dem24]. becomes [Abe13]. Becoming [AGKL^{+}15, Ves96]. bedside [Tim15]. bee [Dur20, SK13]. bee-toxic [Dur20]. Been [Opp79, Ler18]. Beer [Pic02]. Beginnings [Now87, AB80]. Behalf [Ano84]. behavior [Nel13, Pet15]. Behaviour [Cla80, Gar73b, Gor84, Lau90]. Behaviours [Arm09, Arm10, Val10]. Behind [Cas09, PK09]. Being [Ful94a, Kna13, Ste99, Swa89, God15, Ken13, MS14a, PC12, Sah12, Vis15]. Belarus [Nes95]. Belgrade [Jan04]. Belief [Bar79, GM82, Jac94]. Beliefs [LM89]. Believe [Eva99]. bells [JS17]. belonging [RP16]. Below [Wei90]. Ben [FL88, Man93, Wes86]. Ben-David [FL88, Man93, Wes86]. bench [Par25, Tim15]. benchmark [Jat23]. Bendectin [EM00]. Benefit [Hay07, PB84, Sha20]. Benefit-sharing [Hay07]. Benefits [Mir07, Pyc93, RS11]. Benevolent [Jön16]. Benjamin [Egl97].

Benoxaprofen [Abr93]. Benthic [Law81]. Benveniste [FLE97, Pic94]. Berkeley [SG89]. Berlin [Meh80, Mos16]. Berman [Mos79]. Bernard [RD08]. Bertelsmann [Pfe71]. better [Vas15]. Bettie [Pih25]. Between [AH11, Anf21, Ano77e, Bea77, Kwa87, Mel03, Pal02, Rad86, Rii08, Sla07, Ste78a, Wag09, Woo86, dB09, vHVdW07, AD09, Bor17, CERG25a, Dro93, Dug87, Foc16, GKSS24, Gus99, Han86, Ker00, Mat05, MvO23, Mor19, Mül95c, Pih25, Rit16, Smi85, SB18, WC16]. Beverly [Nev80]. Beyond [AA84, Ala20a, Ala20b, Eli00, Foc16, Gor89b, Jas96, Jen07, Moo10, Rem88, RA96, Ure25, AH11, Pri12, SDL^{+}14, SSB24]. Bhadralok [RH96]. Bhaskar [Out76]. Biases [MRL01]. Bibliographical [MC71]. Bibliometric [Ano77n, Hic88, SR77]. Bicycle [Hul92, AQK15, Wie17]. BiDil [Kah08]. Bielefeld [Ano79]. Big [Gre22, KP85, Rem88, Sch94, Fuk16, Met25, HJ23]. Bigelow [Bro89]. biggest [Joh18]. Bijker [Rus86]. Bikes [Ros93a]. Bilingual [ER78]. billed [Hun24]. bin [WL13]. Binding [Rem88]. Bio [Nel08, BFW14]. bio-cultural [BFW14]. biobanking [Aar23, Met25]. Biobanks [STM^{+}08]. biocapital [Rah11]. Biochemists [SM97]. biocultural [Bra14]. Biodiversity [Bow00, Wes24]. Bioethics [Gra16, Hed10, Reu12]. biofuels [WM17]. biographical [Wir15]. Biography [Hyy09, Lat13, Mos79]. Biohierarchies [SFP11]. bioinformatics [SS17a]. Biolegality [Law13]. Biological [AA82, Ams87, JL98, Kwa87, Rad13, Ram02, Leo12, Pap21]. Biologism [Ski06]. Biologistical [Ful08]. Biologists [SM97, Son95]. Biology [AA82, Bar84b, Bix97, Fan09, Fue82, Fue84, Gas90, Mye85, Sin96a, Win76, Wri86, You71, Cal21, Fuk16, Tol19, ZDI^{+}23, Yox84]. biomarkers [Chi24]. Biomedical [BB17, CKSW09, FFMW08, KLBKC11, Lan20, Lev79, STM^{+}08, Avi15, Dav20, FR11, GDLB15, Mur13, Reu12, Sis10, TL18]. Biomedicine [CKM04, FFMW08, RB09, Stu07, Pap21, Str11, TT18]. Biometric [Far75, Kan22]. biomimetic [Hol22]. bionetworking [SFP11]. Biopharming [Mil12]. biopolitical [Bak12, Sis10]. Biopolitics [CSTG23, MBDR24]. Bioscience [Hay07]. biosocial [Chi24]. biosovereignty [vW23]. Biosphere [Kwa05, Yea90]. Biotech [Fau92, Swa07, Tut11, Ree06]. biotechnologies [WG06]. Biotechnology [Bud91, CK85, CLP90, Rah11, Foc16, Twi05]. biovalue [Kow13]. Bioweapons [Vog06]. bird [Lak17]. Birgit [Col07]. Birth [RH83, TL00]. bisexual [Gre16a]. Bit [Pic91]. Black [Lan09, Pin92, Tim03, Yon94, vK01, SvD25]. bleak [Gar73a]. Blend [Lau82]. Blind [Ash93, Del87, Pic99, ST97]. blockchain [MB25b]. Blockmodeling [Len79]. blocks [Edm11]. blogosphere [Rit16]. blogs [Sil11]. Blood [Har96, Mar10, Mor06, BB17, Tan21b]. Bloomington [AB80]. Bloor [Kim94, DeW75, Kus04b, Mer98a, Mer98b, Mey74]. Blue [Tim03, dB09]. Blurred [Sol01b]. Blurring [Pic97]. Board [Bru77, RL21]. Bobby [Rap01]. Bodies [GP09, Hir91, Mia99, MC97, Pre05, Sha01, Ste06, Wal00, FC12, Lee16, Mül18]. Bodily [Bie18, WE15]. Body [Ala09a, Asd08, Car14, Lyn88, Mar04, Mye08, Ala23, Kan22, KRR13, Lap16, Swa24, WLBRS15]. body-environment [Ala23]. Body-in-Interaction [Ala09a]. Body-work [Mye08]. Bog [Yea89a]. Boghossian [Sto08]. Bohemia [Gal24]. Bolts [Fau07]. Bomb [Hol81a, Ral06, Wel04, Zie88, RA22]. Bombay [JS18]. bonding [RG14]. Bones [FS08, Goo10]. Book [Ano77d, Ano77g, Ano77e, Ano77f, Ano77i, Ano77h, Ano77j, Ano77k, Ano88a, Ano88b, Ano89, Ano90a, Ano90b, Ano91a, Ano91b, Ano91c, Ano92, Ash05, Atk05, Bar84a, Bar90, Blo77, Bra04, Car05, Cor94, Dea95, Del05, Elz04, Fis10, For99, Ful94b, Ger95, Gie06a, Gor90, Hag94, Har94b, Hom90, Irw94, Kim94, Kuk94, Lau90, Mor94, OS05a, Old88, Pal94, Ree06, Rei06, Ros10, Rus94, Sch94, Ste90, Twe93, Twi05, Ves96, Wei90, Wei94b, Wes88, Lat13, MS14a]. Books [Day08, Dea95, Mel03, Out76, Ano07c]. Booming [BL87]. border [Hof15, LB19, Pel21]. Borderline [Mül95c]. Borders [KK03]. Boring [SS91]. Boris [Gra85]. Borrowing [Gra75]. Bose [Nan72]. Boston [Blo77, Luk75, dSP75]. Both [RS89b]. Bottle [Geu03]. Boulder [Hol81b]. Boule [Ham82]. Bound [Hyy06]. Boundaries [BV95, BH03, CK05, Jas87, Lam10a, Mel09, PW95, Pic97, Smy01, Sol01b, VB03, Heg16, NM20]. Boundary [Bur08, FLE97, Gus99, Har03, Ma19, PC12, RG16, SG89, Sun07, Bed14, Law13, Rot19]. Boundary-work [RG16]. Bounding [KR11]. Bourdieu [Bre02, Gin10, Mia03]. Bowker [Pra09]. Bowler [Old88]. Box [Lyn01b, SvD25, Car01]. Boxes [Pin92, Yon94]. Boxing [Ric96]. Boycotting [Sla07]. Boyle [Jac78, Sha84]. Boys [CE03]. Brain [Ala08, Bea01, BH91, Gar13, Hol22, Pet17, SZ11, SC18, O’R83]. Brains [Mor86b]. brainwaves [Mar24]. Branding [Bar23]. Brandon [DGC18]. brands [Bia22]. Brasilis [KGDLB^{+}15]. Brazil [CL11, KW15, TL11, VP98, da 22, dMdF98]. Brazilian [AB80, Bot90, Hes87, VK84, Vel90]. BRCA [Bou05]. breach [Ial21]. breakdown [FM17]. breakdowns [Mar17]. Breaking [BS04a, Jas03a]. Breaks [Tur01b]. Breast [Jas02, Par05]. breastmilk [Car14]. Breeder [Kim83]. Breeding [Mar08a]. brew [AABF10]. Brian [Bar90]. Bricolage [Mac03c]. Bridge [Dug87, Pow01]. Bridges [WC99]. Bridging [SAG11]. Brigham [SS91]. Bringing [Ala09b, Rob16]. Bristol [PT18]. Britain [Blo77, GM78, GEJ79, Mac76, Mar08a, McC88, Mor86b, O’R83, Pat83, Wri80, Wei90]. British [Ano77k, Nev80, O’R83, AS99, FG86, HH83, Kir99, Lie87, LW75, Mil84, Mos77, OM76, Ora71, Reu12, San72, Spi97, Wer71, Whi75a, Win02a, Win02b, vdH20]. Broke [MS88b]. Broken [Tou13]. brokerage [SF19]. Brown [Moo10]. Brunel [Buc92]. Bruno [Akr23, Har23, Mar23, NTW^{+}23, Por23, Rii08, Sis22a, Tre13]. Bryant [Bro89]. Bryder [Wei90]. BSHS [O’R83]. Bucks [Mos79, Nev80, O’R83]. budget [Lah21]. Bugs [Van08, Sch15]. Build [Hol81a]. Building [AB80, CH75, Eps09, KGDLB^{+}15, McC92, Muk07b, Ber08, Edm11, Tsa24]. Built [McK99, Ral06]. Bulgaria [KS79]. Bulgarian [Sim95]. Bump [Rii08]. Bureau [Bro89]. bureaucratic [RS11]. bureaucratized [FW23]. Burnet [JL72]. Bush [Red16, dLM00]. bushfire [NM20]. business [Foc16, PW10]. Busted [Joe99b]. Butterfly [Low89]. Button [CE14]. Bynum [Alb95].

C [Ano77i, Del05, HG88, Mar05, Pra09, Ric88]. C. [Mar05, Wyn76]. CA [Col07, Del05, FM17]. Calculation [Blo91a, Hoe19a]. Calibration [Hes00]. Calicivirus [Lan01a]. Calif [Bru77, Nev80]. California [Ano16a, Car12a, Phi16, Sha16, Wak04]. Call [Nel22]. Called [Mac03b, Vic95]. calling [God15]. Cambodia [Gra16]. Cambridge [Alb95, Del05, Fis10, Goo10, Moo10, Pra09, Pri71, Rei06, Ros10, Gas84, Ste90]. Cambrosio [Kle91]. Came [Col96a, Par22]. Camelot [Sol01a]. cameras [Kim16]. Campaigns [Bij01, Hor25b]. Campanario [Nis95]. Campbell [Ros96b]. Campus [Blo77]. Can [BP80, Cri04, Dol75, Gor89b, Jas00, Lau25, Sha01, Skr16, TY85, Vas15, CS17, FW17, Nav11, Ful94b]. Canada [Cur98, Tur01a]. Canadian [Gar73a, LOBB23]. Canal [Car12b]. Cancer [Bou05, Bud78, CC98, Fuj87, KLBKC11, MRL01, Pal02, Par05, RB09, Ric88, Avi15, CGP25, Jat23, KBRB19, SA22, Swa24, TL11]. Cannibalism [Kid88]. Cannot [Joe99b]. Can’t [Mar04, Rap01, RS89b]. Canute [CE03]. capacity [Reu12]. Capital [Bur08, Hon08, Bir23, KHM10]. Capitalism [Wri80]. Capitalist [Ink91]. Caprices [CH75]. carbon [FO22, Lah21, LR20, SBH^{+}20]. Carcinogenic [Abr93, GEJ79]. Card [God97]. cardiac [JS18]. Cardiovascular [Shi05]. cardioverter [Oud15]. Care [AGKL^{+}15, DH15, KH19, Car14, Che23, DPP^{+}20, DA24, Eng20, Hog19, Lee16, Lip23, LB19, MMV15, Mur15, Sch15, ST24, Suz21, SH19, Wat14, dlB11, dlB15, Wri80]. Career [Ano77p, Bur92, Oro77a, Ruf79, Zin76, dC09]. Careers [FS01, HR87, Her07, Rud77, Zin71, Mos16]. careful [DG17]. Caring [Lip23, PPM20, Vis15, Goo18]. Carll [Blo77]. Carlson [Lat92]. Carolina [CC12, Mir07]. Carpenter [Ano77h]. cars [Sti18]. Carving [Ram02, Wyn92a]. cascades [Jan18]. Case [Abr91, Bar85a, BS08, Bud78, Bud85, CK83, CG90, Chu82, CC77, Cri04, Cue90, Del83, Der00, Dun78, Dun91, ER78, Eps97, Far75, FS93, FL88, Gap95, Gar97, GK72, GEJ79, Gor82, Har81b, Hol74, IM83b, Joe94, KSW82, Kwa87, Law73, Mac78b, Mag77, Mar96b, Mar97, Mat05, McC92, Mes88, MB94, Mie98, Mul73, Nes95, Par88, Pic81, Pra07, Pre06, Rea01, RH83, Ruf79, Sha04, Sho05, SS94, Ste78b, SS91, Sut84, Tat86, Tho02, Tur90, Tur87, VB03, WD97, Wes77, Wes78, WSB79, Wyn76, Yea99, dMdF98, vdSvESW98, Bra14, FS23, KHM10, yLL22, iO18, Seg18, SC18, WM17]. Case-Studies [Mag77]. Case-Study [Har81b, Mac78b, Ste78b, SS91, Yea99]. Cases [Mia99, Stu07]. casino [Joh18]. Casting [Sis17a]. Cat [Bru80, Cri10]. Catalogue [dB72]. catchment [CKFS13]. Categorical [Met25]. Categories [Sch13, Nav11]. Category [Kel89, RS89b]. Cathcart [SN89]. Cathedral [Mac90]. Catherine [Gre10]. Cattell [God07]. Caught [Pin99c]. Causal [Hut91, Smi85, WC16]. Cause [Eva99]. caused [Shr14]. Causes [Sle89a]. Caveat [Bud85]. Caveats [Por77]. Celebrating [Sch06]. Celebration [Edg03]. Cell [Pra06, SAG11, SK08, CGP25, RC16, SFP11, SFH12]. Cells [LP11, MBDR24]. Cellulose [Mie98]. Cellulose-Degrading [Mie98]. cements [Sma20]. Census [Lor08, Roc74]. Center [AB80]. centered [HB23, Ros11]. Central [Fue82, Bal95, BFS95a, BFS95b, Sch95]. Centrality [Yox84]. Centre [AB80, CK85, Nes95, BB17]. Centre-Periphery [Nes95]. Centres [HvL08]. Century [Bay83, GHR00, Lau84, Mir99, Pal94, Wei90, Avi15, Car18, Don86, SC18, Alb95, Bud91, Dea92, Gas95, Kra89, Met06, Muk07b, Nur03, Som06a, Wri75]. Cerebral [Mor00]. Certain [JZ12, Sch84, GS11]. Certainty [Bug93, Col87, Mye91, Pin81b, Smy01, BM19]. Cervical [CC98, SM93b, TL11]. Cetina [GP97, Sis93a]. Chadarevian [Mae07]. Chagas [Cou99]. Chains [Hol81b]. Chalfont [Nev80, O’R83]. Challenge [RS89a, SG73, Ves96, JS18]. challenges [Ure14]. Chance [Har01, Wis01]. Chandra [Nan72]. Change [Bar81, Bow96, BW83, BH03, EES96, Eva97, Gas90, HV93, Hof79, Irw94, KC95, Man93, Mar08b, Met06, MO09, Mos79, MF95, Mul98, NHM74, Sin96a, Anf21, Lah21, Mat14, Mcl21, Ple14, RP15a, SH17, Var20, OOO12]. Changed [GB09, Shi03]. Changes [Coh86, Jab95, Lyn12a, McD76, Mül95c, SM93a, Sis12a]. Changing [Bea01, Blo98a, Hes86, KBvLA22, Lah21, Mos79, Nad80, Nes95, Ski06, Oud12, Van08]. Changshan [lL99]. chaos [Law11]. Chapel [Mir07]. Character [Joh85]. Characteristics [Fox05]. Charisma [TS00, TS00]. Charles [Gor10, Mos79, Ano77d, Gor90]. Charm [PN87]. Charting [HR87]. Chaussées [Gra07]. cheating [Joh18]. Checking [Gre24]. cheese [PH14]. Chemical [Buc79, Don86, Lyo18, Man21, SN89]. Chemicals [Oud97]. Chemiker [Ril86]. Chemistry [Rab76, Zin71, Zin76, AO78, Mos79, Opp79, Zin71]. Chemists [Ril86, Whi75b, Mos79]. chequered [Mos16]. Chernobyl [Sch15]. chess [Ens12]. Chicago [Ash07, Doi11, GM78, Gin10, Kim94, Sis10, NK14a]. Chick [Dow07a]. Chicken [Oli25]. Chief [Sis22b]. Children [Sol85, Son09, Lau25, Lau13]. Chilean [CSTG23, KVA21]. Chimerical [Bru80]. China [Buc75b, Dea74, Gu23, SFH12]. Chinese [Gre16b, Hon08, lL99, yLL14, yLL22, Ma19]. Choice [JM01, JLM73, Tod87, Vel90, Wag09, Waj86, SZ11, Rip91]. Choices [Sla07]. Cholesterol [Gar97]. Choosing [Har02b, Pel21]. Choreography [Cus96]. Christmas [Bur75]. chromosome [Gri18]. Chromosomes [Mar04]. Chronobiology [CK83]. Chrysler [Bra92a]. Church [JM01]. cigarette [Bel20]. circa [CC98, Ink88, PS77]. Circle [Bur92]. Circulation [O’C93, Bie18]. Circumstance [Sha84]. Citation [CM75, Coz85, DC76, DB85, Hic87, Hic88, HP91, Len79, Lin80, LMA80, Mor88a, Nad80, Nad83, Por77, RELHR79, Ruf79, Sma77, SWB77a, VK84]. Citations [All97, MM75]. Cite [STGG95, EE14]. Cited [Cam93, NFC83, Sma78]. Cities [Tod87]. Citizen [CC12, EP07, BTS22, BKK21, KH18, KH22, PD19]. Citizen-based [CC12]. citizens [MR17, Sim25, Sis10]. citizenship [RSH18]. City [For99, Gie06b, Hom03, MBDR24]. Civic [Mil04a, RS74]. Civil [Blo98b, Ful94a, Ful94b, Kus10, iO18]. Civilian [Har03, Lic89, OL18]. Claiming [Hen18]. Claims [Mye85, PC84, Col14, Tol19]. Clandestine [Clo01]. Clarendon [Blo74, Sto08]. Clash [Yon94]. Class [Ink88, Wri80]. Classic [Mor88a]. Classical [BS04a]. Classification [Mes10, Mor88a, Tay95, Abe13, Ras20]. Classifications [Rot05, Som06b, Gri18]. Classifying [Som06b, Wat02]. Classroom [Old90a, DH23, KLW19]. Classrooms [Dug87]. Claude [Gre10]. Clergy [Hil86]. Climate [Bak17, HV93, Lah05, Mar08b, OOO12, Sun07, vdSvESW98, Gar73a, Lah21, Mah14, Mat14, Skr16, SBHY15, Var20]. climates [Sch16]. clinic [HJC23, SH19, Rei06]. Clinical [AS99, Arm17, Bou05, Eps08, Ker00, RB09, Bri13, Gar13, KLBKC11, Leo12, Pic13, Tim15, Hed14]. Clinics [Cus96, PH11]. Clio [Del05]. Cloning [Pra06]. Close [Rel24]. closed [Sov10]. Closer [Sim78]. Closing [Kan78a, Uru22]. Closure [Bed91, Car09, Hår94a, Shr14]. cloth [Gre10, Out76]. Club [GT77]. cluster [MS24]. Co [Blo77, Con78, CC12, CL89, Hic87, Hic88, Hol81b, Lau13, Len79, Leo10, Nad83, Pap21, Sma77, SWB77a, Tay95, Whi89, Bea10, Eps21, Gu23, Hed14, LSW24, Mag17, Mon12, Sim25]. Co-Citation [Hic87, Hic88, Len79, Nad83, Sma77, SWB77a]. Co-Construction [Tay95, Hed14]. Co-Evolution [Con78, Leo10]. co-location [Bea10]. Co-participation [Lau13]. co-presence [Bea10]. Co-producing [CC12]. Co-production [Pap21, Eps21, Gu23, Mag17, Mon12]. co-productionist [LSW24, Sim25]. Co-Word [CL89, Whi89]. Coast [Lac16]. Coastal [Bij07, Sar23]. Coauthorship [SH89]. Cod [Mer96, DA24]. Code [Wal01]. coercion [Byc16]. Coffeehouse [Woo89]. Cognition [GM03, Gor97, Mag07]. Cognitive [Blo93, CG90, CG92, Cha99, Gie92, GM03, Gor82, Gor89b, MMB09, Nad80, RJ76, SM93a, Tha89]. cohort [KH19]. Coincidence [Sch95]. Col [Hol81b]. Cold [Lew95, Sim99a, Clo03, Den03, Ger01, Hou01, Pop21, SH12, Sne12, Sol01b, vK01]. colder [Heg16]. Coleridge [Blo83]. Coley [Mos79]. Colin [Mos79]. Collaboration [Cas09, DYS^{+}05, FC79, Gor02, Hac05b, Jef03, LR08, LB05, Lyn95a, Mye91, Pre80, Pyc93, Ras04, EMB^{+}11, Eva10, ELR21, GB16, VFWC15, Wad14]. Collaborations [SCG01, Byc16]. Collaborative [CKM04, CK05, Hef79, SVS21, Ste12]. Collagen [Sma77]. collapse [SK13]. Collecting [Hal98]. Collective [Mac03b, MMB09, MvO23, MS24, PAB15, Rou17]. Collectives [Bou05, BTS22]. Collectivities [Dun85]. Collide [Vog06]. Colliding [Lau87, Mel07, Ste86, Ste87]. Collingridge [Joh85]. Collins [Ash07, Doi11, Hol81b, Chu82, Dre92, Fox94, Hir94, Jas03a, Lau82, Lyn94a, Nic84, RH85, Sis17a, Wyn03]. Collisions [Ful89b]. Colombia [Lyo18, SMW15]. Colombian [SMWCSC15]. Colonial [Hec02, Ink85, MD09, LOBB23, Lan20]. colonialism [Kus10]. Colonization [Ala09b]. Colonizing [MPA87]. colony [HTB20, SK13]. colpocytology [TL11]. Colposcopy [TL11]. Combination [Gra75]. Combustion [Tor92]. Comes [Tur01a, Lip23]. Comfort [Del05]. Coming [Irw06, Lat10, SK13]. commensalism [Ala23]. commensurabilities [JS23]. commensurable [Dix19]. Commensuration [OS05b]. Comment [ABB^{+}08, BL94, Col12, Del87, Fre92, Jas03a, JM73, Lyn95a, Lyn04e, McH09, Pic80a, PM89, Sti87, Woo82, Zan84]. Commentary [RDB03, Bij12, Rif72]. Comments [Dal90, Gus06, Hug06, Lyn01b, NKB86, Pfe79, Sim91]. Commerce [Kin02]. commercial [Coo11]. Commercialization [MV05, Ben12, Ran10, SO17]. commercially [Pin15]. Commission [Bru77, SG74a, Sko73]. Commitment [Nad83, RA96, Zin71]. Committee [Bru77, PC84, Pop21, Shr04]. committees [Hed14]. Commodification [Fis04]. commodities [WC16]. Common [McK99, Sha01, DS20, SMWCSC15]. Communicating [Huu12, Kru13]. Communication [Boh04, CS00, CH75, Gar73b, Lew95, MM86, Nad83, Rib07]. Communications [WG74]. Communist [Hol81b, BFS95b]. Communities [AQL94, KC82c, BTS22, KVA21]. Community [Gas92, Gas95, Mir95, Mos79, Sar23, Ves96, Wei94b, FM17]. Companies [Dae98, Mir07, Foc16, WC18]. Company [Ras04, Jac75]. Company-Medical [Ras04]. comparable [Sch16]. Comparative [AB80, BR79, Elz86, Kra89, Lon71, KGDLB^{+}15]. Compare [Mal98]. Compared [Kir99]. Comparing [Coz85, OS05a, Par05, RB09]. Comparison [Cha99, IM85, Sch16, Tor83, Edm11, Jen20]. comparisons [Sun11]. Competing [Yon94, CKFS13]. Competition [Boh04, Cas09, ZJ09]. Competitive [SF19]. Competitiveness [Lic89]. Complacent [AS99]. Complement [Gor89b]. complements [Mos16]. Complex [Eri71, Mat05, TS00, Col13, Law25]. Complexities [JM01]. Complexity [AABF10, Lan00, Chi24, Lev14]. Component [dMdF98]. Composite [Lin78]. Composition [MM81]. Compounds [GP09, RCB18]. compromising [Huu12]. Computation [Tur82, Rel24]. Computer [CKM04, Gie89, Lag07, LC99, Mel09, Pel99, Sim91, Sle89b, Tur82, WB96, Yea99, Jon17a, Lau13, Püt21]. Computer-Assisted [CKM04]. computer-based [Lau13]. Computers [Aga06, Col89, Hel98, Ken17, Sle89a, RG16]. Computing [Car22, Ger01, Gui98, HHW^{+}19, Lóp22, SvD25]. Concealment [Col07, TL18]. Concept [BH91, DB85, Kel89, Sma78, Dun22, Sla12]. Conception [Bij07]. Conceptions [Bro15, Pel99]. Concepts [Hes86, Rad13]. Conceptual [Ful89b, Hil90, Lew80, Mul74, Pal90, SR73, CLCMW23, SH17]. Conceptualizing [Kle91]. Concerned [Dow88]. Concerning [Mor88a, MG82a]. Concrete [Sim99b]. Conditional [TY85]. Conditions [Wei73]. Conduct [MO09]. Conducting [WG06, FSF24]. Confederation [Cur98]. Conference [MP87]. conferences [CLCMW23]. Confidence [Ano77c, Maz77]. configuration [MB25b]. Configurations [Bou05].

Configuring [HvL08]. Conflict [Far75, MLC06, SCG01, FFB11, Hin14]. Conflicting [AS99, Mes10]. Conflicts [Wad14]. Conformity [Nan72, Whi89]. Confronting [Ram09]. Confronts [Suc06]. Confusion [Mah80]. Congress [CL11, MD20]. Conjectures [BH97]. Conjugating [Hec02]. Connection [Rad86, Sch12]. connections [Pih25, WC16]. Conquered [Hyy09]. Conquers [Hyy09]. consensual [MN23]. Consensus [BP80, Hal06, HI10, KC95, Lew80, Wri86, vdSvESW98]. Consensusing [HI10]. Consensusness [HI10]. Consent’ [LOBB23]. Consequence [FDR08, Gra98]. Consequences [Gar04, Wil91]. Conservation [Yea89a, CC12]. Conservative [Pin97, SN89]. Consider [Hol22]. Considering [WG06]. Consolation [Cam93]. constituent [Pap11]. Constituting [Doe03]. constitution [Pic13]. constitutional [Lau25]. Constitutive [LSS20]. Constraints [Ams87, GS86, GT93, Lan01b, Pic81, Vin95, Hal19]. construct [BSB16]. constructed [Car14]. Constructing [Ake07, Ala09b, BM19, Dix19, Fuj87, Fuk16, GT93, Jas00, Ker00, LS10, LMA80, Mac93, MB25b, Ral06, Rip03, Shi05, Wat02, Wyl15, Yea90, RA22]. Construction [Abr94, Bow85, Bre97, EM00, Elz86, For93, Ful08, Gie92, GS86, Kim07, Mar96a, Mer02, Mie98, Mye85, Nev80, Old90b, Pic90, PB84, Pow01, Ros93a, Rus86, Sin96a, Sør92, Tay95, Tra81, VP98, Wes89, vEG91, AQK15, Bed14, GKSS24, Hed14, PP24, iO18, RP16, Sup14, da 22]. Constructionist [Alb95]. Constructions [MBB00, Sis93b, Wyn92b, Jon17a]. Constructive [Loh00, Ste99]. Constructivism [Cos01, Jen07, KC93, Sto08]. Constructivisms [Tay95]. Constructivist [Dal91, Fre84, Gie82b, KC82a, Kro82, Rad92, Ros96a]. constructs [For94]. Consultancy [Yea92]. Consultants [Kir03]. Containment [Hil07]. Contamination [Mod01, UB25]. Contemporary [All76, Hof79, LM89, Wri80, CS20, Mat13, PC12]. Content [Ano77n, CM75, SR77]. contentious [BM19, Hor25b, Tan21a]. Contents [Sle89a]. Contest [Per04]. Contestation [Car09, McC07, GM23, MB25a]. Contested [Cow95, Edm02, Hai19, Jas87, LJ98, ST16, Sol01b, Hol22, Law20a, PBEW24, Tai19]. Contesting [Tuu05]. Context [Bla76, Coz85, FFMW08, Ham82, Hol81b, Ink85, Kim83, KR00, Liv06, Mor88a, Ore03, SR73, AGKL^{+}15, HH22, Möl16]. Contexts [Ake07, Fra88]. Contextualizations [dMdF98]. Continents [Lau87, Ste86]. Contingencies [VB03]. contingency [EMSL23]. Continuing [Hen90]. Continuities [AD09, Ala23]. Continuity [GT72, Irw94]. Continuum [Ker00]. Contraceptive [Oud97, BM19]. Contract [Bre97, MV05]. Contractual [Sat07]. Contradiction [Rif72, Tou00]. Contradictions [AD09, Mac79]. Contrary [JM73]. Contrasting [Her03a]. Contrastive [Koc10a]. Contribution [Col04a, Lin78, Sch12]. Contributions [Mar07]. Contributorship [LDM^{+}16]. Control [Ano77q, Bar03b, Buc75b, CD84, Hac05a, Muk07b, NHM74, Oud97, Pal90, RDB03, Rot81, SB77, BSB16, Sad24, SV22, Wes22, Del05]. Controlled [Ada02, BSB16]. Controlling [Col85a, Hor04, Joh84, SS17a]. Controversial [Ree01, Iid15]. Controversies [Del83, Mey22, Mye89, RJ76, Pol20]. Controversy [Agr01, Ano77o, Bed91, BH91, Far75, Gar97, GK72, GR08, Gui98, Har76, Har77, Kli02, Mar07, Mar88, MC78, MM84b, MB94, Nel71, Ola96, PM79, Pic94, Pic81, Ram02, Rap05, Ric88, Hin14, Iid15, Law13, PP22, Rit16, SK13, Tol19, Wat14, Waz13, Shr14]. conventionally [PW19]. Conventions [CKSW09]. Convergence [Clo01]. Conversation [Mye89, Woo89]. Conversations [Del05, Vos04]. conveying [RC15]. Convincing [Hal98]. Cookmeyer [Del02]. Cooling [Dow88]. Cooper [Joe99b]. Cooperation [Col95, Jas95, Lub81, Mur13]. Coopmans [CE14]. coordinated [Sun11]. Coordinating [SM25]. Coordination [Hal06, Mor06, LB19]. Coping [SPBG24]. Coproduction [Rea01]. Copula [Blo83, MS14a, MS14b]. Core [Col88, Dow88, Hed06, MB94, SM93a]. Core-Set [Col88]. corn [Bra14]. Corncrake [Lor08]. Corncrakes [Lor08]. Cornell [Bru15]. CORONA [Clo01]. Corporate [Den87, Sch13, Sim10]. Corporation [Bru77]. corporeality [Mül18]. Corpuscular [Fra76]. Corrected [Lin78]. Correction [Ano83a, Mac00, PP21]. correctness [Gre24]. correlative [yLL14]. Corrigenda [Ano86a, Ano90c]. Corrigendum [Ano85, Ano86b, Ano12a, Ano16a]. Cosmological [LM89]. Cosmologies [Mes10]. Cosmology [Kai06]. Cost [Sha20, SK91, BHC16]. Cost-benefit [Sha20]. cost-effectiveness [BHC16]. Costs [Mir07, PI19, RS11]. Couched [Hen90]. Could [Opp79, Wel04]. Council [Rou17, SRM76, Bri09]. Councils [BR79]. Count [Den01, Mar04, Mil01, Sch13]. Countability [BS01]. Counter [CL82, Pic94]. Counter-Enrolment [CL82]. Counterfactual [FR89, Gor91a]. counterpublics [MS24]. Counterweights [Sco91]. Counting [CM75, Ken15, Lor08, Mar04]. Countries [Cla80, Dro93, Fra79, Mor82, Sag73, SK91, Bri14, AB80]. Country [Ves96]. Counts [LMA80, Yai19]. Course [Eas73, Sha80, Spi87, WB78]. Court [CL11, Swa07]. Courtesy [KH19]. Courtroom [Gor06, Car18]. Courts [Hal98]. CoV [Hor25b]. Cover [vK01]. COVID [Sis20a, And21, CLCMW23, Did25, EMSL23, MB25a]. COVID-19 [Sis20a, And21, CLCMW23, Did25, EMSL23]. Cowboys [Olb84]. cows [BD11, FO22]. Crackpots [Rit16]. cradle [Cri10]. Craft [DA01, HH22]. Crafting [DH15]. Crane [Dun91, Pic80a, Pic80b]. craniofacial [Jon23]. Crash [Leo10]. Create [MO09, BM19, Sun11]. Creating [FÖP20, Lez06, RSH18, Ves96]. Creation [Col06, O’C93, PPM20, PBSK25]. Creationism [WO96, Tou91, Tou00]. Creative [RB01]. Creativity [Whi89, Bor17, Wyl15]. Credibility [Bre97, Hil04, KK03, Kir12, Hai19, SH12]. Credible [Cam04]. Credit [Mar07, Del16]. Creek [LR01]. Crime [Ada02, Smi85, CB24]. Crisis [Bed91, Cro91, FM17, Fre92, Hol81b, Lak04, Mas10, Wri80, And21]. Criterion [Dol74]. Critical [Edw94, Fox83, Lyn82, MvR85, Res78, Win76, And21, Cho13, ELR21]. Criticism [Woo81a]. Criticizing [IJJ13]. Critics [Ede06, Kus04a, MI85].

Critique [Bud85, KC82c, KP85, Max17, Woo81a, Hin13, RL16]. critiques [Yai19]. Croom [Wri80]. Crops [Lev01, MLC06, Yea01]. Cross [Jef03, Leo12, Sol92, Kus10]. Cross-Cultural [Sol92, Kus10]. Cross-Disciplinary [Jef03]. Cross-species [Leo12]. Crossing [BH03, FLE97]. Crossovers [Kli02]. crowds [BTS22]. Crucial [Mat05]. Crude [Bow85]. Crushing [Ken00]. Cruz [AB80]. cryo [Kro19]. cryo-kennel [Kro19]. cryobanks [Bra24]. Cryptography [Wes22]. Crystallography [Law73, Mye08]. CSI [Akr23, Ful12]. CT [Rob99]. Cuba [Ste78a]. Cultivated [Eps21]. Cultural [And88, Bar81, Dow88, Hak95, Hol81b, Ink88, Jas95, Mel09, Sha72, Sol92, Wri80, BFW14, Iid15, KGDLB^{+}15, Kus10]. Culturally [Wat90]. Culture [Bij07, Chr88, Col07, Hel98, Hes87, Hir94, Hol81b, Lab95b, Lan01a, Lyn88, Mos79, MG82a, Pra05, Szo92, Tho99, Yai19, CERG25a, HH22, KH18, Sab14, Nev80, Rob99, Dea95]. Cultures [Col95, Dug87, GP09, Goo18, Han86, Jen07, Kir99, Liv99, Tan08, Bak11, Bij12, PJ17, Sis19]. Culturing [LP11]. Curating [Hin13]. curation [GKSS24]. cure [Lev22]. Curie [Pyc93, Wir15]. Curious [Fis10, Goo10, M’c23, Ros10]. Current [Boh04]. curricula [KdR19]. Curriculum [HH83, Mul98]. Curtain [BPBG06]. Cutting [Col13]. cyber [SM25]. Cyberfeminism [Suc06]. Cybernetic [Bow93, PT72]. Cybernetics [Ger01, Hol74, Kli09, Pic02, Col13, Lóp22]. Cyborg [Mir99]. Cyborgs [Kli09, Oud15]. Cycle [Ros93a, Law20b]. Cycles [HR21, Law25, Max17]. Cystic [Ker00]. Czech [Efm92, Mül95c].

D [FOC^{+}71, Bar79, Orl73]. D-7100-8749-7 [Bar79]. Daemmrich [Mir07, Mir07]. Daly [Tur90, Tur91]. Damianos [WZH^{+}25]. Dance [Cri04]. Dancing [Dah24]. danger [Kuc13]. Daniel [Mos79]. Daniels [Str92]. Danish [GKSS24, Hoe19b]. Dar [WB78]. dark [Eng20]. Darwin [Boh91, Chr88, Lau90, Man87]. Dash [Win02b]. Data [Alm08, Chu75, DNTS08, DB85, EMB^{+}11, Gar73b, HAB^{+}23, HR87, Hoe19b, Naf24, SR72, AH11, Arm19, CWvdVdB23, Coo19, DG17, Gre22, Hoe14, Kan22, KH22, May19, Met25, PPM20, PS24, Rei23, Sah25, Sup14, TS24, TL18, Var20, VT21]. data-intensive [PPM20]. Databases [Har02a, Hin06, Jat17, Leo12]. datademic [Did25]. Datadiversity [Bow00]. Datafication [HBP19]. David [Bru77, Kim94, dSP75, Bar03a, Blo03, Bow03, Col03, Elz03, Fox03, FL88, GM78, Gor10, Jas03b, Lat03, Mac03d, Man93, Mey74, Mos79, Pin03, Shr03, Wes86]. Davidson [Ano77h]. Davis [Bol18, Sim79]. Dawson [Hol81b]. Day [Sim07a, Tor92]. DC [Bru77, Del05]. ddC [Eps97]. DDT [Dun78]. de-perimeterisation [SP24]. Dead [Ful90]. Deadly [Hea07]. dealing [HR21]. Deals [OS05b, Yea92]. Dear [Con79, Sim79, Wes79]. Death [BH91, Huf86, RH83, Avi15, GDLB15]. Debate [Abr94, Alb80, Cle86, Col86, Eps97, Kop79, Kus04b, Mul93, Mul94, Mul95a, Pal90, Pfe79, San78, War74, WP74, WO96, Bak11, CL11, KR11]. Debates [Kir99, Tsa24, Del05]. Debating [Hab97]. Deceive [Mir07]. December [Col22, Gie96, Res96]. decide [Reu12]. Deciding [Smi85].

Decision [Bro82, FFMW08, Gra07, Hol81a, Lan01b, VP98, BHC16, Eva22, SFH12]. Decision-Making [Lan01b, VP98, FFMW08, Gra07, BHC16, Eva22, SFH12]. Decisions [Gra07, O’R83, Rie03, Sch94, Zie88]. Declaration [Chu86]. Deconstructing [Pin90, RB90a]. Deconstruction [Got95, MKC97, Old90b]. decoupled [Hin14]. Deductions [Lat08]. Deductive [Gol79]. deep [Gal24, Gar13, dB09]. deep-sea [Gal24]. defanged [Tur12a]. Defects [TL00, SVS21]. Defence [Dol74, Gum90, MG82b, Eri72]. Defending [Tan21a]. Defense [Lic89, Suc23]. Defensive [Sim95]. Defiance [Nan72]. defibrillators [Oud15]. Defines [MR73]. Definition [Fue84, Kuk94, Sis10]. deflating [Lyn13]. deforestation [MR17]. deformed [Sch15]. Degrading [Mie98]. Dehorsification [Lyn94a]. Delhi [Har83a, SG74a]. delineate [Nav11]. Delineating [GL21]. Delphi [Day18]. Demarcating [Mel03]. Demarcation [Pin79, BJ12, Mar20]. dementia [HL19]. Democracy [Car01, Mil01, Mil04a, Moo10, Tho02, Dur11, Lau16, Lau25]. Democratic [SR73, Iid15, LOBB23, Sha20, VSS22]. Democratization [Sim95]. Democratizing [McC07]. Demographics [AW13]. Demography [Gas92, Ram02, Ram09, Lei14, Mer22]. Demonstration [BL94, Sim10]. Demonstrations [Smi09]. dendroclimatology [iO18]. Denmark [Pih25, Sch20]. denunciation [Fis18]. Departing [Hed08b]. Department [Pat83, Suc23, Sim07a, Zin71]. Departments [Bru77]. Depersonalized [Col94a]. depiction [Jon23]. Depoliticizing [Tur89b]. Depth [Goo95, Sec85, Pet17]. depths [Jon23]. Deregulatory [Man21]. Derek [Ano83b, Mac84]. derivatives [Ran10]. Derived [DC76]. Dermatoglyphics [Mil03]. Descartes [Ful89a]. describe [Foc16]. Description [Ple14]. descriptive [DSJ19]. desert [Bar12]. Deserve [Mar07]. Design [Fra97, Gil06, Hår94a, Hsu13, Lam06, LR01, Mul98, Van08, Yan05, CS17, DK12, Eri72, Hed14, LD13, Red16, SS18, Hsu13]. designation [Nav11]. Designing [Bre13, Jat17, Mul98, HB23]. Desire [Bar98, Fis04]. Desk [Mac17]. despite [CKFS13]. Despotism [Jac94]. Destruction [Col81a]. Details [Gre04, Har01]. Detecting [AW19]. Detection [Zie88, MS24, MR17]. Determinants [Kra89, Son95]. Determinism [Bim90, Law81]. Deterministic [AA84]. Deutsch [Mos79]. deutschen [Bay83]. Deutscher [Ril86]. Deutschland [Hom90]. Devastation [Gus06, Hug06]. Developed [AB80, Dro93, Fra79]. Developing [Cla80, DYS^{+}05, Mor82, Bri14, Eng20]. Development [AR02, And88, AB80, Bow87, Bru77, BS08, Buc79, Fue82, GK72, Har03, Hol81b, Kah08, Kir10, Lan81, Law73, Lew72, Lew80, MCBDT00, Met88, Mul73, Oud97, Shr00, SR73, Spi97, Wal75, WB78, Woo76, Zin76, Eri72, FS23, Kna13, Pin21, SDSR25, Sne12, WB78, Twe93]. Deviance [Gor82, Wyn76, Hed14]. Deviant [Par88]. deviation [Hed14]. Device [Jen05, MS14a]. Devices [vdSvESW98, Mar13]. Devicing [Mor19]. Dhirenda [Har83a]. diabetes [DPP^{+}20, Sah12]. Diabolical [Bij01]. diachronic [Dav24]. Diachronous [Ano77p, Oro77a]. Diagnosis [And92, Eps98, KKLL21, Lev22, PJ17, Str11]. diagnostic [PP21]. diagnostics [Eng20, ST24]. Diagrammatic [Vin91]. Diagrams [KIH04]. Dialectical [Cro91, MC79, DSJ19]. Dialectics [Gin97]. Dialogue [KCBT07]. Diana [HDS^{+}98]. diaspora [RA22]. Dichotomies [Gie82a, vHVdW07]. Dick [Ral06]. Dictionary [Kis13]. Did [Aga06, Ber08, CC76]. Diegetic [Kir10]. Dieldrin [GEJ79]. Diesel [Hår94a]. Diesel-Engine [Hår94a]. Dietary [Gar97]. dieting [Mol13]. Difference [Aga06, Eps08, HR87, Mod00, Mor00, Sis10, Ple14, ZDI^{+}23]. Differences [Bij07, DC76, JS89, Mes88, FR11, Lez10]. Different [FR11, MM84b]. Differential [Coh86]. Differentiation [RJ76]. differently [yLL22]. Difficult [Kel89, TL11]. diffuse [Nav11]. Diffusion [BH97]. Digital [Ala08, Wes24, Arm19, CWvdVdB23, CS20, Del16, Fro12, Hoe14, Hoe19a, Jat17, Mar24, Mes17, MB25b, SvD25]. Digitizing [RB99]. Dilemmas [Gel09, LC05, MPA87, Coo19]. Dimension [Mae07, Mar97, Doi11]. dimensional [Mae07]. Dimensions [Col86, Gol94, Pal90, Oud12]. Dingoes [Hea07]. dining [Bor17]. Dinner [Mye89]. Dinosaurs [Cle86, Rie15]. diplomacy [Smi14]. Directions [Tom87]. Dirt [Mod01, Car14]. Dirty [Fis18]. Dis [HFGP17]. disappeared [Smi17]. Disappearing [Bud85, Jan18]. Disarray [Sch72b]. Disaster [All07, Hil07, Sim07b, Bie18, LA19, Shr14, Wyl19]. disasters [eS23]. Discarded [Mos16]. Disciplinary [CK83, CK05, Jef03, SS94, Huu12]. Discipline [Bar03b, BL87, Lyn85, Mil03, AC22, Ell11, Nel13]. Disciplines [AA84, HP91, Whi76, GM78]. Disciplining [HP91, Lei14, Tho02]. Disclosed [SO17]. Disclosing [DK12]. Disconfirmation [FR91, Har80, TDM82]. discontents [And13]. Discourse [AA82, FO86, Hal88, MP87, MG82b, OS05a, Pot87, PM89, Sol92, TJ85, Woo86, JL18]. Discourses [Kwa05, Mer02, Wyn88, Bed14]. Discoveries [BW83, Sch86]. Discovering [EP07]. Discovery [Bra89, Doi11, Dun74, FR89, FR91, Gie89, Mye95, Ore03, Sim91, Sim78, Sle89b, Wes79, Woo76, Woo89, Pol14, TL18]. Discursive [KK03, Lyn98a, Sla07, PBSK25]. Discussant [NKB86]. discussion [PBSK25, Sis25]. discussions [Iid15]. Disease [Cou99, FDR08, Gar97, Hed06, Ker00, Kin02, Lan01a, Mil03, Shi05, Abe13, Sis10]. Dish [Lab95a, Lab95b, Pin95]. disintegration [OOO12]. Disinterested [EP07]. Disinterestedness [Dea92]. Disorder [Smi85, SK13]. Displacement [Mul74]. displacing [LR20]. Display [Sch06, Car18, CS20]. Displays [Col88]. Disposal [Mac03a]. disposals [Tan21b]. dispositifs [Mur13]. dispute [BL12]. Disputes [Cam85, Ger01, Gra98]. Disrupted [BV95]. Dissected [Cos01]. Dissecting [Col94a]. Dissembling [MM84a]. Dissemination [JL98, Eva10]. Dissent [Del08]. Dissident [Lev84]. dissimulation [Smi15]. Distance [Blo91a, Cho13, VSS22]. Distinction [Woo86]. Distinctions [Bur08]. Distinctive [Gor82]. Distortion [Mil09, Som06a]. Distributed [BS08, GM03, Mag07, Mül18]. Distributing [Wet07, Kru13]. Distribution [Ano77h, FNC77, Gun05, Lah05, Rap01, Sim78]. Distributions [BW83]. disturbing [HRI22]. Disunity [Hal06, Mac03c]. Disziplinen [Hom90]. Ditch [LR01]. Divergence [AB91]. divergent [VT21]. diverging [McL15]. Diverse [Wat90]. Diversification [WD97, Ben12]. Diversity [Con78, Rea01, Bur18, CERG25a, CERG25b, JSLS25, Sis25, Tho25]. Divide [Pow01, Shi05]. Divided [Blo77, Rip91]. Division [Har96, Kom22, LDM^{+}16]. DIY [KLW19]. DM [Pfe71]. DNA [Dae98, Der00, Edm11, GM23, Hal98, Hop23, Jas98, Kow13, Kro19, Law13, Ler18, PK09, PBEW24, SMWCSC15, Tsa24, Wri86]. DNA-Typing [Dae98]. Do [Bij12, Col04b, Fuj87, Joe99a, Mae07, Mia99, Mir07, WC99, WC18, AB19, Sma20, UB25]. Do-able [Fuj87]. doable [SAG11]. Dockets [OS05b]. Doctor [HAB^{+}23]. Doctoral [JE75, Tür73]. Doctorate [Tür73]. Doctoring [DA01]. Doctors [lL99]. doctrine [Law11]. document [DA24]. document-practices [DA24]. documented [SAG11]. Documenting [SAG11]. Documents [Sma78]. Does [Dae98, Dal91, GS11, HR87, Mil09, BL12, FW17, KCM97, LSW24, Say14, UB25]. Doing [Bur08, SAG11]. Dolan [Bru77]. Dole [Ber08]. Dollars [Den85]. domains [RHSB19]. Domesticating [EMSL23, PS24]. domesticity [Lac16]. Dominant [Hil90]. Dominated [Mel09]. Domination [Cot75, Hon08, AB80]. Donald [Ros96b]. donated [Car14]. done [God15]. Donna [Del02, Cri10, Har23]. Don’t [Ful90, HM03, Dem24]. Door [Del05]. Doping [vHVdW07]. Dorothy [Ano77l]. dosing [ZDI^{+}23]. Double [Sto82, Tim96]. Doubts [War71]. Douglas [RD08]. Dow [Hsu13]. Down [Ras98, Yan05, SV22, Uru22]. Down-to-Earth [Ras98]. Dr [Wei10]. Drafts [Man87]. dragonfly [TH24]. Drain [LR01]. drama [Möl16, Sor23]. Dramas [Coo19]. Drawing [Fro12]. Dream [Del05, Lyn98a]. Dreams [Nic97, Yon04, Sch20]. Drew [Bru77, Bru77]. Drexler [KR11]. Dreyfus [Col92b]. drift [De 24]. Drifting [Lau87, Ste86, Ste87]. Drilling [Den85]. drive [Pan11]. driven [Pin15]. driving [Sti18]. drone [Cha22]. drones [OL18]. Drosophila [Ens12]. drought [LSW24]. Drug [AS99, AR02, AD09, Eps97, Gel09, Kah08, lL99, Mir07, Ras04, MD20, Pol14]. Drugs [HM03, lL99, Sis10, Len16, Seg18, Mir07]. Drum [Ial21]. Dualisms [Fau00]. Due [Hea07]. Duke [Eps09]. During [Ano77e, Bea77, Sol01b, Fri84, LR01, MB25a, Pop21, SM93a]. Durrant [Pri71]. Dusseldorf [Pfe71]. Dutch [Bij07, HTB20, Kir99, Mol13, Wil91]. Dynamic [Tim96]. Dynamics [Man93, GB16, HHW^{+}19, Oik17, Sun11]. Dynamism [GR08]. Dysfunction [Fis04, Seg18].

e-book [Del05]. e-health [NL18]. E-Infrastructures [VT21]. E. [Ano77e, SS91]. Each [PB84]. Eamon [Dea95]. Earliest [Man87]. Early [Ams87, Ber72, CG90, Dea95, Ger01, Goo10, Joe94, MM81, Zin76, Sil11, Som06a]. Ears [Per04]. Earth [Clo01, Nev80, ÖS24, Bar23, Clo03, Den03, Doe03, Ras98, Row05]. Earthly [Den03]. Earthquake [Sim99b]. Earthquake-Engineering [Sim99b]. Ease [HI10]. East [AB80, HTB20, AB91, IM85]. Eastern [BFS95b, Bal95, BFS95a, Bur18, Res78, Res82, Sch95]. Eat [Van08]. ebk [Hyy09]. Ebola [SAC^{+}20]. Eclipse [Hut91]. eco [Mar13]. eco-show [Mar13]. ecobiopolitics [Car21b]. Ecological [GR08, Pal90, RB99, HRI22]. Ecologies [Kwa05, Jen20, Lyo18]. Ecologists [Nel77, Ano77l]. Ecology [Abr91, Ake07, Hen07, KK03, Kwa87, Lac16, Law81, RB01, Som06b, SG89, FSF24]. Econometrics [WC16]. Economic [Ano77f, Cla80, Fra79, GT93, Hoo17, Inh77a, Ink91, KC82c, Lak04, Mül95c, Ste78a, MB25b, Pin21, Rei13, SV22]. Economics [Bre13, Geu03, Mac03c, Mir99, NK14a, Yon94, AC22]. Economies [Sha01, Bru15, Hai19, SL25]. Economy [Dro93, FG86, Hes00, McC00, Ras04, Len16, Mac17, PS24, Rel24, SM25]. ecosystems [PH14]. ed [AB80, Blo77, Bro95, Hol81b, Nev80, Wri80]. Eden [Ano06c, Gus06, Hug06]. Edgar [KR00]. Edge [Blo03, GM78, Law11, Spa20, Bar03a, Bow03, Col03, Elz03, Fox03, Jas03b, Lat03, Mac03d, Pin03, Shr03]. edges [vdH20]. Edinburgh [Can75, Mor73, Stu07, Wes79]. Edison [CG90, Vin95]. edit [FW17]. Edited [Clo03, Hac05b]. editing [Mey22]. edition [Kim94]. Editor [Ano10, Coz86, Edg95d, Lab04, Sis22b]. Editor-in-Chief [Sis22b]. Editoral [Ano05a]. Editorial [Ano71a, Ano76, Ano86c, Ano05c, Ano05b, Ano06a, Ano07a, Ano07b, Ano08, Ano09a, Ano12b, EM73, EM75, EM78, EM79, EM83a, EM83b, EM84, EM85, EM86, EM87, EM88a, EM88b, Edg91a, Edg91b, Edg92, Edg94a, Edg94b, Edg95a, Edg95b, Edg95c, Edg96a, Edg96b, Edg97b, Edg99a, Edg99b, Edg01a, Edg01b, EL02, Hir10, Lyn02, Lyn03a, Lyn03b, Lyn03c, Lyn03d, Lyn04a, Lyn04b, Lyn04c, Lyn05, Lyn09a, Lyn09b, Lyn10a, Lyn10b, Lyn11b, Sis13]. Editorials [STM^{+}08]. Editors [Ano09a, Con79, EM82, Nel22, Sim79, Swa89, Wes79]. eds [Bar90, Col07, Del02, GM78, Hol81b, Mae07, Nev80, dSP75]. Educating [PS77, Rob96]. Education [Blo77, Bru77, FG86, Rip79, Son09, TG86, Wes86, Zin76, Bed14, Blo77, Bru77]. Educational [Wil91, Dix19]. Edward [Tur95b]. Edwards [Wes88]. Effect [Kan78b, MK84, Mag77, Rud77, Sim84, Gol79, LPKL12, Ros93b]. Effectiveness [Mor82, BHC16]. Effects [Coh86, Gom02, Lan01b, Lin16, Son09, CKFS13, Cra10, Rou17, SPBG24]. Efficiencies [Ale08]. Efficiency [Ale08, Dun22]. Effort [Lew72]. Egypt [Bar12]. Ehrenreich [Wri80]. Ehrlich [Wes88]. Eighteenth [Gas95]. Eighteenth-Century [Gas95]. Einstein [Lat88, Mis85]. Elderly [Hyy06]. Election [Agr01, CFS01, Hil01, Jas01, PP22, Lyn01b, Mil04a]. electoral [PP22]. Electric [Gor77, Tod92, Anf21, Ano77j]. Electrical [Vin95]. Electricity [Mor88b, Win02a, Win02b]. Electrification [Bar84a, Tod87]. Electromagnetic [KSW82, MC97]. Electronic [Gun05, Jen05]. Electronics [CM09]. Elementary [Kri87]. Elements [Dir99, Nur03, ÖS24]. Elga [Del05]. Elichirigoity [Fee00]. elimination [ST24]. Elite [Mul76, Mul72, vHVdW07, Sma20, Bru77]. Elites [HV93, Whi76]. Elitism [Ami74]. Eloquence [Lez10]. Else [Van08]. Elsevier [Wri80]. embedding [MS14a]. Embodied [Bal95, Sor23, Nis11]. embodiment [Ver12]. Embodiments [Mye08]. Embryo [Hab97, HH22, Mul93, Mul94, Mul95b, SK08, CL11]. Embryonic [Pra06]. Embryos [Kir99, Mul95a]. Emergence [Cha99, CG89, Gas84, GM78, Gol04, Lau90, Sho05, Shr00, Fau12, Law11]. Emergency [Dow88, Nur03, MB25a, SS18]. Emergent [Bia22, Gor82]. Emerging [Byc16, KR11, Lyn11d, SSB24]. ‘EMF [Mer02]. Eminent [Del05]. emotions [Did25, SH19]. Empire [Bes20, Nic97, Red02, Lac16]. Empirical [Bur83, Col81b, RH85, Sle89b, CLCMW23, LL13]. Empiricism [Lau82, Ken15, RG14]. Employers [Div90]. enable [Pih25]. Enabling [JS23]. Enacting [BM14, Chi24, MS24, GK16a]. enactment [CWvdVdB23, Lev14]. Enactments [Pap21]. Encounter [Kim07, FLK^{+}14]. encounters [Dah24]. Enculturation [RB01]. End [Ald98, Blo91b, Gil06, Sch86, Win02a, Smi17]. End-to-End [Gil06]. Ending [Ash96, Del05]. Endlessly [Lab95a]. endogenous [Pin21]. Enemies [Col85d]. Energy [Cra80a, Dun85, Dun22, Hod83, IM85, KC95, Mat05, Pic80a, Sol85, Tak24, LR20, RSH18, ST16, SPBG24, Sov10, SBH^{+}20]. Enforcing [Jat23]. Engagement [HvL08, Jas96, AMT11, Bri14, Dem24, IJJ13, LSW24, NL18, Sar23, VA16]. Engagements [Thé09]. Engaging [Cri96]. Engine [Bra92a, CB01, Hår94a, Tor92, Mag17]. Engineered [Hog09]. Engineering [Bij07, Buc79, Div90, Edm02, ER78, Fau00, Fau07, For93, Gil06, GGJ07, Got95, Hor04, Kra89, Law87, Pre08, Sat07, Shi99, Sim99b, Sør93, TG86, Vin91, Bed14, KG12, RvD22, SVS21, WB78, Wyl19]. Engineers [Con89, Fle94, Gra07, Hor04, Por04, Mar20]. Engines [CB01, Zie19]. England [Jac75, O’R83, Div90, Don86, Lau84, Wri75]. English [Kis13, Car18, Far75, RS74, Zin71]. ENGOs [US20]. Enhancing [HvL08, Par25]. enigma [She20]. Enlargement [LS04]. Enlightened [Hed01]. Enlightenment [CS00, Ste90, da 22]. Enquiry [Boh89, CLCMW23]. Enrolling [MB94, Swa24]. Enrolment [CL82]. Enstehung [Hom90]. entangling [HFGP17]. Entering [Hol81a]. Enterprise [Hyy09, Ink85, Ink88, Ma19]. entity [Bre13]. Entrepreneurial [Jon09, Lam10a]. Environment [GB09, MC97, Tür73, Ala23, Lap16]. Environmental [Ala09b, All07, CD84, CBLG23, Doe03, Eps09, Mat05, ML93, McD97, Mer22, MLC06, Nel71, Nev80, Tua71, Ver02, Yea92, Ben12, Car21b, Jeo19, KH22, Naf24, SV22, US20]. Environmentalism [Yea89b, Bro95]. environments [Car22, Tir18]. envirotechnical [Tjh24]. Enzymes [Mie98]. EPA [Man21]. epidemic [Gre16b, Oik17, SAC^{+}20]. epidemiological [AB19]. Epidemiology [Bri14, Rie03, Shi05]. epigenetic [Chi24]. Epigenetics [DSJ19]. Episode [Mor88b]. Epistemic [Bal06, Doi04, FSF24, Kli02, Ley91, Mar04, Nel13, Per04, Rei13, Eva22, Hor25b, Kan21, KH18, LOBB23, Pet17, WM17]. Epistemological [Dal91, Nur03, Sol01a]. Epistemologically [Hen90]. Epistemologies [EM06, Kus10]. Epistemology [Bre97, Gor89a, Jas96, Kri78, Mil04a, DP21, Rel24]. Eponymy [dBB76]. Epstein [Sis10, Pic97]. Equation [Mac03c]. Equatorial [MD09]. Equivocal [ST24]. equivocation [FLK^{+}14]. Era [Ale08, Eps97, KLBKC11, Rei21, WH21]. Erasers [Bar11]. erasures [Bar11]. Eri [dSP75]. ERIC [Aar23]. Ernest [Blo77, Wes88]. Erratum [Ano91d, Ano96, Ano99a, Ano99b, Ano09b, Ano09c, Ano09d, Ano12c]. Error [Der00, Jas01, Sti87, Ial21]. errors [AW19]. Escape [Das92, Eli00]. Essay [Alb80, All76, AB80, Bar79, Blo71, Blo74, Blo77, Bru77, Buc75a, GM78, GW74, Gin10, Gre10, Haw73, Hol73, Hol81b, Kea73, Lat08, Mir07, Mos79, Nev80, Pin11, Sch72b, Sis10, Sto08, Tua71, Wri80, Doi11, Edm11, Hor25a, LP11]. Essays [Day08, Nev80]. Essential [Hac05a]. essentialism [Sch20]. Establishing [Hod83]. estate [Gre22, Har83a]. Estimates [Bow85, Den85]. Estrangeirados [CS00]. Estrogen [MM84b]. Etcetera [Hug86]. Ethanol [Car09]. ethic [Elg22]. Ethical [MO09, Sch72b, Kow13, Zie19]. Ethics [Bri09, FS23, Gor06, Lan20, Cra10, DSJ19, Hed12, Hed14, HJ13, HFGP17, Hol22, Jac19, LP11]. Ethnicity [STM^{+}08, Bur18, Lei14]. Ethnobotanical [GP09]. ethnobotany [She20]. Ethnographic [Vau00, Vos04]. Ethnography [CLP90, CLP91, GP97, KCM97, Lyn94a, Ano16b, Bea10]. ethnomethodological [Lyn11a]. Ethnomethodology [Bar85b, MS00, Mul73, Woo81a]. Etiquette [Gre04]. Eugenic [GGJ07]. Eugenics [All76, Bix97, God07, Kim83, Mac76, Ram02, Ram09]. EUR [Ash07, Col07, Del02, Del05, Doi11, Fis10, Gin10, Goo10, Gre10, Hyy09, Mae07, Moo10, Pra09, Rei06, Ros10, Sis10, Sto08]. EUR88.50 [Ash07]. Europe [Bal95, BFS95a, BFS95b, GM23, Lev01, Mag17, Sch95, SR72, Wat02]. European [Law20a, MLC06, Aar23, Bij01, Fau12, Law20a, da 22]. Eva [ABB^{+}08]. Evaluating [KHM10, KP85, Kru13, MI85, MvR85, Bru77]. Evaluation [AD09, Bru77, Eps97, Har81b, Ric88, Sol85, Son95, SWB77a, Ber72, Cho13, MM82, Ros10]. Evaluative [Mah79]. Evaluators [MI85]. Evans [Jas03a, Sis17a, Wyn03]. event [Abe13]. events [Cra10, HJ23]. Ever [Jas00, Lyn98b]. Everett [Blo77]. Every [Car18]. Everyday [Hea04, McD97, Joh13, ST16]. Everyman [Jas98]. Everyone [Gie06a, FW17]. Everything [Lyn98b, BJ12]. Evidence [Bar90, Bor17, Col98, Dae98, FDR08, Gre10, RB09, TDM82, HL10, Kna13, Kru13, May19, WC16]. Evidence-based [Bor17, Kna13]. Evidential [Bak11, Pin85]. evidentiary [Lyo18]. Evolution [Baz84, Con78, JL18, KSW82, RH96, WO96, Dun22, Leo10, SS17b, Wei73, Old88]. Evolutionary [Blo74, Lau90, You71]. exactness [Püt21]. Examination [GW74, Sha72]. examinations [Nis11]. examining [BH97, WC16]. Example [All97, Sha01, Sta83]. except [BL13]. exception [Leo12]. exceptional [Kro19, PW19]. Excess [Kid88]. Exchange [Ano77e, Bea77, Gun05, Hay07, Bru15, Ken17]. Excluding [lL99]. Exclusion [Mar10]. exclusivity [Pin11]. Excursion [Vau00]. Executive [HB23]. Executive-centered [HB23]. Exemplars [Mac03c, Pic80a].

Exemplary [Sto82]. exemplifying [Lyn12b]. Exercises [Jen23]. exhibit [Lau13]. existence [Lat13]. exobiology [RB22]. exome [Tim15]. Exoticism [Lyn94a]. Expansion [Car21b, Bar12, SPBG24]. Expectation [BV95]. Expectations [HM03, Hed06, Hed10, Rei06, VT21, Bir23, Mil12, PW10]. Experience [Ano77l, Blo85, CB97, CE02, Hea04, Ink85, Nel77, OS05b, Rud96, Shi05, AH11]. Experiences [Bix97, Gor06, Sup14]. Experiencing [Dav21]. Experiment [Jen05, Rad86, WSB79, Wir15]. Experimental [Asd08, Baz84, Gor90, Har81b, Hil86, Kan78b, KMH83, Mag77, ST97, SFP11, AC22, DA24, FLK^{+}14, Kim14, Lau16, Suz21]. experimentalized [Mar13]. Experimentation [MB94]. Experimenting [DA24]. Experiments [Blo91b, Col88, KC95, Bol18, Lau16]. Expert [All07, Col98, CD84, Dae98, Ham86, Lyn06a, vEG91, Kum24, MB25a, SG21, Tak24, Bar90]. Expertise [AS99, Blo00, Bri08, CE02, Gor02, Joe94, Lew01, LC05, Moo10, Nel75, Rip03, Smi85, BL12, Car18, CB14, Dav24, Hoo17, Huu12, Kum24, Lau25, Lyn14, NM20, Pet15, PBEW24, RC15, RL16, Sne12, Tai19]. Experts [Cam85, Dre92, Hal98, KCBT07, Sle92, Tho02, Tur01c, Dem24, SK13, WLBRS15]. experts-or [Dem24]. Explaining [SMW15]. Explanation [Gol79, Koc10a, yLL22, RB90b, Woo81b, Car21a]. Explanations [Pfe79, Pin79]. Explicit [Doi11]. Exploited [Swa89]. Exploiting [Kah08]. Exploration [Ver12, ÖS24, Rei21]. Exploratory [Cra80a, Lew80, NFC83]. Exploring [CKFS13, LSS20, MvO23, Ras20, Sco92]. Exporting [Fit86]. Exposure [Gra16, Rie03, Ure25]. expression [Car18]. Expulsion [Ham82]. extend [AH20]. Extending [MG92, PH11]. Extension [She06]. External [Ano77o, Har77, Man87]. Externalists [All76]. Externality [Pin85]. extinction [Hun24]. Extra [Mes17]. Extra-terra [Mes17]. Extracting [MD09]. Extraction [Yan05, PT18]. extradisciplinary [TS24]. Extraordinary [Fau98]. Eye [CWvdVdB23, Jas98]. Eyeballing [CB14].

F [Alb95, Bug93]. F-4 [Bug93]. Fabricating [Fox94]. fabrication [HL19]. Fabulous [Mil04b]. Face [MM87, Sti87, Hop23, MvO23, Plá23, Coo11]. Faces [Bim90]. facework [Hed12]. facial [Del23, Ple14]. Facilitating [SK08]. facilities [Law20b]. Fact [Mel03, JL18, Sim99a, SS17b, ZDI^{+}23]. Factor [JLM73, Man87]. Factors [GT93, Har76, Har77, Ste78a, Ano77o]. Factory [Kis13]. Facts [LW82, Lew95, Mye98, PB84, BM19]. faculty [KG12]. Fadlon [Pic97]. Fail [CC76, MB81]. Failed [Law77, LT92, Nic84, Sco91]. Failing [Eli00]. fails [Rei21]. Failure [Bar84b, Fre84, Hac92, Jas01, Tod92, Tor92, Wol95, Cal21, Fuk16, Heg16]. failures [MB22]. Fall [LM80, Rip91, San78, Sim07b, Arm22, MBDR24]. fallible [GS11]. falling [HLRK14]. Fallout [Kop79, BFW14]. False [HI10, Lam10b]. Falsification [Eva97, FR91]. Family [Fox05, FFB11]. Famous [Mis85]. Fan [Dow07a]. Fantasy [Wat02]. Farce [Lyn06b, Pic94]. Fare [KCM97]. farm [Hin14]. Farrington [Str92]. Farris [Gor91a]. Faster [Mod12]. Faster-than-light [Mod12]. Fat [Gar97, Sch13]. fatbergs [Mic20]. fate [AW13]. fauxtomation [CWvdVdB23]. favourites [Jen23]. Fax [Lew95]. FDA [Fro17]. Fear [Ful90, Mul93, SH19, Sto08]. Fears [Jas97, SAC^{+}20]. Features [Sha01]. February [Sim79]. federal [Hoo17, CL11, Gil86, Huf86, PKS88, RDB03, Tat86]. Federation [Blo00]. feeding [SS18]. Female [AB91, Del05, Fis04, Seg18]. Feminine [RS89a]. Feminism [Sin96b]. Feminist [Del02, Loh00, Mat13, Has12, Mur15, Del02]. feminization [Ple14]. Feminized [Sør92]. Fermilab [Hod83]. Ferrets [CW87]. Fertile [JM73]. Fertility [Hud73, HH22]. Fetish [Blo91a]. fetus [Nis11]. Fever [Ste78a]. Few [CKSW09, Nis95]. Feynman [KIH04]. Fibbing [RB01]. Fibrosis [Ker00]. Fiction [Mel03]. Fictional [Kir03]. Fictions [PI19, Mat14, ÖS24]. Field [Bix97, Del05, FFMW08, FS01, GK16a, Gie06b, Hen00, Hon08, LL13, Ree01, Ree07, RB01, Wat02, AS16, AH11, Bor17, GK16b, Lac16, PBSK25, SF19, TH24, Ure25]. Field-Sites [Gie06b]. Field-Workers [Bix97]. Fields [MC97, KG12, Pet17]. Fieldwork [RB99]. Fifty [Dea12, Sis12b]. Fighting [Dow07b, Wei10]. Filling [KdR19]. Films [Kir03, Kir10]. filter [Sil11]. Final [Pin99a]. Finalization [Pfe79]. Finance [Jab95]. Financial [Mac03c, Pre06, Del21, MB22, WC16]. financing [Bir23]. Finding [Mor86a]. Findings [Gil76, JM73, MK84]. Findlay [SN89]. fine [Bor17]. Fingerprinting [Col98]. Finland [Jau15]. Fire [Ano06c, Gus06, Hug06, Ans06]. fires [SAC^{+}20]. Firing [Ver02]. Firm [San72, Tuu05]. Firmament [Pri71]. First [Col96a, Eas73, Zie88, Zin71, NTW^{+}23, WB78, Mac79, SG73]. First-Year [Eas73, Zin71]. Fish [Sch10, Car21b, DA24]. Fisheries [Bro89]. Fission [Hug80]. five [Sis12b]. Fixation [Str92]. fixes [LR20]. flashlight [KLW19]. Flawed [ABB^{+}08]. Fleck [ABB^{+}08, For94, FL88, Har86, Hed08b, Wei10]. Fleet [MS88a, MS88b]. Fleron [Hol81b]. Flexible [FG86]. flibanserin [Seg18]. Floating [De 24]. Flood [Wet07, Jen20, Shr14]. floodgates [Sha20]. Floor [Ore03]. Flopping [vHVdW07]. Florida [Agr01]. Floundering [KCM97]. Flounders [GP97]. flow [BB17]. flows [VSS22]. flu [Lak17]. Fluid [Red16, dLM00]. Fluidity [Mor00]. Fluids [ML94]. Fluoridation [Mar88]. fluorinated [RCB18]. Flush [HJ23]. Flying [Mod00]. FM [DH08]. fMRI [Ala08]. focus [Arm10, Pet17, Val10]. Focused [Mar05]. folded [Hop23]. Folkestone [Hol81b]. Follow [Sch12]. Following [AB19, Eva22, Sha88, Mül95c]. Follows [Ash07]. Folly [Dis04, Lam06]. Food [Kli02, MM96, Yea01, Fro17]. Foot [Ful91]. Footsteps [Rob99]. Force [Cur94]. Forecasting [Ano71b, Mil76, Roc81, Dix19]. Forensic [Bie18, GM23, Hal98, LJ98, PK09, Edm11, GDLB15, Hop23, Jon23, Kru13, LW10, Law13, MvO23, Pot11, SMWCSC15]. Foreskin [Nor17]. Forest [KHM10]. Forestry [Muk07a]. forests [Mat14]. Forget [Wel04]. Forgotten [Fra97]. Form [Cur94, Lyn85, SMWCSC15]. Formal [Kuc13, MM86, Nad83, Gre10]. Formalization [SBHY15]. formalizing [Lam10b]. format [Sil11]. Formation [Gra75, KC95, Shi05, Bak17, Bre13, Car12a, LD13]. Formatting [Kan22]. Former [Nes95]. Forms [Ash96, Col92b, Col94a, Wat90]. Formula [Bal06, MS14b]. Forsythe [HDS^{+}98, Fle94]. Fortes [Ves96]. Forum [SP24]. Fossil [Col96a, Goo10]. fossilization [Hig21]. Foucault [Jah72]. Foundation [Abr91, Lin16, Yox84, AA82, Bar84b, Cue90, Fit86, Fue84, GKM75]. Foundations [Blo87, Len79, McD76, Wat90]. fourth [Sch20]. fragile [Lak17]. Fragmentation [Law80, Mil81]. Fragmented [Wal00]. Frame [Ano77h, Sch88b, Smi09]. Frames [Dor89b, CKFS13]. Framework [BH03, Gor02, McC07, Mil78, SC18]. frameworks [SH17]. Framing [RDB03]. France [Gre10, Ald98, BL87, BR79, Cha99, Fre90, Kra89, Muk07b, Pri12]. Francis [Ano77h]. Francisco [AB80, MBDR24]. Franco [Ano77j, Cro76, Gor77]. Franco-Prussian [Cro76, Gor77]. Frankenscience [Rob99]. Frankenstein [Rob99]. Franklin [Hic88]. Fraternity [Con89]. Fraud [Har81a, Pin79]. Fraught [Tol19]. Frederic [Hol81b]. Freedom [Yea85]. Freire [Cro92]. French [Gre10, Ste90, Cas09, CG89, Gra07, Hec96, Pop21, Pri12, Shi99, Som06a, dC09]. frequencies [Sup14]. frequency [Mac17]. friction [EMB^{+}11]. frictions [PI19]. Friends [Ano77e, Bea77, Bro95, Col85d]. Frigate [Low89]. Frogs [Ano77k, Mos77]. Frolicking [KCM97]. Front [Sch72b, GS11]. frontier [Rie15]. frozen [Kro19]. Fuel [Col96a, dMdF98, Law20b]. fuel-cycle [Law20b]. fuels [PBSK25]. Fuerst [AA84]. Fuhrman [Pot87]. Fukushima [Lin16, LA19, Pol20, Tak24]. Fuller [Lam06, Wei94a, Wol94]. fun [Col13]. Function [MM75]. Funding [Bar84b, Coh86, Gil86, Har03, Huf86, dSP71, Tat86, Rei21, SSB24]. funeral [JS17]. fungible [MB25b]. Further [Lau84, TDM82, WO96]. Furze [O’R83]. Fusion [Lew95, Sim99a]. Futile [Col96b]. Future [GT00, Jen05, Kir10, LM75, Sch20, CS20, Gal24, GS25, Mah14, Mir18, Mor19, Rei13, SDSR25, SBH^{+}20]. Future-Generating [Jen05]. Futures [Ski06, Dix19, LR20, Sha20, Tut11]. futurity [dlB15]. Fuzzy [Lam10a, NM20].

G [Mos79, PCL98, SS91, Wyn76]. G. [Ano77m]. Ga [Win02b]. Gabriel [BS04b, Lyn04e]. Gains [Elz04]. Galápagos [Hen18]. Gale [KCM97]. Galileo [FM17, Gol94, Mul95a]. Galison [Blo91b]. Galton [God07]. gambling [Hsu13]. Game [Bru77, Jon09, Mar78, Wra03, RL21, Sok23]. Gaming [GT00, Zie19]. Gap [Ver08, Zin71, FW17]. gaps [KdR19]. Garfinkel [Lyn11c]. Gas [Bal06, Bra92a, Whi75b]. Gascoigne [Ste90]. Gasoline [Car09]. Gathering [Law25, vK01]. Gaussian [MS14a, MS14b]. gay [Gre16a]. Geeks [DH08]. Gemini [Hac92]. Gender [Del02, Dug92, DH08, Fau00, Fau07, FS01, Fox05, Kel89, Mel09, RS89a, Rie03, Son09, Waj00, BL13, FW17, GB16, KG12, Lag12, Lóp22, Mon12, RS11, VFWC15]. Gender-Troubled [Fau07]. gendered [OL18, Sør92, Ree06]. Gene [Rob02, RH83, Wal01, vHVdW07, Hal19, Mey22]. Genealogical [Arm09]. Genealogies [Jas12, Cha22, Del23]. Genealogy [Nel08, Kim14]. General [Osb93, SM93b, HAB^{+}23]. Generating [Jen05, Kir10, Pan11]. Generation [Cam03, RH83, TJ85, DG17, Del02]. Generations [Pin93]. generative [dLDV21]. Generification [PWD07]. Generous [M’c23]. Genetic [BFW14, Del05, Ful08, GGJ07, Got95, Hel98, Ker00, Nel08, PD19, Par05, Sah12, FR11, TT18]. Genetically [Kli02, Lez06, MLC06, Yea01, Elg22]. geneticists [GM23]. Geneticization [Hed01]. Genetics [All76, Bou05, FDR08, Gas90, KW15, Kim83, Pal02, RB09, Ram09, Rob99, Vic95, Wei94b, AABF10, GDLB15, Iid15, Nav11, Nel13, Pan11, SMWCSC15, SMW15, Smi17]. Genius [Lyn14, Woo74]. Genoma [KGDLB^{+}15]. Genome [Bal96, OS05a, Rea01, Bur18]. Genomic [Nav11, SS17a, Tal13, WLBRS15, Ken13, KGDLB^{+}15, KLBKC11, SDL^{+}14]. genomics [CGP25, HJC23, KBRB19]. Genre [Gor90]. Genres [Sin96a]. Gentlemanly [ST97]. Gentler [Lyn04e, Sto04a]. Geoffrey [Pra09, Pri71]. geography [Hun24, Kim14, Mil12]. Geological [Rud96]. Geology [Mac03a, OM76, AH11, Dam25, WZH^{+}25]. geopolitics [Sne12]. Geosphere [Kwa05]. geothermal [SPBG24]. Gerard [GM78, Lat07]. Germ [KC85]. German [Ril86, Wei94b, Bay83, Got95, OS05a, Sol92, SR73, Wol95, Yai19]. Germanies [SR72]. Germany [Mir07, Pfe79, PS77, Rob02, Tod92]. Gerontology [Bol18]. Gerrylynn [Mos79]. Gesellschaft [Ril86]. Gestural [Ala08]. Gestures [Col96b]. get [Jat17]. Getting [CB01, Gar99]. Ghost [Lyn04d, AQK15]. Ghosts [Bia22, McH09, Sis09a, Sis09b]. Gieryn [Han84, Kro82]. Gilbert [Ano77m, Hal88]. Gilded [Rie15]. Giles [Nev80, O’R83]. Gingras [Pic99]. Give [Har01, Wis01]. Giving [Hay07]. Glasgow [SN89]. Global [Hår94a, Kim07, Kin02, Kwa05, Mat05, ML93, PWD07, Pol14, Smi17, Abe13, Ben12, Cra10, CBLG23, Law20b, Sar23, ST24, TT18, WM17]. Globalization [Lak04, Row12]. GM [Kli02, Lev01]. GMO [BL12]. Goals [FOC^{+}71]. Goes [Rus83]. Goethe [BL94, Jac94]. Going [MS88b]. gold [Car14]. Golden [Per04, RH90, Wil91]. Good [Geu03, Son95, AC22, IJJ13, ST24]. Gooding [Gor10]. Goodman [Old93]. Gopakumar [Row12]. gordos [GDLB15]. gordos/país [GDLB15]. Gorman [Ano77j, FR91, Lat92]. got [VI17]. Gothic [Atk05]. Gould [She02]. Governance [BD11, Irw06, SFH12, ÅSS21, CBLG23, Did25, FS23, Gus14, KH22, RG23, Sad24, SS17a, Sch13, SZ11, SSB24, Sti18, WM17]. Governing [FFMW08, Pol20, Sha20, SSB24, Thé09, Hen18]. Government [Hol81b, Hut71, Lez06, Lon71, DG17, SV22, Sma20]. Governmental [Shr00, US20]. Govind [Row12]. Graduate [Blo77, Bru77, FG73]. graffiti [DP21]. Grafton [Mos79]. grammar [Lez10]. Grant [GGJ07, Lan01b, Wes88, PW19, Abr91, Jab95]. Grants [CG89]. graphene [Yan22]. Graphing [SG74b]. Graphs [CG92, LMT92, SBS^{+}00]. Grasping [Dor89a]. Gravitational [Ash07, Col99c]. gravity [GL21, Ash07]. Great [Del05, Mul93, Pri71, Wri80, GEJ79, Jan18, Muk07a]. Greek [Lat08]. Greene [Sis10]. Grid [WN22]. Grid-scale [WN22]. grids [TH24]. Gross [Hol81b]. Ground [McK99, Jat17]. grounds [DS20]. groundwater [UB25].

Groundwork [Jat23]. group [Arm22, DH15, Ste12, Whi75a]. Groups [Sch90, Rou17]. Growth [Coh86, Gil86, Inh77a, MC71, Ram02, San78, Urb82, Bar79, Ano77f]. guarantee [LSW24]. guessing [Sok23]. Guest [Clo03, Hac05b, Shr05]. Guest-Edited [Clo03, Hac05b]. Guidance [MS88a, MS88b]. Guide [Del05, Dun78, MC71, Lor17]. Guideline [Cas09, Kna13]. Guidelines [MO09, Fro12]. guides [TH24]. Guinea [MD09]. Guinean [Str11]. GWP [CBLG23]. gynecology [BM19].

H [Ano77e, Del02, Mia99]. H. [Jas03a]. habitat [Car21b]. Habituating [AS16]. Hackers [Max17]. Hague [GM78]. Half [Pin99b, Red02, KRR13]. Half-Life [Red02]. half-lives [KRR13]. Halfpenny [O’R83, PM89]. hallucinogenic [She20]. Hand [Ash05, Doi04]. Handberg [GH84]. Handle [CB01]. hands [Fis18]. Hans [Vin99]. Happen [Wil88]. happened [Tur12b]. Happily [Jas00]. Haraway [Cri10]. Harbor [Car21b]. Hard [Cam93]. hardback [Blo74, Out76]. Hardcore [Wak04]. hardcover [Luk75]. Hardest [Chu82]. Hardin [Pin81a]. harmonization [JSLS25]. Harold [Lyn11c]. Harry [Doi11, Ash07]. Hart [Har83a]. Harvard [Del05, Fis10, Goo10, Ros10]. Harvest [Mil84]. Harvester [Rip91]. harvesting [HJ13]. Harvill [Hol81b]. Harwood [Ano77o]. Hatice [Del02]. haunted [De 24]. Haven [Rob99]. Hawking [Lyn14]. Hazard [Wet07]. Hazards [Kop79, Ray86, Wri86, Cur22, LA19]. Haze [RH90]. hbk [Ash07, Col07, Cri10, Del05, Doi11, Eps09, Fis10, Gin10, Goo10, Hyy09, Mae07, Moo10, Pra09, Rei06, Ros10, Sis10]. Healing [LA19]. Health [Arm09, Blo91a, Eps09, God97, Kin02, MS24, Mul72, SN89, Wri80, vEG91, Arm10, CSTG23, Cra10, Eps21, HBP19, Hoe19b, Hog19, KBRB19, Mur15, NL18, Pic13, Pih25, Pol14, Sar23, ST24, Val10, VT21, ZDI^{+}23]. Health-related [Arm09, Arm10, Val10]. Healthcare [BL13, Goo18, ZJ09, JZ12, Oud12, Pic19]. Healthy [LR01]. Healy [NK14b]. Heart [BH91, Gar97, JS18]. heart-lung [JS18]. Hearts [Hug24]. Heating [Rus94]. Heavy [Wak04]. Hedfors [ABB^{+}08]. Hedgecoe [Rei06]. Hegemony [Van08, Wyn03]. Heidegger [Koc10b, Rii08, Sch12]. Heights [Gin99]. Heinemann [Mos79]. Helen [Ken15]. Helix [Shi02, Sto82]. Helm [Wri80]. Help [EES96, dV07, AH20, Skr16]. Helping [LW10]. Hemel [Rip91]. Hempstead [Rip91]. Henrietta [LP11]. Henry [Del05, Bro89, Mer98b, MR73]. HEP [PN87]. Heralds [Law20b]. Herbert [Ano77f]. Here [Har02a, Ler18]. Heritage [Egl97]. Herts [Rip91]. Hessen [Gra85]. Heterodox [Hes87]. Heterogeneity [Mor06, Nic96, SDL^{+}14]. Heterogeneous [Aar23, Del08]. Hicks [Fra88]. hidden [Cha22]. Hierarchies [Fau00]. Hierarchy [SBS^{+}00]. Higber [Blo77]. High [Cra80a, Hod83, IM85, KC95, Pic80a, Pri71, Foc16, Gu23, Heg16, Mac17, MB22, SDSR25, Tan21b]. High-Energy [IM85, KC95]. high-frequency [Mac17]. high-tech [Foc16, SDSR25]. high-temperature [Gu23]. Higher [Bru77, Heg16, Wes86]. Highest [Blo77]. Highly [Cam93, NFC83].

Highly-Cited [Cam93]. Hill [Blo77, Mir07]. Hills [Nev80]. Hindsight [Kre71, Bra24]. Hiroshima [Lin16]. Hirschauer [Lyn94a]. Hirsh [Han84]. Historian [She02]. Historians [Car05]. Historical [Boh89, BW83, Bru80, Elk78, FG86, Gas92, Ink85, Nev80, Old90a, dSP75, Thé04, Wri80]. Histories [Den87, Wal01, Sah12]. historiographical [Mar17]. Historiography [Boh91, Nic96, Wei74a, Wei74b]. History [AA84, All76, AB80, Ano09a, dBB76, Boh99, Bow96, Buc75a, CM09, Del05, Den87, Dug87, Eas73, Eps98, FG73, God98, Goo10, Gra85, Har83b, Hun82, Jas00, Jen05, Kuk94, Mac78a, Mac81, Mar05, Mar04, Meh80, Mil84, Mil04b, Mir72, Mor88b, Mos79, Nev80, Nic97, O’R83, Ola96, Old88, Pan80, Rei80, Roc81, Sec85, Sha80, SS99a, SS94, Tom87, Twe93, Wei74a, Wei74b, Wei90, Woo76, AQK15, Cur22, Ens12, Gar13, Gu23, Joh18, Mar17, Sil11]. hitchhiking [CB24]. hitherto [Jön16]. Hits [Dow07a]. HIV [Cra10, Dav20, Eng20, Gre16a, MS24, Mon12]. hoc [Lyn11a, Sis16b]. Holds [Dow07b]. Holmes [Wes88]. Holocaust [ABB^{+}08, Hed08b, Hed08a]. home [GK16b, Lac16, Rob16]. homes [Mar13]. Homo [Dav20, KGDLB^{+}15]. Homogeneity [SDL^{+}14]. honey [SK13]. Honeybee [Cri04]. Honoring [Car05]. Honouring [Old93]. Hope [Mul93, Pol14]. Hopkins [Sis10]. Hopwood [Mae07]. Horae [RG16]. Horizon [Kan21]. horizons [Mcl21]. Hormone [Mat13]. Hormones [Oud90]. Horses [Ful90]. hospital [Rob16, Str16]. Hospitals [Ray86]. Hot [Hin14]. House [McC92, Mir07, Pin99b, Ral06]. House-Building [McC92]. Houses [McK99]. HP8 [O’R83]. Hubert [Col92b]. Hud [AB80]. Hughes [Bar84a, Car05]. Hugo [Ano77g]. Human [Bal96, Blo92, Fee00, Gie89, Gie91, Goo10, Ham82, Jah72, Kir99, Mar04, MM98, Mul94, Mul95a, Nol92, OS05a, Pet15, Pra06, Twi05, BSB16, Dah24, Eps21, FR11, HRI22, Hol22, Hoo17, Joh13, Jon17a, Lip23, Mey22, Nel13, Pih25, Pop21, Rad13, Reu12, Smi17, Str16, Rea01]. human-machine [Lip23]. humanitarian [Red16, SS18]. Humans [Vos04, FÖP20, Leo12]. Hundred [Dea74]. Hungary [MF95]. Hurricane [Shr07, Sim07a, Shr14, Sim07a]. Hurt [Mod01]. Hybrid [Tuu05, Yan22]. Hybridization [JM01]. Hybridizing [CGP25]. hybrids [KLBKC11]. hydraulics [Pri12]. hydrocapitalism [Pri12]. hydroimperialism [Pri12]. Hydrothermal [Ore03]. Hypatia [Del02]. Hype [Rei06]. Hyper [Edw94]. Hypertension [Edw94]. Hypo [Tir18]. Hypo-interventions [Tir18]. Hypothesis [Oro77b]. Hysterectomy [Woo06].

I. [Buc92]. ian [Sin98, Twe93]. IAPT [Pic19]. ICD [Eps21]. Ice [OOO12, Var20]. Idea [Dun74, Ful89a, Nol90, Arm17, Iid15, Old88]. Ideal [Por92b, Thé04]. Idealism [Blo96]. Ideals [Whi75b]. Ideas [MK84, Yon94, Del21, PW19, PP24, Pin21]. Identification [Col98, Pel21, Bie18, Plá23]. Identifying [Bri13, SG74b]. Identities [Fau07, Hec96, LJ98, Tim96]. Identity [Dow88, DH08, Hac05a, Kan78b, Mar10, MD09, Nan72, Ski06, Wyn96, Hor25b, Kan22, Ken13, PD19, PAB15]. Ideological [Ger01, Kli02]. Ideologies [Hec96, Kin02]. Ideology [Fri84, Hil86, Lei75, Lyn94b, MG92, Man87, McC88, You71]. idiom [Sim25]. Idylls [Lyn95b]. Ignorance [Dur20, Jeo19, Pin15]. ignoring [Sok23]. II [Ano77o, Ala20b, Bea77, Bud78, Bug93, FS93, GSSD74, Har77, MS88b, Res82, Sis12a, Wei74b, Ano77e]. Illness [Wri80]. Illumination [Ano77j, Del05, Gor77]. Image [Som06a, dG75, Fro12, Jat17, Law25]. Imagery [JS89, Kim94]. Images [Ala08, Joy05, Lyn85, Hoe19a, Nev80]. Imaginaries [Hyy06, Suc23, Hug24, Kan22, LR20, PT18, Sis20b, ST16, PJ17]. Imaginary [BS01, GS25, Mag17, RvD22, SDSR25, Sch20, Sma20]. Imagination [BD11, Law20b, SDSR25]. Imaginations [Atk05, JM73]. imagined [ÅSS21]. Imaging [Clo01, Joy05, Pra05, Pra07, Coo11]. Imagining [RA22, SBH^{+}20, Jun23]. Imaginings [Mac03b]. Immigrant [Tor83, Tor94]. Immigration [Ala09b]. immobile [Rei23]. immobilization [Oli25]. Immortal [LP11]. immuno [CGP25]. immuno-oncology [CGP25]. Immunology [SS94]. immunotherapy [CGP25, Jat23, Swa24]. Impact [AA84, DC76, Hol81b, LB05, Lic89, Mcl21, NHM74, Nel75, SL91, Vin91, KG12, LOBB23, LSW24]. Impairment [MMB09]. Impartiality [Yea92]. Imperfect [KKLL21, TL11]. imperialism [BFW14]. Implant [Jas02]. implantable [Oud15]. Implications [GR08, GHR00, HH83, Mul79, PT72, Roc74]. Importance [Hen90, Ken13, Sov10]. Imports [JS18]. Impossible [Coo19]. Imposture [Cal99]. Impractical [Wyn88]. Improving [Har02a]. improvised [Tak24]. Improvising [Suz21]. imputation [Wad14]. Ina [Ano77n]. incl [O’R83]. Include [Lag07]. Inclusion [Eps08, Sis10, Sis10]. inclusive [Par22]. incognita [Mes17]. Incoherent [Wal00]. Income [McC92]. Incommensurability [Nad83, Zan84, Lez10]. Incompleteness [Gum20]. Independent [MB25a, Sim78, Wes79]. Indeterminacy [Sch10]. Index [Ami74, Ano02, Ano04, Ano05d, Ano06b, Lin78, Bru77]. India [Row12, Bha76, GS02, Hai19, Har83a, JS18, Pra05, Red75, Sah25, SFP11]. Indian [Bes20, Cho85, Mah14]. Indiana [AB80]. Indians [Jac75]. Indicator [Nad83, SR72]. Indicators [SWB77b]. Indices [DC76]. Indies [HTB20]. Indifference [Dun74]. indigeneity [Tal13]. indigenization [Ler18]. Indigenous [Gar99, Ken13, KRR13, Kow13, LOBB23, Mar08b, She06, Tsa24, JS18]. indigestion [Bes20]. individual [GKSS24, MvO23]. Individualism [Hut91, Sat07]. Indonesia [Smi14]. Inductivism [Blo77]. Industrial [Bar71, Eri71, Hol81b, Lic89, MF95, Twe93, Ver88, Oli25, Pih25, Sch20]. Industrialisation [Hol81b]. Industrialized [McC92]. Industry [And88, Bur75, Don86, Eva10, Lam10a, MK84, MM98, Mül95c, Ros93a, San72, Win02a, Win02b, Bes20, Byc16, Gar73a, HH22, VI17, Wad14]. ineffable [SC18]. Inequality [All80, SL91]. inequity [RS11]. Infant [Huf86]. Infanticide [Ree01]. infection [BSB16]. Inference [Blo92, Nol92]. Inferences [TY85]. Infertility [Cus96, Ker00]. Influence [Doe03, Hil74, Hut71, Son09, SH89, Vel90, Tor94]. influencers [Bia22].

Influences [Hea04]. Influential [HSP02]. influenza [MD20]. Informal [Mar20, WG74, Zin71]. Informatics [Bea01]. Information [Blo91a, Har02a, Mac93, Smi09, Fro17, Rie15, Sok23, SP24, TL18, VFWC15, Wes22, SAC^{+}20]. informational [Fro17, Jan18]. informs [AH11]. Infrastructural [Cal21, Mac03b, HJ23]. Infrastructure [Pin21, RP15b, Sim07a, Sim07b, Str16, Aar23, Ben12, Car12b, FW17, Fro17, Ken17, Mos16, RP15a, SA22]. Infrastructures [Cha22, VT21, Gum20, Ure14]. Infrastructuring [Aar23, LD13, VSS22, Mic20]. Ingenious [MR73]. Ingrid [Del02]. Inhaber [Ano77f]. inherently [Has12]. Initiative [Lic89, Pic19]. Initiatives [McD97]. injury [Hol22]. Inkster [Twe93]. Innovating [VA16]. Innovation [AR02, Bal95, BS04a, Byc16, CKM04, GH83, Gra75, Hac92, Hol74, LT92, Nor78, Rud96, Tor92, Bir23, Bra14, CGPP24, Gar13, God15, JS18, KLBKC11, Par22, PI19, PJ17, PW10, Rot19, SSB24, VSS22]. Innovations [Leo10, Nic96, Sch90, Ver88, BD11, CS20, Pih25]. Innovator [Nis95]. inquiries [Goo18, LW10]. Inquiry [Lyn82, Mer02, O’R83, Rem88, Yea89a, Lat13]. Inscription [SBS^{+}00]. Insect [Cri04]. insecticides [SK13]. Inside/ [Ash05]. insides [Dán20]. insight [Ell11]. Instability [Mil04a]. Institute [LRC87, Mos79]. Institutes [Gra75, Lau84, SB77, Ano77q]. Institution [Mos79, Ray86, CERG25a, Ber72, Ber08]. Institution-building [Ber08]. Institutional [Abr93, GB09, Han86, Hun82, Man93, Pal90, SK91, SR73, SG89]. institutionalization [Byc16]. Institutions [Blo04, Cla80, Moo10]. Instrumental [Blo00, DSJ19, Sla12]. Instrumentalized [Kus10]. Instruments [AB80, BS04a, CB01, Hin06, Sta02, dBtMV21]. insurance [Jau15, Sad24]. Integrated [CT89, Tur89c, Vis15]. Integration [Ake07, Aar23]. Integrative [BH03, PP21]. Integrity1 [MO09]. Intellect [Ink88]. Intellectual [Kra72, Tor94, Woo76]. Intellectual-influence [Tor94]. Intelligence [Ano77o, Bra89, Col92b, CL89, For93, Har76, Har77, Kuk94, Wes77, Wes78, vK01, Ens12, SM25, Suc23, SS91]. Intelligent [Gui98, Lam06]. Intelligentsia [Hol81b]. Intended [Wil91]. intensification [Smi17]. Intensive [Bro89, Car14, PPM20]. Interaction [Ala08, Ala09a, Ver12]. interactional [RL16]. Interactions [SWB77b]. Interactivity [HvL08]. Intercourse [Wag09]. Interdependence [WSB79]. Interdisciplinary [Nic97]. Interest [Haw73, Wad14]. Interests [Abr93, Abr94, CL82, Chu86, Gie89, Gie91, Lau87, Mac78b, Mac81, Mer02, Ste86, Ste87, Woo81b, Wyn92b, dMdF98]. interface [Leo12, MvO23, ZDI^{+}23]. Interfaces [Col07, Plá23]. Interfacing [Sor23, Lip23]. Intergovernmental [OOO12]. Interim [Kah08]. Intermediaries [MM98]. Intermediary [Oud97]. Internal [Tor92]. Internal-Combustion [Tor92]. Internalists [All76]. International [AR02, AB80, FC79, Gra75, Hal06, Hol81b, Kwa87, Lub81, Rad13, Ran17, SM97, SG73, SRM76, KVA21, RC16, Kwa05, SG74a, Sko73]. Internationale [Sal71]. Internationalism [Hod83]. Internationalized [Pop21]. Internet [DYS^{+}05, Gil06, Gre22, Shr05]. Interoceanmetal [Gal24]. Interplay [Mac03a, Ste78a]. interpret [GB16]. Interpretation [AdSPG^{+}76, AH11, KdR19]. Interpretations [Mul94, Pin97]. Interpretative [MP87]. Interpreting [Ano77q, Bar90, SB77, Yea01, SBHY15]. Interrogating [Gar99, Mil04a, Oik17]. Interscience [GM78]. Intersection [Mel09, Tol19]. Intersections [Ful89b]. intersects [Lei14]. intersex [Gri18]. Interstitial [Sta02]. intersubjectivity [Ala20a, Ala20b]. intervention [Plá23]. interventions [Tir18]. interview [Tre13]. Interwar [Mod00, Ras04]. intimacies [Sch15]. Intimate [Hea04, Tir18]. Intra [Kum24]. Intra-mediary [Kum24]. Intrinsic [Rad86]. Introduction [And02, BFS95b, BA12, Bra92b, Coz86, Dán20, Eas73, FDR08, Gas90, Hac05b, HBP19, JE75, LMR10, Lyn11a, Por92a, Sis16b, Pic19]. Introductory [CKSW09, Smy01]. Iñupiat [Lan20]. Inventing [Gus06, PW95, Sab14]. Invention [CG90, Gor97, Mye95, Pra07, Wes79, Dav24]. Inventions [Pat83]. Inventors [Pat83]. Inverted [Mil76]. Investigating [Lyn13, Wat90]. Investigation [KMH83, NFC83, PC84, Lez10]. Investigations [Hir94, Lyn82]. Investment [Law80, Lic89, MS14b]. invisibility [HR21, LSS20]. Invisibilizing [Jeo19]. Invisible [MC97, Kuc13, Lee16, Nis11, SMW15]. Involvement [Mar07]. Ionospheric [Gil86]. IPCC [Lah21]. IQ [Har83b]. Iran [Sha20]. Irish [OS05a]. Iron [BPBG06]. irreconcilable [Lez10]. irreduction [Jen23]. irresponsibility [Mey22]. Irvine [Bud85, KP85, MvR85]. Isaac [IM83b, IM83a, Smi83]. Isaiah [Fan09]. Isar [Tjh24]. ISBN [Alb95, AB80, Ash07, Bar79, Blo77, Bro95, Bru77, Col07, Cri10, Del02, Del05, Doi11, Eps09, Fis10, GM78, Gin10, Goo10, Gre10, Hol81b, Hsu13, Hyy09, Kim94, Mae07, Mir07, Moo10, Mos79, Nev80, O’R83, Out76, Pra09, Rei06, Rip91, Rob99, Ros10, Sis10, Sto08, Wes88, Wri80]. ISBN-0-906450-00-4 [Nev80]. Island [Her03a, Bur75]. Islands [Hen18]. Isolation [Dun74, SM97]. Isomorphism [SK91]. Isotope [Hon08]. Israel [Lei14, LS10, PAB15, Pra06, Tor83, Tor94]. Israeli [Yai19]. issledovania [Rab76]. Issue [Clo03, Hac05b, Joh84, Thé09, AH20, BA12, HBP19, Sis20b]. issue-politics [AH20]. Issues [CK85, Gum90, Mar07, ML93, dB09]. Italy [Pan80]. Itera [Hul92]. iteratively [Dur20]. itineraries [Mur15]. Itself [Dae98]. IV [Rud82]. Ivan [DGC18]. Ivory [Hun24, Lac16, Lam10a]. Ivory-billed [Hun24]. ix [Rob99]. Izd [Rab76].

J [AB80, GM78, Gor08, Hol81b, Lau90, Mos79, Nev80, Old88, Pri71, Wri80, Wyn76]. J. [Ano77h, TS00]. Jackson [BL94, Mar05]. Jacqueline [Ves96]. Jacques [Fan09]. Jagadis [Nan72]. Jahrhunderts [Bay83]. James [Par88]. January [Blo03, Bre02, Col22, PCL98, Tur95b, Wes79]. Japan [dSP75, Fuk16, Hod83, Hug80, Iid15, Ink91, KIH04, LA19, Low89, Nak12, Pol20, Sab14, SFP11, SFH12]. Japanese [Suz21]. Jaroslav [Efm92]. Jay [She02]. Jeremy [Sis10]. Jericho [SP24]. Jerry [Mir07]. Jewish [RA22].

Jizz [Ell11]. Job [CC98, Mos79, Pal94, TL11]. Joe [Ash07, Zie88]. Joe-1 [Zie88]. Johannsen [Kim91]. John [Ano83b, Bur75, Mac84, Ano77i, Ger95, Ste90]. Johnny [Mir07]. Johns [Sis10]. Johnston [Col85a]. Joint [Ano09a, Gal24]. Joking [MG82a, Sco92]. Jon [Rob99]. Jonathan [Ano77o]. Jordanova [Nev80]. Joseph [Del05, FL88, MR73, Res96, Tor92, Wes86]. Josephson [For99]. Journal [Gun05, SM93a, SR72, SL91, Fro12, KBvLA22]. Journalist [MBB00]. Journalistic [Kid88]. Journals [Ano09a, Gor84, Har02a, Lev79, STGG95, KVA21]. Journeys [Fau98]. joy [Ell11]. Jr [Blo77, Hol81b]. Judge [Tai19]. Judgements [STGG95]. judgment [Cho13, Fis10, Fis10, Ros10]. Judgments [Hir10, Rou17]. July [EHB^{+}97, Meh80, Tur95b]. Jumping [Dor89b]. June [EHB^{+}97, Sis22a]. Jungle [Wyn92a]. Junior [LMA80]. Junior-Authored [LMA80]. Jupiter [Gol94]. jurisdiction [WM17]. Just [Kel89, Say14, Wel04, SS17b, Wie17]. Justice [All07, Lyo18]. Justification [Mah80].

K. [Blo74]. Kampala [HHW^{+}19]. Kantor [Mag78]. Kaplan [BCM76, Sal76]. Karen [Pin11]. Karl [Bim90, Nor78]. Karpf [Bru77]. Kass [Bri09]. Kathryn [Del05]. Katrina [Shr14, Sim07a]. kava [Bak11]. Kay [Bro95]. Keech [CC77]. Keele [Nev80]. Keeping [Gre16a, de 02]. Kegan [Bar79, Blo77, Luk75]. Keith [FPM03, LcaM82]. KEK [Hod83]. Keller [RS89b]. Kendall [Hol81b]. kennel [Kro19]. Kenya [Sar23]. Kenyan [GP09, PP22]. kernel [RP15a]. Kevles [Mos79]. Keynes [Mos79]. Keywords [Whi89]. khimii [Rab76]. kicks [HH22]. killed [MS14b]. Killer [Sch10]. Kind [Hil86]. Kinder [Sto04a, Lyn04e]. King [CE03]. kitty [Ial21]. Klapping [vHVdW07]. Kleinman [CLP91]. Knight [Mos79]. Knopf [Mir07, Mos79]. Knorr [GP97, Sis93a]. Knorr-Cetina [GP97, Sis93a]. Know [Col04b, Gie06a, Lyn98b, Van08, BL12]. knowers [MR17]. Knowing [Bud98, Fle94, Par22, PK09, Stu07, M’c23, Str11]. Knowledge [AA82, Ake07, All07, Bal06, Blo71, Blo74, Blo92, Blo96, Coh86, Col98, Col74, Col82a, Col96b, Col01, Del83, DA01, Dol71, EM00, Eps08, For93, Fre84, FL88, GP09, Gie92, Gil76, Gil86, Gor90, Gor06, Gus06, Har81b, Har86, Haw73, HP91, Hor04, HI10, Hug06, Jas97, Jas02, Jen07, Joy05, Kim94, Koc10a, Kus04b, Lin78, Luk75, Lyn94b, Lyn98a, Mac78a, MS00, McD97, Mey77, Mil78, Mod01, Mul79, Mye85, Old90b, Oud90, Pal02, Pic90, PC84, Rad86, Sch72a, Sha01, Shi02, Sin96a, Sin96b, Sør93, Sto08, TJ85, Tur90, Tur91, Vau99, Woo06, Yea90, Yox87, de 02, Bea10, BB17, BSB16, Col14, CC12, DU15, Doi11, Dur20, Elg22, Eva10, FLK^{+}14, Foc16, FC12, Hed12, HH22, Hof15, Hun24, LDM^{+}16, MN23, Mül18, PT18]. knowledge [Pic13, Pol14, iO18, Rei13, SS17a, SMW15, SH12, SPBG24, Tan21a, Tur12b, US20, VSS22, da 22, Sim91, Doi11, Kim94, Nev80, Ano77g, Sha04]. Knowledge-Making [Mod01, Pol14]. Knowledge/Practice [Tur90]. Knowledge1 [Col89]. knowledges [LOBB23]. Koertge [Col99a]. Korea [Kim16, SFH12]. Korean [Kim07]. Kraken [Mac99b, Mac00]. Kranzberg [Gie96]. Krause [Wri80]. Kuhn [EHB^{+}97, Col12, Ful12, Lyn12b, Nak12, Pic12, Pin97]. Kuhnian [Bri72, Cra80a, Lam06, McD76, Pic80a]. Kuhnians [Pin97]. Kurt [AB80]. Kusch [Blo04, Sha04].

L [Col92b, Hol81b, dSP75]. Lab [Hon08, Leo10, AC22, Cal21, PP24, Ros11, Yan22, Doi04]. label [Fro17]. labeling [MD20, Sch13]. Labelling [Kli02]. Labinger [Col95, Ful95, Hak95, Kei95, Lyn95a, Mar95, Sto95]. Labor [Doi04, Lóp22, LDM^{+}16]. Laboratories [Den87, Gie06b, Hut71, BL13, Wyl15, Wyl19]. Laboratory [Asd08, Har94b, Kim07, Lyn82, Lyn88, Sco91, Sim99b, Sim05, Woo82, dB09, Ala23, Dem24, Hen18, PPM20, SO17, Suz21, Tou13, Ala08, Mar05, Sut84, VP98, vK01]. Labour [And88, Har96, Str16]. labs [Sub25]. Lacks [LP11]. Ladd [Blo77]. ladenness [Tho25]. Lady [Goo10]. Laetrile [PM79]. Lamont [Fis10, Ros10]. Lampert [Hol81b]. Lamto [Lac16]. Lancers [Har83a]. Land [GGJ07, PT18, Pic99, Rei80]. Land-Grant [GGJ07]. Landlords [Gre22]. Landowners [Ver02]. Landscape [SK08, Tjh24]. landscapes [BFW14]. Lane [O’R83]. Language [BV95, Col11, Cri04, Mye91, Por04, Rib07]. Large [MB81, McC00, KHM10]. Large-Scale [MB81]. Larissa [Ves96]. Larry [Bar79]. Laser [CH75]. Last [Gie82a, Lat86]. late [Kus10, Pop21, Ale08, Don86, Lau84]. Latent [Col98, Rad13]. Later [Cam93, BJ12]. Latin [Bri14, Cue90, Ves87, WLBRS15]. Latina [Lóp22]. Latina/o/e [Lóp22]. Latour [Sis22a, Akr23, Ano10, CG92, Ful12, Koc10b, Lyn12b, Mar23, Mia12, NTW^{+}23, Por23, Rii08, Row05, Sch12, Sch88b, Sco92, Tre13]. Latourian [Jen23, SBS^{+}00]. Laudan [Bar79, Ste87]. Laughter [FLE97]. launch [Sim10]. Law [dBB76, CB24, EM00, Edm02, Fau12, Ger95, LJ98, Pot11, Ram02, Van83, Yea85, Coo19, DSJ19, Edm11, Pic13, AdSPG^{+}76, MM82, Bar90]. Law-Science [EM00]. Lawrence [Sut84]. Laws [Far75]. Lay [KCBT07, Shi05, Dav24]. Layer [Har03]. Lead [CD84]. Leadership [Har03, MM81]. Leading [Zie88]. leaf [Sch15]. Learn [Dol75, CS17]. Learning [Sol85, Tra81, Sti18, US20]. Leave [Lat86]. Lectures [Blo91c]. Lederman [Del02]. Leeds [Out76]. Left [Ash05]. Left-Hand [Ash05]. Legacies [MD09, Hal19, KKLL21]. Legacy [FS93, Man93, Sol01b, HRI22]. Legal [Edm02, Gol04, Swa07, Joh18]. Legalization [Ful06, Sol92]. Legends [WC99, Rek14]. legislation [Fau12]. Legitimacy [Agr01, McL15]. Legitimate [Cho13]. legitimating [Jeo19]. Legitimation [RH96, SM97, GM23, Sud22]. Legitimization [Mel07]. Leicester [Pri71]. Leigh [Bal10, Del05]. Lemaine [GM78]. Lenin [Hol81b]. Lens [Boh04]. Less [AB80, HSP02, Rav99]. Lesser [Dro93, Fra79]. Lesson [Luc01, Swa07]. Lessons [yLL14, Wei74b]. Let [Bro95, MN23]. Letters [Sil74]. letting [dBtMV21]. leukemia [Lev22]. Levees [Muk07b]. Level [HH83, MM86].

Levels [Gor02]. Levers [Sco91]. l’évidence [Gre10]. Lewin [Pic97]. Lewin-Epstein [Pic97]. Lewis [Hol81b]. Liberal [Tho02, LOBB23]. Liberation [Eas74]. Library [Bru15]. Licensing [OS05b]. Life [Ash05, Col92b, Col94a, Cri96, EM00, Kro19, LP11, OS05a, Pal94, Pin81b, Pra06, Red02, Thé09, Vin99, Wat90, Arm19, Bie18, GK16a, GK16b, Hen18, Hoo17, Jau15, Mor19, Rad13, Rah11, Sad24, Wes88]. Lifecourse [Rie03]. LifeStraw(R) [Red16]. Light [ABB^{+}08, Fra76, Hed08b, Hed08a, Sha01, VP98, Mod12]. Lighting [Vin95]. Like [Col85d, Mae07, Bre13, VI17, Ver12]. Liminal [Rud96, McL15]. Limitations [Hic87, Hic88, Naf24]. Limits [Ano77g, Mey77, San78, HR21, Lak17, She20, Ure14]. Limoges [Kle91]. Lincoln [Bru77]. Linda [Hol81b, Wei90]. Line [Kim91, Fro12]. Linguistic [Sha01]. Linking [Len79, Mat14, RG14, Sat07]. links [Rit16]. Lipset [Blo77]. liquid [Car14]. Lisa [Del02]. list [FSF24]. Listening [EP07, Hun24, Wat14, Sup14]. Listens [Bar87]. Listing [Rud77]. Lists [Har02b]. Literacy [Fou97]. Literary [Sha84, Sin96a]. Literature [Ano77p, GW74, Gun05, Oro77a, PN87, PC84, Win76]. Literatures [BH97, GSSD74, SG74b]. Litigation [EM00, Jas02, Kir12]. litter [Ial21]. Little [Mod01, AH20, Avi15, BL13, Mar05]. Live [Jas00, Bij12]. Livermore [Sut84]. Lives [Del05, Fau98, Tim96, KRR13, Lau25, LB19, Pet15, Sis19, ST16]. Living [Blo87]. Lizards [RB99]. Loathing [Ful90]. Local [CK85, Hår94a, Kwa05, MC78, TB97, Tur91, Hed12]. Location [Hut91, Bea10, Eri72]. locative [SAC^{+}20]. Loeb [Fan09]. Logic [KMH83, Lab04, Vin95, Gre10, RHSB19, Gre10]. Logics [Sad24, Wes24]. Logo [AWN06]. Lomnitz [Ves96]. London [AB80, Ash07, Bar79, Blo77, Bro95, Del02, Hol81b, Hyy09, Jah72, Kim94, Luk75, Mos79, Nev80, Wes88, Wri80, Mic20, Ver08]. Long [BPBG06, CM09, HRI22, Mar17]. Long-Term [BPBG06, HRI22]. Longitudinal [AQL94, Sma77]. Look [Lau84, Mae07, Sim78, GS11]. loop [Jon17a, OL18]. loops [Dur20]. loose [Col13]. Lord [Gui98]. Loss [Sil74, Woo74]. Losses [Elz04]. Lost [Ink91, Joe94]. Lotka [AdSPG^{+}76, MM82]. Louis [Ful90]. Love [Mar08a]. Low [DH08, FO22, McC92, Law13, LR20, SBH^{+}20]. Low-carbon [FO22, LR20, SBH^{+}20]. Low-Income [McC92]. Low-power [DH08]. Loyalties [SG73]. Ltd [Jah72]. Lubrano [Hol81b]. Luck [DP23]. Ludmilla [Nev80]. Ludwik [ABB^{+}08, FL88, Har86, Hed08b]. Lung [Blo00, JS18]. Lynch [Kus04b]. Lynn [Ano06c, Gus06, Hug06]. Lysenko [Iid15].

M [Boh91, Jas03a]. M.I.T [dSP75]. MA [Del05, Fis10, Goo10, Moo10, Pra09, Ros10]. Mårup [JM01]. Machine [Col92b, Gar73b, Ger01, Hsu13, Lyn98a, McH09, Pre05, Sis09a, Sis09b, Sti18, Tur01b, Ken17, Lip23, Phi16]. Machine-Readable [Gar73b]. Machinery [And88, Sma20]. Machines [Ano77i, Bud98, Mar77, Vos04, Hof15, JS18, Pot11]. MacKenzie [Sul99]. Maclaurin [Gra98]. MacLeod [GM78]. Macmillan [Hol81b]. Macro [GSSD74, Sat07]. Macro- [GSSD74]. MacRoberts [Sti87]. Macroeconomic [Eva97]. Maddox [Pic94]. Made [Ink91, Por96, MRL01]. Madness [Tom87]. Madrid [LS04]. Magic [Col07, Wei90, Smi15]. Magnetic [Joy05, Pic81, Pra05, Pra07]. Magnum [Ash07]. Mahoney [Gor08]. Main [AB80]. Maintaining [CGPP24, Cho13]. Maintenance [DP21, HJ23, RB22]. Make [Aga06, HR87, CGPP24]. Makes [MB81, Son95, Jon17b]. Making [Ald98, Ano16a, Bar71, Boh99, CLP91, Car12b, CC98, Col07, Den85, Gre16b, HV93, Has12, Hen00, HLRK14, JS18, KH22, KVA21, Lan01b, Law84, Lez06, Mar05, MB25a, Met25, Mil01, Mod01, Mon12, Oud90, Pal02, PAB15, Ras20, Rot05, Rot19, Sch84, Sch16, Sha16, Sim25, Sim99a, ST24, Thé04, VP98, Vin99, Vos04, Wya07, ZDI^{+}23, dlB15, BHC16, Che23, Eva22, FSF24, FFMW08, Gra07, Ham19, HL19, Hun24, Jac19, Kis13, Lag12, LSS20, Mey22, PP22, PS24, Pol14, Pot11, Sah25, SFH12, Wag13, dBtMV21, Col07, Elz04]. Malaria [Dal90, Tur89a, Tur90, BSB16]. Malarial [lL99]. Malaysia [Mel09]. Malcontents [Cor94]. Male [Ker00]. Malthus [GT77]. Malthusianism [Mer22]. Man [MRL01, Wra03, OL18, Sud22]. Man-made [MRL01]. Management [Bug93, Her03a, Jab95, KHM10, Ray86, Sch94, CKFS13, DH15, HLRK14, Hog19, Kna13, Law20a, MB22]. Managers [Hol81b]. Managing [Bal96, Cam03, Pet17, Püt21, Rap05, Rei23, Suz21, Tol19, Car12b, Gre16b, Ras20]. Maneuvers [Sis04]. Mangin [NK14b]. Mangle [Nic96, Pic02]. Manpower [RS74, SR72]. Manufacture [Hir91]. Manufacturing [Bug93, Fis04]. many [Ben12, LB19]. Map [Ran17, Shi03, Ver08]. Mapping [Bal96, CKM04, Yea01]. Maps [CG92, LS10, Mes17]. Maralee [Del02]. Marcellin [Ham82]. March [Con79, Res96]. Marcia [Mir07]. Margin [And08]. Marginal [Sim84]. Marginalia [GH84]. Marginality [GH83, Sim84]. Marginalization [Waz13]. Marginalized [JSLS25]. Margins [Nev80]. Marianne [Goo10]. Marie [Wir15, Pyc93]. Marine [Van83, dB09]. Mark [Ano77h, Hol81b, WL15, Ano77p, Moo10]. Marker [Eps97]. Market [Lan09, Bre13, CS20, Del21, Ken17, LR20, SO17, WH21]. market-like [Bre13]. marketplace [Par22]. Markets [Pre06, ZJ09, MB22]. Marking [Smy01]. marriage [KG12]. Marriages [Wag09]. Mars [Ver12]. Marshall [Ano77i]. Marta [Del02]. Martian [Mes17, RB22]. Martin [Blo77, Wri80, Blo77, Blo04, Bud85, KP85, MvR85, Rii08, Sha04]. martyred [Smi17]. Marx [Bim90]. Marxism [Gra85]. Mary [Del02, Del05, RD08]. Mass [Blo77, EM00, Kai06]. Mastering [DA01]. Masters [Her03a]. Material [Blo92, Lyn85, Sim05, Bak12, Bel20, Mac17, NL18, PT18, Rel24, RSH18, vW23]. Materialism [Cro91]. Materialisms [HFGP17]. materialities [De 24, Swa24]. Materiality [Eps09, Pot11, Sil11]. materialization [WE15]. Materializations [Bri20]. Materializing [Bes20]. Materials [Mod01, Oud90, Spa20]. maternal [Lap16]. Math [Leo10].

Mathematical [Gas84, Kra72, Mac78a, Mac99b, Wag09, Gre24, Ger01]. Mathematicians [Pop21]. Mathematics [Blo87, Gas84, LC99, Mac78a, Mac93, Meh80, Ste78b, VB03, Wag09, Wat90, GS11, KG12, Lyn11a, Ste12]. Matilda [Ros93b, LPKL12]. Matter [Ste06, Lau16, Lip23, Oud12, Pet15, Wat14]. Matters [Den03, FDR08, Law81, Lyn04d, dlB11]. Matthew [Gol79, Ros93b]. Mauguin [CG92]. Max [AH20]. Maximization [Gor84]. May [Ano79, Ros96b, EES96]. Maya [Wat14]. Mayberry [Del02]. Mazur [Ano77c]. McClenon [Par88]. McClintock [Del05]. McGraw [Blo77]. McGraw-Hill [Blo77]. McHenry [Sis09b]. McKinlay [Hal89]. McKinney [Pin99a, Pin99b]. Me [Den01, Sin96b, PBEW24]. Meadows [Pri71]. Mean [Mil09, Say14]. Meaning [Der00, Mor88b, Pin21, Sil11, Twi05]. meaninglessness [FW23]. meanings [Aar23]. means [Tou13]. Measure [Dir99, Div90, Bie18]. Measurement [Der00, MM87, Sti87]. Measurements [Har96]. Measures [Lin80, MM86, JSLS25]. meat [DK12, Jön16]. meats [Jön16]. Mechanic [Ano77q, Lau84, SB77]. Mechanical [Jas01, Bij01]. Mechanics [Bru80, Har81b, MR73, Tou00, dLM00, Pin11]. Mechanizing [Woo89]. Media [Hin14, Lew01, OS05a, Del21, EMSL23, KR11, Law25, Oik17, Ros11]. Media-speak [Hin14]. mediary [Kum24]. Mediating [Kwa87, Mul76]. mediation [Ano10]. mediators [Kus10]. Medical [Ams87, ABB^{+}08, And92, And08, Ber95, Cue90, Gol04, Hed08b, Hed08a, Joh07, Law97, Nic96, Nur03, Pal02, Ras04, Sis09a, Sis10, Smi14, Sol92, Ste78a, Tan21b, TB97, Wri80, de 02, Car18, Coo11, yLL14, Mur13, Nav11, Ras20, Str11, WB78, AB80]. medically [Ras20]. Medicine [Ada02, Alb95, Ano09a, Asd08, Eps98, Ful08, Kim07, Mir07, Par05, Rei06, Wal00, Wri80, Fau12, GKSS24, Hoe19b, Jau15, JSLS25, yLL22, LA19, RC16, SF19, WH21, Wri80, Har94b]. Medicines [Mir07]. Medico [Sch72b]. Medieval [Dea95]. Medium [Hoe19a]. Medvedev [Hol81b]. Meet [Cri10]. Meeting [SG74a]. Meetings [SS94, PP24, Püt21]. Meets [CE03, Mir99, Ros11]. megaproject [Rot19]. Meier [Wes88]. Mekong [Jen20]. Mel [Ano77j]. Melvin [Gie96]. members [O’R83]. Membrane [vLR98]. Memoriam [Bar03a, Bow03, Col03, Elz03, Fox03, Jas03b, Lat03, Mac03d, Pin03, Sal76, Shr03, Sis22a, Van89]. memories [Bra24]. Memory [Pra09, SH19, Tou13, Pra09]. Men [God07, dC09, FFB11, Gre16a, WE15]. Mendel [Bra79, Far75]. Mendelian [Vic95]. Mendelssohn [AB80]. Menlo [FS23]. Mental [Smi85, Pic13, Pic19]. Mephistopheles [Fan09]. Mere [Pic90]. Merge [GM03]. Meritocracy [Gas84]. Mermin [Bar98, Blo98a]. Merton [Col04a, Gar04, Gie04, Kea73, Oro77b, PN04, Ric04]. Mertonians [Pin97]. Merz [GP97]. Mesoscale [Jan04]. Meta [DH08, Wir15]. Meta-Geeks [DH08]. metabolic [Lev14]. metabolism [FO22, Oli25]. Metadata [May19, EMB^{+}11]. Metal [Wak04, Rel24]. Metaphor [Chr88, Kuk94, Tsa24]. metaphorize [Vas15]. Metaphors [Mar77, Swa24, Ano77i]. Metaphysics [Gin99]. Metascientists [Lyn06a].

Meteorites [Wes78]. meteorology [Ran10, Sun11]. Meter [Per04]. meters [Kim16]. Methadone [Gom02]. Methanol [dMdF98]. Method [AO78, Buc75b, Ham86, Har96, Hil86, Mer02, MM96, RS89a, Day18]. Methodological [Ake07, Cla80, SR72, Bij12, De 24]. Methodology [Bed14, Bud85, Hir94, KP85, MvR85, Mor86a, Wei74b, Say14]. Methods [Fra88, HL10, M’c23, SC18, SS17b, VA16]. metrics [CBLG23]. Metrological [Cur98, Mal98]. Metrology [O’C93]. Mexican [LRC87, Mat14, dG75]. Mexicano [KGDLB^{+}15]. Mexico [Ves96, BFW14, Fit86, GDLB15, LC99, RG23]. Meyer [Col07]. Meynell [Ano77g]. Michael [AB80, GM78, Rip91, Doi11, Gor08, Osb93, Van89]. Michel [Jah72]. Michele [Fis10, Ros10]. Michelle [Eps09]. Michiel [Rip91]. Microbial [Som06b, FSF24, FO22, PH14]. Microbiology [Sch10]. Microeconomic [Dix19]. Microelectronics [CM09]. Microorganisms [Som06b]. Microprocessor [God97]. Microstructure [GSSD74]. microworlds [Red16]. Mid [Mir99, Mor73, SC18]. Mid-Century [Mir99, SC18]. Mid-Victorian [Mor73]. Middle [Bur18]. Midst [Roc81]. Might [dV07, PB84]. Migration [Mul74, vdH20]. Migrations [Jan04]. Mild [MMB09]. militarism [Cha22]. Military [Doe03, Eri71, Har03, Mac86, Pat83]. Military-Industrial [Eri71]. Mill [O’R83]. Millenarianism [JL72]. Milton [Bro95, Mos79]. Mind [Car18, Gor97, Lau90, Mol13, MPA87, Ver08, Pet17, SvD25, SC18, Del05]. mind-brain [SC18]. minded [HL19]. Minds [GP09, Mar77, Hug24, Ano77i]. Miner [Blo00]. Miners [Blo00]. mining [Gal24, Rie15]. Minnesota [Cri10]. Minority [GS25]. Minute [Rie03]. Mirror [Som06a]. misalignment [Met25]. misconduct [HR21]. Misparadigmatic [Lip74]. Missile [MS88a, MS88b]. Missing [WL15, Mia12, Smi17]. Mission [Ami74, Ver12, Gol94]. missionary [HTB20]. Mistaken [Kan78b]. Mistakes [Sch94]. Mister [Mia99]. Mistrust [Sar23]. Misuse [Mil76]. Mitigation [Wet07, Car21b]. Mixin [Mir07]. Moas [Mac90]. Mobile [Rei23, Gum20]. Mobility [Urb82, Anf21, Dav21, SBH^{+}20]. Mobilization [McC07, BL12]. Mobilizing [Osb93]. modal [Uru22]. mode [Bir23]. Model [Bar82, Bra89, Gor84, Gor92, Ink85, Kah08, MP87, Sma77, vdH20, And21, Joh13, LSW24, PH14, Sok23, PJ17]. model-based [LSW24]. Modeling [BH97, Mye08, Nel13]. Modelling [Eva97, Hic88, EMSL23, MS14b, vdH20]. Models [Blo85, Blo93, CD84, Dur11, Fra97, God15, HH83, KC82c, Lah05, Mae07, Pet00, TJ85, Yea99, BHC16, EMSL23, FC12, Jen20, Pot11, Mae07]. Modern [Baz84, Col07, Dea95, FG86, Ham86, Mir07, Mos79, Pra05, dSP75, RH96, Res78, Res82, Tou91, Wri80, Pot11, Sah25, Bra04]. modernen [Hom90]. Modernity [Col07, Eli00, Now87, Ger95]. Modernizing [BB17, Jas02]. Modes [JZ12, Lat13, M’c23]. Modified [Kli02, Lez06, MLC06, Yea01, AH20, Mos16]. modifying [Lah21]. modifying-work [Lah21]. modularity [Law20b]. Molasses [Gra98]. Molecular [AA82, Bar84b, BFW14, CM09, Fra97, Fue82, Fue84, Gas90, JL98, Mye08, SM97, Som06b, Win76, Wri86, Yox84, Cal21, MS24, Nor17].

Molecularization [Sho05]. Moment [Ver02]. moments [Mar23]. momentum [Uru22]. Money [Cur98, Mar08a, Yai19]. Mongolism [Mil03]. Monica [Bru77]. monitoring [Cal18, CC12]. Monograph [Nev80]. Monographs [Bro95, Nev80, O’R83]. Monopole [Pic81]. Monsieur [Sco92]. Monster [Ric96]. Monthly [Wri80]. Montreal [HJ23]. Moore [Blu09]. Moral [Blo83, Cur98, McC00, Mul84, RH96, Ras04, SK08, DK12]. Moravcsik [AB80, Cla82, Van89]. more-than-human [Hol22]. More-than-seeing [CWvdVdB23]. Morocco [Tan21b]. Morris [Mos79]. Moseley [Ano77k]. Moses [WC99]. Moskovskogo [Rab76]. Moskva [Rab76]. mote [Edg97a]. Motherhood [Kyv90, Dun22]. Motion [CG90]. motivated [Elg22]. Motivation [Ore03]. motive [Wad14]. motives [EE14, PP21]. Mountain [Ros93a, Wei90, Mac03a]. Mouse [Mar05, Nel13]. Mouse-Focused [Mar05]. Mouth [Ful91]. Mouton [GM78]. move [AC22]. Movement [Bay83, Yea89b, Hor25a, Kom22]. Movements [LM80, McC07]. Moving [Ala09a, SDL^{+}14]. MRI [Pra07]. Mrs [CC77]. MSM [Mar10]. muertos [GDLB15]. Mulkay [GM78, Hal88]. Mullins [HG88]. Multidisciplinary [Gor02, DSJ19]. multigenerational [Ran17]. multimodal [AMT11, Plá23]. Multinationals [GT77]. Multiple [BW83, Lin80, Tim96, Hof15]. Multiples [Con78]. Multiplicity [Som06b]. multiplied [Pap21]. multistability [eS23]. Multivariate [Lev14, Nad80]. municipal [SB18]. Munitions [Pat83]. Murderous [Per89]. Muriel [Del02]. Murphy [Eps09]. Murray [Bur75]. Museum [SG89]. Museums [HvL08]. Music [PB04, Thé04]. Musical [BS04a]. Must [Ash96, Mar13]. Mutant [Lak17]. mutating [KRR13]. Mutual [Pre05, GKSS24, Jau15, Pic13, Pin11]. My [Ede06, Lab95a, Sle89a, Bar20, Kus04a]. Mysticism [Res78, Res82]. Mystification [Blo74]. Myth [RS89a].

n.p [AB80]. Nadel [Zan84]. Nairobi [Gum20, HHW^{+}19]. Naïveté [Kro82]. Nakayama [dSP75]. Nancy [AB80]. Nanotechnology [CM09, VFWC15, KR11, MK10, Vis15]. Naples [dB09]. Narin [Ano77h]. Narrating [Bur18]. Narrative [Cur94, Hed01, Ree01]. Narratives [FFMW08, CB24, SO17]. narrativity [SP24]. NASA [Rei21, Rot19]. Natasha [Hsu13]. Nathan [AB80]. Nathaniel [Del05]. Nation [MD09, Tsa24, WLBRS15, GDLB15, KGDLB^{+}15, SMWCSC15]. Nation-building [Tsa24]. National [Abr91, AC88, Blo91a, Bru77, CL11, Fra79, GKM75, Lub81, ML93, MD09, VP98, Bur18, LS10, PAB15, RP16, LC99]. Nationalism [Hod83]. nationalists [PD19]. Nations [HI10]. Native [Ler18, Sah12]. Natural [dBB76, Bru15, Bur83, Col82b, Dol71, Rob96, Sch86, Sch88b, Sec85, Wat90, vHVdW07, Hen18, Nev80, Kim94]. Naturalism [Jac78]. Naturalist [Cri96, Ell11]. Naturalization [Nev80]. Naturalized [Pic91]. Nature [Ano77q, Car12b, Chr88, Dea95, Eas74, Gin97, Kwa87, Mos79, SB77, Wat02, Wei74a, Yea85, Tjh24, RL21]. natures [PH14]. Naturwissenschaft [Bay83]. Naturwissenschafts [Bay83]. Naturwissenschafts-Rezeption [Bay83].

Naukometricheskie [Rab76]. Naval [Smi14, vK01]. Navarro [Wri80]. Navigation [MS88a, MS88b]. Navy [Ore03]. Nazi [Per89]. NC [Mir07, And88]. Neanderthal [Hig21, Som06a]. Neanderthals [Ham82]. near [Var20]. near-real-time [Var20]. Neb [Bru77]. Need [Chu75, Dic00]. Needham [Res96]. Needle [Mar96b]. Needs [Col85d, Eas74]. Neeley [Del05]. Negational [MM84a]. negative [Ran17]. Neglected [Lan81, dlB11]. Negotiating [Dow88, Mac93, Mar10, Sat21, Pra06, TT18, Tjh24]. Neil [Hyy09]. Nelkin [Ano77l]. neoantigen [Jat23]. neoantigens [CGP25]. neoantigens-based [CGP25]. Neoliberal [Gre16b, LMR10, NK14a, Rah11]. Neonatal [MM96, Car14]. Nerve [Bal06]. Nerves [Hea04]. Nervous [Law84]. net [Sis17a]. Netherlands [BR79, RB75]. Network [Akr23, Ala09b, KHM10, Say14, Thé04, FM17, Lag12, PI19]. Networking [lL99]. Networks [BH97, CS00, Col74, EES96, Gar97, GR08, Ley91, Mie98, Mil03, ML94, SM93b, Tor94, WB96, Joh18, KHM10, Ran17, Smi17, Bar84a, Rus94]. Netz [Lat08]. Netz-Works [Lat08]. neuro [BM14]. Neuroeconomics [SZ11]. Neuropharmacology [Coz85]. Neuropsychedelia [Cas15]. Neuroscience [Bea01, Sta83, FLK^{+}14, Tol19]. Neurosciences [Lyn88, Mat13]. Neurosurgery [Mor06]. Neurosurgical [Mor00]. Neutrality [RA96, Yea92]. Never [Del05, Mod01]. News [MBB00, ML93]. Newton [IM83b, IM83a, Smi83]. Newtonian [Jac94]. Next [BFS95a, Cam03]. Nexus [Sol01a]. Nice [Gin10, BHC16]. Nicholas [Hol81b, HG88]. Nick [Mae07]. Nigel [Ano77m]. Nineteenth [Alb95, Don86, Lau84, Car18, Kra89]. Nineteenth-Century [Lau84, Don86, Car18]. Ninety [Cou99]. Nitrogen [Str92]. Nixon [Fri84]. NJ [Gre10, Hol81b, Hsu13]. No [Bro95, Dow07b, JS17, Nev80, O’R83, Sha04]. No-Holds-Barred [Dow07b]. Nobel [Opp79, AO78, IP76]. Noel [Mos79]. Noise [Bij01, Lyn01a, Wak04, Kim16]. nomenclature [Hal19]. Non [Dug87, Edm02, Mel03, RCB18, SK91, Shr00, DU15, Dur20, KG12, Kna13, MB25b, O’R83, SPBG24, US20]. non-evidence [Kna13]. non-fungible [MB25b]. Non-Governmental [Shr00, US20]. non-knowledge [DU15, Dur20, SPBG24]. non-members [O’R83]. Non-Science [Mel03, Edm02]. non-SEM [KG12]. Non-stick [RCB18]. noncontentious [BM19]. nonhumans [Say14]. Nonlinear [Law11]. nonproliferation [Law25]. Nonviolent [Mar97]. Nonwestern [Han86]. Norm [Woj74]. Normal [Hen07, MB22, Pin79, Rei80, War74, WP73, CKFS13, Gu23]. Normalizing [Ure14]. Norman [BCM76, Sal76]. Normative [Cur94, Kim94, Rad92]. Normativity [Pin15]. Norms [Col82a, GB09, MO09, Sut84, Woj74, Fis10]. North [Mir07, Bij01, CC12, De 24, Dun91, FG73, Pri12]. Norton [Hol81b]. Norwegian [And88, Anf21, RP16]. not-knowing [Str11]. Not-Last [Gie82a]. Note [Bea10, CK85, EM73, EM82, EM83b, EM84, LMT92, Lat07, Lau82, Law77, LSW95, Nis95, Ano10]. Notes [Ano88a, Ano88b, Ano89, Ano90a, Ano90b, Ano91b, Ano91c, Ano92, CFS01, Edg95d, Mag77, Sil74, Ano91a, LL13, Lyn12b]. Nothing [Sma20]. Notion [Wes79]. Notions [Bea01]. Novel [PH14]. November [Gie96, HDS^{+}98, Ros96b]. NSF [MC79]. Nuclear [Bar03b, BP80, Del83, Dow88, Gus06, Har83a, Hec96, Hec02, Hol81a, Hug06, Mac03a, MS88a, MS88b, MC78, MM98, O’R83, Spi87, Spi97, Tak24, Win02a, Hai19, Ial21, Law20b, Law25, Pol20, SH12, eS23]. Number [Wat90]. Numbers [BS01, Did25, Lag07, Wis01, Lam10b, ÖS24, Sis10]. Numerical [Har03]. Nutrition [SN89, Fro17]. Nuts [Fau07].

O [GM78, EM83b, Lóp22, Doi04]. Obesity [GDLB15, Gre16b]. Obituary [Bal10, Blo03, Blu09, FPM03, Gor08, Gor10, RD08]. Object [Ala20b, Lez06, Luc01, Lyn88, Mie98, Hop23, Sor23]. Objectification [Cus96]. Objective [Blo74, Mey74, Ano16a, Sha16, Blo74]. Objectivity [And13, Bea01, CKSW09, Das92, Dol74, Hog09, MMB09, GL21, Kis13]. Objects [Kai06, SG89, Sun07, CSTG23]. obligation [GKSS24]. Obscure [Nis95]. Observation [Pin85]. Observational [Pin85, TH24]. Observations [Cho85, Ink85, Jef03]. Observer [Sch88b]. Observing [Ree01]. Obtainable [O’R83]. Obvious [Gra07]. Occupational [RJ76]. Ocean [MB81, Leh18, De 24]. Oceanographers [PKS88]. Oceanographic [Ore03]. oceanography [Sun11]. Ochre [Goo10]. October [Lyn11c, Sis22a]. OECD [Tak24]. Oehler [Pot87]. Off [Bur75, Wyn96]. Off-and-On [Wyn96]. Office [Eps09, Gus99]. Officer [Rii08]. Official [Ola96]. Oil [Bow85, Con89, Tan21a, UB25]. OJ [Lyn98a]. Olby [AA84]. Old [Geu03]. olfactory [Ala20a, Ala20b]. Olga [Blu09]. Omnibus [DU15]. omniscience [Suc23]. once [Wie17]. oncogene [SA22]. Oncology [Cas09, CGP25, Mur13]. One [Ben12, Dea74, Hir94, Mul84]. One-Act [Mul84]. Ones [HSP02]. online [BTS22, Hor25b, KH18, Sil11]. Ontography [Lyn13]. Ontological [Cus96, Lan00, Mor00, Pow01, Sis15, RP15a]. Ontologies [Sch10, Bri13, Pap11, vdH20]. Ontology [CLP91, KC95, Asp15, LL13, Lyn13, Mar13, Pin21, WL13, WL15, vHLW13]. ontonorms [Mol13]. Open [Mos79, Dav24, DG17, Mir18, Sov10]. open-source [Dav24]. Opening [Car01, Moo10, Pin92, SvD25, Uru22]. Operational [Ial21]. Operationalization [Zan84]. Operations [FS93, Mir99, Day18]. Opinions [Jab95]. Oppenheimer [TS00, Tho02]. Opportunities [Ree06, Sch95]. Opportunity [GGJ07]. Opposition [LM80, Ros81]. optical [PP24]. Optics [Fra76, Tur87]. optimization [Zie19]. Opus [Ash07]. oracle [Day18]. Oransky [DGC18]. Order [Buc75b, Hes86, Jah72, Nev80, Rud77, Sim05, Sim07a, DP21, LOBB23]. Ordered [Bow87]. Ordering [Hin06, Lau16, Rap05, Rot05, SH17]. Orderly [Tur95a]. Ordinary [Blo92, Fau98, Nol92]. organ [HJ13]. organisations [Hed12]. organism [Joh13]. Organisms [Lez06]. Organization [Dae98, Dun91, Gus99, Hyy09, MV05, Mos79, Nad80, Oud97, Rot81, TS00, Vau99, WD97, Yea90, Dem24, Nis11, RS11, BL12, Gal24]. Organizational [BS98, CK05, KK03, Kis13, Leo10, PWD07, Sut84, Wol95, Hed14, MS14a, MB22, SV22]. Organizations [Bal95, Gus06, Hug06, Hyy09, MO09, SM97, SG73, Sch94, Shr00, Pan11, PC12, PW10, US20]. Organized [Hut71]. Organizing [KK03, RP15a, Ger95]. orientation [WE15]. orientations [FSF24]. Origin [Mar96b, Ros81, dB72]. Originality [Dir99]. Originally [Gre10]. Origins [Arm09, CM09, Hol81b, Kop79, Mos79, Nor78, ST97, Wil91, Hor25b, Ple14, SA22]. Orleans [Sim07a, Wet07]. Oromaner [Ano77p]. Orphan [Kow13]. Orthodoxy [KK03]. Oswaldo [AB80]. Other [Her07, Law81, Mar95, PB84, PK09, Pra05, Pra09, CS20, Lyn04d, Sok23, WC99]. Ottawa [AB80]. Otters [GR08]. Our [MI85, Roc81, Jat17]. Outcome [Lan01b]. Outcomes [Vin91, Sim25]. Outer [Bur92, Red02]. Outline [Her07]. Outsider [Del87, Wer71]. outsides [Dán20]. Ovario [Woo06]. Ovario-Hysterectomy [Woo06]. Overflow [Hil07]. Oversimplification [Bar98]. Overthrown [Win02b]. Ovonic [GK72]. Owen [Bar03a, Blo03, Bow03, Col03, Elz03, Fox03, Jas03b, Lat03, Mac03d, Pin03, Shr03]. Own [Hir94]. Oxford [Blo74, Col07, Sto08, Kis13]. Ozturk [Del02].

P [Ano77h, Bar84a, Jah72, Mil09]. P&L [MS14a]. p.&p [O’R83]. pace [dlB15]. pacemakers [Oud15]. Pacific [De 24, Gal24]. Packaged [Hyy09]. Packages [PWD07]. packaging [Bel20]. packet [Bel20]. pads [Par22]. Pagan [Jac78]. Page [Wis01, Ste12, Mos79]. Pain [Dow07b]. País [GDLB15]. Pakal [Wat14]. palace [Col13]. palatability [SB18]. paleoclimatology [Sch16]. paleontology [Wyl15]. Panacea [PJ17]. Panama [Car12b]. pancreas [Dav24]. pandemic [CLCMW23, MB25a]. pandemics [Abe13]. Pandora [Lyn01b]. Panel [OOO12, Rou17]. Pap [CC98]. Paper [ABB^{+}08, Hed08b, KC93, Lat07, Mah79, Mul74, Blo74]. paperback [Hol81b, Luk75, Out76, Wri80]. Papers [Cam93, Coz85, LMA80, NFC83, Ano10, Del16, dB72]. Papua [Str11]. Para [KR11]. Para-scientific [KR11]. Parade [Wyn96]. Paradigm [FL88, McD76, Rei80]. Paradigm-Land [Rei80]. Paradigms [Blo71, CW87, Cra80a, Lam06, Nad84, Pic80a, Tur87, WP73]. Paradox [DYS^{+}05, Mye89, Pin11]. Paradoxes [Res78, Res82]. Paraglider [Hac92]. Parallelism [Res82]. Parallels [Han86, Res78, Res82]. Parameterizations [Sun07]. Paranormal [Nic84, Pin79, PC84]. Parapsychology [Gor82, Har81a, Par88, Pin79]. Parent [Woo74]. Parental [Sil74, Son09]. parenthood [KG12]. Parenting [HL10]. Parents [Son09]. PAREX [Mac78a]. Paris [GM78, Gre10, DP21]. Parity [WSB79]. Parliament [Asd08]. Parliamentary [Kir99, Mul94, Mul95a, Mul95b, Bar20, Dán20]. parliaments [AH20]. Paro [Che23]. Part [WB78, Bar20, Sis12a]. Partial [McK99]. Participation [FFMW08, SS94, Lau13, Mar13, Moo10, NL18, Rei13, RSH18, SFH12, VA16]. participatory [CKFS13, LD13]. Particle [KSW82, Kai06]. Particles [Kri87, Del16]. particularity [KH22]. Particulars [O’C93]. Partnerships [Row12]. Parts [Rav99, KRR13]. Party [Roc81]. Pask [Pic02]. Passive [MBB00]. Past [Jas00, Mac99a]. Pasteurization [Wir15]. Patent [Bal06, Pot11]. Patenting [Ber08]. Patents [Met06, Mos79, Mye95, CS20]. Path [MS88b, PBSK25, Ste99]. pathology [CWvdVdB23]. Paths [Thé04, dC09]. Patient [God97, Jen05, Pan11, HAB^{+}23, JZ12]. patient-provided [HAB^{+}23]. Patients [Bou05, Pal02, NL18]. Paton [SN89]. Patrick [Bru77]. Patron [PKS88]. Patronage [Mor73, Sol01a]. pattern [Ell11]. Patterns [Bur92, Del05, FFB11, LR08, LRC87, Nad83, SM93a]. Paul [Bar79, Blo77, Luk75, For99, Sto08, Wes88]. Pauling [Olb84]. Paviland [Goo10]. Pavitt [FPM03]. Pax [Ful95]. Pays [Bur75]. pbk [Alb95, Ash07, Col07, Cri10, Del02, Del05, Doi11, Eps09, Gin10, Gre10, Kim94, Mae07, Moo10, Nev80, Pra09, Rei06, Rip91, Sis10, Nev80]. Peaceful [Sch06]. Pearson [Nor78]. Pedagogy [RH90]. Peer [Abr91, Hir10, Lan01b, MC79, Son95, Gre24, Huu12, KdR19]. Pels [Col07]. Pelton [Con78]. Penalties [KG12]. Penicillin [Lie87]. People [Fau07, KCBT07, Por82, BHC16, NK14b, PC12]. Peoples [Mar08b]. perception [Spa20]. Perceptions [PKS88, Sol92, Yea92]. Perceptrons [Ola96]. perfect [IJJ13]. Perfection [BS04a]. perform [AC22]. Performance [Bur92, Hor04, Ink91, IM85, MvR85, SCG01, SL91, Joh13]. Performances [Cam04]. Performative [God98, Jen05]. performativities [Fau12]. Performativity [Cam04, Mac03c, CBLG23]. Performing [Asp15, Kan21]. perils [PH14]. perimeterisation [SP24]. period [Lei14]. Periods [Rav99]. Peripheral [Fue82, Sch87]. Periphery [Cou99, Nes95, SM97]. permanently [LTC18]. Permission [Mah80]. Permissive [AD09, MD20]. permit [Jan18]. perplexed [Lor17]. Persistence [Mil03, CKFS13, Jon23]. Person [Cal99, Lat86]. Personal [Cho85, CB24, Gor06, KC93, Tur82]. Personalised [Rei06]. personality [Pic13]. Personalized [Rei06, Hoe19b, Jat23, WH21]. Personane [Ano84]. Personhood [Ree07, Jon17a]. persons [HL19]. Perspective [Bru80, Das92, FLE97, Res82, Ste86, Yea85, HJ13, KGDLB^{+}15, LSW24]. Perspectives [Del02, Ham86, Nur03, Sma20, GM78]. Persuasion [Ano77m, Gil77, LW82]. Persuasive [AQL94]. pessimism [Tut11]. Pest [Pal90]. pesticides [Bri14]. pet [Kro19]. Peter [Col07, GM78, Old88, PCL98, Pfe71]. Petri [Lab95a, Lab95b, Pin95]. Petroleum [Con89, Den85, AH11]. Pfiesteria [Sch10]. phallometric [WE15]. Phantom [Bug93, MM98]. Phantomatic [Sch10]. pharma [BM19]. pharma-contraceptive [BM19]. Pharmaceutical [Gre04, Mil12, Mir07, Sis04, NK14a, Pol14]. Pharmacogenetics [HM03, Hed06, Rei06]. Pharmacogenomics [Kah08]. Pharmacopolitics [Mir07]. Pharmacovigilance [Lan09]. Phenomenon [Ash93, Col81a, Wyn76, Arm10, Ler18, Val10]. phenotypically [Nav11]. phenotyping [GM23, Hop23]. philosopher [Lat10]. Philosophers [Nic90]. Philosophical [Blo75, Elk78, Ful90, Sha72, AQL94]. Philosophism [Lyn04e]. Philosophy [Boh89, Bru80, Col99a, Hut91, Pic91, PM89, Ric04, RH85, Sch86, Kim14, Cor94]. Phoney [Ash93]. Phosphate [Bur75]. photo [Edm11]. Photographic [JS89]. Physical [AA82, Col81a, Del23, SMW15, Baz84]. Physician [Wad14]. Physician-industry [Wad14]. Physicians [de 02]. Physicist [Efm92, Mer98b]. Physicists [Cal99, GP97, MKC97, Mos79, PS77, Tür73]. Physics [Ano77k, Baz84, Cra80a, Cro91, Fra76, Gil86, IM85, Kan78b, KC95, KCM97, KSW82, Mag75, Mag77, Mel03, Mos77, Nad80, Pic80a, Pin85, Res78, Res82, Wyn76, PP24, RG14, Fre92, Can75, Mos79]. Physik [Hom90]. Physiological [Tur87]. Physiology [Cue90, SN89]. Pickering [Cra80b, Gie92, Gin99, Old90b]. Picking [Old90b]. Pickup [LT92]. picture [CE16]. Pictures [CG90]. Piece [Ash93, Wyl15]. Pierre [Bre02, Gin10, Mia03]. Pig [DK12]. pigs [Pih25]. Pillars [Shi99]. Pinch [Col22, Bar11, McK99, Nic84, Rus86]. Pinnick [KCM97, SS99a, SS99b]. Pioneers [Hec96]. pipeline [WC18]. piscicida [Sch10]. Pitfalls [Hun82, Pyc93]. Pittsburgh [Con79]. Place [Chr88, Hen00, Pal71, War71]. placebo [Cra10]. Places [Pol14, Oud12]. Plan [Bha76]. Planarian [Tra81]. Planck [LSW95]. planet [FO22, Row05]. Planetary [Mas10, Rei21, Tat86]. Planets [Mes10]. Planning [Bra24, Sis09a, SA22]. Plants [GP09, Van08, She20]. Plastic [Hul92, Sat21]. Plate [Mes88, Mol13]. Plausibility [Har81b]. Play [Mul84, Mur13, Bar20, Hoe19a, Sor23]. Please [Fee00, Lat86, RA96, SS17b]. Plight [Nis95]. Ploughs [Tod92]. plumes [SPBG24]. Plumpy’nut(R) [SS18]. Plural [Ray86]. Pluralism [DeW75]. pluralization [RC16]. Pluto [Mes10]. Pocket [Ash93]. Point [KC93, Sag73, Eng20, ST24]. point-of-care [Eng20, ST24]. Points [Cla80, MC71, Arm19]. Poison [Whi75b]. Poisson [Sim78]. Poland [Jab95, Wal75]. Polanyi [Doi11, Tur12a]. Polar [GK16b]. Polaris [MS88a]. Polemics [And92]. Police [Col99b, Rii08, Sim07a]. Policy [AA82, AA84, Abr94, AB80, BFS95a, Bar84b, BR79, CLP90, CLP91, CB97, Col85c, Cue90, Elz04, Fra88, Fue84, Gum90, Hal06, HV93, Hic87, Jas87, Jas95, Kre71, Kre73, Kwa87, Mac03a, Mar10, MC79, Ron78, SG74a, Sko73, SR73, SRM76, TG86, Vog06, WG74, Wyn92b, vdSvESW98, CGPP24, Day18, EMSL23, Gar73a, McL15, MS24, Par25, Sma20, SBHY15, ZDI^{+}23]. Policy-Making [CLP91]. Policy-Relevant [Jas87]. Policymaking [CD84]. Polio [Mar96b]. Politeness [Mye91]. Politic [Sin98]. Political [Bak12, Cal18, CLP91, Ezr71, GT93, Hil90, JL72, Jon17a, Lau16, Muk07a, Nel75, Pow01, SR73, Tan08, Wil73, Wri80, dV07, Bri20, Bri14, CK85, Eva22, Gra85, LSS20, Lei14, Mac17, Mar13, Moo10, PJ17, Pol20, Rel24, SV22, SM25]. Politicization [Bri09]. Politicizing [Bro15]. Politics [AR02, Ala09b, Alb80, Bar03b, Dea74, DNTS08, Doi04, Eps97, Eps08, Eps09, Fra76, Got95, Gus99, Hag99, Hay07, Hol81b, Irw06, Jas96, JM01, Joe99a, KK03, Kle91, Lan00, Lat07, Lau84, Lub81, Lyn01a, McD97, MM96, Mos79, Muk07b, Oud97, Pel96, PM79, Rad98a, Rad98b, Ree07, Rei06, Ric88, RA96, Rip91, Ros96a, Ros81, Sin96b, Sis10, Ski06, Sla07, Sol01a, Ste90, SS91, VSS22, Waj86, Wer71, Wri86, Wyn92a, AH20, Bar20, BFW14, Bre13, Bro15, CEW17, DS20, Dán20, Fis18, FLK^{+}14, Fro17, GM23, Gra16, Has12, Hen18, Ial21, Jan18, Jeo19, Joh18, Ken13, KW15, LS10, LTC18, MMV15, MvO23, Nor17, OL18, Pap11, PLPS22, Rah11, SS17a, SvD25, Sim25, SB18, Spa20, WN22, dV07, Blo77, RL21]. Pollock [Hyy09]. polluted [LTC18]. pollution [Cal18]. Polygenic [KBRB19]. Polytechnics [LW75]. Polytheistic [Whi76]. Ponts [Gra07]. Ponts-et-Chaussées [Gra07]. Popovsky [Hol81b]. Popper [Blo74, Blo74, DeW75, Kri78]. Popular [Ale08, Cur94, Kid88, Kir10, Mel03, Ree07, Rob99, She02]. Popularization [Hil90, Mil09]. Popularized [JS89]. Popularizer [She02]. Population [GKSS24, Hal06, Oud97, Ram02, Ram09, Arm17, Hog19]. populations [Mor19, SMWCSC15]. Populist [RG23]. pork [Pih25]. porosity [BM14]. Porter [Gre10, Nev80, Hag99, Sch02]. Portraits [JS89]. Portrayals [Cri96]. Portugal [CS00]. Poseidon [MS88a]. position [LSS20]. position-making [LSS20]. Positions [KCBT07]. Positivism [Mac81]. Possibilities [Col85c, WHBM02]. Possibility [Woo89, Bar12]. Possible [Chu82]. Post [BFS95b, Edw94, Fee00, Hed08b, Lan09, Mos79, Pol20, Sis17b, Tür73, JS17, LA19, Lyn17, Oli25, SDL^{+}14]. post-3.11 [LA19]. Post-Black [Lan09]. Post-Communist [BFS95b]. Post-Doctorate [Tür73]. post-Fukushima [Pol20]. post-genomic [SDL^{+}14]. Post-Human [Fee00]. post-industrial [Oli25]. Post-political [Pol20]. Post-Structuralist [Edw94]. Post-truth [Sis17b, JS17, Lyn17]. Post-war [Hed08b]. Postal [Har01]. Postcolonial [Ada02, And02, Hu25, Kin02, Ver02, Gra16, Hin13, KRR13, Kow13, yLL22, Pol14]. postcolony [Lac16]. postgenomics [Chi24]. Postmodernism [Sin96b]. Postmodernity [Ros93a]. Postparadigmatic [Lip74]. postphenomenology [dBtMV21]. Postscript [Edg99b, Mye89, Thé09]. Postwar [Ber95, Doe03, Gre04, Iid15]. potability [SB18]. Potential [Bal95]. Potter [FO87, Hal89]. Pottery [Sha72]. Poultry [Mar08a]. poverty [Bri14, Tol19]. Power [AA82, Bar84a, Bar87, BP80, Bru77, Dow88, FFMW08, GR08, Hec02, MC97, Rad86, Tur89b, Win02a, CKFS13, DH08, GB16, Hai19, KLW19, Kim16, Uru22, NK14b, Wri80]. Powerful [Mir07]. Pox [Ful95]. pp [Alb95, AB80, Ash07, Bar79, Blo77, Bro95, Bru77, Col07, Cri10, Del02, Del05, Doi11, Eps09, Fis10, GM78, Gin10, Goo10, Gre10, Har83a, Hol81b, Hsu13, Hyy09, Kim94, Luk75, Mae07, Moo10, Mos79, Nev80, O’R83, Out76, Pfe71, Pra09, Rab76, Rei06, Rip91, Rob99, Ros10, Sis10, Sto08, Wes88, Wri80, Bru77]. Practical [Wyn88]. Practice [Aga98, Alb95, Ber95, Del08, Fou97, Gor97, Hår94a, Hal06, Hyy06, JE75, Joh07, Ken00, Kir03, Kra89, Lyn88, LJ98, Mor00, Mos79, Nic96, Pal90, Pal02, Pic90, Tur90, Woo06, AMT11, BM14, Col11, FÖP20, HAB^{+}23, Hoe14, Hoe19a, IJJ13, Joh13, Jon23, yLL14, M’c23, MvO23, Mcl21, RL16]. Practice-Bound [Hyy06]. Practices [AD09, Bou05, Pra09, RB09, Rot05, Sim99b, SBS^{+}00, Som06b, VK84, BB17, Dán20, DA24, KH19, KBvLA22, Lam10b, Lyo18, MS14b, Mar20, Mur15, Plá23, Rad13, Sim10, SH19, Tan21a]. Practising [Cho85, Han86, Ree01]. Practitioners [Joh07, Osb93, SM93b]. Praeger [Hol81b]. Pragmatic [Hal19, Hog09]. Pragmatist [Mar07]. Praise [Col96b]. Prankster [Ash93]. Praxeology [Hir10]. Praxis [TJ85, Woo86, Gre16a]. Pre [Cur98, Gra16, Mul94, Lei14]. Pre-Confederation [Cur98]. Pre-Embryo [Mul94]. Pre-Exposure [Gra16]. pre-state [Lei14]. Precambrian [Yea90]. Precarious [She06, Var17]. Precaution [LT12]. Precautionary [Lev01]. Precise [Sti87]. precision [CGP25, GKSS24, JSLS25]. Precog [GS25]. Preconceptions [Mer98a]. Predicament [Mir07]. Predicted [Opp79]. Predicting [AO78, GS25]. Prediction [Har03, Jan04, Arm17, Jat23, Mac17, NM20]. Predictive [Abr91, HB23, Mah14, eS23]. predictors [FFB11]. Preface [Bru77]. Prefatory [Mul74]. Preferences [FS01]. pregnancy [Rob16]. Prejudice [EP07]. premiums [KG12]. prenatal [Nis11]. Prepackaged [Shi02]. Preparadigmatic [Lip74]. Preparing [Cam03]. Preprint [Gun05, Del16]. Prescribing [Sis10]. Prescription [Mir07, Ras03, Ple14]. presence [Bea10]. Present [Buc79, Har23, Jas00, LM75, Wal75]. Presentation [Eri04, Pin81b, Shr04]. Presentations [LR01]. presented [Bru77]. Preservation [Dun91, Rad13, Tan21a]. Presidential [Lyn01b, Mil04a, PP22]. Press [Alb95, Ash07, Blo74, Col07, Cri10, Del05, Doi11, Eps09, Fis10, Gin10, Goo10, Gre10, Hol81b, Hsu13, Kim94, Mae07, Mir07, Moo10, Mos79, OS05a, Pra09, dSP75, Pri71, Rei06, Rob99, Ros10, Sis10, Som06a, Sto08, Wri80, Ano07c, Cle86]. Presses [Gre10]. Pressure [Har96, Mor06, Tan21b]. Prestigious [Sha01]. Presuppositions [Abr94]. prevention [Dav20, Gre16a, Mon12]. preventive [Law11]. price [CS20, Ken17, Mac17, Ano83b, BH97, Mac84]. Prices [O’R83, WC16]. Primatologists [AB91]. Primatology [Ree07]. PRIMES [Mil09]. Princeton [Gre10, Hol81b, Hsu13]. principal [Car21a]. principal-agent [Car21a]. Principle [AD09, Gil06, LSW95, Bar11]. Principles [dSP71, McL15]. Print [Onu84]. Prior [PW95]. Prism [Ash93]. prisms [GB16]. Privacy [RvD22]. Private [Jas97, MO09, PC84]. Privatization [Win02a]. Privatizing [LDR10]. Prize [CG89]. prizes [LPKL12, IP76]. Prizewinner [Opp79]. Prizewinners [AO78]. probabilistic [PBEW24]. Problem [AdSPG^{+}76, Arm09, Bre97, BS08, Dun85, Eps09, Ham86, Lin80, LMA80, MBB00, Mes10, Pin79, Tur01c, Vel90, Avi15, Eri72, Lau25, Tan21b, WM17]. Problematic [Wei74a]. Problematizing [Mor19, Ros11]. Problems [Bar79, DYS^{+}05, EES96, Fuj87, Hil90, Hun82, Mal98, Mat05, Mir72, Ram09, Rap05, Rav99, WHBM02, Hop23]. procedures [Sim25]. Proceed [Bro95]. Proceedings [DU15]. Proces [Hea07]. Process [And88, Cam03, CG90, FR89, Jef03, Shr04, Str92, Tay95, VP98, Wol95, PBSK25]. Processes [Hår94a, HMH80, Joh18, Lan01b, RDB03, Sat07, BHC16, Waz13]. processing [Fro12, Jat17]. Prodist [Wri80]. produce [Rei13]. produced [Dur20]. producers [Rei23]. produces [Kan21]. Producing [Del23, Dow07b, SDL^{+}14, CC12]. Product [Sim10]. Production [Fra79, Hil01, Joy05, Lin80, Lyn98a, MMB09, Oud90, PWD07, Shi02, Vau99, Waj86, Wri86, Eps21, Foc16, Gu23, HH22, JS23, LDM^{+}16, Lee16, Lyn13, Mag17, Mon12, Oli25, Pap21, Pih25, Ros11, SB18, Str11, Str16]. productionist [LSW24, Sim25]. Productivity [All80, Bur92, DYS^{+}05, Fox83, Fox05, GT72, Kyv90, LB05, Tür73, Tur87, HL10, MM82]. Profession [Kuk94, Mos79]. Professional [Ano77l, Bla76, Har76, HJ13, Nel77, Oro77b, PT72, Sut84, Tor94, Van08]. Professional-Support [Tor94]. Professionalism [Mac79]. Professionalization [Ano77k, Bot90, Mos77, OM76, Por04]. Professionals [Pal02, Rot81, SG89]. Professors [Blo77, Fis10, Ros10, Ham19]. Profiling [Der00, Edm11, Law13, PBEW24]. profits [LW10, Ran10]. Program [Bug93, Lan01a, Pel99, Rad13]. Programme [Ano71b, Bro82, Chu82, Col81b, Fre84, Kim94, KC82a, Koc10a, Kro82, KR00, Kwa87, Nol92, Ros81, SM93b, Sle89b, Sle91, Yox84, Kwa05]. Programmes [Bru77, Gie82b, Gor89b, Nad84]. Programming [Püt21]. Programs [KH18, MS24]. progres [Hop23]. Progress [AR02, Bot90, Nol90, Bar79]. Progressive [Ale08]. Prohibitions [Rap05]. Project [Bal96, BS08, Fuk16, Mye91, KH18, SSB24, AB80, Car21b, Mat05, Rea01, Sol01a]. Projecting [dSP71]. projection [Bol18]. projections [OOO12]. Projects [HH83, MB81, Bur18, CSTG23, Ran17]. Proliferation [Vog06, Pic19]. Promethean [Ash93]. Prominence [Pyc93]. promiscuity [And21]. promise [BM14, Hoe19b]. Promises [PI19, Jön16, PH14]. Promising [Tut11]. promissory [PW10]. Prompted [VB03]. Pronovost [Kle91]. Proof [Hei03, Mac93, Mac99b, Sim10, Ste12]. properties [SDL^{+}14]. Prophecy [CC76, Law77]. Prophylaxis [Gra16]. proposal [Sis25]. Proposals [Abr91, Huu12, PW19]. Propositional [KMH83]. Propositionalism [Wyn03]. Prosopographical [Gas95, SS94]. Prospecting [Rie15]. Prospects [FS01, Mir72, Wal75, Gar73a, Mil12]. Protein [Fuk16, Law73, Mye08]. protest [Kim16]. Protestant [BL94]. Protocols [TB97]. Prototypes [Kir10]. Provenance [PWD07]. Proverbial [Sha01, SL25]. provided [HAB^{+}23]. Providing [Har01]. Provincial [Ora71, Sha72]. Proving [Liv99, Liv06]. Provision [McC92]. proxy [Var20]. Prussian [Ano77j, Cro76, Gor77]. Pseudo [Mag77, Kan78b]. Pseudo-Effect [Mag77, Kan78b]. Psychiatric [RB09, AABF10]. Psychiatry [Ras03, Tom87, SvD25]. Psychoanalysis [Osb93]. Psychological [Gol94, Han86, McD76, Sch90, Pet15, Pic19]. Psychology [AB95, Bri72, Cro00, Gor91b, HSP02, Lip74, Mah79, Pal71, STGG95, Smy01, Tur82, War71, WP73, Wei74a, Wei74b]. psychometrics [Sta18]. psychophysics [Ala20a, Ala20b]. Public [Ano77c, Bal06, BPBG06, Col87, Col88, Dic00, FFMW08, Jas97, JS17, KCBT07, LM75, LM80, Mar07, Maz77, Mer02, MO09, Mul72, PC84, Rob99, RDB03, SMW15, SFH12, SN89, TG86, Vog06, Wyn88, Yea89a, Yea92, Yea99, BKK21, CB24, DP21, HB23, HLRK14, HBP19, Hoe19b, KBRB19, Kum24, LR20, MB25a, MN23, Mon12, Par25, SMWCSC15, Sma20, VT21, VA16]. Publication [Fox83, Fox05, LRC87, Rie03, Sis09a, VK84, Jac19, KBvLA22]. Publications [AB80, Col99c, DC76, Gun05, Jah72, Mos79, Nev80, Oro77b]. Publics [Irw94, LD13, WLBRS15]. Publish [Cam93, Inh77b]. published [Gre10]. Publishers [Har83a]. PubPeer [DGC18]. Pump [Red16, Sha84, dLM00]. Pumping [Bar12]. Puncturing [WC18]. Pure [BS01, Kim91]. Puritanism [Sch88a]. Pursuit [Nel08]. Push [Dal90, Tur89a]. Put [Ful91]. Putting [LW82, Vin99, Wis01]. Puzzled [Del87]. Pyramids [Mil76].

Q [Col01]. Quality [Bru77, IP76, Lin78, MM75, Pre80, KH19, MN23, Mor19]. Quantification [Hag99, Por92b]. Quantitative [Ell75, Fra88, GW74, KSW82, Len79, MM86, PN87, Ste86, WD97]. quantity [Mor19]. Quantum [Bru80, Cro91, Har81b, GL21, Pin11]. Quarks [RB90a]. Quarterly [SS17b]. Quasi [KC82c]. Quasi-Economic [KC82c]. Quebec [CLP90, ER78]. Queries [Por77]. Query [Bar82]. Question [Mac86, MM84b, MG82b, WL15]. Questioning [Bri08, Hil74]. Questions [Dea74, FDR08, MM96]. Quiet [Mac99a].

R [Blo74, For99, FOC^{+}71, Orl73]. R&D [Lic89]. Rôle [Gus99, Vau99]. Rabbit [Lan01a, Mar08a]. Race [And08, Ano77o, Del23, FS08, FDR08, Ful08, Har76, Har77, Hol81a, Hop23, Kah08, Mel09, Shi05, STM^{+}08, FR11, Jon23, KW15, Lóp22, M’c23, MvO23, Plá23, SMWCSC15, SMW15, SDL^{+}14, WLBRS15]. Race-Intelligence [Ano77o, Har77]. Racial [Eps08, Shi05]. Racialized [Ski06, OL18]. racism [Elg22]. radar [Eri72]. Radder [Sin98, Wyn98]. Radiation [Lin16, LA19, Ray86, Kuc13]. Radical [HI10, Pin97, Sim95, Wri80]. Radicalism [Ash96, SBH^{+}20]. Radio [DH08, Eri72, GM78]. Radio-location [Eri72]. Radiology [Bur08, RIA^{+}11]. radium [UB25]. rails [PH11]. Railway [Buc92]. Raise [Sco91]. Raised [Gum90]. Ramchandra [RH90]. Rand [Bru77]. Randall [Col07]. Random [Mir07]. Randomized [Ada02]. rank [GB16]. Ranking [Bru77]. Rapid [MCBDT00, CB24, OOO12]. rare [ÖS24]. Rate [Coh86]. Ratio [Lin78]. Rational [BS01]. Rationalities [Ray86]. Rationality [Cot75, Har80, Hel98, Kim94, Mah80, TDM82, Woj74, Kus10, O’R83]. Rationalization [OS05b, Ros77, Tou00, Ano77d]. Rationally [TY85]. Rawification [DG17]. Ray [Hes00, Law73]. rays [Gol04]. rBST [Tur01a]. Re [Aud01, BH97, Ber72, For94, GK16a, Ker00, LR01, lL99, MM82, FSF24, Jun23, TL18, EP07]. Re-Discovering [EP07]. Re-enacting [GK16a]. Re-evaluation [Ber72, MM82]. Re-examining [BH97]. re-imagining [Jun23]. Re-Networking [lL99]. re-orientations [FSF24]. Re-Presentations [LR01]. Re-Thinking [Aud01]. re-use [TL18]. Re/Design [LR01]. Reach [Eli00]. Reaction [Kro82]. Reactions [Hel98]. Reactor [Del83]. Readable [Gar73b, HSP02]. Reader [Del02, Del02]. Readers [Har01, Per04, Wis01]. Reading [Tut11, Var17, Mod12, NTW^{+}23]. Readings [Woo81a]. Readjusting [TH24]. Reads [Mer98b]. Reaffirming [Pow01]. Reagency [Shr05]. Reagent [CKM04]. Real [BS01, Fee00, Kir10, Kre71, Vin95, Gre22, Nis11, Var20]. real-time [Nis11].

Real-World [Vin95, Kir10]. Realism [Elk78]. Realist [Out76, UB25]. Realistic [For89]. Realities [FS01, Hof15]. Reality [Ano77d, Bow87, Ros77, vLR98]. Really [Wra03]. Realness [Rob02]. Reaping [Pyc93]. Reappraisal [AA82]. Reason [Agr01, JS17]. Reasoning [And92, LW10]. Reassembling [Lag12]. reassembly [Sha20]. Reassessing [AA84, Kim94]. Rebels [Hec96]. Reception [Bay83, Gol04, Kim91, Mes88, Oro77b]. reclamation [Bar12]. Recognition [Bla76, Hef79, Vos04, Del23, Ell11, Ham19, LOBB23]. Recombinant [Wri86]. Recombination [Hel98]. Recommendations [MG82a]. Reconceptualizing [Ber95]. reconfiguration [Mül18, Reu12]. reconfigurations [DPP^{+}20]. Reconfiguring [Hoe19b, MCBDT00, ZJ09]. Reconnaissance [Res78]. Reconsidering [Met06]. Reconstituting [Joh07]. Reconstructing [Hun82, Jas00, MM81, Jun23]. Reconstruction [Hak95]. Reconstructions [TJ85]. Reconstructivist [WHBM02]. record [BE00, Edg97c, Jac19, Str11]. Recorder [Lat86]. recorders [Bru15]. Recording [Hor04, Por04]. Records [Jen05]. Recovering [Bel20, CC76, Law77]. recruiting [WC18]. recruitment [Ham19]. Recruitmentology [Eps08]. recumbent [AQK15, Wie17]. recuperation [Max17]. recurrence [HRI22]. Recursion [Dor89b]. Red [Goo10, RS11]. Redefining [Rip91]. Redemption [TL00]. Rediscovering [RIA^{+}11]. rediscovery [Tsa24]. Redivivus [Mil92]. Reductio [Mey74]. Reductionism [Fue82, Hel98]. Redundance [Gie82b]. Redundancy [MM98]. Redux [FO87]. Reengaging [Col13]. REF [MN23]. Reference [BH97, Har02b, MM84a, SM93a, Sch88b, CKFS13]. References [AQL94, CM75, Wis01]. Referencing [Ano77m, Gil77, LRC87]. referential [Jat17]. Referring [All97]. reflect [Reu12]. Reflected [SM93a]. Reflections [Ans06, dBB76, Dea74, Hal88, Nis95, Por23, RH96, Ree07, Waj00, CS17]. Reflexions [FG86, Rad92]. Reflexive [Bir23, HJC23, HP91, Pot87, CBLG23, DSJ19, VA16]. Reflexivity [Dor89a, Dor89b, FO86, FO87, Ham92, Mul84, PM89, Gin10, Hoe14]. Refocusing [SSB24]. Reform [Cur98, McC88, Muk07a, Row12, Wil91]. Reformation [BL94]. Reforming [Col85b]. Reforms [CB97, Dix19]. Refutation [Nol92, Sle89b]. regenerative [Fau12, RC16, SF19]. Regime [Ald98]. Regimes [Boh04, Ver02]. Regions [ML94]. Regress [Alb95, Ken00]. regular [PP24]. Regulating [Dun78, Lev01, WH21]. Regulation [AS99, AD09, Eps97, Mir07, Pra06, RDB03, Yox87, MD20, SF19, ZDI^{+}23]. Regulatory [AR02, CKSW09, Leo10, MMB09, MLC06, SF19, Tur01a, Wyn92a, BHC16, Hed12, Hed14, Hoo17, KLBKC11, RvD22]. regulatory-scientific [KLBKC11]. Reification [Bra79]. Reinforcement [Hed10]. Rejected [Nev80]. Rejecting [Col14]. Rejoinder [WP74]. Related [EES96, Arm09, Arm10, Val10]. Relation [Chu75, Mul95b]. Relational [Var20, Ken15, Str16]. relationality [Plá23]. Relations [Mul98, Nes95, PKS88, Sol92, BKK21, Gra16, Leh18, MvO23, VI17, vW23]. Relationship [Bar82, Dro93, Lev79, Lew80, Mat05, Mor19]. relationships [PT18, VFWC15]. Relativism [Col82b, Dal91, Sto08, Col81b, Elk78, Jas96, Jen07, KC82b, Lau82, PB86, RH83]. Relativist [Gie82b, Kro82]. Relativist/Constructivist [Gie82b, Kro82]. Relativistic [Lat88]. Relativity [BH97, CC76, CC77, Lat88, Waz13, Law77]. Relax [Wel04]. Relenza [MD20]. Relevant [Dun85, Elz04, Jas87]. reliability [MB22]. Religion [Col07, Hes87, Mul95a, Ste90]. remains [Cal21]. Remaking [Car21b, Goo18]. Remarks [CKSW09, MvR85, MG82b]. remedial [RIA^{+}11]. remembrance [Lyn11c]. reminder [Lyn11c]. remodeling [Car22]. Remote [Law20b, AH11]. Removing [DP21]. renaturation [Tjh24]. Rendering [Tur95a]. Renormalization [Met06]. repair [Jac19, Ma19, PD19, Püt21]. Repairing [SH12, Ure14, Hen07]. repeated [Sab14]. Repertoires [MP87]. Replacement [MM84b]. Replicating [Tra81, JS18]. Replication [Mul84, Tra81, PP21]. Replies [Rad98b, Sle89a, Sto04b]. Reply [AA84, Blo04, BS04b, CLP91, Cla82, Col99a, Cra80b, Cro92, FR91, For94, FO87, Ful94a, Gie82a, GH84, Gin99, Hal89, Hic88, Joe99b, Joh85, KCM97, Kus04a, Lab95a, Lat92, Mac99c, MM87, Mag78, MI85, McK99, Mer98a, Nad84, Pic97, Pic80b, Pin81a, PB86, Pin99c, RS89b, RH85, RB90b, SS99a, Sin98, Sis95, Sis09b, Ste87, Tur90, Tur91, Wyn98, NK14b, Sis17a, Arm10]. Report [AB80, Baz84, Bru77, GS25, FS23]. Reports [Pin85]. Repositioning [SH12]. Repository [MC78]. Representation [Ake07, Ald98, BPBG06, Gar99, Tur01b, Moo10]. Representations [Don86, Edm02, Hyy06, Kwa87, Mar08b, Rob99, Ver08, FÖP20, Kuc13]. Representativeness [Dow07a]. Representatives [Mar08b]. Representing [CLP90, Wyn92b]. Reproducibility [Dow07a]. Reproducing [Dow88]. Reproductive [BV95, FC12]. Republic [Jac94, Mül95c, SR73]. Republics [Nes95]. Request [Onu84]. Request-A-Print [Onu84]. Require [Dal91]. Requiring [Hea07]. Research [AA84, Ada02, Ams87, AB80, Ano71b, Ano77d, Bal95, BFS95b, Bar71, Bea10, Bru77, Bud78, CK85, Cla80, CG89, Cue90, CK05, Den87, Dun85, DYS^{+}05, Eps08, FS93, FC79, Fra79, Fuj87, Gap95, Gil76, Goo10, Gra75, Hab97, Hac05a, Ham82, HMH80, Har03, HV93, Hef79, Hol81b, Huf86, Hug80, IP76, IM83b, IM85, Jab95, Jef03, Joh89, KC82c, KR00, LR08, LB05, Lew72, Lin16, LRC87, LC99, MK84, Mag75, MS00, MB81, Mel07, Mie98, Mir99, MV05, MO09, Mor82, MF95, Mul71, Mul94, Mul95a, Mul95b, Mül95c, OS05a, Ore03, Oud90, Pra06, Pra05, Pre80, Ron78, Ros77, Rou17, San72, Sch95, Sch90, Shr00, Sma77, Smi14, Som06b, Sta83, Ste78a, SK08, Wei74b, WG06, Whi75b, dC09, vK01, Avi15, BM14, CL11]. research [DH15, DP23, Dav20, Day18, Dem24, FFB11, FR11, GB16, GL21, Gu23, HHW^{+}19, Hed12, Hed14, Heg16, HRI22, Hof15, HL10, Huu12, Jac19, JSLS25, KHM10, LSW24, Lan20, Leo12, May19, McL15, Mon12, Pan11, PC12, PH11, PW19, PP24, PPM20, Reu12, RP15a, RCB18, RS11, Sar23, SFH12, Sov10, SS17b, WLBRS15, Yan22, Sis10]. researcher [AGKL^{+}15, VFWC15]. Researchers [MRL01, Foc16, Rei23, Rit16]. Researching [Jun23, Yea90]. Resentment [Yea85]. Reservations [War71]. Reserve [Den85]. Reserves [Bow85]. reservoir [AH11]. reshapes [PH11]. Resistance [Dun74, RC16, Sma20]. resistances [MS24]. Resolution [Rap01, HJC23]. Resolving [CERG25a]. Resonance [Joy05, Pra05, Pra07]. Resource [PT18, RS89a, Iid15, Sne12]. Resources [Ezr71, Fra79, GT93, Mar88, Mer02]. responded [Dah24]. Respondents [Sle89a]. Responders [Lab95a]. Responding [Sch10]. Response [Ano09a, Bar81, Bar98, Blo98a, CG92, Col85a, Col92a, Col95, Dal91, Dre92, Ede06, Fee00, Fle94, Fox94, Fra88, Ful95, GP97, Gie92, Gor91a, Hak95, Han84, Hir94, Kan78a, Kei95, Kel89, Kle91, Lau87, Mar95, Pin99a, Pin99b, Pot87, Rad98b, Rus86, Sch88b, SS99b, Sha04, Sim07a, Sis93a, Sto95, Sul99, Tay95, WZH^{+}25, Wei94a, Wol94, Wyn03, CE14, MS24]. Responses [CE03, Elg22, Yea01, HAB^{+}23, McL15]. Responsibilities [Wet07]. Responsibility [Ano77l, GB09, MK10, Nel77, RB75, Sch10, Kru13, Mey22]. Responsibilization [KBRB19]. responsible [Elg22, da 22]. Restoration [GR08, Hun82, Jac78, MM81, Ste90, LDR10]. Restraint [Kid88]. result [AW19]. Results [BH97, MM75, RB01, Ste87, Sun11]. Resuscitation [Tim96]. rethink [Skr16]. Rethinking [Hou01, Sol01a, Ure25, Wra05, Zie19]. Retraction [DGC18]. retractions [AW19]. Retreat [Gie82b]. Retreats [KC82a]. Retrieval [Wei74a, Wei74b]. retrospect [Lan20]. return [Tjh24]. Revealing [Mer96]. Revelation [Col07]. Revenge [Del02]. reversed [Cho13]. Review [Abr91, Alb80, Alb95, All76, AB80, Ano77d, Ano77g, Ano77e, Ano77f, Ano77i, Ano77h, Ano77j, Ano77k, Ano06c, Ash05, Ash07, Atk05, Bar79, Bar84a, Bar90, Blo71, Blo74, Blo77, Blo91b, Bra04, Bri08, Bro95, Bru77, Buc75a, Car05, Col07, Cor94, Cri10, Day08, Dea74, Dea95, Del03, Doi11, Edm11, Elz04, Eps09, Fis10, For99, Fox83, Ful94b, Gar73a, Gel09, Ger95, Gie06a, GM78, GW74, Gin10, Goo10, Gor90, Gre10, Hag94, Har94b, Har83a, Hir10, Hol81b, Hom90, Hyy09, Irw94, Jah72, Jen07, Kim94, Kuk94, Lan01b, Lat08, Lau90, LP11, Luk75, Mae07, Mah79, Mir07, MC79, Moo10, Mor94, Mos79, Nev80, O’R83, Old88, Pal94, Pin11, Pra09, Pre08, dSP75, Rav74, Ree06, Rei06, Rip91, Rob99, Ros10, Row12, Rud82, Rus94, Sch94, Sch72b]. Review [Sis10, Ste90, Sto08, SS91, Tan08, Twe93, Twi05, Ves96, Wal82, Wei90, Wei94b, Wel03, Wes88, Wri80, Wya07, DSJ19, Gre24, Hed12, Hor25a, Huu12, KdR19, Mar17, Tai19, Baz84]. Reviewing [Day08, Wie17]. Reviews [Del05, Haw73, Hol73, IG02, Kea73, Out76, Pfe71, Pri71, Rab76, SS91, Tua71]. revision [Eps21]. Revisited [CC77, Col88, For99, Jen07, Kea73, Kus04b, LSW95, Mil09, PN87, Szo92, CERG25b, Hu25]. Reviving [AQK15]. Revlin [Gor91a]. Revolution [Fra76, Har94b, Nic84, Pal71, Sol01a, Tha89, War71, Wis01, Sch20, Ste90]. Revolutionary [Gra75, Sch20, WSB79]. Revolutions [Sis12b, Lyn12b]. Reward [Gor84]. Rewriting [Mye95]. Rezeption [Bay83]. RFID [Sim10]. Rhetoric [Alb80, God97, Mer02, Mod00, Seg18]. Rhetorical [GHR00]. Rhetorics [Mul93, vLR98]. Rhizomic [US20]. Richard [Gin10, She20]. Richards [Lau90, Kel89]. Right [CC98, Pal94, Sah25, Jon17a, Pel21]. Righteous [Mia03]. rights [Eps21, Pop21, Smi17]. Rigoni [Avi15].

Rigoni-Stern [Avi15]. Rigor [VB03, Mur13]. Rise [Arm22, Eps08, LM80, Mos79, Pyc93, San78, Sch90, vLR98, MBDR24]. Risk [Abr93, AS99, Bij07, Dav20, GEJ79, Hac05a, Irw94, Rob02, Sim07b, TL00, Yox87, dMdF98, AB19, Jan18, KBRB19, Lak17, Man21, MB22, Nor17, Sah12]. risk-scores [AB19]. risk-stratified [KBRB19]. Risks [Mir07, MLC06, Wet07, vEG91]. Risky [Hen07]. Ritual [Lyn88, O’R83]. Road [Ber08, Leo10]. Roads [Del93]. Robert [Hol81b, Mos79, TS00, Boh91, Chr88, Gar04, Jas03a, Lau90, PN04, Ric04, Sha84]. Roberts [Mos79]. Robertson [Wri80]. Robin [Hyy09]. Robot [CS14, AMT11, Che23, Jon17b]. robotics [AMT11, Sab14]. Robots [Ala09a, CGPP24, FÖP20, Leh18, Lip23]. Roby [LcaM82]. Rochdale [Ink88]. Rockefeller [AA82, Bar84b, Cue90, Fit86, Fue84, Yox84]. Rocking [Lab95a]. Roger [Bar90]. Rogue [Bal06]. Role [BH97, Bla76, Blo91a, Cle86, Fre84, Fue82, Fue84, Gun05, Hin06, Kir10, Mul76, RJ76, Row12, Rud96, WG06, Gar13, Hed12, Hug24, RG14, SMWCSC15, TL18]. Roles [FL88, Her03a]. Roman [FM17]. Romantic [Hel98]. Rome [GT77]. Ronald [Bru77]. Roots [Gra85, Zil00]. Rosenberg [Ano77d]. Rosental [Gre10]. Rosing [Ano77n]. Roth [Pin90]. Rothschild [Wil73]. rough [vdH20]. roundabouts [Sis15]. rounded [FSF24]. roundtable [DGC18]. Routing [Ful89b]. Routinization [JL98, Dav24]. Routledge [Bar79, Blo77, Bro95, Del02, Hyy09, Luk75]. Rover [Ver12]. Roy [GM78, Nev80, Out76, Sch02]. Royal [AQL94, Mos79, Ber72, Jac75, MM81, dB72]. Rs [Har83a]. Rudd [Blo77]. Rudolf [Hom90]. Ruhr [Tod87]. Rule [Blo04, FR91, Kus04b, Sha04]. Rule-Scepticism [Blo04, Kus04b, Sha04]. Rules [Col82a, Hac84, Jon09, Wyn88]. Rupture [Hec02]. Rupture-Talk [Hec02]. Ruse [Lyn06b]. Russell [Ano77k, PB86]. Russia [Byc16, Gap95, GB09, Zie88]. Russian [Hol73, Mir95, Tor94].

S [Bru77, EHB^{+}97, Nev80, Che23, Met25, Mir18, Sin98]. S&TS [Del02, Pin95, WC99]. Sacred [JL72]. Sacrifice [Lyn88]. Safe [Sim05, Gre16a]. Safety [Del83, Hyy06, Bak11, BM19, Car14, JZ12]. Saga [Lew95, Tim94]. Sagas [OS05b]. Sagasti [AB80]. Sage [Nev80, Eva22, MB25a]. Said [Gor91c]. Salaam [WB78]. salaries [KG12]. Sale [Yea92]. Sales [Lak04]. Salesman [Gre04]. Same [Ald98, MM84b]. samples [Kow13]. sampling [Sah25]. samplism [Ure25]. San [MBDR24]. Sandy [Pin99c]. sanitary [Par22]. Santa [Bru77, Del05]. SAP [Hyy09]. Sapphire [Col01]. SARS [Hor25b]. SARS-CoV-2 [Hor25b]. satellite [Ben12, Law25]. satellite-based [Ben12]. Satisfaction [Her03b]. Satisfactions [LM75]. Sauce [RA96]. Savannah [Car21b]. save [Lau25, Par22]. Saves [Row05]. Saving [Tim96]. say [BL13, Dah24, Say14]. Says [Ful91]. SBPC [Bot90]. scaffolds [Nel13]. scalable [Sta18]. Scale [LS04, MB81, Pot11, Sch94, WN22]. Scaling [Yan05, PLPS22]. Scams [Joe99b]. Scandal [Lyn95a, Goo18]. Scandalous [Jen07]. Scandals [Mat14].

Scans [Ala08]. Scarlet [Doi04]. Scenarios [Gum20]. Scene [Cho85, Col94b, Lor17]. scenes [DH23]. Scepticism [Blo04, Kus04b, Yox84, Sha04]. Schaffer [Pin99c]. Scheduled [Hac92]. Schism [Pow01]. Schizophrenia [Hed01]. Schlumberger [Bow87]. Scholarly [Boh04, Kid88, War74, Fis10, Sok23]. Scholarship [WP74]. Scholastic [Hut91]. School [And08, HH83, JLM73, McC88, NK14a, Sol85, Lau13, Far75, SN89]. Schools [All97, ER78]. Schüll [Hsu13]. Schultes [She20]. Schuster [Kel89]. Schwarz [Rip91]. Science [Abr94, AR02, Abr91, All76, AB95, ABB^{+}08, AB80, And92, Ano77c, Ano77e, Ano79, Ano05c, Ano09a, Ano16b, Atk87, Aud01, Bak17, BFS95a, Bar82, Bar85b, Bar87, Bar90, BA12, BPBG06, Bay83, Bea77, Ber95, Blo75, Blo93, Boh89, Bot90, BL87, Bre97, BR79, Bru77, Bur83, Bur75, CLP90, CLP91, Cam04, CS00, CB01, CB97, Cha99, CT89, Cho85, Chu75, Col04a, Col85b, Col82a, Col85c, Col87, Col99a, Col99b, CE02, Col07, Con89, Cor94, Cow95, Coz85, Cro76, Cur94, Day08, Dea74, DNTS08, Del87, Del02, Del05, Den87, Dic00, Dir99, Dol71, Don86, Dor89b, Dro93, Dug87, Dun78, Eas73, EM00, Edm02, Edw94, EMB^{+}11, ER78, Elz04, Eva97, ELR21, Ezr71, Far75, For89, FNC77, Fra76, Fra88, FG73, Fre84, Fre90, Fri84, Ful94b]. Science [Ful06, Geu03, Gie82a, Gie82b, GH83, GW74, God07, Gol79, Gor82, Gor91a, Gor91b, Gra85, GSSD74, GKM75, GHR00, Gum90, Gus99, Hag99, Hak95, Ham92, Han86, HMH80, HH83, Har92, HvL08, Hed08b, Hed08a, Hed06, Hen90, Hen00, Hes87, Hic87, HP91, Hil04, Hil74, Hol73, Hol74, Hol81b, Hor25b, Hou01, Hug86, Hun82, Ink88, Ink91, JE75, Jan04, Jas87, Jas95, Jas96, Jas00, Jas02, Jas03a, Joh89, Joh84, JLM73, Jon09, KS79, KW15, Kim07, Kir03, Kle91, KC82a, KC82c, Kra72, Kre71, Kre73, Kri78, KP85, Kro82, Kwa87, Kwa05, Lab95b, LS04, Lag07, Lat86, Lau84, LMR10, Law73, Law97, Lei75, Len79, Lev84, Lin80, LC99, Lon71, Lor08, Low89, LW75, LJ98, Lyn04e, LC05, Ma19, Mac86, Mac99a, MM86, Mae07, Mag07, Mah14]. Science [MBB00, Mar08a, Mar96a, Mar96b, Mar97, Maz77, McC07, McC88, McC92, MC71, Mel03, Mel09, Mer96, Mer98b, MKC97, Mia99, Mil84, Mil04b, Mir95, Mir72, Mod01, MR73, Mor73, Mos79, Mul74, Mul84, MPA87, Mul95a, MM81, Mul73, Mye90, Nan72, Nel08, Nev80, Nis95, O’R83, Old90a, Pan80, Par88, Per89, PM79, PC84, PB84, PB86, Pin92, Pin95, Por92b, Pra05, dSP71, RELHR79, Rad92, RH96, Ram02, Ras98, Ree01, Ree07, Rei80, Rem88, RS89a, Ric04, Rip79, Rip03, RP16, Rob99, Roc74, Ron78, Ros93b, Rot05, Row05, Sag73, Sal71, Sch88a, SG74a, Sch88b, ST97, Sch87, Sha72, SB77, Sha80, Sha01, SK91, She06, She02, Shi05, Shi02, Shr00, Shr04, Shr05, Sim99a, Sim78, Sim79, Sim05, Sis12b, Sko73, Sla07, STM^{+}08, SS94, Sol01a]. Science [Sol01b, Som06a, SR73, SRM76, SR77, Ste99, Ste78a, Ste78b, Ste90, Sto95, Sto04a, SS17b, SWB77a, SWB77b, Sut84, Swa07, Tan08, TG86, Tim94, Tur89c, Tur03, Urb82, Van83, Vas15, Vel90, Ver02, Ves87, Vog06, Wal75, War74, WP73, Wes78, Wes79, Wes86, WSB79, Whi75a, Whi89, WB78, Wil91, WG74, WHBM02, Woo81b, Wra03, Wri75, Wyn92a, Yea85, Yea89a, Yea89b, Yea99, Yon94, Zil00, Zin71, Zin76, vK01, vdSvESW98, ÅSS21, Ano16a, Bar23, BTS22, BL13, BKK21, Bor17, Bro15, CS17, Col14, CEW17, CERG25a, Cra10, Dav21, DH23, DGC18, Doi11, Dur11, Edm11, Elg22, Eva22, FW23, FFB11, Gar73a, GK16b, Gre16a, Gre16b, Gu23, HHW^{+}19, Hof15, Hoo17, Hor25a, Iid15, Joh13, KBvLA22, KH18, KH22, KG12, KVA21, Kow13, Kum24]. science [Lap16, Law11, LPKL12, Lyn11a, Man21, MB25a, Mir18, NK14a, Pan11, PD19, Par22, Par25, Pet17, PP21, Pol14, Pop21, Püt21, Ran17, RG23, RCB18, RS11, RP15b, Sab14, SV22, SS17a, SMWCSC15, Sha16, SDL^{+}14, Skr16, Sma20, Smi14, SSB24, SBHY15, TS24, Tho25, Ure25, Waz13, WL13, Yai19, da 22, dBtMV21, Mer02, Alb95, AB80, Ano77n, Ano77q, Bay83, Bha76, Bru77, Dea95, Del05, Elz80, Hol81b, Mae07, Moo10, Ora71, Out76, dSP75, Twe93, Ano77h, Con79, For99, Gor90, Dán20, Hag94, Irw94]. Science- [Kra72]. science-in-the-making [dBtMV21]. Science-Lay [Shi05]. Science-state [Bak17]. Science-Technology [Bar82]. Science-the [Hol74]. Sciences [Ada02, AC88, Bru77, Clo01, Clo03, Coz85, Den03, Doe03, Hud73, Rud82, Sis09a, SBS^{+}00, Smy01, Day18, FC12, Jeo19, Reu12, Sha12, SC18, CG89, Jah72, Ma19, Nev80, Pra09, Sim95, Pal94]. Scientific [AA84, Abr93, AS99, AB91, All80, Ami74, AB80, AC88, Bar79, dBB76, Bla76, Blo71, BW83, Bra89, Buc75b, Cle86, Col74, Col89, Col96b, CLCMW23, Cue90, Del08, DC76, Doi04, Dol74, DB85, Dun74, DYS^{+}05, Eps98, Eri04, Fou97, Gas92, Gas95, Gas90, GK72, Gie92, GM78, Gil76, GM82, GR08, GSSD74, Gun05, Hac05b, Ham86, Ham82, HV93, Haw73, Her07, HP91, Hin06, Hof79, Hon08, Ink85, IM85, Irw06, JS89, Kim91, Kim07, Kir03, KC82c, Koc10a, Kop79, Kus04b, Kyv90, Lan81, LW82, LB05, Lev79, Lin78, Lub81, Lyn82, Lyn85, Lyn88, Lyn94b, MM82, Man93, MS00, McD97, MD09, Mer02, Mes88, MV05, Mir95, MvR85, MO09, Mos79, Mul76, MG82a, MG82b, Nev80, Nor78, Now87, Old90b, Pal71, Pal90, Pic94, Pin81b, PC84]. Scientific [Pin85, Pra05, Pyc93, RJ76, Roc81, Roc74, RS74, Ron78, Ruf79, SM97, Sch86, SG73, Sch90, Sha04, She02, Shr84, SCG01, Shr04, Sim91, Sin96a, Sin96b, Sis12b, Sle89b, SG74b, Sma77, SBS^{+}00, SS94, Sol92, SR72, Sta83, Sta85, SL91, Tho02, TJ85, Tou91, Vau99, Ves96, WD97, WB96, War71, Wei74b, Wei73, Wei94b, Whi76, Woo74, Woo76, Wra05, Wri86, Yea90, Yea92, dB72, vK01, Aar23, Bia22, BSB16, BL12, Bru15, Car21a, Cho13, Dav21, Dem24, Eva10, Fro12, GK16a, HR21, Hoe19a, KR11, KLBKC11, LC13, May19, Mcl21, RA22, Rit16, Str16, Sup14, Tsa24, VFWC15, dBtMV21, Luk75]. Scientism [Lyn04e]. Scientist [Ami74, Cam93, Ell75, KMH83, Lam10a, Mah79, MP87, Mul84, Ras04, She02, Son95, Ves96, Wes88, dG75, ÅSS21, AW13, BL13, Vis15]. Scientists [And08, Ano05c, Bot90, Cam03, Del05, FS93, Fox83, FS01, Fox05, Ful94a, GT72, Gar73b, GB09, GR86, GS02, Her03a, Her03b, Hil74, Hol81b, Hut71, Inh77a, Inh77b, JS89, Lam10a, Lev79, Mac79, MJM77, MR17, Nel71, Nel77, Pat83, RB75, Rud77, Sha88, Sil74, Son09, Tor83, Tor94, TY85, VK84, Ver02, Wer71, AS16, Cra10, Dem24, Law13, Lin16, SV22, Ano77f, Ano77l, Dow88]. Scientometric [Rab76]. Scientometrics [God07, SM93a]. Scope [Vin91]. scores [AB19]. Scotland [BJ12]. Screening [CC98, SM93b, KBRB19, TL11]. Screw [Pot87]. Scripts [Tim96, WE15]. Sea [Ore03, Van83, CC12, Gal24, Var20, dB09]. Sea-Floor [Ore03]. seabed [Gal24]. Seamless [Hug86]. Search [Ash07, Bru77, Col99c, Del05, Gom02, Ley91, Mag17, Lev22, Sok23, Zie19]. seas [PH14]. Seasick [Wyn03]. Seawalls [Muk07b]. Second [Rad98b, Roc81, Kim94, WB78, Wil91]. Secondary [HH83]. Secrecy [Bal06]. Secret [Bal06]. Secrets [Bea77, Dea95, Ano77e]. section [Dán20, Lyn11a, Sis16b]. Sector [Cas09, Wri80, HB23]. securitized [Pel21]. Security [Kin02, ELR21, Kom22, SP24, TL18]. Sedimentary [HRI22]. Sedimentation [DB85]. Sedimentology [Law80, Mil81]. Seductive [Lah05]. seed [Cur22]. Seeing [Goo95, KLW19, Ver12, BFW14, CWvdVdB23, Ell11, SVS21]. Seeker [Cha22]. Seen [Boh04]. Seismology [Bar03b]. Select [Gor84]. Selected [dSP75]. Selection [GR86, KSW82]. Selective [Mar78, Sol92]. Self [Eri04, Gre10, Lyn12b, PP21, Shr04, WD97, DPP^{+}20, Möl16, Püt21, Sti18, Gin10]. self- [DPP^{+}20]. Self-analysis [Gin10]. Self-correction [PP21]. self-driving [Sti18]. Self-Evidence [Gre10]. Self-exemplifying [Lyn12b]. Self-Organization [WD97]. self-repair [Püt21]. Selling [Bar23]. Selye [Vin99]. SEM [KG12, KG12]. Semantic [Sin96a]. Semantics [Gie91]. Seminar [Eas73]. Semiotic [Wag09, AMT11]. semiotics [dBtMV21]. Senegalese [Tou13]. Sense [Phi16, Sha01, Sim99a, FW23, SMWCSC15]. Sensing [Oud15, SVS21, Cal18, KLW19, Law20b, Leh18]. Sensitivity [vdSvESW98]. Sensory [SB18, Kim16]. senti [Lyo18]. senti-actuando [Lyo18]. separation [SBHY15]. September [Blo03]. sequencing [Tim15]. Serbian [Jan18]. Serious [Gar99, ÖS24]. Servants [Her03a]. Service [Blo91a, Mul72]. sessions [WC18]. Set [Col74, Col88, MB94, Pin81b]. Sets [Hed06]. Setting [Fro12, AR02]. Settings [Har02b, Eng20]. Settle [Mal98]. settlers [Ler18]. Seven [Col81a, Mar23]. Seventeenth [Dea92, Wri75]. Seventeenth-Century [Wri75]. Sewage [Van08]. sewer [CGPP24]. sewing [Jun23]. Sex [Gri18, Oud90, Seg18, Avi15, Pap21, Ple14, Sud22, ZDI^{+}23]. sex-assignment [Sud22]. sex-difference [ZDI^{+}23]. Sexes [Col81a]. sexism [BL13]. Sexual [Eps21, Fis04, Seg18]. sexualities [WE15]. Sexuality [JM73]. Seymour [Blo77]. Shadow [Ash07, Spa20]. Shallow [Pin99c]. shape [Del16, NL18, PW10]. shaped [Lev22]. Shapin [Ano77q, Nev80, Pin99c]. Shaping [AWN06, Ano77d, Cle86, Fau12, Hea04, HM03, MS88a, MS88b, MG92, Ros77, Ver88, Vin95, Yea90, Jau15, Gor90]. Shared [Fre84]. sharing [Eva10, Hay07]. Sharma [Har83a, Har83a]. Sheet [OOO12]. Sheriff [Olb84]. Shibboleths [Mar20]. Shifting [Gri18, KCBT07, Möl16, Mul98, Eng20]. Shifts [Bea10, McD76, MO09, Sin96a]. Shigeru [dSP75]. Shils [Tur95b]. Shipbuilding [And88]. ships [Leh18]. shipwreck [HMA14]. Shoal [Pin99c]. Shoot [Ful90]. Shop [Bre97]. Short [BL87]. Shortcut [TT18]. shortsighted [SZ11]. Should [Gie06a]. Shoulders [MJM77]. show [Mar13].

Shutting [Por82]. Siberian [For99]. Sick [Eps09]. Side [Ash96, Ash05]. signatures [KLBKC11]. Significance [Pin85]. Signing [Wyn96]. Signing-Up [Wyn96]. Signs [Bel20, Sha97]. Silence [Sub25]. Silent [Gol04]. Similarities [JS89]. Simon [Col91, Ful91, Gor91c]. Simonton [GH84]. Simple [Col92b]. Simplification [Sta83]. Simply [Hea07]. Simpson [Jas98, Lyn98a]. Simulated [Joh07]. Simulating [GT00]. Simulation [Blo85, NM20, Pre05, Smi15]. simulationists [Sun11]. Simulations [Gor91a, Joh07, Lah05]. Simulators [Joh07]. Simultaneous [Wes79]. Since [GGJ07, Mos79, Pic12]. sincerity [Dam25, WZH^{+}25]. sinew [Sne12]. Single [Ray86, Law20a]. Singleton [Rad98b]. Singular [Alm08, HMA14]. Sir [Bur75]. Sirs [Sim79, Wes79, Con79]. Sismondo [KC93, McH09, Tay95]. Sites [Gie06b, Lau16]. Situ} [She06]. situated [JS23]. Situating [AB95, HHW^{+}19]. Situation [Hen07, Mir95, DP23]. situational [MN23, SDL^{+}14]. Situations [Alm08]. Sivin [AB80]. Six [Gom02]. Sixty [RCB18]. Skepticism [Blo98b, iO18]. Sketch [Gin10]. Skill [And92, Law97]. Skills [Yea92]. skulls [Jon23]. Sky [Law20a]. Slaying [Mac99b, Mac00]. sleep [HLRK14]. Slezak [Col91, Col92a, Ful89b]. Slippery [LL13]. Sloterdijk [Skr16]. Small [Fau92]. Smalley [KR11]. Smear [CC98]. smell [Spa20]. Smith [Bar90]. Smithsonian [dB72]. Smoking [MBB00]. Smoothing [Jef03]. snake [Tan21a]. sociability [Pan11]. Social [AWN06, Ala09a, AB95, And08, Ano71b, Ano79, BA12, Blo85, Bow85, BL87, Bre97, Bro95, Cam03, CT89, Col81a, Cro00, Dem24, Den87, Dol74, Eas73, Edw94, Elz86, Eva97, Fou97, Gar97, GB09, GM03, Goo10, Ham82, Hec96, Hil74, Hof79, Hol81b, Hou01, Ink88, Joh89, Joh84, Joh85, KS79, Kel89, Kim94, KC95, Kra89, Kra72, Lat86, Law87, LM80, Lei14, Low89, LW75, Mac78a, Mac78b, MG92, MBB00, Mar08a, Mar96a, Mar97, MS00, McD76, Meh80, Mil84, ML94, MPA87, MM81, Mul73, Mye85, Nev80, Nor78, Old90b, Old90a, PW95, Pan80, PT72, PB84, Pow01, RS89a, RS89b, RB75, Rip91, Ron78, Ros93a, Ros10, Rus86, SM97, Sat07, Sau04, Sch88b, Sha80, Sha01, Sim79, Sim05, Sis93b, Sol85, SH89, Tra81, Ver88, Ves87, Vin91]. Social [Wat90, Wei90, Wes77, Wes78, Wes79, Wes89, Woo81b, Yea89b, dMdF98, vLR98, AQK15, AMT11, Arm19, Arm22, DPP^{+}20, Del21, Dur11, EM06, Ens12, Gu23, Hin14, Jon17b, KHM10, Leh18, PP24, iO18, Sab14, Sol01a, Sti18, Vis15, Mos79, Con79, Dán20]. Socialism [Bay83, Hol81b]. sociality [Tol19]. Socialization [Wyl19, CLCMW23]. Socially [Gor82]. Societal [Mir95, Yea01]. Societies [Lev79]. Society [All76, Aud01, Bot90, Con89, Edm02, ER78, Ell75, FL88, Gap95, GP09, Hag99, Hol73, Hol81b, Lau84, Nev80, O’R83, Pin92, dSP75, Ril86, Rip79, dB72, DSJ19, Jac75, MM81, Sha72, Bar84a, AQL94, Irw94]. Socio [CK85, CG92, Car09, EES96, Gra85, Hal06, Hed10, Hol81b, LMT92, Pre06, Sch90, Sco92, Ste78a, Fuk16, Tai19]. Socio-Cultural [Hol81b]. Socio-Economic [Ste78a]. Socio-political [CK85, Gra85]. Socio-Psychological [Sch90]. Socio-Technical [CG92, EES96, Hal06, LMT92, Pre06, Sco92, Car09, Hed10, Fuk16, Tai19].

Sociobiology [Alb80]. Sociohistory [Mac99b]. Sociological [Ano77o, Ano77p, Fre90, Ful90, GM78, Gin10, GHR00, Har76, Har77, Law81, Luk75, Mul71, Nev80, Ola96, Oro77a, Oro77b, Res82, RB90b, Whi75a, Woo89, Zil00, Jas95]. Sociological/ [Jas95]. Sociologist [KC93, Wes86]. Sociologists [Nic90]. Sociology [AA84, Ano77g, Blo92, Blo96, Boh04, Col04a, Col82a, Col89, Col96b, Coz85, Del87, Der00, Dol71, Dor89b, Dug92, Fre84, FL88, Gie82a, Gie82b, Hag94, Har86, Hen90, Her07, HP91, Kim91, Kim94, KC82a, Koc10a, Kro82, Kus04b, LR08, Len79, Lin80, Lyn94b, Lyn98a, Mac78a, Mac93, MS00, Mey77, Mil78, Mor88b, Mul79, Mul84, Mye90, PB84, PB86, Rud82, Sha04, Sim91, Sin96b, Ste78b, Wri80, Gre10, Whi75a]. Sociotechnical [Bow96, LR20, Sis20b, GS25, Hug24, Joh18, Sch20, Sma20, ST16, SH17, Uru22]. Software [BS08, Edw94, Hyy09, Mac03b, PWD07, Coo11, PBEW24, Hyy09]. Soil [AC88, FSF24, dlB15]. Solar [Dun85, Jan18, PBSK25]. soldier [Wag13]. Solla [Ano83b, Mac84]. Solomon [Hol81b]. Solution [RB01]. Solve [EES96]. Solving [BS08, Tan21b]. Somaticization [HL19]. Some [Ano77k, Bow93, Cla80, DC76, FG86, Hal88, HH83, Jen23, Lew72, Mag77, MM96, MM75, Mos77, Mul71, MG82a, Pfe79, Pin97, Sha01, Sis93b, Ste87, Wei74b, Wil73, Mal98, Gra98, MG82b]. Somerville [Del05]. Sometimes [Cam93]. Sommer [Goo10]. Son [Col81a, AB80]. Sonic [Hun24]. sonification [Sup14]. Soothsaying [Eva97]. Soraya [Mae07]. Sort [KCM97]. Sorting [Sis16b, Lei14]. Sound [Bij01, Hor04, PB04, Por04]. Sound-Recording [Por04]. Sounding [ML93]. Sounds [Bru15, Bra04]. Source [HR87, Dav24, LC13]. Sources [MK84, Vel90, AH11, dSP75]. South [Blo00, Cha22, Kim16, Pol14, SFH12]. Southeast [Sne12]. Southern [Wak04]. sovereignty [Lan20]. Soviet [Ano77e, Hol81b, Bea77, Eri72, Ger01, Gra75, Gra85, Hof79, Hol73, Hol74, Hol81a, Hol81b, KIH04, Lev84, Lew72, NFC83, Nes95, Roc81, Sch06, Sol92, Tor83]. Sozialismus [Bay83]. Space [Mel07, Red02, Ver08, dB09, RP16, Rot19]. Spaces [Tho99, Foc16, McL15]. Spacing [Yan22]. Spain [SM97]. Spanish [MD09]. Spare [SK08]. Sparks [Mor88b]. Spatial [AMT11, BH97, Bal06, Oud12]. Spatial-Epistemic [Bal06]. Speak [Cri04, Dae98, Hin14, dBtMV21]. Speaking [Por04]. Special [BH97, Clo03, CC77, Col82b, Hac05b, Hol81b, Thé09, BA12, Dán20, HBP19, Lyn11a]. Specialization [JM73, LR08, OM76, WD97, Wra05]. Specialties [GM78, Law73, Mul73, Shr84, SG74b, SH89]. Specialty [Lan81, Nad80, Sma77, SWB77b]. Species [Cri10, Bol18, Leo12]. Specific [LM80]. Specificity [CB01]. Spectroscopic [Baz84]. Spectrum [Jac94]. speculation [WC16]. Speculative [Jun23, CSTG23, ÖS24]. Speech [Vos04]. Speed [Rii08, Wie17]. spheres [Skr16]. Spiegel [Ano77n]. Spiegel-Rosing [Ano77n]. spirit [Rah11]. Spoken [Ful94b]. spokespersons [Pel21]. Spontaneous [RH83]. Sport [vHVdW07]. Spot [Gie06b]. Spots [Del87]. Spouses [MJM77]. Spray [Rap01]. spread [Abe13]. Spreading [KIH04]. Spreads [FLE97]. squares [Skr16]. SSafránek [Efm92]. SSK [Aga98, Col99a, Gor89b, Pin93, RA96, Sin98, Wyn92b, Wyn96]. SST [Ros81].

St [Nev80, O’R83]. St. [Ful90]. Stability [Met06, Mil04a, Anf21]. stabilization [Dav24, SMWCSC15]. Stabilizing [Gus99, Has12]. Stable [Hes86]. Staff [Zin71, LSS20]. Staffs [Nev80]. Stage [Hil04, GK16a, Smi15]. Stages [Col81b, Zin76]. Staging [Che23, RB09, Tak24, BL12]. Stake [CK83]. Stakeholder [LSW24]. Stalin [Hol81b]. Stalking [GP97]. Stand [Fee00, Rip91]. Standard [AR02, Hed14]. Standard-setting [AR02]. Standardization [Hog09, JL98, STM^{+}08, TB97, HJC23, RC16]. Standardize [Mal98]. Standardized [Alm08]. Standardizing [TL00]. Standards [Abr93, AD09, Ale08, Hal19, MLC06, Thé09, Tur91, War74, WP74, Yea89a, Sun11, Tim15, WM17]. Standing [Mir95]. Stanford [Col07, Mae07]. Star [Bal10, Ken00, CBLG23, Sla07]. Starring [Ash93]. Start [Ber08]. starters [PBSK25]. startup [Foc16]. State [Bot90, Buc79, Gui98, Hol81b, McC92, Muk07b, PT72, SA22, Sol01b, SWB77b, Waj00, Wal75, Woo82, Wri80, Bak17, Car12a, GKSS24, Lei14, Mah14, RP15b]. Statehood [Pre08]. States [GKM75, Bar23, BR79, Buc75b, Dun78, GEJ79, Orl73, Ste78a, Sud22]. station [GK16a, Lac16]. Stations [dB09, GK16b, HTB20, Heg16]. Statistical [Jas02, Mac78b, Sah25]. Statistics [MC71, Nor78, Del23, Lev14, Phi16]. Staying [Lan00]. Steam [Con78, Ink88]. Stell [DGC18]. Stem [Pra06, SAG11, SK08, RC16, SFP11, SFH12]. Step [Ful06]. Stepan [AB80]. Stephen [She02]. Stern [Avi15]. Steve [Wei94a, Wol94]. Steven [Nev80, Sis10, Ano77q]. Stewardship [Muk07b]. Stewart [Lau87]. Stichweh [Hom90]. stick [Has12, RCB18]. Sticking [Mar96b]. Sticky [SS18]. Stigler [MM87]. Stigma [Ram09]. Still [Lyn11d]. stimulation [Gar13]. Stock [Ken17, Pre06]. Stocktaking [TG86]. Stolzenberg [BS04b, Lyn04e]. stone [Wyl15]. Stop [Sto95]. storage [WN22]. Stored [Pel99]. Stored-Program [Pel99]. Stories [Mia99, Shr07, Wyl19]. Story [Cou99, Cur94, Del05, Lan01a, Mar05, Shr00]. Story-Telling [Cur94]. Straight [LC13]. Strangeness [Hir94]. Strata [Pin86]. Strategic [Lic89, MD20]. Strategies [AC88, Bow93, FR91, Kwa05, Lag07, Sim95, WH21, de 02, SAG11]. Strategy [Bud78, Edg03, FR89]. Stratification [HMH80, WH21]. stratified [KBRB19]. stratigraphic [Dam25, WZH^{+}25]. Strawsonian [Yea85]. stream [LDR10]. Street [MS14b]. Strength [Lag07]. Stress [Chi24, Vin99]. Stroke [Tor92]. Strong [Gor89b, Kim94, KC93, Koc10a, Nol92, Sle89b, Sle91]. Strongest [And92]. Structural [MF95]. Structuralist [Edw94]. Structure [BH97, Buc79, Dea12, GSSD74, Mul72, Nak12, Ril86, Sis12b, SG74b, SH89, Tod87, RS11]. Structures [Bow96, FFMW08, GGJ07, Kra72, Kri87, Mar88]. Struggle [Hon08, Wri80]. struggles [Bak17]. STS [AGKL^{+}15, CEW17, Edg03, For94, Gor06, Gum90, Hag99, HH83, Hom03, Jas12, LMR10, yLL14, yLL22, Lyn17, Mia12, Moo10, Rad98a, Rad98b, SL25, Sis16a, dV07, vHLW13, Gie06a]. Stuck [Vas15]. Students [Atk87, Blo77, Cam03, Eas73, FG73, JE75, Zin71, WB78]. Studie [Elz80].

Studies [Ano77n, BFS95a, BA12, BL87, Cam04, CT89, CE02, Con79, Dán20, Day08, Del02, Div90, Eas73, Edw94, EES96, Elz04, Fre90, Ful94b, Ful06, Gie06b, Ham92, Hol81b, Jas00, Jas03a, Joh84, KS79, Lat86, Lev84, Loh00, Low89, LW75, LC05, Mac86, Mag77, Mir72, Nev80, Old90a, Pan80, PB04, Rab76, RS89a, Rip03, RS74, RH83, RH85, Ron78, Ros96a, Row05, STGG95, Sim79, SR73, SRM76, SR77, Ste99, Vas15, Ver02, Vog06, Waj00, Wes79, WHBM02, Woo82, Bro15, CS17, DH23, Doi11, Dur11, FSF24, Hor25a, KBvLA22, yLL22, Lyn11a, Pin15, Ros10, RP15b, Wat14, WL13, da 22, Del02, SG74a, Sko73, WB78, Jas03a]. Studio [Thé04, Mar20]. Study [ABB^{+}08, Arm09, Blo77, Bro89, Bru77, Bud78, CL89, Cra80a, Elz86, Eps97, Far75, FL88, Gas95, GK72, GM78, GW74, Har81b, Hil74, Hol81b, Hou01, Jas95, Kan78b, KSW82, LW82, Mac78b, MM86, Mar07, Ola96, Oro77a, Pal90, Pic80a, Rea01, Ruf79, Sch90, Sho05, Sma77, Sol85, Sol92, SR72, Ste78b, SS91, Sut84, Tat86, Tor92, Tur82, Ves87, VB03, Vos04, WD97, Whi75a, Woo81b, Wyn76, Yea99, Bea10, Bra14, KH19, PW19, PBSK25, Rou17, Sch12, SC18, Zin71, Ano77p]. Studying [BH03, CK85, Rav99]. stuff [Bra24]. Styers [Col07]. Style [HSP02, de 02, Gre16b]. Styles [Har92, MM81, Sov10, Wei94b]. Sub [dV07]. Sub-politics [dV07]. subcultures [Pet17, RG14]. Subfields [Lew80]. Subject [Ala20a, JM73, KCBT07, Pet15, Reu12, Sta18]. Subjected [Asd08]. Subjectifying [GL21]. Subjective [GB09, Tur82, Ano16a, Sha16]. Subjectivity [Mia99, Phi16, Sha12]. Subjects [JLM73, Gra16, KH18, Sis10]. Sublime [Sup14]. Submarine [MM98]. Submission [Gor84]. Submissions [Mer02]. Subordinates [Hef79]. Subramaniam [Del02]. subsistence [PT18]. Substance [Gom02, HSP02]. subtle [BL13]. Subverting [Wyn03]. Succeed [MB81]. Success [Bar84b, Cou99, Lie87, Wol95, Heg16, LC13, TT18]. Successful [Tor92]. Sudden [Huf86]. Suggestions [Mul71]. Suitable [CC77]. Sullivan [Mac99c]. Sum [Koe99, Mac99c, Rav99]. summarizing [SBHY15]. sun [Jan18, Pin81b]. Sun-Set [Pin81b]. Sunset [RH90]. ‘Sunshine’ [Mat05]. superconductivity [Gu23]. Superorganisms [KC95]. Superpowers [Ink91]. Supply [Row12, Win02a, Win02b]. Support [Olb84, Tor94]. suppression [SV22]. Supreme [CL11]. Surfing [Law11]. Surgeons [Joh07]. Surgery [Col94a, Fox94, Hir91, Lyn94a, Woo06, JS18, Ple14, Sud22]. Surgical [Joh07, Pre05, Lee16]. Surrogate [Eps97]. surrogates [Mos16]. Surveillance [Lan09, Raa97, Ben12, Kim16, MS24]. Survey [Bur83, FG73, Jab95, MS00, SBS^{+}00, Muk07a]. Surveying [Raa97]. Survivors [Lin16]. Susan [Hol81b, Bal10]. sustainability [WM17]. Sustainable [KHM10, SBH^{+}20]. Sustaining [Oud15]. Svalbard [RP16]. Swain [dSP75]. Sweden [Elz80]. Swedish [BB17, KH22, Rou17]. Switches [GK72]. Sydney [Sch02, Bed91]. symbiotic [FO22]. Symbolic [All97, Bur08, Cam85, Len16]. Symbolism [Bij01]. Symbols [Sma78]. Symmetry [Koe99, Pel96, Rii08, Lyn17]. Symphony [Bij01]. Symposium [Ano06c, Sim91, Wei73]. symptom [HAB^{+}23, Ras20]. Synchrotron [VP98].

Syndrome [Eps09]. syndromes [Gri18]. Syntax [Gie91]. Synthetic [CS17]. System [Car09, CG89, Gap95, Hil74, Hyy09, Jab95, Jas01, Law84, LT92, MS88a, MS88b, Tod92, Vin95, Bar23, Ben12, Pri71]. systematic [SBHY15]. Systemic [MB22]. Systems [BFS95b, Blo91a, Hom90, Mac03b, Sat07, Sch95, Shr84, Wes89, Win02b, BHC16, HB23, Mos16, Uru22].

T [Ros96b]. Table [Har81a]. Table-Turning [Har81a]. Tacit [Col74, Col01, Doi11, Hor04, Doi11]. Tactics [NL18]. Tadpoles [Mos77, Ano77k]. Taiwan [SFH12]. Taiwanese [Tsa24]. Taken [Del93]. Taking [DH23, Hay07, Mey22, Pal71, War71, Woo06, Wya07]. Tale [Sto82, Tod87]. Tales [Tur01a]. Talk [Hec02, Irw06, MP87, Mye98, Nic90]. Talking [Hal88, Sto95]. Talks [Mul84]. tandems [BSB16]. Tangled [Del05]. tank [McL15]. tannins [Bes20]. Tantalus [Col99c]. Tanzania [GK16a, WB78]. Tape [Lat86, RS11]. Tape-Recorder [Lat86]. Task [GR86, Mag07]. taste [Ano16a, Phi16, Sha16]. tastes [GL21]. Tavistock [Jah72]. taxonomy [EE14, LT12]. Taylor [Sis95]. tea [Bes20, Col74, CH75]. Teacher [Gie04]. Teachers [Bru77, Bur83, Blo77]. Teaching [And08, Atk87, CT89, GGJ07, HH83, JLM73, LW75]. Team [Ral06, PP24, Lyn98a]. tech [Foc16, SDSR25]. Technical [Ale08, Bru77, CG92, Del83, EES96, Fan09, GT93, Hal06, Hol81b, LMT92, Lyn82, Nel75, Pre06, Sco92, Shr84, Vin95, Car09, Fuk16, Hed10, Lyn11a, Smi17, Sne12, Tai19]. Technician [Tim03]. Technicians [Hol81b, Lyn98b, BB17]. Technique [JL98, Rii08, FC12, Sch13]. Techniques [Dow07b, HTB20, SDSR25]. Techno [DPP^{+}20, Vau99, LR20, MB25b, RvD22, SvD25]. techno-economic [MB25b]. techno-market [LR20]. techno-politics [SvD25]. techno-regulatory [RvD22]. Techno-Scientific [Vau99]. Techno-social [DPP^{+}20]. Technocratic [Hec96, McC88]. Technological [And88, Ano71b, Bij07, Bim90, BH03, BS98, Con78, Dow07a, Fou97, Gar13, Hof79, Kir10, Leo10, LC99, MM98, NHM74, Pet00, Thé04, Waj86, Wes89, Max17, Mil12, Mül18]. Technologies [BV95, Dow07b, Fau98, LM75, LP11, MD09, PB04, PK09, Smi15, Vos04, Arm19, MB25b, Mos16, Oud12, PI19, Red16, RIA^{+}11, SS18]. Technologists [Gar73b]. Technology [AB80, Ano09a, BFS95a, Bar82, Bha76, Bij01, Boh04, Buc75a, Col85a, Cot75, Del02, Dro93, Dug92, Dun78, EES96, Elz04, FFMW08, Ful08, God97, Gor92, Gum90, Gus99, Hår94a, Ham92, Hol81b, Hou01, Hug86, Ink91, Joe94, Joh84, Joh85, Law87, Loh00, LW75, LC05, Mac86, MS88a, MS88b, Mac93, MG92, Mar97, Mat05, McC92, Met88, Mil03, Mul98, PB84, PB86, Pin92, Rad86, Rad92, Rap01, Red75, Rip91, Rob02, Ron78, Ros93a, Ros96a, Row05, Sag73, Sha84, Shi02, Sim78, Smi09, Sør92, Spi87, Ste99, Szo92, Vas15, Vin95, Waj00, WB78, WG74, Wri86, Wyn88, Yox84, dLM00, vLR98, AQK15, Bro15, CS17, ELR21, Hol22, Hor25a, KBvLA22, KR11, KLW19, Lag12, LW10, Mag17, Max17, Möl16, Mon12, PP22, PW10, RP15b, Sab14]. technology [Sad24, Sim10, Sma20, SSB24, eS23, WL13, WC18, da 22, vW23, Hol81b]. technopolitical [Hug24]. Technoscience [And02, Eps09, Gor92, Lan01a, Hu25, Jun23, KRR13, Lóp22, MMV15, Oik17, Pap11, dlB11]. Technoscientific [Möl16, Tur90, dlB15, BD11, Car12a]. Technoscientization [VA16]. technosexual [WE15]. technotopias [Jön16]. Techonology [Twe93]. Tectonics [Mes88]. Teil [CG92]. Telecare [Oud12]. Telephone [Gor97]. Telescope [IM83b, IM83a, Smi83]. Telescopes [McC00]. Television [Efm92]. Television1 [Col87]. Telling [Cur94, Tur01a]. temperature [Gu23, SPBG24]. template [Law13]. temporalities [Ial21, KRR13, Sch15]. temporality [Ran17, SBH^{+}20]. tempos [Tou13]. Tendencies [Bay83]. Tendenzen [Bay83]. tension [CERG25a]. Tensions [Anf21, Hac05a, Loh00, Pin15]. Term [BPBG06, HRI22]. Termination [Ros81]. Terms [Irw06]. terra [Mes17]. Terrain [Cow95]. territories [LS10]. territory [Hen18, Mah14]. Test [Col92b, Gor84, Oro77b, Zie88, Rob16, ST24, WE15]. Testing [Bar03b, Bug93, Fra88, Gel09, Hal98, Kuk94, Leo10, Nel08, Par05, Sim99b, Spi97, KKLL21, PD19]. Tests [Dow07a]. Texas [Con89]. Text [Edw94]. Textbook [Woo06]. Textbooks [Day08, Gas90, Smy01, BM19]. Texts [Mye85, Ree01, AC22, Oik17]. Textual [BL94, Oik17]. TFTR [Bro82]. Thalidomide [TL00]. Thames [AB80]. Theater [Pra05]. Theatre [Ash93, Sim10, Smi09]. Their [Hin06, Lan01b, Sch90, Son09, Whi76, CL82, CS00, HH83, Lev79, MK84, NL18]. Them [MB81, God15]. Theme [Mar04]. themed [Sis20b]. Theoretical [Cra80a, GP97, Ken00, KCM97, MKC97, Rud96, GL21]. Theoretician [Ken00]. Theories [BW83, Hof79, Lau90, Mes88, RH96, SH17]. Theorists [MR73]. Theorizing [Sim25]. Theory [Aga98, Akr23, Bar79, Cro91, Edw94, For89, Fra76, JL72, JE75, KIH04, Kai06, Kim91, KSW82, Kre73, Mac78b, Man87, Mar78, Mes88, MM96, Out76, Pal90, Ros96a, Say14, Sin96b, Vin99, WSB79, Del21, Lag12, Mcl21, Moo10, SA22, Ste12, SS17b, Waz13, Luk75]. Therapeutic [Eps97, Mar24, Ric88, Pet15]. therapies [CGP25]. therapy [Pic19, SFP11]. There [BP80, Van08, OL18]. thermal [KLW19]. Thermometer [Cur98]. These [Mac03b, BL13, Lyn11d]. Thesis [Koe99, McD76, KR00, Spi97]. Thing [Vic95, De 24, LP11]. Things [Ald98, Bow96, Jah72, Mac03b, Sim07b, Woo06, Bel20, BL13, Lyn13, PC12, dlB11]. Think [Fis10, McL15, Ros10]. Thinking [Aud01, Elk78, Hof15, yLL22, MKC97, Mye98, Shi02, dLDV21]. Third [Cal99, CE02, CE03, Dic00, Jas03a, Mae07, Rip03, Tur03, Wyn03, Hig21, She06]. Thomas [Car05, Bar84a, CG90, EHB^{+}97]. Thompson [Rip91]. Thought [All97, SN89, Wei94b, Wel04, de 02]. Thoughts [Pic99, Rad98b, VB03]. Thousand [Dea74, CE16]. Threat [Ano09a, SM25]. Three [Bim90, CD84, CFS01, Del87, Lew80, Mae07, STGG95, Tod87, PJ17]. Three-dimensional [Mae07]. Throw [Sha01]. Ticker [Pre06]. Tier [Elk78]. Time [Wya07, And21, Mar24, MB25a, Nis11, SZ11, Var20, dlB15]. Times [Sha97, Eva22, GK16b, Var17]. Timing [BJ12]. Tinkering [Wak04]. ’tis [Dis04]. tissue [Jon23, Rad13, Str16]. Tissues [Hog09]. Titanic [Bie18]. Titles [Whi89]. Tobacco [RDB03]. Today [BS04a, Mir95, Sch06, Hol81b]. Together [LW82, Hoe14]. Tokamak [GT93]. tokens [MB25b]. Tokyo [dSP75]. Tolerance [Ald98]. Tomorrow [Sch06]. Tongue [Ful94b]. Tool [Bal06, CC98, Fra97, Hic87, Has12]. Tools [CB01, KIH04, Pal94, Chi24, Gar13, TL11]. topic [Doi11]. Topography [MC97]. Topology [Coh86, ML94, RB99]. Tort [EM00]. Total [Sad24]. Tower [Lam10a]. towers [DK12]. Toxic [Cal21, EM00, LTC18, Rie03, Sch10, Dur20, Fis18, Tir18]. Toxic-Exposure [Rie03]. toxicity [Bak11]. Toxicogenomics [Sho05]. toxicology [Dem24]. Traceable [Lez06]. Tracing [DS20, Jön16, Mur13, Spi87, Swa24]. Trade [BL12, Mac79]. Trading [Bru15, Gor02, HTB20, Mac17, MB22]. Tradition [Bak11]. Traditional [Tuu05]. Traditionalist [Lam10a]. traditions [Mos16]. traffic [Law20a]. train [Day18]. Training [DH23, Tür73]. Trajectories [And88]. trame [Gre10]. Trans [Kum24, Sch13, Wes88]. Trans-science [Kum24]. transaction [PI19]. Transactions [AQL94]. Transantiago [Ure14]. Transcultural [HTB20]. Transdisciplinary [TS24, FLK^{+}14]. Transepistemic [KC82c]. Transfer [Gor06, Law87, Yox84, vW23, Gus99]. Transformation [BFS95b, Bar03b, CL82, DeW75, Gap95, Gil76, Har96, Lyn88, Mül95c, Sch95, BB17, Pic19]. Transformed [GM78]. Transforming [Row12]. Transgenes [Del08, BFW14]. transgenic [BD11]. Transgressions [Del08]. Transgressive [HJ13]. transit [CB24]. Transition [Boh04, Wol95, Anf21, Gum20]. Transitional [Gap95]. Transitioning [Tim15]. transitions [ST16, SBH^{+}20]. Translated [Gre10, Gin10]. translating [Por23]. Translation [Mor00, SG89, Kom22, Pel21]. Translational [BM14]. Translations [LR01]. Transmissions [Sis17c]. Transnational [SG73, HH22, Mur15]. transnationalities [Law20a]. Transonic [VB03]. transparency [HR21, Law20b, PP22]. Transporting [Tur90]. Transposing [FC12]. trauma [PAB15]. traumatic [Hol22]. Travel [Rud96, KH22]. Traveling [PS24]. travels [Pih25]. treatable [Pic13]. Treatise [RH90]. Treatment [CC77, Kid88, Van08, Wyn76, Swa24]. Trends [BPBG06, Bia22]. trespassers [Hor25b]. Trevor [Col22]. Trial [Hen00, Jas98, LC05, Bri13, Gra16]. Trials [Pic94, Yan05, BSB16, Hed14, KLBKC11, LW10, Lev22]. Triazolam [AS99]. tribulations [LW10]. trickle [SV22]. trickle-down [SV22]. Trident [MS88b, Spi87]. Triple [Shi02]. Tripos [Gas84]. Triumph [Mul94]. Tropical [AC88, GK16b]. Trouble [DH08, Sch95, MvO23]. Troubled [Fau07]. Troubles [Ala20a, Ala20b, Mat13]. Troubling [LP11, Mur15]. Trow [Blo77]. True [Lan00]. Trump [Man21]. truncated [KKLL21]. Trust [BHC16, Col01, Hag99, Hed12, KH22, McD97, ÖS24, ST97, SCG01, Tim15, BSB16, Par25, RG14, SP24]. Trusted [Chu75]. Truth [Bar98, Dea92, Gie06b, Mey74, Mir07, Rei06, JS17, Lyn17, Sis17b]. Truth-Spot [Gie06b]. truthing [WE15]. truths [Jat17]. Tuberculosis [Wei90, Eng20]. tug [SB18]. tug-of-war [SB18]. tuning [Oud15]. Turbine [Bra92a]. Turbines [Con78]. Turbulence [LB19, Tur95a]. turbulent [Day18]. Turkish [Tür73]. Turn [Gar99, Lat86, Pot87, CGP25, CKFS13, Fro17, Tre13, Vas15, WL13, WL15]. Turnbull [Dal90, Dal91]. Turney [Rob99]. Turning [AH20, BTS22, Ber95, Har81a, Lat07, Rii08, vHLW13, vHLW13]. turnoffs [Sis15]. turns [Sis15]. turtle [CC12]. TWAP [CFS01]. Twentieth [Pal94, Wei90, Bud91]. Twentieth-Century [Pal94, Wei90]. twenty [Sis12b]. twenty-five [Sis12b]. Twice [Kuc13]. twins [Wes24]. Twist [WO96]. twitter [Bru15]. Two [Ano79, Ano10, Ash96, Blo71, Col95, ER78, Elk78, Elz86, Ink91, Jen07, Mye85, Mye95, Pel99, RH83, SR72, Tor92, Woo81a, Mia99]. Two-Stroke [Tor92]. Two-Tier-Thinking [Elk78]. types [Del23]. Typhus [Wei10]. typification [KBvLA22]. Typing [Dae98]. Typographic [Har02b]. Typologies [SH17]. Typology [Mag75, SDSR25].

U.S. [Mul72]. UFOs [Wes77]. Ugandan [Cal21]. UK [Hol81b, Nev80, O’R83, dSP75, Abr93, Blo91a, Ell11, EMSL23, Joh18, LR20, Lor08, MB25a, MM84b, MN23, Par05, SM93b, STM^{+}08, SAG11]. Ultimate [MG82b]. Ultracentrifuges [Elz86]. ultrasound [Nis11]. Umbrella [Whi76]. Un/common [DS20]. unaligning [VI17]. Unanticipated [Gar04]. Unboxing [SDSR25]. Uncertain [Mat05, CWvdVdB23]. Uncertainties [dMdF98, JZ12, Pol20, VT21]. Uncertainty [Cam85, DA01, Eps09, Eva97, Lev01, Lyn98a, MM87, MMB09, Smy01, Sta85, Bar11, Eva22, HJC23, Rou17, SM25, Str11, Lah05]. Unclearing [Naf24]. Unconventional [PW19]. Undead [Sim99a]. Underdeveloped [Sag73, Ves96]. Underdevelopment [Sag73]. Undergraduate [Bru77]. Underground [Ver08]. Underlying [Mac03a]. understand [CS20]. Understanding [CG90, Col87, Gus14, KMH83, McL15, MLC06, Sau04, Sha01, Tur89c, Wyn88, Yea99, dBtMV21, Kum24]. undisclosed [SO17]. unexpected [Naf24]. unexplained [Ras20]. unfortunate [Bar11]. Unification [KSW82, Wol95]. uniform [Hal19]. unimaginable [Jön16]. Unintended [Gar04, Wil91]. uninvention [Ste12]. Union [Ano77e, Hol81b, Bea77, Dow88, Fau12, Hol74, KIH04, Lew72, MLC06, Sch06]. Unionism [Mac79]. Unique [Jon23]. uniqueness [TT18]. Unit [Smi14]. United [Rip91, Bar23, BR79, Buc75b, Dun78, GEJ79, GKM75, Orl73, Ste78a, Sud22]. Unity [Swa07]. Universal [Bow93, MM87]. Universality [O’C93, TB97]. Universitaires [Gre10]. Universitätsverlag [Pfe71]. Universiteta [Rab76]. Universities [Ber08, Div90, GGJ07, RS74, FFB11]. University [Alb95, Ash07, Bru77, Col07, Cri10, Del05, Doi11, Eps09, Fis10, Gin10, Goo10, Gre10, Hol81b, Hsu13, Kim94, Lam10a, LC99, Mae07, Met06, Mir07, Mos79, Nev80, OS05b, PKS88, dSP75, Pri71, Rei06, Rob99, Ros10, Sis10, Vel90, Byc16, PC12, Tou13, Zin71, Bur83, FG86, MK84, Mor73, Tuu05, WG06, WB78]. university-industry [Byc16]. Unknowable [Hil01, Coo19]. Unknown [Wal01, Wag13]. unmaking [HL19, Wag13]. Unnatural [Rob96]. Unpacking [SK08, MB25b]. Unruly [Lee16, Wyn88]. Unsettling [Mur15]. Unthinkable [GT00]. upon [AA84, Coh86, CBLG23]. Urban [Gie06b, LS04, Row12, Ver08, WC99, Gum20, MBDR24, Rek14]. Uros [Ken13]. US$24.95 [Mir07]. US$27.50 [Mir07]. US$34.95 [Mir07]. US-USSR [Lub81]. USA [Ale08, Car09, CC98, Cha99, CC12, Doe03, Gol04, KIH04, MM84b, Par05]. Use [Blo00, DB85, Fra97, Hyy06, Mil76, Muk07a, PC84, Smi09, Sol92, Woo76, Bea10, Car14, FR11, MD20, TL18]. Usefully [Dol75]. Usefulness [Mar78, Mye89]. User [MCBDT00, Mul98, Ver08, Eng20, FÖP20, Ros11]. user- [Eng20]. user-centered [Ros11]. User/Using [Mul98]. Uses [Del83, Hil90, Sha72, Sol85]. Using [Fra88, FL88, GB16, Gor92, MCBDT00, Mul98, SR72]. USSR [Hol81b, Lub81, Mir72]. Usual [Mal98, AH20]. Utility [Mul79, KLBKC11]. Utilization [Onu84]. Utopian [Mar97].

V [Rud82, Wri80, Rab76]. V. [Wri80]. Vaccine [Dal90, Tur89a, Tur90, SA22]. Vaccines [Mar96b]. vagueness [Püt21]. Valence [SBH^{+}20]. Validity [Bow85, Kim91]. Vallgårda [Arm10]. valuation [Bie18, Bir23, Hoo17]. Value [EP07, Hog19, PPM20, Coo11, Hoo17, Mar24, PS24, RL16, Rie15, Str16, Tho25]. Value-based [Hog19]. value-ladenness [Tho25]. value-making [PS24]. Values [And88, Del83, Mil92, RG23, Sat07, Sør92, Mag17]. Variable [Eng20, Pap21]. Variation [Ams87]. variations [Gri18]. Various [Ake07, Pin86]. Vedic [WO96]. Vegas [Hsu13]. vehicles [TS21]. Velikovsky [Dol75]. Velocity [Eli00]. vendettas [RG23]. Venezuela [Roc74]. Vents [Ore03]. venture [Bir23]. venues [Sis16a]. Verbal [Ful89b]. verbs [Sch12]. Verfahrenstechnik [Buc79]. Verran [Ken15]. Versions [FFMW08, DA24]. Verstehen [Cri96]. Versus [Rie03, Car09, FR11, Jac94, Kim94]. Vertebrate [SG89]. Very [Ful89a, Nol90, Sto08]. Veterinary [Woo06]. Vey [Col92a]. vi [Col07]. via [Row05]. Vicissitudes [Bar79]. Victim [Jas02, MF95]. Victorian [Chr88, Ham86, Mor73]. View [Bro95, Dal91, Ell75, Fre90, Hil90, Jas95, KC93, Kra89, Lab95b, Lev84, Sag73, She02, Thé09, Coo11]. Viewpoint [Red75]. Views [Ano79, GM23, Met06, Sol85, Car22]. viii [Col07]. Violation [WSB79]. violence [Lan20]. Virtual [CERG25a, CERG25b, RG14, Sis25, Tho25]. Virtuosity [Col88, Ell11]. viscosity [SS18]. Visibility [Lyn85, RELHR79]. Visible [Por96, Nis11, SMW15]. Vision [McC88, RB99, DGC18, Plá23]. Visions [FFMW08, Mag17, Nic97, Rob96, Ben12, CS14, GS25, SBH^{+}20, TS24]. Visual [Law25]. Visualization [Ver12]. Visualizing [Var20]. vitae [KdR19]. Vitamin [Ric88]. vitro [DK12, Jön16, She06]. Vocabularies [Yea85]. Vocabulary [AB91]. voice [Kan22]. Voices [Bix97, Mac99a, Mar95]. Volume [Ano02, Ano04, Ano05d, Ano06b]. Volumes [Rud82]. Voting [Hir10]. Voxels [Bea01]. Vries [Lat07]. vs [Hor25b]. vulnerability [PAB15].

W [Bru77, Hol81b]. Wait [Rie03]. Waiting [Zie88]. Wajcman [Suc06]. Wake [All07, SPBG24]. Wales [Blo00, Div90]. Wall [Som06a, MS14b]. Wallis [Nev80]. Wanted [Lyn98b, Kis13]. War [Ano77e, Ash96, Bea77, Bud78, Cal99, Cro76, Fan09, FS93, GT00, Gor77, Mel07, Tur03, Whi75b, Hed08b, Law11, SB18, Clo03, Den03, Ger01, Hou01, Mac79, Pop21, SG73, SH12, Sne12, Sol01b, vK01, Ano77j]. Warfare [KC85, Lyo18]. warming [CBLG23]. Warped [Sto82]. Warranting [GM82]. Warren [WP74]. Wars [Mac99a, Sto04a, Sla07, Lyn04e]. Wartime [Hug80]. Was [TS00, Kis13]. Washington [Bru77, Del05, Jan04]. Wasserman [Del05]. Waste [Mac03a, MC78, Van08, HLRK14, Ial21, LR20]. waste-energy [LR20]. wastewater [UB25]. Watch [DGC18]. watchdog [DGC18]. watched [Car18]. Water [Ala09b, Bak12, BA12, Car12a, Con78, Row12, Bij12, Sne12, SB18]. Waters [Jef03]. Watershed [LR01, Car12b]. waterways [Sha20]. Wave [CE02, CE03, Fra76, Jas03a, Rip03, Wyn03]. Waves [Col99c, Jas03a, Ash07]. Way [Del93, Dic00, Gol94, Mod00]. Wayne [Del02]. Ways [PK09, RS89b]. Weak [KSW82, SWB77b]. Weak-Electromagnetic [KSW82]. Weapon [MS88a, MS88b]. Weapons [Gus06, Hug06, Rap05, Spi87, Spi97, SH12]. wearable [Jun23]. Weasel [Del02]. Weather [Har03, Jan04, Mas10, Muk07b, Ran10]. Weaver [Olb84]. Weaving [Gre10]. web [Sil11, Zie19, Hug86]. Weber [AH20]. Weighting [RB09]. Weinberg [Ful94a]. Weinel [Sis17a]. Weingart [GM78, Pfe71]. Welcome [Bro95, Tha89, Mos16]. Welfare [Sch72a]. Well [Bow87, FSF24]. well-rounded [FSF24]. West [Meh80, Cho85, Han86, IM85, OOO12, ST16]. Western [AB80, Ano77e, Bar84a, Mos79, Bea77, Cho85, Dug87, Han86, SK91]. Westview [Hol81b]. We’ve [Ler18]. Whaley [Del05]. Whatever [Tur12b]. Wheatsheaf [Rip91]. Wheels [Con78]. Where [GM03, Har01, Inh77b, Kli09, Mia12, Vog06]. Which [Col96a]. While [Pyc93]. White [Ler18, vK01, OL18, PD19]. Who [Col85d, Eas74, Son09, TS00, Oud97, SAC^{+}20]. Whole [Ano06c, Ans06, Gus06, Hug06]. Whooping [Dun91]. Whose [Cal99, Kai06, Yon04]. Widening [Zin71]. wider [CERG25a, Sis17a]. Wikipedia [FW17]. Wild [Hea07, Muk07b, ZJ09]. Wildlife [Dun91]. Wiley [GM78]. Wiley-Interscience [GM78]. Wilhelmine [Tod92]. Will [Fee00, Lat86, Wil88, Bar20, MD20]. William [Bru77, Alb95, Dea95, Mia99]. Williams [Hyy09]. Wimereux [dB09]. Winch [PCL98, Lyn94a]. wind [Hin14, SPBG24]. Windscale [O’R83]. Wine [Geu03, Ano16a, Phi16, Sha16]. Winner [WC99]. Wintry [Gar73a]. WIPP [Ial21]. Wisconsin [Str92]. Wise [Dis04]. wish [CBLG23]. wissenschaftlicher [Hom90]. Wissenschaftslobby [Pfe71]. witche [AABF10]. with/in [Vas15]. Within [Shr07, SH89, CG89, Mir95]. Without [EP07, Mye90, Ano05c, FR91, Jan18, Mia12]. Witness [Gol04]. Witnesses [Dae98]. Witnessing [Col98, Col06, Ham86, Jas98]. Wittgenstein [Blo91c, Hac84]. Wolpert [Ful94a]. woman [Sud22]. Women [Bix97, Bur83, Cow95, Del05, Eps09, GS02, Lag07, Mel09, dC09, FFB11, WC18, ZDI^{+}23, Del02, Del05]. Wood [Ash93, Del93]. Woodpecker [Hun24, Hol22]. Woolgar [Bar81, Joe99b]. Word [CL89, Kis13, Whi89]. Words [Gie82a, CE16]. Wordsworth [Pri71]. Work [Bix97, Bur08, BS98, CKM04, Eps09, FFB11, GP09, HM03, Hes00, Hin06, Lyn82, MKC97, Mie98, Pal94, PT72, PWD07, Rot05, Sim05, Sta83, Sta85, Tor83, WB96, Yea90, Bed14, Bri20, CGPP24, Edm11, FC12, GK16a, God15, HJ13, JS18, KH19, Lah21, LSS20, Law13, Mye08, Püt21, RG16, SVS21, Sok23, UB25, Zie19]. Workers [Bay83, Bix97, Mor94, Eps09]. Working [AB91, Ala08, Bor17, Hoe14, Rot19, BB17]. Workplace [Hec96]. Works [Lat08, STGG95]. Workshop [Mac78a, Rif72, Tak24, Meh80]. Workshops [CKM04]. worksite [RIA^{+}11]. World [Ano77e, Ano77h, Ano06c, Ans06, BS04a, Clo01, Fis10, FNC77, Gor97, Gus06, Hug06, Kre71, Mac79, PW95, Ran17, Ros10, SG73, Sco91, Shi03, Vin95, Ano16a, Cho13, Col13, Kir10, KVA21, Leh18, LTC18, Par22, Phi16, Pic12, Sha16, Bea77, BL12, Bud78, Col07, FS93, Hig21, Hyy09, She06, Whi75b]. Worlds [CT89, Gar97, Mac03c, Mar08a, Mel07, BA12, DS20, DK12]. Worms [Tra81]. Worn [Gie82a]. Worn-Out [Gie82a]. worth [CE16]. would [Mia12, VI17]. wouldn’t [BL13]. Wrath [Mia03]. Writers [Har01]. Writing [AC22, Bri14, Col07, God98, Hal88, Lyn04d, Mye95, Woo76]. Writings [PN04]. Written [Gor90]. Wrong [Ash96, Sah25, WL13]. Wyer [Del02]. Wynne [Bar90, O’R83, Rad98b].

X [Alb95, Bar79, Blo74, Blo77, Bro95, Cri10, Del05, Eps09, Gin10, Hol81b, Luk75, Mos79, O’R83, Gol04, Hes00, Law73, Mia99, Sto08]. X-Ray [Hes00, Law73]. X-rays [Gol04]. xi [Kim94, Mos79]. xii [Hol81b]. xiii [Del05, Doi11, Hol81b]. xiv [Del05, Hol81b]. xv [Blo77, Hol81b]. xvi [Alb95, Bru77, Del05, Wes88]. xvii [Bru77]. xx [Nev80]. xxii [Blo77]. xxv [Mos79].

Yagi [dSP75]. Yale [Rob99]. Year [CFS01, Eas73, Zin71, WB78]. Yearbook [Rud82]. Years [Ber72, Cou99, Edg03, Fri84, GT72, Dea12, Ler18, Lyn11d, RCB18, Sis12b]. Yellow [Ste78a]. Yen [dSP75]. York [AB80, Blo77, Bro95, Col07, Del02, GM78, Hol81b, Ken17, Mir07, Mos79, Rip91, Wes88, Wri80]. Young [FS01, Wra03, AW13, Boh91, Chr88]. Youth [Nic97, Day18]. You’ve [Col04b]. Yoxen [AA84]. Yucca [Mac03a].

Zandvoort [Nad84]. Zapatista [Bra14]. Zealand [BD11, CB97]. Zero [Koe99, Mac99c, SP24]. Zero-Sum [Koe99, Mac99c]. Zhores [Hol81b]. Zika [KKLL21, Oik17]. Zilsel [KR00]. Zimbabwe [dLM00]. Zionist [Ala09b]. Zipper [Sto82]. Zombie [Ran17]. zone [Hof15, Hsu13]. Zones [Gor02, HTB20]. Zoology [SG89]. Zoonati [Hor25b]. Zuckerman [Oro77b, Spi97]. Zur [Hom90].

References

Abir-Am:1982:DPP

[AA82] Pnina Abir-Am. The discourse of physical power and biological knowledge in the 1930s: a reappraisal of the Rockefeller Foundation’s ‘policy’ in molecular biology. Social Studies of Science, 12(3):341–382, August 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631282012003001.

Abir-Am:1984:BDS

[AA84] Pnina Abir-Am. Beyond deterministic sociology and apologetic history: Reassessing the impact of research policy upon new scientific disciplines (reply to Fuerst, Bartels, Olby and Yoxen). Social Studies of Science, 14(2):252–263, May 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/030631284014002007. See [Fue84, Bar84b, Olb84].

Arribas-Ayllon:2010:CAW

[AABF10] Michael Arribas-Ayllon, Andrew Bartlett, and Katie Featherstone. Complexity and accountability: The witches’ brew of psychiatric genetics. Social Studies of Science, 40(4):499–524, August 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312710363511.

Aarden:2023:IES

[Aar23] Erik Aarden. Infrastructuring European scientific integration: Heterogeneous meanings of the European biobanking infrastructure BBMRI-ERIC. Social Studies of Science, 53(4): 572–598, August 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127231162629.

Anderson:1980:ERS

[AB80] Mary B. Anderson and Peter Buck. Essay review: Scientific development: The development of science, science and development, and the science of development: Kurt Mendelssohn, Science and Western Domination (London: Thames and Hud son, 1976), 224 pp., £4.95. Michael J. Moravcsik, Science Development: The Building of Science in Less Developed Countries (Bloomington, Indiana: International Development Research Center, 1975), 291 pp., n.p. Francisco Sagasti, Science and Technology for Development: Main Comparative Report of the Science and Technology Policy Instruments Project (Ottawa: International Development Research Centre, 1978), 111 pp., n.p. ISBN 0-88936-174-6. Nancy Stepan, Beginnings of Brazilian Science: Oswaldo Cruz, Medical Research and Policy, 1890–1920 (New York: Science History Publications, 1976), 225 pp., £9.75. ISBN 0-88202-032-3. Nathan Sivin (ed.), Science and Technology in East Asia (New York: Science History Public. Social Studies of Science, 10(2):215–230, May 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278001000205.

Adams:1991:SVD

[AB91] Elizabeth R. Adams and G. W. Burnett. Scientific vocabulary divergence among female primatologists working in East Africa. Social Studies of Science, 21(3):547–560, August 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631291021003006.

Allwood:1995:SSP

[AB95] Carl Martin Allwood and Jan Bärmark. Situating the social psychology of science. Social Studies of Science, 25(3):600–608, August 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631295025003009.

Amelang:2019:FAH

[AB19] Katrin Amelang and Susanne Bauer. Following the algorithm: How epidemiological risk-scores do accountability. Social Studies of Science, 49(4):476–502, August 1, 2019. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312719862049.

Amsterdamska:2008:MSL

[ABB^{+}08] Olga Amsterdamska, Christian Bonah, Cornelius Borck, Johannes Fehr, Michael Hagner, Marcus Klingberg, Ilana Löwy, Martina Schlünder, Florian Schmaltz, Thomas Schnelle, Antke Tammen, Paul Weindling, and Claus Zittel. Medical science in the light of a flawed study of the Holocaust: a comment on Eva Hedfors’ paper on Ludwik Fleck. Social Studies of Science, 38(6):937–944, December 1, 2008. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/0306312708098609. See [Hed08b] and reply [Hed08a].

Abeysinghe:2013:WSD

[Abe13] Sudeepa Abeysinghe. When the spread of disease becomes a global event: The classification of pandemics. Social Studies of Science, 43(6):905–926, December 1, 2013. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713492559.

Abrams:1991:PAP

[Abr91] Peter A. Abrams. The predictive ability of peer review of Grant proposals: The case of ecology and the US national science foundation. Social Studies of Science, 21(1):111–132, February 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631291021001006.

Abraham:1993:SSI

[Abr93] John Abraham. Scientific standards and institutional interests: Carcinogenic risk assessment of benoxaprofen in the UK and US. Social Studies of Science, 23(3):387–444, August 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312793023003001.

Abraham:1994:IPS

[Abr94] John Abraham. Interests, presuppositions and the science policy construction debate. Social Studies of Science, 24(1): 123–132, February 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631279402400108.

Arvanitis:1988:NSS

[AC88] Rigas Arvanitis and Yvon Chatelin. National scientific strategies in tropical soil sciences. Social Studies of Science, 18(1):

113–146, February 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631288018001005.

Asdal:2022:WGE

[AC22] Kristin Asdal and Béatrice Cointe. Writing good economics: How texts ‘on the move’ perform the lab and discipline of experimental economics. Social Studies of Science, 52(3):376–398, June 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/03063127221079600.

Abraham:2009:DEP

[AD09] John Abraham and Courtney Davis. Drug evaluation and the permissive principle: Continuities and contradictions between standards and practices in antidepressant regulation. Social Studies of Science, 39(4):569–598, August 1, 2009. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312709103480.

Adams:2002:RCC

[Ada02] Vincanne Adams. Randomized controlled crime: Postcolonial sciences in alternative medicine research. Social Studies of Science, 32(5–6):659–690, December 1, 2002. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631270203200503.

Allison:1976:LLP

[AdSPG^{+}76] Paul D. Allison, Derek de Solla Price, Belver C. Griffith, Michael J. Moravcsik, and John A. Stewart. Lotka’s Law: a problem in its interpretation and application. Social Studies of Science, 6(2):269–276, May 1, 1976. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277600600205.

Agar:1998:STP

[Aga98] Jon Agar. SSK in theory and practice. Social Studies of Science, 28(4):649–655, August 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631298028004008.

Agar:2006:WDD

[Aga06] Jon Agar. What difference did computers make? Social Studies of Science, 36(6):869–907, December 1, 2006. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706073450.

Atkinson-Graham:2015:CCB

[AGKL^{+}15] Melissa Atkinson-Graham, Martha Kenney, Kelly Ladd, Cameron Michael Murray, and Emily Astra-Jean Simmonds. Care in context: Becoming an STS researcher. Social Studies of Science, 45(5):738–748, October 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312715600277.

Agre:2001:LRF

[Agr01] Philip E. Agre. Legitimacy and reason in the Florida election controversy. Social Studies of Science, 31(3):419–422, June 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631201031003005.

Almklov:2011:BBD

[AH11] Petter G. Almklov and Vidar Hepsø. Between and beyond data: How analogue field experience informs the interpretation of remote data sources in petroleum reservoir geology. Social Studies of Science, 41(4):539–561, August 1, 2011. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312711403825.

Asdal:2020:MIT

[AH20] Kristin Asdal and Bård Hobæk. The modified issue: Turning around parliaments, politics as usual and how to extend issue-politics with a little help from Max Weber. Social Studies of Science, 50(2):252–270, April 1, 2020. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720902847.

Akera:2007:CRE

[Ake07] Atsushi Akera. Constructing a representation for an ecology of knowledge: Methodological advances in the integration of knowledge and its various contexts. Social Studies of Science, 37(3):413–441, June 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706070742.

Akrich:2023:ANT

[Akr23] Madeleine Akrich. Actor network theory, Bruno Latour, and the CSI. Social Studies of Science, 53(2):169–173, April 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231158102.

Alac:2008:WBS

[Ala08] Morana Alac. Working with brain scans: Digital images and gestural interaction in fMRI Laboratory. Social Studies of Science, 38(4):483–508, August 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312708089715.

Alac:2009:MAS

[Ala09a] Morana Alac. Moving android: On social robots and body-in-interaction. Social Studies of Science, 39(4):491–528, August 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312709103476.

Alatout:2009:BAE

[Ala09b] Samer Alatout. Bringing abundance into environmental politics: Constructing a Zionist network of water abundance, immigration, and colonization. Social Studies of Science, 39(3):363–394, June 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312708101979.

Alac:2020:BIOa

[Ala20a] Morana Alac. Beyond intersubjectivity in olfactory psychophysics I: Troubles with the subject. Social Studies of Science, 50(3):440–473, June 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312720915645.

Alac:2020:BIOb

[Ala20b] Morana Alac. Beyond intersubjectivity in olfactory psychophysics II: Troubles with the object. Social Studies of Science, 50(3):474–502, June 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312720915646.

Alac:2023:BEC

[Ala23] Morana Alac. On body-environment continuities from a laboratory commensalism. Social Studies of Science, 53(2):242–270, April 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
abs/10.1177/03063127221136556.

Albury:1980:ERP

[Alb80] W. R. Albury. Essay review: Politics and rhetoric in the sociobiology debate. Social Studies of Science, 10(4):519–536, November 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631278001000409.

Albury:1995:RCR

[Alb95] W. Randall Albury. Review: Constructionist’s regress: William F. Bynum, Science and the Practice of Medicine in the Nineteenth Century (Cambridge: Cambridge University Press, 1994), xvi + 283 pp., £40.00 / $54.95, £12.95 / $15.95 pbk. ISBN 0-521-25109-5 (0-521-27205-X pbk). Social Studies of Science, 25(2):393–396, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631295025002021.

Alder:1998:MTS

[Ald98] Ken Alder. Making things the same: Representation, tolerance and the end of the Ancien Regime in France. Social Studies of Science, 28(4):499–545, August 1, 1998. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631298028004001.

Alexander:2008:EBT

[Ale08] Jennifer Alexander. Efficiencies of balance: Technical efficiency, popular efficiency, and arbitrary standards in the Late Progressive Era USA. Social Studies of Science, 38(3):323–349, June 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312707083531.

Allen:1976:ERG

[All76] Garland Allen. Essay review: Genetics, eugenics and society: Internalists and externalists in contemporary history of science. Social Studies of Science, 6(1):105–122, February 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277600600107.

Allison:1980:ISP

[All80] Paul D. Allison. Inequality and scientific productivity. Social Studies of Science, 10(2):163–179, May 1, 1980. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278001000203.

Allen:1997:RST

[All97] Bryce Allen. Referring to schools of thought: an example of symbolic citations. Social Studies of Science, 27(6):937–949, December 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631297027006004.

Allen:2007:EJE

[All07] Barbara L. Allen. Environmental justice and expert knowledge in the wake of a disaster. Social Studies of Science, 37 (1):103–110, February 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706069431.

Almklov:2008:SDS

[Alm08] Petter G. Almklov. Standardized data and singular situations. Social Studies of Science, 38(6):873–897, December 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312708098606.

Amick:1974:ISE

[Ami74] Daniel J. Amick. An index of scientific elitism and the scientist’s mission. Social Studies of Science, 4(1):1–16, January 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277400400101.

Amsterdamska:1987:MBC

[Ams87] Olga Amsterdamska. Medical and biological constraints: Early research on variation in bacteriology. Social Studies of Science, 17(4):657–687, November 1, 1987. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631287017004004.

Alac:2011:WRS

[AMT11] Morana Alac, Javier Movellan, and Fumihide Tanaka. When a robot is social: Spatial arrangements and multimodal semiotic engagement in the practice of social robotics. Social Studies of Science, 41(6):893–926, December 1, 2011. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312711420565.

Andersen:1988:TTC

[And88] Håkon With Andersen. Technological trajectories, cultural values and the labour process: The development of NC machinery in the Norwegian shipbuilding industry. Social Studies of Science, 18(3):465–482, August 1, 1988. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631288018003003.

Anderson:1992:RSP

[And92] Warwick Anderson. The reasoning of the strongest: The polemics of skill and science in medical diagnosis. Social Studies of Science, 22(4):653–684, November 1, 1992. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631292022004005.

Anderson:2002:IPT

[And02] Warwick Anderson. Introduction: Postcolonial technoscience. Social Studies of Science, 32(5–6):643–658, December 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631270203200502.

Anderson:2008:TRM

[And08] Warwick Anderson. Teaching ‘race’ at medical school: Social scientists on the margin. Social Studies of Science, 38 (5):785–800, October 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312708090798.

Anderson:2013:OD

[And13] Warwick Anderson. Objectivity and its discontents. Social Studies of Science, 43(4):557–576, August 1, 2013. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712455732.

Anderson:2021:MCH

[And21] Warwick Anderson. The model crisis, or how to have critical promiscuity in the time of Covid-19. Social Studies of Science, 51(2):167–188, April 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312721996053.

Anfinsen:2021:BSC

[Anf21] Martin Anfinsen. Between stability and change: Tensions in the Norwegian electric mobility transition. Social Studies of Science, 51(6):895–913, December 1, 2021. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/03063127211022842.

Anonymous:1971:E

[Ano71a] Anonymous. Editorial. Social Studies of Science, 1(1):1–2, January 1, 1971. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277100100101.

Anonymous:1971:PRS

[Ano71b] Anonymous. A programme of research into social and technological forecasting. Social Studies of Science, 1(2):243–244, April 1, 1971. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277100100210.

Anonymous:1976:EA

[Ano76] Anonymous. Editorial announcement. Social Studies of Science, 6(3–4):279, September 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277600600301.

Anonymous:1977:Aa

[Ano77a] Anonymous. Abstracts. Social Studies of Science, 7(2):134–137, May 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277700700201.

Anonymous:1977:Ab

[Ano77b] Anonymous. Abstracts. Social Studies of Science, 7(3):268–269, August 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631277700700301.

Anonymous:1977:AMP

[Ano77c] Anonymous. Allan Mazur, ‘public confidence in science’. Social Studies of Science, 7(1):iii, February 1, 1977. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700106.

Anonymous:1977:BRC

[Ano77d] Anonymous. Book review: Charles E. Rosenberg, ‘Rationalization and Reality in the Shaping of American Agricultural Research, 1875–1914’. Social Studies of Science, 7(4):398, November 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277700700401.

Anonymous:1977:BRHb

[Ano77e] Anonymous. Book review: E. H. Beardsley, ‘Secrets Between Friends: Applied Science Exchange Between the Western Allies and the Soviet Union During World War II’. Social Studies of Science, 7(4):399, November 1, 1977. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700403.

Anonymous:1977:BRHc

[Ano77f] Anonymous. Book review: Herbert Inhaber, ‘Scientists and Economic Growth’. Social Studies of Science, 7(4):400, November 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277700700407.

Anonymous:1977:BRHa

[Ano77g] Anonymous. Book review: Hugo Meynell, ‘On the Limits of the Sociology of Knowledge’. Social Studies of Science, 7(4): 399–400, November 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277700700405.

Anonymous:1977:BRJb

[Ano77h] Anonymous. Book review: J. Davidson Frame, Francis Narin and Mark P. Carpenter, ‘The Distribution of World Science’. Social Studies of Science, 7(4):400, November 1, 1977. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700406.

Anonymous:1977:BRJa

[Ano77i] Anonymous. Book review: John C. Marshall, ‘Minds, Machines and Metaphors’. Social Studies of Science, 7(4):399, November 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277700700404.

Anonymous:1977:BRM

[Ano77j] Anonymous. Book review: Mel Gorman, ‘Electric Illumination in the Franco–Prussian War’. Social Studies of Science, 7(4):400, November 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277700700408.

Anonymous:1977:BRR

[Ano77k] Anonymous. Book review: Russell Moseley, ‘Tadpoles and Frogs: Some Aspects of the Professionalization of British Physics, 1870–1939’. Social Studies of Science, 7(4):398, November 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631277700700402.

Anonymous:1977:DNS

[Ano77l] Anonymous. Dorothy Nelkin, ‘Scientists and Professional Responsibility: The Experience of American Ecologists’. Social Studies of Science, 7(1):ii, February 1, 1977. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700103.

Anonymous:1977:GNG

[Ano77m] Anonymous. G. Nigel Gilbert, ‘referencing as persuasion’. Social Studies of Science, 7(1):iii, February 1, 1977. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700105.

Anonymous:1977:ISR

[Ano77n] Anonymous. Ina Spiegel-Rosing, Science Studies: Bibliometric and Content Analysis. Social Studies of Science, 7(1):ii, February 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277700700104.

Anonymous:1977:JHR

[Ano77o] Anonymous. Jonathan Harwood, ‘The Race-Intelligence Controversy: a Sociological Approach: II — “External” Factors’. Social Studies of Science, 7(1):i, February 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700101.

Anonymous:1977:MOC

[Ano77p] Anonymous. Mark Oromaner, ‘The Career of Sociological Literature: a Diachronous Study’. Social Studies of Science, 7 (1):iii, February 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277700700107.

Anonymous:1977:SSB

[Ano77q] Anonymous. Steven Shapin and Barry Barnes, ‘Science, Nature and Control: Interpreting Mechanics’ Institutes’. Social Studies of Science, 7(1):i, February 1, 1977. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700102.

Anonymous:1979:SAS

[Ano79] Anonymous. The social assessment of science (Bielefeld, 26– 28 May 1978): Two views. Social Studies of Science, 9(1): 125–133, February 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277900900108.

Anonymous:1983:C

[Ano83a] Anonymous. Correction. Social Studies of Science, 13(1):i, February 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631283013001001.

Anonymous:1983:DJS

[Ano83b] Anonymous. Derek John de Solla Price (1922–83). Social Studies of Science, 13(4):ii, November 1, 1983. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631283013004002.

Anonymous:1984:BP

[Ano84] Anonymous. On behalf of the personane. Social Studies of Science, 14(2):283, May 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631284014002009.

Anonymous:1985:C

[Ano85] Anonymous. Corrigendum. Social Studies of Science, 15(1): 201–203, February 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631285015001011.

Anonymous:1986:Cb

[Ano86a] Anonymous. Corrigenda. Social Studies of Science, 16(3):568, August 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631286016003012.

Anonymous:1986:Ca

[Ano86b] Anonymous. Corrigendum. Social Studies of Science, 16(2):384, May 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312786016002012.

Anonymous:1986:E

[Ano86c] Anonymous. Editorial. Social Studies of Science, 16(4):571– 572, November 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631286016004001.

Anonymous:1988:BNa

[Ano88a] Anonymous. Book notes. Social Studies of Science, 18(3): 565–571, August 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631288018003012.

Anonymous:1988:BNb

[Ano88b] Anonymous. Book notes. Social Studies of Science, 18(4): 755–763, November 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631288018004008.

Anonymous:1989:BN

[Ano89] Anonymous. Book notes. Social Studies of Science, 19(2):381– 384, May 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631289019002013.

Anonymous:1990:BNa

[Ano90a] Anonymous. Book notes. Social Studies of Science, 20(1): 184–191, February 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631290020001011.

Anonymous:1990:BNb

[Ano90b] Anonymous. Book notes. Social Studies of Science, 20(3): 565–576, August 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631290020003009.

Anonymous:1990:C

[Ano90c] Anonymous. Corrigenda. Social Studies of Science, 20(1):192, February 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631290020001012.

Anonymous:1991:BNa

[Ano91a] Anonymous. Book notes. Social Studies of Science, 21(1): 191–192, February 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631291021001017.

Anonymous:1991:BNb

[Ano91b] Anonymous. Book notes. Social Studies of Science, 21(2):377– 384, May 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631291021002010.

Anonymous:1991:BNc

[Ano91c] Anonymous. Book notes. Social Studies of Science, 21(3): 585–593, August 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631291021003012.

Anonymous:1991:E

[Ano91d] Anonymous. Erratum. Social Studies of Science, 21(1):6, February 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631291021001002.

Anonymous:1992:BN

[Ano92] Anonymous. Book notes. Social Studies of Science, 22(2):429– 432, May 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631292022002016.

Anonymous:1996:E

[Ano96] Anonymous. Erratum. Social Studies of Science, 26(3):574, August 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631296026003003.

Anonymous:1999:Ea

[Ano99a] Anonymous. Erratum. Social Studies of Science, 29(2):317–318, April 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631299029002017.

Anonymous:1999:Eb

[Ano99b] Anonymous. Erratum. Social Studies of Science, 29(4):i, August 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631299029004001.

Anonymous:2002:IV

[Ano02] Anonymous. Index to volume 32, 2002. Social Studies of Science, 32(5–6):939–942, December 1, 2002. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631270203200510.

Anonymous:2004:IV

[Ano04] Anonymous. Index to volume 34, 2004. Social Studies of Science, 34(6):957–959, December 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704050652.

Anonymous:2005:Eb

[Ano05a] Anonymous. Editoral. Social Studies of Science, 35(6):835– 836, December 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312705060704.

Anonymous:2005:Ea

[Ano05b] Anonymous. Editorial. Social Studies of Science, 35(3):331, June 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312705055560.

Anonymous:2005:ESS

[Ano05c] Anonymous. Editorial — science without scientists? Social Studies of Science, 35(4):499–501, August 1, 2005. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312705057049.

Anonymous:2005:IV

[Ano05d] Anonymous. Index to volume 35, 2005. Social Studies of Science, 35(6):959–960, December 1, 2005. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312705060734.

Anonymous:2006:E

[Ano06a] Anonymous. Editorial. Social Studies of Science, 36(2):171–172, April 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312706064669.

Anonymous:2006:IV

[Ano06b] Anonymous. Index to volume 36, 2006. Social Studies of Science, 36(6):959–960, December 1, 2006. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706073447.

Anonymous:2006:RSL

[Ano06c] Anonymous. Review symposium on Lynn Eden, whole world on fire. Social Studies of Science, 36(4):627, August 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706068476. See [Ans06, Hug06].

Anonymous:2007:Ea

[Ano07a] Anonymous. Editorial. Social Studies of Science, 37(1):5–6, February 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/0306312706075791.

Anonymous:2007:Eb

[Ano07b] Anonymous. Editorial. Social Studies of Science, 37(2):171–172, April 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312707077985.

Anonymous:2007:NBP

[Ano07c] Anonymous. New books from The MIT Press. Social Studies of Science, 37(6):967–968, December 1, 2007. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127070370060602.

Anonymous:2008:E

[Ano08] Anonymous. Editorial. Social Studies of Science, 38(1):5–6, February 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/0306312706088562.

Anonymous:2009:EJU

[Ano09a] Anonymous. Editorial: Journals under threat: a joint response from history of science, technology and medicine editors. Social Studies of Science, 39(1):6–9, February 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/03063127090390010702. Declaration by the editors of 27 history journals that their journals must be excluded from the proposed European Reference Index for the Humanities (ERIH), an initiative originating with the European Science Foundation.

Anonymous:2009:Ea

[Ano09b] Anonymous. Erratum. Social Studies of Science, 39(1):163, February 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/0306312708101566.

Anonymous:2009:Eb

[Ano09c] Anonymous. Erratum. Social Studies of Science, 39(3):485, June 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312709336791.

Anonymous:2009:Ec

[Ano09d] Anonymous. Erratum. Social Studies of Science, 39(4):643, August 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312709340455.

Anonymous:2010:ENT

[Ano10] Anonymous. Editor’s note: Two papers on Latour and mediation. Social Studies of Science, 40(4):577, August 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312710375492.

Anonymous:2012:C

[Ano12a] Anonymous. Corrigendum. Social Studies of Science, 42(5): 793–794, October 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712460483.

Anonymous:2012:Eb

[Ano12b] Anonymous. Editorial. Social Studies of Science, 42(3):323, June 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312712447267.

Anonymous:2012:Ea

[Ano12c] Anonymous. Erratum. Social Studies of Science, 42(1):3, February 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312711433378.

Anonymous:2016:CTS

[Ano16a] Anonymous. Corrigendum to a taste of science: Making the subjective objective in the California wine world. Social Studies of Science, 46(4):646, August 1, 2016. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716663859.

Anonymous:2016:SEA

[Ano16b] Anonymous. Science, ethnography, art. Social Studies of Science, 46(6):961–981, December 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312716678490.

Anspach:2006:RWW

[Ans06] Renee Anspach. Reflections on Whole World on Fire. Social Studies of Science, 36(4):628–633, August 1, 2006. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706064817. See [Ano06c, Hug06].

Ashton:1978:MPN

[AO78] S. V. Ashton and C. Oppenheim. A method of predicting Nobel Prizewinners in Chemistry. Social Studies of Science, 8 (3):341–348, August 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277800800306.

Ahmed:2015:RGH

[AQK15] Hassaan Ahmed, Omer Masood Qureshi, and Abid Ali Khan. Reviving a ghost in the history of technology: The social construction of the recumbent bicycle. Social Studies of Science, 45 (1):130–136, February 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312714560640.

Allen:1994:PCL

[AQL94] Bryce Allen, Jian Qin, and F. W. Lancaster. Persuasive communities: a longitudinal analysis of references in the Philosophical Transactions of the Royal Society, 1665–1990. Social Studies of Science, 24(2):279–310, May 1, 1994. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631279402400204.

Abraham:2002:PIR

[AR02] John Abraham and Tim Reed. Progress, innovation and regulatory science in drug development: The politics of international standard-setting. Social Studies of Science, 32(3):337–369, June 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312702032003001.

Armstrong:2009:OPH

[Arm09] David Armstrong. Origins of the problem of health-related behaviours: a genealogical study. Social Studies of Science, 39 (6):909–926, December 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312709104258.

Armstrong:2010:FHR

[Arm10] David Armstrong. Is the focus on health-related behaviours a new phenomenon? Reply to Vallgårda (2010). Social Studies of Science, 40(4):615–617, August 1, 2010. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312710367843. See [Val10].

Armstrong:2017:CPI

[Arm17] David Armstrong. Clinical prediction and the idea of a population. Social Studies of Science, 47(2):288–299, April 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716685926.

Armstrong:2019:SLD

[Arm19] David Armstrong. The social life of data points: Antecedents of digital technologies. Social Studies of Science, 49(1):102– 117, February 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312718821726.

Armstrong:2022:RFS

[Arm22] David Armstrong. Rise and fall of the (social) group. Social Studies of Science, 52(4):618–634, August 1, 2022. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127221096389.

Abraham:1999:CCS

[AS99] John Abraham and Julie Sheppard. Complacent and conflicting scientific expertise in British and American drug regulation: Clinical risk assessment of Triazolam. Social Studies of Science, 29(6):803–843, December 1, 1999. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029006001.

Alcayna-Stevens:2016:HFS

[AS16] Lys Alcayna-Stevens. Habituating field scientists. Social Studies of Science, 46(6):833–853, December 1, 2016. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716669251.

Asdal:2008:SPL

[Asd08] Kristin Asdal. Subjected to Parliament: The laboratory of experimental medicine and the animal body. Social Studies of Science, 38(6):899–917, December 1, 2008. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706098607.

Ashmore:1993:TBS

[Ash93] Malcolm Ashmore. The theatre of the blind: Starring a Promethean prankster, a phoney phenomenon, a prism, a pocket, and a piece of wood. Social Studies of Science, 23 (1):67–106, February 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631293023001003.

Ashmore:1996:EWS

[Ash96] Malcolm Ashmore. Ending up on the wrong side: Must the two forms of radicalism always be at war? Social Studies of Science, 26(2):305–322, May 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631296026002005.

Ashmore:2005:BRL

[Ash05] Malcolm Ashmore. Book review: The Life Inside/ The Left-Hand Side. Social Studies of Science, 35(5):827–830, October 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312705056053.

Ashmore:2007:RMO

[Ash07] Malcolm Ashmore. Review: Magnum opus: When Harry met Joe, and what follows: Harry Collins, Gravity’s Shadow: The Search for Gravitational Waves (University of Chicago Press, Chicago & London, 2004), 864 pp., £63.00 / $100.00 / EUR88.50 (hbk), £25.00 / $39.00 / EUR 38.80 (pbk). ISBN 0-226-11377-9 (hbk), 0-226-11378-7 (pbk). Social Studies of Science, 37(1):159–166, February 1, 2007. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706069951.

Aspers:2015:PO

[Asp15] Patrik Aspers. Performing ontology. Social Studies of Science, 45(3):449–453, June 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312714548610.

Aam:2021:ISS

[ÅSS21] Heidrun Åm, Gisle Solbu, and Knut H. Sørensen. The imagined scientist of science governance. Social Studies of Science, 51(2):277–297, April 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312720962573.

Atkinson:1987:TSA

[Atk87] Paul Atkinson. Teaching students about science. Social Studies of Science, 17(1):189–192, February 1, 1987. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631287017001012.

Atkinson:2005:BRG

[Atk05] Paul Atkinson. Book review: Gothic imaginations. Social Studies of Science, 35(4):653–664, August 1, 2005. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312705054592.

Audetat:2001:RTS

[Aud01] Marc Audétat. Re-thinking science, re-thinking society. Social Studies of Science, 31(6):950–956, December 1, 2001. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031006006.

Aviles:2015:LDR

[Avi15] Natalie B. Aviles. The little death: Rigoni-Stern and the problem of sex and cancer in 20th-century biomedical research. Social Studies of Science, 45(3):394–415, June 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312715584402.

Arbesman:2013:DFY

[AW13] Samuel Arbesman and K. Brad Wray. Demographics and the fate of the young scientist. Social Studies of Science, 43 (2):282–286, April 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712470218.

Andersen:2019:DER

[AW19] Line Edslev Andersen and K. Brad Wray. Detecting errors that result in retractions. Social Studies of Science, 49(6): 942–954, December 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312719872008.

Agalianos:2006:SSL

[AWN06] Angelos Agalianos, Geoff Whitty, and Richard Noss. The social shaping of Logo. Social Studies of Science, 36(2):241–267, April 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312706053809.

Barnes:2012:WWI

[BA12] Jessica Barnes and Samer Alatout. Water worlds: Introduction to the special issue of social studies of science. Social Studies of Science, 42(4):483–488, August 1, 2012. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712448524.

Baker:2011:TTE

[Bak11] Jonathan D. Baker. Tradition and toxicity: Evidential cultures in the kava safety debate. Social Studies of Science, 41 (3):361–384, June 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312710395341.

Bakker:2012:WPB

[Bak12] Karen Bakker. Water: Political, biopolitical, material. Social Studies of Science, 42(4):616–623, August 1, 2012. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712441396.

Baker:2017:CSS

[Bak17] Zeke Baker. Climate state: Science-state struggles and the formation of climate science in the US from the 1930s to 1960s. Social Studies of Science, 47(6):861–887, December 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312717725205.

Balazs:1995:IPE

[Bal95] Katalin Balázs. Innovation potential embodied in research organizations in Central and Eastern Europe. Social Studies of Science, 25(4):655–683, November 1, 1995. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631295025004004.

Balmer:1996:MMH

[Bal96] Brian Balmer. Managing mapping in the human genome project. Social Studies of Science, 26(3):531–573, August 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631296026003002.

Balmer:2006:SFR

[Bal06] Brian Balmer. A secret formula, a rogue patent and public knowledge about nerve gas: Secrecy as a spatial-epistemic tool. Social Studies of Science, 36(5):691–722, October 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706063786.

Balka:2010:OSL

[Bal10] Ellen Balka. Obituary. Susan Leigh Star (1954–2010). Social Studies of Science, 40(4):647–651, August 1, 2010. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312710376010.

Barnes:1971:MIR

[Bar71] S. B. Barnes. Making out in industrial research. Social Studies of Science, 1(2):157–175, April 1, 1971. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277100100203.

Barnes:1979:ERV

[Bar79] Barry Barnes. Essay review: Vicissitudes of belief: Larry Laudan, Progress and its Problems: Towards a Theory of Scientific Growth (London: Routledge and Kegan Paul, 1977), x + 257 pp., £5.95. ISBN D-7100-8749-7. Social Studies of Science, 9(2):247–263, May 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277900900205.

Barnes:1981:HWC

[Bar81] Barry Barnes. On the ‘hows’ and ‘whys’ of cultural change (response to Woolgar). Social Studies of Science, 11(4):481– 498, November 1, 1981. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631278101100404.

Barnes:1982:STR

[Bar82] Barry Barnes. The science-technology relationship: a model and a query. Social Studies of Science, 12(1):166–172, February 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631282012001013.

Barnes:1984:BRT

[Bar84a] Barry Barnes. Book review: Thomas P. Hughes, Networks of Power: Electrification in Western Society. Social Studies of Science, 14(2):309–314, May 1, 1984. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014002012.

Bartels:1984:RFF

[Bar84b] Ditta Bartels. The Rockefeller Foundation’s funding policy for molecular biology: Success or failure? Social Studies of Science, 14(2):238–243, May 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631284014002004. See re-

ply [AA84].

Barnes:1985:CA

[Bar85a] Barry Barnes. A case of amnesia? Social Studies of Science, 15 (1):175–176, February 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631285015001007. See re-

ply [RH85].

Barnes:1985:ES

[Bar85b] Barry Barnes. Ethnomethodology as science. Social Studies of Science, 15(4):751–762, November 1, 1985. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631285015004008.

Barnes:1987:PLS

[Bar87] Barry Barnes. Power listens to science. Social Studies of Science, 17(3):555–564, August 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631287017003007.

Barnes:1990:BRR

[Bar90] Barry Barnes. Book review: Roger Smith and Brian Wynne (eds), Expert Evidence: Interpreting Science in the Law. Social Studies of Science, 20(4):757–760, November 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020004009.

Barnes:1998:ODT

[Bar98] Barry Barnes. Oversimplification and the desire for truth: Oversimplification and the desire for truth: Response to Mermin. Social Studies of Science, 28(4):636–640, August 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/030631298028004006. See [Mer98b] and reply [Mer98a].

Barr:2003:MDO

[Bar03a] Stephen Barr. In memoriam: David Owen Edge. Social Studies of Science, 33(2):194–195, April 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127030332013.

Barth:2003:PSN

[Bar03b] Kai-Henrik Barth. The politics of seismology: Nuclear testing, arms control, and the transformation of a discipline. Social Studies of Science, 33(5):743–781, October 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312703335005.

Barad:2011:EEP

[Bar11] Karen Barad. Erasers and erasures: Pinch’s unfortunate ‘uncertainty principle’. Social Studies of Science, 41(3):443–454, June 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312711406317.

Barnes:2012:PPA

[Bar12] Jessica Barnes. Pumping possibility: Agricultural expansion through desert reclamation in Egypt. Social Studies of Science, 42(4):517–538, August 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712438772.

Barry:2020:AWP

[Bar20] Andrew Barry. Afterword: ‘I will play my part!’, or parliamentary politics in action. Social Studies of Science, 50(2):335–342, April 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312720911151.

Barton:2023:BES

[Bar23] Jenifer Barton. Branding the Earth: Selling Earth system science in the United States, 1983–1988. Social Studies of Science, 53(1):49–80, February 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

Bayertz:1983:NST

[Bay83] Kurt Bayertz. Naturwissenschaft und Sozialismus: Tendenzen der Naturwissenschafts-Rezeption in der deutschen Arbeiterbewegung des 19. Jahrhunderts. (German) [Science and socialism: Tendencies of the science reception in the 19th Century German workers movement]. Social Studies of Science, 13 (3):355–394, August 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631283013003002.

Bazerman:1984:MEE

[Baz84] Charles Bazerman. Modern evolution of the experimental report in physics: Spectroscopic articles in Physical Review, 1893– 1980. Social Studies of Science, 14(2):163–196, May 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014002001.

Berner:2017:MFB

[BB17] Boel Berner and Maria Björkman. Modernizing the flow of blood: Biomedical technicians, working knowledge and the transformation of Swedish blood centre practices. Social Studies of Science, 47(4):485–510, August 1, 2017. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312717693464.

Bershady:1976:NK

[BCM76] Harold J. Bershady, Diana Crane, and Norman Miller. Norman Kaplan 1923–1975. Social Studies of Science, 6(2):277–278, May 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277600600206.

Bloomfield:2011:ITI

[BD11] Brian P. Bloomfield and Bill Doolin. Imagination and technoscientific innovations: Governance of transgenic cows in new Zealand. Social Studies of Science, 41(1):59–83, February 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312710385372.

Bloor:2000:R

[BE00] David Bloor and David Edge. For the record. Social Studies of Science, 30(1):158–160, February 1, 2000. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030001009.

Beardsley:1977:SBF

[Bea77] E. H. Beardsley. Secrets between friends: Applied science exchange between the Western Allies and the Soviet Union during World War II. Social Studies of Science, 7(4):447–473, November 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277700700411.

Beaulieu:2001:VBN

[Bea01] Anne Beaulieu. Voxels in the brain: Neuroscience, informatics and changing notions of objectivity. Social Studies of Science, 31(5):635–680, October 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631201031005001.

Beaulieu:2010:RNC

[Bea10] Anne Beaulieu. Research note: From co-location to co-presence: Shifts in the use of ethnography for the study of knowledge. Social Studies of Science, 40(3):453–470, June 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312709359219.

Beder:1991:CCS

[Bed91] Sharon Beder. Controversy and closure: Sydney’s beaches in crisis. Social Studies of Science, 21(2):223–256, May 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631291021002003.

Beddoes:2014:MDB

[Bed14] Kacey Beddoes. Methodology discourses as boundary work in the construction of engineering education. Social Studies of Science, 44(2):293–312, April 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312713510431.

Bell:2020:STP

[Bel20] Kirsten Bell. Signs, things and packaging: Recovering the material agency of the cigarette packet. Social Studies of Science, 50(1):30–49, February 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312719876995.

Benson:2012:OIM

[Ben12] Etienne Benson. One infrastructure, many global visions: The commercialization and diversification of Argos, a satellite-based environmental surveillance system. Social Studies of Science, 42 (6):843–868, December 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712457851.

Berman:1972:EYR

[Ber72] Morris Berman. The early years of the Royal Institution 1799– 1810: a re-evaluation. Social Studies of Science, 2(3):205–240, July 1, 1972. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277200200301.

Berg:1995:TPS

[Ber95] Marc Berg. Turning a practice into a science: Reconceptualizing postwar medical practice. Social Studies of Science, 25 (3):437–476, August 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631295025003002.

Berman:2008:WDU

[Ber08] Elizabeth Popp Berman. Why did universities start patenting? Institution-building and the road to the Bayh–Dole Act. Social Studies of Science, 38(6):835–871, December 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312708098605.

Besky:2020:EIM

[Bes20] Sarah Besky. Empire and indigestion: Materializing tannins in the Indian tea industry. Social Studies of Science, 50 (3):398–417, June 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312720915780.

Balazs:1995:STS

[BFS95a] Katalin Balázs, Wendy Faulkner, and Uwe Schimank. Science and technology studies and policy in Central and Eastern Europe: What next? Social Studies of Science, 25(4): 873–883, November 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631295025004013.

Balazs:1995:TRS

[BFS95b] Katalin Balázs, Wendy Faulkner, and Uwe Schimank. Transformation of the research systems of post-Communist Central and Eastern Europe: an introduction. Social Studies of Science, 25 (4):613–632, November 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631295025004002.

Bonneuil:2014:GFB

[BFW14] Christophe Bonneuil, Jean Foyer, and Brian Wynne. Genetic fallout in bio-cultural landscapes: Molecular imperialism and the cultural politics of (not) seeing transgenes in Mexico. Social Studies of Science, 44(6):901–929, December 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312714548258.

Brante:1991:BHC

[BH91] Thomas Brante and Margareta Hallberg. Brain or heart? The controversy over the concept of death. Social Studies of Science, 21(3):389–413, August 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631291021003001.

Baldi:1997:REP

[BH97] Stéphane Baldi and Lowell L. Hargens. Re-examining Price’s conjectures on the structure of reference networks: Results from the special relativity, spatial diffusion modeling and role analysis literatures. Social Studies of Science, 27(4):669–687, August 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631297027004004.

Bruun:2003:CBI

[BH03] Henrik Bruun and Janne Hukkinen. Crossing boundaries: an integrative framework for studying technological change. Social Studies of Science, 33(1):95–116, February 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312703033001178.

Bhaneja:1976:IST

[Bha76] Balwant Bhaneja. India’s Science and Technology Plan, 1974– 79. Social Studies of Science, 6(1):99–104, February 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277600600106.

Brown:2016:TRP

[BHC16] Patrick Brown, Ferhana Hashem, and Michael Calnan. Trust, regulatory processes and NICE decision-making: Appraising cost-effectiveness models through appraising people and systems. Social Studies of Science, 46(1):87–111, February 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312715609699.

Biagioli:2022:GBI

[Bia22] Mario Biagioli. Ghosts, brands, and influencers: Emergent trends in scientific authorship. Social Studies of Science, 52 (3):463–487, June 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/03063127221095046.

Bier:2018:BCM

[Bie18] Jess Bier. Bodily circulation and the measure of a life: Forensic identification and valuation after the Titanic disaster. Social Studies of Science, 48(5):635–662, October 1, 2018. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312718801173.

Bijsterveld:2001:DSM

[Bij01] Karin Bijsterveld. The diabolical symphony of the Mechanical Age: Technology and symbolism of sound in European and North American noise abatement campaigns, 1900–40. Social Studies of Science, 31(1):37–70, February 1, 2001. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031001003.

Bijker:2007:ADC

[Bij07] Wiebe E. Bijker. American and Dutch coastal engineering: Differences in risk conception and differences in technological culture. Social Studies of Science, 37(1):143–151, February 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706069437.

Bijker:2012:DWL

[Bij12] Wiebe Bijker. Do we live in water cultures? A methodological commentary. Social Studies of Science, 42(4):624–627, August 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312712441690.

Bimber:1990:KMT

[Bim90] Bruce Bimber. Karl Marx and the three faces of technological determinism. Social Studies of Science, 20(2):333–351, May 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631290020002006.

Birch:2023:REI

[Bir23] Kean Birch. Reflexive expectations in innovation financing: an analysis of venture capital as a mode of valuation. Social Studies of Science, 53(1):29–48, February 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

Bix:1997:EVE

[Bix97] Amy Sue Bix. Experiences and voices of eugenics field-workers: ‘women’s work’ in biology. Social Studies of Science, 27(4): 625–668, August 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631297027004003.

Beynon-Jones:2012:TED

[BJ12] Siân M. Beynon-Jones. Timing is everything: The demarcation of ‘later’ abortions in Scotland. Social Studies of Science, 42 (1):53–74, February 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312711426596.

Blacker:2021:WCS

[BKK21] Sarah Blacker, Aya H. Kimura, and Abby Kinchy. When citizen science is public relations. Social Studies of Science, 51 (5):780–796, October 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/03063127211027662.

Bowker:1987:BDS

[BL87] Geof Bowker and Bruno Latour. A booming discipline short of discipline: (social) studies of science in France. Social Studies of Science, 17(4):715–748, November 1, 1987. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631287017004006.

Bjelic:1994:GPR

[BL94] Dusan Bjelic and Michael Lynch. Goethe’s ‘Protestant Reformation’ as a textual demonstration: Comment on Jackson. Social Studies of Science, 24(4):703–724, November 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631279402400404. See [Jac94].

Bonneuil:2012:HDW

[BL12] Christophe Bonneuil and Les Levidow. How does the World Trade Organization know? The mobilization and staging of scientific expertise in the GMO trade dispute. Social Studies of Science, 42(1):75–100, February 1, 2012. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312711430151.

Bevan:2013:WSS

[BL13] Valerie Bevan and Mark Learmonth. ‘I wouldn’t say it’s sexism, except that … it’s all these little subtle things’: Healthcare scientists’ accounts of gender in healthcare science laboratories. Social Studies of Science, 43(1):136–158, February 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712460606.

Blau:1976:SRA

[Bla76] Judith R. Blau. Scientific recognition: Academic context and professional role. Social Studies of Science, 6(3–4):533– 545, September 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277600600311.

Bloor:1971:ERT

[Blo71] David Bloor. Essay review: Two paradigms for scientific knowledge? Social Studies of Science, 1(1):101–115, January 1, 1971. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277100100108.

Bloor:1974:ERP

[Blo74] David Bloor. Essay review: Popper’s mystification of objective knowledge: K. R. Popper, Objective Knowledge: an Evolutionary Approach. Oxford: Clarendon Press, 1972. x + 380. £4.50 hardback, £1.50 paper. Social Studies of Science, 4 (1):65–76, January 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277400400104.

Bloor:1975:PAS

[Blo75] David Bloor. A philosophical approach to science. Social Studies of Science, 5(4):507–517, November 1, 1975. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277500500410.

Bloor:1977:ERI

[Blo77] Celia Bloor. Essay review: Inductivism on the campus: Martin Trow (ed.), Teachers and Students Aspects of American Higber Education (New York: McGraw-Hill Book Co., 1975), xxii + 419 pp. $17.50 (£11.90). ISBN 0-07-010070-5. Everett Carll Ladd, Jr. and Seymour Martin Lipset, The Divided Academy Professors and Politics (New York: McGraw-Hill Book Co., 1975), xv + 407 pp. $17.50 (£11.90). ISBN 0-07-010112-4. Ernest Rudd, The Highest Education: A Study of Graduate Education in Britain (London and Boston, Mass.: Routledge and Kegan Paul, 1975), x + 198 pp. £3.90. ISBN 0-7100-8307-6. Social Studies of Science, 7(1):133–138, February 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700115.

Bloor:1983:CMC

[Blo83] David Bloor. Coleridge’s moral copula. Social Studies of Science, 13(4):605–619, November 1, 1983. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631283013004006.

Bloomfield:1985:ASE

[Blo85] Brian P. Bloomfield. Anomalies and social experience: Backcasting with simulation models. Social Studies of Science, 15 (4):631–675, November 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631285015004003.

Bloor:1987:LFM

[Blo87] David Bloor. The living foundations of mathematics. Social Studies of Science, 17(2):337–358, May 1, 1987. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631287017002009.

Bloomfield:1991:RIS

[Blo91a] Brian P. Bloomfield. The role of information systems in the UK national health service: Action at a distance and the fetish of calculation. Social Studies of Science, 21(4):701–734, November 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631291021004004.

Bloor:1991:RGH

[Blo91b] David Bloor. Review of Galison, how experiments end. Social Studies of Science, 21(1):186–189, February 1, 1991. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631291021001016.

Bloor:1991:WL

[Blo91c] David Bloor. Wittgenstein’s lectures. Social Studies of Science, 21(2):374–376, May 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631291021002009.

Bloor:1992:OHI

[Blo92] David Bloor. Ordinary human inference as material for the sociology of knowledge. Social Studies of Science, 22(1): 129–139, February 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312792022001008.

Bloor:1993:CMS

[Blo93] David Bloor. Cognitive models of science. Social Studies of Science, 23(4):743–757, November 1, 1993. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631293023004005.

Bloor:1996:ISK

[Blo96] David Bloor. Idealism and the sociology of knowledge. Social Studies of Science, 26(4):839–856, November 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631296026004005.

Bloor:1998:CAR

[Blo98a] David Bloor. Changing axes: Response to Mermin. Social Studies of Science, 28(4):624–635, August 1, 1998. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/030631298028004005. See [Mer98b] and reply [Mer98a].

Bloor:1998:CS

[Blo98b] David Bloor. A civil skepticism. Social Studies of Science, 28 (4):655–665, August 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631298028004009.

Bloor:2000:SWM

[Blo00] Michael Bloor. The South Wales Miners Federation, miners’ lung and the instrumental use of expertise, 1900–1950.

Social Studies of Science, 30(1):125–140, February 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030001005.

Bloor:2003:ODO

[Blo03] David Bloor. Obituary: David Owen Edge (4 September 1932– 28 January 2003). Social Studies of Science, 33(2):171–176, April 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/03063127030332003.

Bloor:2004:IRS

[Blo04] David Bloor. Institutions and rule-scepticism: a reply to Martin Kusch. Social Studies of Science, 34(4):593–601, August 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/

10.1177/0306312704044910. See [Kus04b] and reply [Kus04a].

Blume:2009:OOA

[Blu09] Stuart Blume. Obituary: Olga Amsterdamska Moore (1953– 2009). Social Studies of Science, 39(6):957–959, December 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312709350403.

Brosnan:2014:ENP

[BM14] Caragh Brosnan and Mike Michael. Enacting the ‘neuro’ in practice: Translational research, adhesion and the promise of porosity. Social Studies of Science, 44(5):680–700, October 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312714534333.

Bertotti:2019:CCN

[BM19] Andrea M. Bertotti and Skye A. Miner. Constructing contentious and noncontentious facts: How gynecology textbooks create certainty around pharma-contraceptive safety. Social Studies of Science, 49(2):245–263, April 1, 2019. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312719834676.

Bohlin:1989:PSH

[Boh89] Ingemar Bohlin. Philosophy of science and historical enquiry. Social Studies of Science, 19(2):360–367, May 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019002009.

Bohlin:1991:RMY

[Boh91] Ingemar Bohlin. Robert M. Young and Darwin historiography. Social Studies of Science, 21(4):597–648, November 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631291021004001.

Bohlin:1999:MH

[Boh99] Ingemar Bohlin. Making history. Social Studies of Science, 29(3):459–480, June 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631299029003007.

Bohlin:2004:CRC

[Boh04] Ingemar Bohlin. Communication regimes in competition: The current transition in scholarly communication seen through the lens of the sociology of technology. Social Studies of Science, 34(3):365–391, June 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312704041522.

Bolman:2018:HEA

[Bol18] Brad Bolman. How experiments age: Gerontology, beagles, and species projection at Davis. Social Studies of Science, 48 (2):232–258, April 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312718759822.

Borkenhagen:2017:EBC

[Bor17] Chad Borkenhagen. Evidence-based creativity: Working between art and science in the field of fine dining. Social Studies of Science, 47(5):630–654, October 1, 2017. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312717725204.

Botelho:1990:PBS

[Bot90] Antonio José Junqueira Botelho. The professionalization of Brazilian scientists, the Brazilian Society for the Progress of Science (SBPC), and the state, 1948–60. Social Studies of Science, 20(3):473–502, August 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631290020003004.

Bourret:2005:BPC

[Bou05] Pascale Bourret. BRCA patients and clinical collectives: New configurations of action in cancer genetics practices. Social Studies of Science, 35(1):41–68, February 1, 2005. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312705048716.

Bowden:1985:SCV

[Bow85] Gary Bowden. The social construction of validity in estimates of US crude oil reserves. Social Studies of Science, 15(2):207–240, May 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631285015002001.

Bowker:1987:WOR

[Bow87] Geof Bowker. A well ordered reality: Aspects of the development of Schlumberger, 1920–39. Social Studies of Science, 17 (4):611–655, November 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631287017004003.

Bowker:1993:HUS

[Bow93] Geof Bowker. How to be universal: Some cybernetic strategies, 1943–70. Social Studies of Science, 23(1):107–127, February 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631293023001004.

Bowker:1996:HTC

[Bow96] Geoffrey C. Bowker. How things change: The history of sociotechnical structures. Social Studies of Science, 26(1): 173–182, February 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631296026001009.

Bowker:2000:BD

[Bow00] Geoffrey C. Bowker. Biodiversity datadiversity. Social Studies of Science, 30(5):643–683, October 1, 2000. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030005001.

Bowker:2003:MDO

[Bow03] Geoffrey C. Bowker. In memoriam: David Owen Edge. Social Studies of Science, 33(2):189–190, April 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127030332007.

Bickerstaffe:1980:CTC

[BP80] Julia Bickerstaffe and David Pearce. Can there be a consensus on nuclear power? Social Studies of Science, 10(3):309–344, August 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631278001000303.

Bauer:2006:LTT

[BPBG06] Martin W. Bauer, Kristina Petkova, Pepka Boyadjieva, and Galin Gornev. Long-term trends in the public representation of science across the ‘Iron Curtain’: 1946–1995. Social Studies of Science, 36(1):99–131, February 1, 2006. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312705053349.

Brickman:1979:SPA

[BR79] Ronald Brickman and Arie Rip. Science policy advisory councils in France, The Netherlands and the United States, 1957–77: a comparative analysis. Social Studies of Science, 9(2):167–198, May 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277900900202.

Brannigan:1979:RM

[Bra79] Augustine Brannigan. The reification of Mendel. Social Studies of Science, 9(4):423–454, November 1, 1979. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277900900403.

Brannigan:1989:AIA

[Bra89] Augustine Brannigan. Artificial intelligence and the attributional model of scientific discovery. Social Studies of Science, 19 (4):601–613, November 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631289019004002.

Braun:1992:CAG

[Bra92a] Hans-Joachim Braun. The Chrysler automotive gas turbine engine, 1950–80. Social Studies of Science, 22(2):339–351, May 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631292022002009.

Braun:1992:I

[Bra92b] Hans-Joachim Braun. Introduction. Social Studies of Science, 22(2):213–230, May 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631292022002002.

Braun:2004:BRM

[Bra04] Hans-Joachim Braun. Book review: Modern Sounds. Social Studies of Science, 34(5):809–817, October 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704044907.

Brandt:2014:ZCC

[Bra14] Marisa Brandt. Zapatista corn: a case study in biocultural innovation. Social Studies of Science, 44(6):874–900, December 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312714540060.

Braun:2024:SMP

[Bra24] Veit Braun. The stuff of memories: Planning hindsight in animal cryobanks. Social Studies of Science, 54(5):728–748, October 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127241252081.

Breslau:1997:CSE

[Bre97] Daniel Breslau. Contract shop epistemology: Credibility and problem construction in applied social science. Social Studies of Science, 27(3):363–394, June 1, 1997. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631297027003001.

Breslau:2002:PBA

[Bre02] Daniel Breslau. Pierre Bourdieu (1 August 1930–23 January 2002). Social Studies of Science, 32(4):631–635, August 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312702032004006.

Breslau:2013:DML

[Bre13] Daniel Breslau. Designing a market-like entity: Economics in the politics of market formation. Social Studies of Science, 43 (6):829–851, December 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312713493962.

Briskman:1972:KAA

[Bri72] L. B. Briskman. Is a Kuhnian analysis applicable to psychology? Social Studies of Science, 2(1):87–97, January 1, 1972. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277200200103.

Briggle:2008:RQE

[Bri08] Adam Briggle. Review: Questioning expertise. Social Studies of Science, 38(3):461–470, June 1, 2008. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706093668.

Briggle:2009:KCP

[Bri09] Adam Briggle. The Kass Council and the politicization of ethics advice. Social Studies of Science, 39(2):309–326, April 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312708101048.

Brives:2013:IOC

[Bri13] Charlotte Brives. Identifying ontologies in a clinical trial. Social Studies of Science, 43(3):397–416, June 1, 2013. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712472406.

Brisbois:2014:EDC

[Bri14] Ben W. Brisbois. Epidemiology and ‘developing countries’: Writing pesticides, poverty and political engagement in Latin America. Social Studies of Science, 44(4):600–624, August 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312714523514.

Brichzin:2020:MTP

[Bri20] Jenni Brichzin. Materializations through political work. Social Studies of Science, 50(2):271–291, April 1, 2020. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312719891626.

Bromberg:1982:TAP

[Bro82] Joan Lisa Bromberg. TFTR: The anatomy of a programme decision. Social Studies of Science, 12(4):559–583, November 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631282012004004.

Brosco:1989:HBB

[Bro89] Jeffrey P. Brosco. Henry Bryant Bigelow, the US Bureau of Fisheries, and intensive area study. Social Studies of Science, 19(2):239–264, May 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631289019002002.

Brown:1995:RWF

[Bro95] Paul Brown. Review: Welcome friends: Let’s proceed. Kay Milton (ed.), Environmentalism: The View from Anthropology, Association of Social Anthropologists Monographs, No. 32 (London & New York: Routledge, 1993), x + 240 pp., $17.95 / £14.99. ISBN 0-415-094755. Social Studies of Science, 25(1):

185–190, February 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631295025001009.

Brown:2015:PSC

[Bro15] Mark B. Brown. Politicizing science: Conceptions of politics in science and technology studies. Social Studies of Science, 45(1):3–30, February 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312714556694.

Brush:1977:ERS

[Bru77] Stephen G. Brush. Essay review: The search for quality in university research programmes: Drew, David. E.: Science Development: An Evaluation Study: A Technical Report presented to the National Board on Graduate Education (Washington, DC National Academy of Sciences, 1975), xvi + 182 pp. ISBN 0-309-02329-7. $5.75. Drew, David E., and Karpf, Ronald S. Evaluating Science Departments: A New Index (Santa Monica, Calif.: The Rand Corporation, 1975), 37 pp. Dolan, W. Patrick: The Ranking Game: The Power of the Academic Elite. Preface by William A. Arrowsmith (Lincoln, Neb. Evaluation of Higher Education Committee of the Study Commission on Undergraduate Education and the Education of Teachers, 1976), xvii + 108 pp. Social Studies of Science, 7(3):395–400, August 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700309.

Brush:1980:CCP

[Bru80] Stephen G. Brush. The chimerical cat: Philosophy of quantum mechanics in historical perspective. Social Studies of Science, 10(4):393–447, November 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL http://www.

jstor.org/stable/284918;   https://journals.sagepub.
com/doi/pdf/10.1177/030631278001000401.

Bruyninckx:2015:TTA

[Bru15] Joeri Bruyninckx. Trading twitter: Amateur recorders and economies of scientific exchange at the Cornell Library of Natural Sounds. Social Studies of Science, 45(3):344–370, June 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312715580404.

Button:1998:OAT

[BS98] Graham Button and Wes Sharrock. The organizational accountability of technological work. Social Studies of Science, 28(1):73–102, February 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631298028001003.

Bowker:2001:PRR

[BS01] Geoffrey C. Bowker and Susan Leigh Star. Pure, real and rational numbers: The American imaginary of countability. Social Studies of Science, 31(3):422–425, June 1, 2001. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031003006.

Bijsterveld:2004:BWP

[BS04a] Karin Bijsterveld and Marten Schulp. Breaking into a world of perfection: Innovation in today’s classical musical instruments. Social Studies of Science, 34(5):649–674, October 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312704047171.

Bricmont:2004:RGS

[BS04b] Jean Bricmont and Alan Sokal. Reply to Gabriel Stolzenberg. Social Studies of Science, 34(1):107–113, February 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312704040491. See [Sto04a, Sto04b].

Bruun:2008:DPS

[BS08] Henrik Bruun and Seppo Sierla. Distributed problem solving in software development: The case of an automation project. Social Studies of Science, 38(1):133–158, February 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312707077366.

Bijker:2016:CHM

[BSB16] Else M. Bijker, Robert W. Sauerwein, and Wiebe E. Bijker. Controlled human malaria infection trials: How tandems of trust and control construct scientific knowledge. Social Studies of Science, 46(1):56–86, February 1, 2016. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312715619784.

Baudry:2022:TCC

[BTS22] Jérôme Baudry, Élise Tancoigne, and Bruno J. Strasser. Turning crowds into communities: The collectives of online citizen science. Social Studies of Science, 52(3):399–424, June 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/03063127211058791.

Buchanan:1975:ERT

[Buc75a] Angus Buchanan. Essay review: Technology and history. Social Studies of Science, 5(4):489–499, November 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277500500408.

Buck:1975:OCS

[Buc75b] Peter Buck. Order and control: The scientific method in China and the United States. Social Studies of Science, 5 (3):237–267, August 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277500500301.

Buchholz:1979:VCE

[Buc79] Klaus Buchholz. Verfahrenstechnik (chemical engineering) — its development, present state and structure. Social Studies of Science, 9(1):33–62, February 1, 1979. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277900900102.

Buchanan:1992:ARK

[Buc92] R. A. Buchanan. The atmospheric railway of I. K. Brunel. Social Studies of Science, 22(2):231–243, May 1, 1992. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631292022002003.

Bud:1978:SAC

[Bud78] R. F. Bud. Strategy in American cancer research after World War II: a case study. Social Studies of Science, 8(4):425– 459, November 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277800800402.

Bud:1985:CDC

[Bud85] Robert Bud. The case of the disappearing caveat: a critique of Irvine and Martin’s methodology. Social Studies of Science, 15 (3):548–553, August 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631285015003008. See

[IM85, MI85].

Bud:1991:BTC

[Bud91] Robert Bud. Biotechnology in the Twentieth Century. Social Studies of Science, 21(3):415–457, August 1, 1991. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631291021003002.

Bud:1998:KM

[Bud98] Robert Bud. Knowing machines. Social Studies of Science, 28(2):349–351, April 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631298028002009.

Bugos:1993:MCT

[Bug93] Glenn E. Bugos. Manufacturing certainty: Testing and program management for the F-4 Phantom II. Social Studies of Science, 23(2):265–300, May 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631293023002002.

Burstyn:1975:SPS

[Bur75] Harold L. Burstyn. Science pays off: Sir John Murray and the Christmas Island phosphate industry, 1886–1914. Social Studies of Science, 5(1):5–34, February 1, 1975. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277500500102.

Burrage:1983:WUT

[Bur83] Hilary F. Burrage. Women University teachers of natural science, 1971–72: an empirical survey. Social Studies of Science, 13 (1):147–160, February 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631283013001008.

Burrows:1992:IOC

[Bur92] Andrea C. Burrows. Inside the outer circle: Productivity, performance and career patterns. Social Studies of Science, 22 (3):581–585, August 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312792022003007.

Burri:2008:DDB

[Bur08] Regula Valérie Burri. Doing distinctions: Boundary work and symbolic capital in radiology. Social Studies of Science, 38 (1):35–62, February 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312707082021.

Burton:2018:NED

[Bur18] Elise K. Burton. Narrating ethnicity and diversity in Middle Eastern national genome projects. Social Studies of Science, 48 (5):762–786, October 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312718804888.

Bloomfield:1995:DBN

[BV95] Brian P. Bloomfield and Theo Vurdubakis. Disrupted boundaries: New reproductive technologies and the language of anxiety and expectation. Social Studies of Science, 25(3):533–551, August 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631295025003005.

Brannigan:1983:HDM

[BW83] Augustine Brannigan and Richard A. Wanner. Historical distributions of multiple discoveries and theories of scientific change. Social Studies of Science, 13(3):417–435, August 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631283013003004.

Bychkova:2016:ICE

[Byc16] Olga Bychkova. Innovation by coercion: Emerging institutionalization of university-industry collaborations in Russia. Social Studies of Science, 46(4):511–535, August 1, 2016. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716654768.

Callon:1999:WIP

[Cal99] Michel Callon. Whose imposture?: Physicists at war with the third person. Social Studies of Science, 29(2):261–286, April 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631299029002011.

Calvillo:2018:PAM

[Cal18] Nerea Calvillo. Political airs: From monitoring to attuned sensing air pollution. Social Studies of Science, 48(3):372–388, June 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312718784656.

Calkins:2021:TRI

[Cal21] Sandra Calkins. Toxic remains: Infrastructural failure in a Ugandan molecular biology lab. Social Studies of Science, 51 (5):707–728, October 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/03063127211011531.

Campbell:1985:USA

[Cam85] Brian L. Campbell. Uncertainty as symbolic action in disputes among experts. Social Studies of Science, 15(3):429–453, August 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631285015003002.

Campanario:1993:CSS

[Cam93] Miguel Campanario. Consolation for the scientist: Sometimes it is hard to publish papers that are later highly-cited. Social Studies of Science, 23(2):342–362, May 1, 1993. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631293023002005. See reflections [Nis95].

Campbell:2003:PNG

[Cam03] Robert A. Campbell. Preparing the next generation of scientists: The social process of managing students. Social Studies of Science, 33(6):897–927, December 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312703336004.

Campbell:2004:CPP

[Cam04] Nancy D. Campbell. Credible performances: The performativity of science studies. Social Studies of Science, 34(3):433–442, June 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312704043693.

Cantor:1975:APE

[Can75] G. N. Cantor. The Academy of Physics at Edinburgh 1797– 1800. Social Studies of Science, 5(2):109–134, May 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277500500201.

Carson:2001:ODB

[Car01] John Carson. Opening the democracy Box. Social Studies of Science, 31(3):425–428, June 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631201031003007.

Carroll:2005:BRH

[Car05] Patrick Carroll. Book review: Honoring Thomas Hughes, Appealing to Historians. Social Studies of Science, 35(2):313–321, April 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312705048717.

Carolan:2009:EVG

[Car09] Michael S. Carolan. Ethanol versus gasoline: The contestation and closure of a socio-technical system in the USA. Social Studies of Science, 39(3):421–448, June 1, 2009. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312708101049.

Carroll:2012:WTS

[Car12a] Patrick Carroll. Water and technoscientific state formation in California. Social Studies of Science, 42(4):489–516, August 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312712437977.

Carse:2012:NIM

[Car12b] Ashley Carse. Nature as infrastructure: Making and managing the Panama Canal watershed. Social Studies of Science, 42 (4):539–563, August 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712440166.

Carroll:2014:BDL

[Car14] Katherine Carroll. Body dirt or liquid gold? How the ‘safety’ of donated breastmilk is constructed for use in neonatal intensive care. Social Studies of Science, 44(3):466–485, June 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312714521705.

Carson:2018:EEW

[Car18] John Carson. ‘Every expression is watched’: Mind, medical expertise and display in the nineteenth-century English courtroom. Social Studies of Science, 48(6):891–918, December 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312718820160.

Carro:2021:WSA

[Car21a] Remo Fernández Carro. What is a scientific article? A principalagent explanation. Social Studies of Science, 51(2):298–309, April 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312720951860.

Carse:2021:EEM

[Car21b] Ashley Carse. The ecobiopolitics of environmental mitigation: Remaking fish habitat through the Savannah Harbor Expansion Project. Social Studies of Science, 51(4):512–537, August 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312721992541.

Carlsson:2022:CVR

[Car22] Moa Carlsson. Computing views, remodeling environments. Social Studies of Science, 52(2):227–252, April 1, 2022. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/03063127211048943.

Castel:2009:WBG

[Cas09] Patrick Castel. What’s behind a guideline?: Authority, competition and collaboration in the French oncology sector. Social Studies of Science, 39(5):743–764, October 1, 2009. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312709104435.

Casper:2015:N

[Cas15] Stephen T. Casper. Neuropsychedelia. Social Studies of Science, 45(4):620–622, August 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312715598398.

Cartner:1997:SPR

[CB97] Monica Cartner and Tim Bollinger. Science policy reforms: The new Zealand experience. Social Studies of Science, 27 (5):775–803, October 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631297027005003.

Carroll-Burke:2001:TIE

[CB01] Patrick Carroll-Burke. Tools, instruments and engines: Getting a handle on the specificity of engine science. Social Studies of Science, 31(4):593–625, August 1, 2001. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031004005.

Coopmans:2014:EE

[CB14] Catelijne Coopmans and Graham Button. Eyeballing expertise. Social Studies of Science, 44(5):758–785, October 1, 2014.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312714531472. See response [CE14].

Cole:2024:LAP

[CB24] Simon A. Cole and Alyse Bertenthal. Law’s artefacts: Personal rapid transit and public narratives of hitchhiking and crime. Social Studies of Science, 54(5):749–776, October 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127241229071.

Cusworth:2023:WYW

[CBLG23] George Cusworth, Jeremy Brice, Jamie Lorimer, and Tara Garnett. When you wish upon a (GWP) star: Environmental governance and the reflexive performativity of global warming metrics. Social Studies of Science, 53(1):3–28, February 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

Collins:1976:RRD

[CC76] H. M. Collins and Graham Cox. Recovering relativity: Did prophecy fail? Social Studies of Science, 6(3–4):423–444, September 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277600600307. See

note [Law77].

Collins:1977:RRM

[CC77] Harry Collins and Graham Cox. Relativity revisited: Mrs Keech — a suitable case for special treatment? Social Studies of Science, 7(3):372–380, August 1, 1977. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700306.

Casper:1998:MPS

[CC98] Monica J. Casper and Adele E. Clarke. Making the Pap smear into the ‘right tool’ for the job: Cervical cancer screening in the USA, circa 1940–95. Social Studies of Science, 28(2):255–290, April 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631298028002003.

Cornwell:2012:CPC

[CC12] Myriah L. Cornwell and Lisa M. Campbell. Co-producing conservation and knowledge: Citizen-based sea turtle monitoring in North Carolina, USA. Social Studies of Science, 42(1): 101–120, February 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312711430440.

Collingridge:1984:TMP

[CD84] David Collingridge and Jenny Douglas. Three models of policymaking: Expert advice in the control of environmental lead. Social Studies of Science, 14(3):343–370, August 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014003002.

Collins:2002:TWS

[CE02] H. M. Collins and Robert Evans. The third wave of science studies: Studies of expertise and experience. Social Studies of Science, 32(2):235–296, April 1, 2002. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/0306312702032002003. See comments [Jas03a, Wyn03].

Collins:2003:KCM

[CE03] H. M. Collins and Robert Evans. King Canute meets the Beach Boys: Responses to the third wave. Social Studies of Science, 33(3):435–452, June 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/03063127030333007.

Collins:2014:AAR

[CE14] Harry Collins and Robert Evans. Actor and analyst: a response to Coopmans and Button. Social Studies of Science, 44 (5):786–792, October 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312714546242.  See

[CB14].

Collins:2016:TWW

[CE16] Harry Collins and Robert Evans. A thousand words is worth a picture. Social Studies of Science, 46(2):312–324, April 1, 2016.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716637437.

Collins:2025:VDRa

[CERG25a] Harry Collins, Robert Evans, and Luis Reyes-Galindo. Virtual diversity: Resolving the tension between the wider culture and the institution of science. Social Studies of Science, 55(2):262–287, April 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127241263609.

Collins:2025:VDRb

[CERG25b] Harry Collins, Robert Evans, and Luis Reyes-Galindo. Virtual diversity revisited. Social Studies of Science, 55(2):316–324, April 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
abs/10.1177/03063127251330545.

Collins:2017:SSP

[CEW17] Harry Collins, Robert Evans, and Martin Weinel. STS as science or politics? Social Studies of Science, 47(4):580–586, August 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312717710131. See reply [Sis17a].

Collins:2001:WTT

[CFS01] Harry Collins, Sam Finn, and Patrick Sutton. What is TWAP: Three notes on the American election in the year 2000. Social Studies of Science, 31(3):428–436, June 1, 2001. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031003008.

Crosland:1989:ERG

[CG89] Maurice Crosland and Antonio Gálvez. The emergence of research grants within the prize system of the French Academy of Sciences, 1795–1914. Social Studies of Science, 19(1):71– 100, February 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631289019001002.

Carlson:1990:UIC

[CG90] W. Bernard Carlson and Michael E. Gorman. Understanding invention as a cognitive process: The case of Thomas Edison and early motion pictures, 1888–91. Social Studies of Science, 20(3):387–430, August 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631290020003001.

Carlson:1992:STG

[CG92] W. Bernard Carlson and Michael E. Gorman. Socio-technical graphs and cognitive maps: a response to Latour, Mauguin and Teil. Social Studies of Science, 22(1):81–91, February 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/0306312792022001004. See [LMT92] and reply [Lat92].

Chiapperino:2025:PIO

[CGP25] Luca Chiapperino, Nils Graber, and Francesco Panese. A precision immuno-oncology turn? Hybridizing cancer genomics and immunotherapy through neoantigens-based adoptive cell therapies. Social Studies of Science, 55(4):565–590, August 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127241303720.

Cuevas-Garcia:2024:MIH

[CGPP24] Carlos Cuevas-Garcia, Federica Pepponi, and Sebastian M. Pfotenhauer. Maintaining innovation: How to make sewer robots and innovation policy work in Barcelona. Social Studies of Science, 54(3):352–376, June 1, 2024. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231207082.

Collins:1975:BTL

[CH75] H. M. Collins and R. G. Harrison. Building a TEA laser: The caprices of communication. Social Studies of Science, 5(4): 441–450, November 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277500500404.

Chamak:1999:ECS

[Cha99] Brigitte Chamak. The emergence of cognitive science in France: A comparison with the USA. Social Studies of Science, 29 (5):643–684, October 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631299029005001.

Chandler:2022:ADI

[Cha22] Katherine Chandler. Apartheid drone: Infrastructures of militarism and the hidden genealogies of the South African Seeker. Social Studies of Science, 52(4):512–535, August 1, 2022. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127221105748.

Chevallier:2023:SPC

[Che23] Martin Chevallier. Staging Paro: The care of making robot(s) care. Social Studies of Science, 53(5):635–659, October 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127221126148.

Chiapperino:2024:EBC

[Chi24] Luca Chiapperino. Enacting biosocial complexity: Stress, epigenetic biomarkers and the tools of postgenomics. Social Studies of Science, 54(4):598–625, August 1, 2024. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231222613.

Choudhuri:1985:PWS

[Cho85] Arnab Rai Choudhuri. Practising Western science outside the West: Personal observations on the Indian scene. Social Studies of Science, 15(3):475–505, August 1, 1985. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631285015003004.

Chong:2013:LJA

[Cho13] Phillipa Chong. Legitimate judgment in art, the scientific world reversed? Maintaining critical distance in evaluation.

Social Studies of Science, 43(2):265–281, April 1, 2013. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712475256.

Christie:1988:RYD

[Chr88] J. R. R. Christie. Robert Young, Darwin’s metaphor: Nature’s place in Victorian culture. Social Studies of Science, 18(1): 187–189, February 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631288018001009.

Chubin:1975:TAS

[Chu75] Daryl E. Chubin. Trusted assessorship in science: a relation in need of data. Social Studies of Science, 5(3):362–367, August 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277500500307.

Chubin:1982:CPH

[Chu82] Daryl E. Chubin. Collins’s programme and the ‘hardest possible case’. Social Studies of Science, 12(1):136–139, February 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631282012001008.

Chubin:1986:DI

[Chu86] Daryl Chubin. A declaration of interests. Social Studies of Science, 16(3):549–557, August 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631286016003009.

Cambrosio:1983:DSC

[CK83] Alberto Cambrosio and Peter Keating. The disciplinary stake: The case of chronobiology. Social Studies of Science, 13(3): 323–353, August 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631283013003001.

Cambrosio:1985:SBR

[CK85] Alberto Cambrosio and Peter Keating. Studying a biotechnology research centre: a note on local socio-political issues. Social Studies of Science, 15(4):723–737, November 1, 1985.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631285015004006.

Cummings:2005:CRA

[CK05] Jonathon N. Cummings and Sara Kiesler. Collaborative research across disciplinary and organizational boundaries. Social Studies of Science, 35(5):703–722, October 1, 2005. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312705055535.

Cook:2013:PNC

[CKFS13] Brian R. Cook, Mike Kesby, Ioan Fazey, and Chris Spray. The persistence of ‘normal’ catchment management despite the participatory turn: Exploring the power effects of competing frames of reference. Social Studies of Science, 43(5):754–779, October 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312713478670.

Cambrosio:2004:MCW

[CKM04] Alberto Cambrosio, Peter Keating, and Andrei Mogoutov. Mapping collaborative work and innovation in biomedicine: a computer-assisted analysis of antibody reagent workshops. Social Studies of Science, 34(3):325–364, June 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704043767.

Cambrosio:2009:BCR

[CKSW09] Alberto Cambrosio, Peter Keating, Thomas Schlich, and George Weisz. Biomedical conventions and regulatory objectivity: a few introductory remarks. Social Studies of Science, 39(5): 651–664, October 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312709334640.

Callon:1982:ITT

[CL82] Michel Callon and John Law. On interests and their transformation: Enrolment and counter-enrolment. Social Studies of Science, 12(4):615–625, November 1, 1982. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631282012004006.

Courtial:1989:CWS

[CL89] Jean-Pierre Courtial and John Law. A co-word study of artificial intelligence. Social Studies of Science, 19(2):301–311, May 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631289019002005.

Cesarino:2011:ERD

[CL11] Letícia Cesarino and Naara Luna. The embryo research debate in Brazil: From the national Congress to the Federal supreme court. Social Studies of Science, 41(2):227–250, April 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312710386637.

Clark:1980:EBR

[Cla80] Norman Clark. The economic behaviour of research institutions in developing countries — some methodological points. Social Studies of Science, 10(1):75–93, February 1, 1980. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278001000104.

Clark:1982:RM

[Cla82] Norman Clark. Reply to Moravcsik. Social Studies of Science, 12(1):147–150, February 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631282012001011. See

[Mor82].

Collins:2023:SCS

[CLCMW23] Harry Collins, Willow Leonard-Clarke, and Will Mason-Wilkes. Scientific conferences, socialization, and the Covid-19 pandemic: a conceptual and empirical enquiry. Social Studies of Science, 53(3):379–401, June 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127221138521.

Clemens:1986:ADR

[Cle86] Elisabeth S. Clemens. Of asteroids and dinosaurs: The role of the Press in the shaping of scientific debate. Social Studies of Science, 16(3):421–456, August 1, 1986. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631286016003002.

Cloud:2001:IWB

[Clo01] John Cloud. Imaging the world in a barrel: CORONA and the clandestine convergence of the earth sciences. Social Studies of Science, 31(2):231–251, April 1, 2001. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312701031002005.

Cloud:2003:SGE

[Clo03] John Cloud. Special guest-edited issue on the Earth sciences in the Cold War. Social Studies of Science, 33(5):629–633, October 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312703335001.

Cambrosio:1990:RBE

[CLP90] Alberto Cambrosio, Camille Limoges, and Denyse Pronovost. Representing biotechnology: an ethnography of Quebec science policy. Social Studies of Science, 20(2):195–227, May 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/030631290020002001. See comments [Kle91] and reply [CLP91].

Cambrosio:1991:ASP

[CLP91] Alberto Cambrosio, Camille Limoges, and Denyse Pronovost. Analyzing science policy-making: Political ontology or ethnography?: a reply to Kleinman. Social Studies of Science, 21(4): 775–781, November 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631291021004008. See

[Kle91].

Chubin:1975:CAR

[CM75] Daryl E. Chubin and Soumyo D. Moitra. Content analysis of references: Adjunct or alternative to citation counting? Social Studies of Science, 5(4):423–441, November 1, 1975.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277500500403.

Choi:2009:LHM

[CM09] Hyungsub Choi and Cyrus C. M. Mody. The long history of molecular electronics: Microelectronics origins of nanotechnology. Social Studies of Science, 39(1):11–50, February 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312708097288.

Cohn:1986:EFC

[Coh86] Steven F. Cohn. The effects of funding changes upon the rate of knowledge growth in algebraic and differential topology, 1955– 75. Social Studies of Science, 16(1):23–59, February 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631286016001003.

Collins:1974:TST

[Col74] H. M. Collins. The TEA set: Tacit knowledge and scientific networks. Social Studies of Science, 4(2):165–185, April 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277400400203.

Collins:1981:SSS

[Col81a] H. M. Collins. Son of seven sexes: The social destruction of a physical phenomenon. Social Studies of Science, 11(1): 33–62, February 1, 1981. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631278101100103.

Collins:1981:SEP

[Col81b] H. M. Collins. Stages in the empirical programme of Relativism. Social Studies of Science, 11(1):3–10, February 1, 1981. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278101100101. See note [Lau82].

Collins:1982:KNR

[Col82a] H. M. Collins. Knowledge, norms and rules in the sociology of science. Social Studies of Science, 12(2):299–309, May 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631282012002005.

Collins:1982:SRN

[Col82b] H. M. Collins. Special relativism — the natural attitude. Social Studies of Science, 12(1):139–143, February 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631282012001009.

Collingridge:1985:CTR

[Col85a] David Collingridge. Controlling technology (response to Johnston). Social Studies of Science, 15(2):373–380, May 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/030631285015002006. See [Joh84] and reply [Joh85].

Collingridge:1985:RS

[Col85b] David Collingridge. Reforming science. Social Studies of Science, 15(4):763–769, November 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631285015004009.

Collins:1985:PSP

[Col85c] H. M. Collins. The possibilities of science policy. Social Studies of Science, 15(3):554–558, August 1, 1985. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631285015003009.

Collins:1985:ELW

[Col85d] H. M. Collins. With enemies like this, who needs friends? Social Studies of Science, 15(1):176–177, February 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631285015001008. See reply [RH85].

Collins:1986:DD

[Col86] H. M. Collins. Dimensions of debate. Social Studies of Science, 16(3):557–563, August 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631286016003010.

Collins:1987:CPU

[Col87] H. M. Collins. Certainty and the public understanding of science: Science on television1. Social Studies of Science, 17(4): 689–713, November 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631287017004005.

Collins:1988:PED

[Col88] H. M. Collins. Public experiments and displays of virtuosity: The core-set revisited. Social Studies of Science, 18(4): 725–748, November 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631288018004006.

Collins:1989:CSS

[Col89] H. M. Collins. Computers and the sociology of scientific knowledge1. Social Studies of Science, 19(4):613–624, November 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631289019004003.

Collins:1991:SS

[Col91] H. M. Collins. Simon’s Slezak. Social Studies of Science, 21 (1):148–149, February 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631291021001009.

Collins:1992:AVR

[Col92a] H. M. Collins. AI-Vey!: Response to Slezak. Social Studies of Science, 22(1):201–203, February 1, 1992. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312792022001011. See [Sle92].

Collins:1992:HDF

[Col92b] Harry Collins. Hubert L. Dreyfus, forms of life, and a simple test for machine intelligence. Social Studies of Science, 22(4): 726–739, November 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631292022004008.

Collins:1994:DSF

[Col94a] H. M. Collins. Dissecting surgery: Forms of life depersonalized. Social Studies of Science, 24(2):311–333, May 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/030631279402400205. See [Hir91] and reply [Hir94].

Collins:1994:SA

[Col94b] H. M. Collins. Scene from afar. Social Studies of Science, 24(2):369–389, May 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631279402400209.

Collins:1995:CTC

[Col95] H. M. Collins. Cooperation and the two cultures: Response to Labinger. Social Studies of Science, 25(2):306–309, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631295025002005. See [Lab95b].

Cole:1996:WCF

[Col96a] Simon A. Cole. Which came first, the fossil or the fuel? Social Studies of Science, 26(4):733–766, November 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631296026004002.

Collins:1996:PFG

[Col96b] H. M. Collins. In praise of futile gestures: How scientific is the sociology of scientific knowledge? Social Studies of Science, 26(2):229–244, May 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631296026002002.

Cole:1998:WIL

[Col98] Simon A. Cole. Witnessing identification: Latent fingerprinting evidence and expert knowledge. Social Studies of Science, 28 (5–6):687–712, October 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631298028005002.

Collins:1999:PSS

[Col99a] H. M. Collins. Philosophy of science and SSK: Reply to Koertge. Social Studies of Science, 29(5):785–790, October 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029005006. See [Koe99].

Collins:1999:SP

[Col99b] H. M. Collins. The science police. Social Studies of Science, 29(2):287–294, April 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631299029002012.

Collins:1999:TAP

[Col99c] H. M. Collins. Tantalus and the aliens: Publications, audiences and the search for gravitational waves. Social Studies of Science, 29(2):163–197, April 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631299029002001.

Collins:2001:TKT

[Col01] H. M. Collins. Tacit knowledge, trust and the q of Sapphire. Social Studies of Science, 31(1):71–85, February 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031001004.

Collins:2003:MDO

[Col03] Harry Collins. In memoriam: David Owen Edge. Social Studies of Science, 33(2):185–187, April 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127030332006.

Cole:2004:MCS

[Col04a] Stephen Cole. Merton’s contribution to the sociology of science. Social Studies of Science, 34(6):829–844, December 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312704048600.

Collins:2004:HDY

[Col04b] Harry Collins. How do you know you’ve alternated? Social Studies of Science, 34(1):103–106, February 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704041342.

Cole:2006:WC

[Col06] Simon A. Cole. Witnessing creation. Social Studies of Science, 36(6):855–860, December 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/0306312706067901.

Collins:2007:RMM

[Col07] Samuel Gerald Collins. Review: Making magic, writing culture: The magic of the magic of modernity: Randall Styers, Making Magic: Religion, Magic, & Science in the Modern World (Oxford & New York: Oxford University Press, 2004), vi + 298 pp., £52.00 / EUR 60.92 / $85.00 (hbk), £16.00 / EUR 25.90 / $27.50 (pbk). ISBN 0-19-515107-0 (hbk), 0-19-516941-7 (pbk). Birgit Meyer and Peter Pels (eds), Magic and Modernity: Interfaces of Revelation and Concealment (Stanford, CA: Stanford University Press, 2003), viii + 392 pp., £47.50 / EUR 69.85 / $93.83 (hbk), £16.95 / EUR 14.93 / $24.95 (pbk). ISBN 0-8047-4463-7 (hbk), 0-8047-4464-5 (pbk). Social Studies of Science, 37(3):489–494, June 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706071792.

Collins:2011:LP

[Col11] Harry Collins. Language and practice. Social Studies of Science, 41(2):271–300, April 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312711399665.

Collins:2012:CK

[Col12] Harry Collins. Comment on Kuhn. Social Studies of Science, 42(3):420–423, June 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712436571.  See

[Lyn12b].

Collins:2013:CLF

[Col13] Samuel Gerald Collins. Cutting loose in the fun palace: Reengaging cybernetics in a complex world. Social Studies of Science, 43(2):314–318, April 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712461712.

Collins:2014:RKC

[Col14] Harry Collins. Rejecting knowledge claims inside and outside science. Social Studies of Science, 44(5):722–735, October 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312714536011.

Collins:2022:TPJ

[Col22] Harry Collins. Trevor Pinch (1 January 1952–16 December 2021). Social Studies of Science, 52(1):144–146, February 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/03063127221074075.

Constant:1978:DCE

[Con78] Edward W. Constant II. On the diversity and co-evolution of technological multiples: Steam turbines and Pelton water wheels. Social Studies of Science, 8(2):183–210, May 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277800800202.

Constant:1979:ESS

[Con79] Edward Constant. To the Editors, Social Studies of Science Pittsburgh, 13 March 1979 Dear Sirs. Social Studies of Science, 9(4):510, November 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277900900410.

Constant:1989:SSP

[Con89] Edward W. Constant III. Science in society: Petroleum engineers and the oil fraternity in Texas, 1925–65. Social Studies of Science, 19(3):439–472, August 1, 1989. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019003003.

Coopmans:2011:FVN

[Coo11] Catelijne Coopmans. ‘Face value’: New medical imaging software in commercial view. Social Studies of Science, 41(2): 155–176, April 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312710389226.

Cool:2019:IUA

[Coo19] Alison Cool. Impossible, unknowable, accountable: Dramas and dilemmas of data law. Social Studies of Science, 49(4): 503–530, August 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312719846557.

Corones:1994:BRP

[Cor94] Anthony Corones. Book review: Philosophy of Science and its Malcontents. Social Studies of Science, 24(1):139–142, February 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631279402400111.

Costelloe:2001:CD

[Cos01] Timothy M. Costelloe. Constructivism dissected. Social Studies of Science, 31(3):469–472, June 1, 2001. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031003017.

Cotgrove:1975:TRD

[Cot75] Stephen Cotgrove. Technology, rationality and domination. Social Studies of Science, 5(1):55–78, February 1, 1975. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277500500104.

Coutinho:1999:NYC

[Cou99] Marilia Coutinho. Ninety years of Chagas disease: a success story at the periphery. Social Studies of Science, 29(4): 519–549, August 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631299029004003.

Cowan:1995:WSC

[Cow95] Ruth Schwartz Cowan. Women and science: Contested terrain. Social Studies of Science, 25(2):363–370, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631295025002017.

Cozzens:1985:CSC

[Coz85] Susan E. Cozzens. Comparing the sciences: Citation context analysis of papers from neuropharmacology and the sociology of science. Social Studies of Science, 15(1):127–153, February 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631285015001005.

Cozzens:1986:EI

[Coz86] Susan E. Cozzens. Editor’s introduction. Social Studies of Science, 16(1):9–21, February 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631286016001002.

Crane:1980:ESK

[Cra80a] Diana Crane. An exploratory study of Kuhnian paradigms in theoretical high energy physics. Social Studies of Science, 10 (1):23–54, February 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631278001000102. See

comment [Pic80a] and replies [Cra80b, Pic80b].

Crane:1980:RP

[Cra80b] Diana Crane. Reply to Pickering. Social Studies of Science, 10 (4):502–506, November 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631278001000405. See

[Cra80a, Pic80a, Pic80b].

Crane:2010:AEP

[Cra10] Johanna Crane. Adverse events and placebo effects: African scientists, HIV, and ethics in the ‘global health sciences’. Social Studies of Science, 40(6):843–870, December 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312710371145.

Crist:1996:NPA

[Cri96] Eileen Crist. Naturalists’ portrayals of animal life: Engaging the Verstehen approach. Social Studies of Science, 26(4): 799–838, November 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631296026004004.

Crist:2004:CIS

[Cri04] Eileen Crist. Can an insect speak?: The case of the honeybee dance language. Social Studies of Science, 34(1):7–43, February 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312704040611.

Crist:2010:RCC

[Cri10] Eileen Crist. Review: Cat’s cradle with Donna Haraway [When Species Meet (University of Minnesota Press, 2008), x + 360 pp., £15.50/18.22/$24.95 (pbk), £46.50/54.68/$75.00 (hbk). ISBN: 978-0-8166-5046-0 (pbk), 978-0-8166-5045-3 (hbk)]. Social Studies of Science, 40(4):641–646, August 1, 2010. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312710365588.

Crosland:1976:SFP

[Cro76] Maurice Crosland. Science and the Franco-Prussian war. Social Studies of Science, 6(2):185–214, May 1, 1976. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277600600202.

Cross:1991:CPD

[Cro91] Andrew Cross. The crisis in physics: Dialectical materialism and quantum theory. Social Studies of Science, 21(4):735–759, November 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631291021004005. See comment

[Fre92] and reply [Cro92].

Cross:1992:RF

[Cro92] Andrew Cross. Reply to Freire. Social Studies of Science, 22 (4):742–744, November 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631292022004010.

Crozier:2000:SP

[Cro00] Ivan Crozier. Social psychology. Social Studies of Science, 30 (4):633–639, August 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631200030004006.

Carneiro:2000:ESP

[CS00] Ana Carneiro and Ana Simões. Enlightenment science in Portugal: The Estrangeirados and their communication networks. Social Studies of Science, 30(4):591–619, August 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030004004.

Castaneda:2014:RV

[CS14] Claudia Castañeda and Lucy Suchman. Robot visions. Social Studies of Science, 44(3):315–341, June 1, 2014. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713511868.

Calvert:2017:WCS

[CS17] Jane Calvert and Pablo Schyfter. What can science and technology studies learn from art and design? Reflections on ‘synthetic aesthetics’. Social Studies of Science, 47(2):195–215, April 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312716678488.

Cochoy:2020:BFD

[CS20] Franck Cochoy and Bastien Soutjis. Back to the future of digital price display: Analyzing patents and other archives to understand contemporary market innovations. Social Studies of Science, 50(1):3–29, February 1, 2020. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312719884643.

Castillo-Sepulveda:2023:BSO

[CSTG23] Jorge Castillo-Sepúlveda, Francisco Tirado, and Ana Gálvez. Biopolitics and speculative objects in Chilean health projects.

Social Studies of Science, 53(2):194–212, April 1, 2023. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127221136804.

Chambers:1989:SWI

[CT89] David Wade Chambers and David Turnbull. Science worlds: an integrated approach to social studies of science teaching. Social Studies of Science, 19(1):155–179, February 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019001008.

Cueto:1990:RFM

[Cue90] Marcos Cueto. The Rockefeller Foundation’s medical policy and scientific research in Latin America: The case of physiology. Social Studies of Science, 20(2):229–254, May 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020002002.

Curtis:1994:NFN

[Cur94] Ron Curtis. Narrative form and normative force: Baconian story-telling in popular science. Social Studies of Science, 24 (3):419–461, August 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631279402400301.

Curtis:1998:MTM

[Cur98] Bruce Curtis. From the moral thermometer to money: Metrological reform in pre-Confederation Canada. Social Studies of Science, 28(4):547–570, August 1, 1998. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631298028004002.

Curry:2022:HSB

[Cur22] Helen Anne Curry. The history of seed banking and the hazards of backup. Social Studies of Science, 52(5):664–688, October 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127221106728.

Cussins:1996:OCA

[Cus96] Charis Cussins. Ontological choreography: Agency through objectification in infertility clinics. Social Studies of Science, 26 (3):575–610, August 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631296026003004.

Clark:1987:PF

[CW87] Tim Clark and Ron Westrum. Paradigms and ferrets. Social Studies of Science, 17(1):3–33, February 1, 1987. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631287017001001.

Carboni:2023:EAM

[CWvdVdB23] Chiara Carboni, Rik Wehrens, Romke van der Veen, and Antoinette de Bont. Eye for an AI: More-than-seeing, fauxtomation, and the enactment of uncertain data in digital pathology. Social Studies of Science, 53(5):712–737, October 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231167589.

daCostaMarques:2022:AEE

[da 22] Ivan da Costa Marques. Anthropophagy, European enlightenment, science and technology studies, and responsible knowledge construction in Brazil. Social Studies of Science, 52(6):812–828, December 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

Delamont:2001:DUM

[DA01] Sara Delamont and Pauln Atkinson. Doctoring uncertainty: Mastering craft knowledge. Social Studies of Science, 31(1): 87–107, February 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631201031001005.

Druglitro:2024:ECC

[DA24] Tone Druglitrø and Kristin Asdal. Experimenting with care and cod: On document-practices, versions of care and fish as the new experimental animal. Social Studies of Science, 54(5):706–727, October 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
abs/10.1177/03063127231223904.

Daemmrich:1998:EDS

[Dae98] Arthur Daemmrich. The evidence does not speak for itself: Expert witnesses and the organization of DNA-typing companies. Social Studies of Science, 28(5–6):741–772, October 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631298028005004.

Dahlin:2024:SAR

[Dah24] Emma Dahlin. And say the AI responded? Dancing around ‘autonomy’ in AI/human encounters. Social Studies of Science, 54(1):59–77, February 1, 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127231193947.

Daly:1990:CTP

[Dal90] John A. Daly. Comments on Turnbull’s ‘push for a malaria vaccine’. Social Studies of Science, 20(2):371–379, May 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020002008. See [Tur89a].

Daly:1991:DCV

[Dal91] John A. Daly. Does a constructivist view require epistemological relativism?: a response to Turnbull. Social Studies of Science, 21(3):568–571, August 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631291021003009. See

[Tur90, Tur91].

Damianos:2025:AAA

[Dam25] Alexander Damianos. Anthropocene angst: Authentic geology and stratigraphic sincerity. Social Studies of Science, 55(3):444–464, June 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127241282309.

Danyi:2020:IOP

[Dán20] Endre Dányi. The insides and outsides of parliamentary politics: Introduction to a special section of Social Studies of Science on parliamentary practices. Social Studies of Science, 50 (2):245–251, April 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312720908935.

Daston:1992:OEP

[Das92] Lorraine Daston. Objectivity and the escape from perspective. Social Studies of Science, 22(4):597–618, November 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631292022004002.

Davis:2020:HAR

[Dav20] Clay Davis. Homo adhaerens: Risk and adherence in biomedical HIV prevention research. Social Studies of Science, 50(6): 860–880, December 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312720933943.

Davies:2021:ASE

[Dav21] Sarah R. Davies. Atmospheres of science: Experiencing scientific mobility. Social Studies of Science, 51(2):214–232, April 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720953520.

Davis:2024:RLE

[Dav24] Clay Davis. The routinization of lay expertise: a diachronic account of the invention and stabilization of an open-source artificial pancreas. Social Studies of Science, 54(4):626–652, August 1, 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127231214237.

Daye:2008:REA

[Day08] Christian Dayé. Review essays: ‘according to the books … ’: Reviewing science studies textbooks. Social Studies of Science, 38(2):303–318, April 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312707088127.

Daye:2018:HTY

[Day18] Christian Dayé. How to train your oracle: The Delphi method and its turbulent youth in operations research and the policy sciences. Social Studies of Science, 48(6):846–868, December 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312718798497.

deBBeaver:1972:SOR

[dB72] Donald deB. Beaver. The Smithsonian origin of the Royal Society Catalogue of Scientific Papers. Social Studies of Science, 2(4):385–393, October 1, 1972. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277200200407.

Dolman:1985:SSC

[DB85] Han Dolman and Henk Bodewitz. Sedimentation of a scientific concept: The use of citation data. Social Studies of Science, 15 (3):507–523, August 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631285015003005.

deBont:2009:BLD

[dB09] Raf de Bont. Between the laboratory and the Deep Blue Sea: Space issues in the marine stations of Naples and Wimereux. Social Studies of Science, 39(2):199–227, April 1, 2009. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312708097325.

Beaver:1976:RNH

[dBB76] Donald de B. Beaver. Reflections on the natural history of eponymy and scientific law. Social Studies of Science, 6(1): 89–98, February 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277600600105.

deBoer:2021:USM

[dBtMV21] Bas de Boer, Hedwig te Molder, and Peter-Paul Verbeek. Understanding science-in-the-making by letting scientific instruments speak: From semiotics to postphenomenology. Social Studies of Science, 51(3):392–413, June 1, 2021. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720981600.

Dieks:1976:DIS

[DC76] Dennis Dieks and Hans Chang. Differences in impact of scientific publications: Some indices derived from a citation analysis. Social Studies of Science, 6(2):247–267, May 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277600600204.

deCheveigne:2009:CPW

[dC09] Suzanne de Cheveigné. The career paths of women (and men) in French research. Social Studies of Science, 39(1): 113–136, February 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312708097656.

deCamargo:2002:TSP

[de 02] Kenneth Rochel de Camargo, Jr. The thought style of physicians: Strategies for keeping up with medical knowledge. Social Studies of Science, 32(5–6):827–855, December 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631270203200501.

DeWolff:2024:FTM

[De 24] Kim De Wolff. ‘Floating things’ and methodological drift: Accounting for haunted materialities in the North Pacific Ocean. Social Studies of Science, 54(4):536–556, August 1, 2024. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127241226829.

Dean:1974:RSP

[Dea74] Genevieve Dean. Review: Science and politics in China: Reflections on one hundred thousand questions. Social Studies of Science, 4(1):93–96, January 1, 1974. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277400400106.

Dear:1992:TDS

[Dea92] Peter Dear. From truth to disinterestedness in the Seventeenth Century. Social Studies of Science, 22(4):619–631, November 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631292022004003.

Dear:1995:BRW

[Dea95] Peter Dear. Book review: William Eamon, Science and the Secrets of Nature: Books of Secrets in Medieval and Early Modern Culture. Social Studies of Science, 25(2):388–393, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631295025002020.

Dear:2012:FYS

[Dea12] Peter Dear. Fifty years of structure. Social Studies of Science, 42(3):424–428, June 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712437237.

DelSesto:1983:UKV

[Del83] Steven L. Del Sesto. Uses of knowledge and values in technical controversies: The case of nuclear reactor safety in the US. Social Studies of Science, 13(3):395–416, August 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631283013003003.

Delamont:1987:TBS

[Del87] Sara Delamont. Three blind spots? A comment on the sociology of science by a puzzled outsider. Social Studies of Science, 17 (1):163–170, February 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631287017001008.

Delamont:1993:RTW

[Del93] Sara Delamont. Roads not taken, or ‘the way through the woods’. Social Studies of Science, 23(1):217–222, February 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631293023001008.

Delamont:2002:HRF

[Del02] Sara Delamont. Hypatia’s revenge?: Feminist perspectives in S&TS. Muriel Lederman and Ingrid Bartsch (eds), The Gender and Science Reader (London & New York: Routledge, 2001), 505 pp., £18.99 / EUR 32.49 / $29.95 (pbk). ISBN 0-415-21358-4 (pbk). Maralee Mayberry, Banu Subramaniam and Lisa H. Weasel (eds), Feminist Science Studies: A New Generation (London & New York: Routledge, 2001), 354 pp., £15.99 / EUR 27.35 / $23.95 (pbk). ISBN 0-415-92696-3 (pbk). Mary Wyer, Donna Cookmeyer, Mary Barbercheck, Hatice Ozturk and Marta Wayne (eds), Women, Science and Technology: A Reader in Feminist Science Studies (London & New York: Routledge, 2001), 376pp., £15.99 / EUR 32.41 / $27.95 (pbk). ISBN 0-415-92607-6 (pbk). Social Studies of Science, 32(1): 167–174, February 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312702032001007.

Delamont:2003:R

[Del03] Sara Delamont. Review. Social Studies of Science, 33(2): 315–322, April 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/03063127030332019.

Delamont:2005:BRL

[Del05] Sara Delamont. Book reviews: Lives of the great women scientists: The never ending story? Nathaniel C. Comfort (2001) The Tangled Field: Barbara McClintock’s Search for the Patterns of Genetic Control (Cambridge, MA: Harvard University Press), x + 337 pp., £9.00 / EUR 15.95 / $17.95 (pbk), £26.95 / EUR 36.50 / $41.00 (hbk). ISBN 0-674-01108-2 (pbk), 0-674-00456-6 (hbk). Kathryn A. Neeley (2001) Mary Somerville: Science, Illumination and the Female Mind (Cambridge: Cambridge University Press), xiii + 263 pp., £17.99 / EUR 22.59 / $24.99 (pbk), £39.02 / EUR 61.90 / $70.00 (hbk). ISBN 0-521-62672-2 (pbk), 0-521-62299-9 (hbk). Elga Wasserman (2000) The Door in the Dream: Conversations with Eminent Women in Science (Washington, DC: Joseph Henry Press), xiv + 254 pp., £10.01 / 17.95 / $19.95 (pbk). ISBN 0-309-08619-1 (pbk). Leigh Ann Whaley (2003) Women’s History as Scientists: A Guide to the Debates (Santa Barbara, CA. Social Studies of Science, 35(3):491–496, June 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312705054517.

Delborne:2008:TTS

[Del08] Jason A. Delborne. Transgenes and transgressions: Scientific dissent as heterogeneous practice. Social Studies of Science, 38 (4):509–541, August 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312708089716.

Delfanti:2016:BPP

[Del16] Alessandro Delfanti. Beams of particles and papers: How digital preprint archives shape authorship and credit. Social Studies of Science, 46(4):629–645, August 1, 2016. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716659373.

Delfanti:2021:FMI

[Del21] Alessandro Delfanti. The financial market of ideas: a theory of academic social media. Social Studies of Science, 51(2):259–276, April 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312720966649.

Delgado:2023:RSF

[Del23] Abigail Nieves Delgado. Race and statistics in facial recognition: Producing types, physical attributes, and genealogies. Social Studies of Science, 53(6):916–937, December 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127221127666.

Demortain:2024:HSB

[Dem24] David Demortain. How scientists become experts-or don’t: Social organization of research and engagement in scientific advice in a toxicology laboratory. Social Studies of Science, 54 (3):405–428, June 1, 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127231204578.

Dennis:1985:DDM

[Den85] Michael Aaron Dennis. Drilling for dollars: The making of US petroleum reserve estimates, 1921–25. Social Studies of Science, 15(2):241–265, May 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631285015002002.

Dennis:1987:ARN

[Den87] Michael Aaron Dennis. Accounting for research: New histories of corporate laboratories and the social history of American science. Social Studies of Science, 17(3):479–518, August 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631287017003005.

Dennis:2001:CM

[Den01] Michael Aaron Dennis. Count me out. Social Studies of Science, 31(3):436–439, June 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631201031003009.

Dennis:2003:EMC

[Den03] Michael Aaron Dennis. Earthly matters: On the Cold War and the Earth sciences. Social Studies of Science, 33(5):809–819, October 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312703335007.

Derksen:2000:TSM

[Der00] Linda Derksen. Towards a sociology of measurement: The meaning of measurement error in the case of DNA profiling. Social Studies of Science, 30(6):803–845, December 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030006001.

DeWitt:1975:BTP

[DeW75] Larry W. DeWitt. On Bloor’s transformation of Popper’s pluralism. Social Studies of Science, 5(2):201–209, May 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277500500206.

deGomezgil:1975:MAI

[dG75] Maria Luisa Rodriguez Sala de Gómezgil. Mexican adolescents’ image of the scientist. Social Studies of Science, 5(3):

355–361, August 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277500500306.

Denis:2017:RCG

[DG17] Jérôme Denis and Samuel Goëta. Rawification and the careful generation of open government data. Social Studies of Science, 47(5):604–629, October 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312717712473.

Didier:2018:NWV

[DGC18] Emmanuel Didier and Catherine Guaspare-Cartron. The new watchdogs’ vision of science: a roundtable with Ivan Oransky (Retraction Watch) and Brandon Stell (PubPeer). Social Studies of Science, 48(1):165–167, February 1, 2018. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312718756202.

Dunbar-Hester:2008:GMG

[DH08] Christina Dunbar-Hester. Geeks, meta-geeks, and gender trouble: Activism, identity, and low-power FM radio. Social Studies of Science, 38(2):201–232, April 1, 2008. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312707082954.

Davies:2015:CGC

[DH15] Sarah R. Davies and Maja Horst. Crafting the group: Care in research management. Social Studies of Science, 45(3):371–393, June 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312715585820.

Decker:2023:TST

[DH23] Kris Decker and Christoph Hoffmann. Training scenes: Taking science studies to the classroom. Social Studies of Science, 53(3):402–426, June 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127231164583.

Dickson:2000:SPN

[Dic00] David Dickson. Science and its public: The need for a ‘third way’. Social Studies of Science, 30(6):917–923, December 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631200030006004.

Didier:2025:NEG

[Did25] Emmanuel Didier. Numbers and emotions in the governance of the Covid-19 datademic. Social Studies of Science, 55 (2):238–259, April 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127241262457.

Dirk:1999:MOE

[Dir99] Lynn Dirk. A measure of originality: The elements of science. Social Studies of Science, 29(5):765–776, October 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029005004.

Disco:2004:TFW

[Dis04] Cornelis Disco. ’tis folly to be wise. Social Studies of Science, 34 (1):133–138, February 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312704038745.

Divall:1990:MAE

[Div90] Colin Divall. A measure of agreement: Employers and engineering studies in the universities of England and Wales, 1897– 1939. Social Studies of Science, 20(1):65–112, February 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020001003.

Dix:2019:MFC

[Dix19] Guus Dix. Microeconomic forecasting: Constructing commensurable futures of educational reforms. Social Studies of Science, 49(2):180–207, April 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312719837364.

Driessen:2012:PTV

[DK12] Clemens Driessen and Michiel Korthals. Pig towers and in vitro meat: Disclosing moral worlds by design. Social Studies of Science, 42(6):797–820, December 1, 2012. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712457110.

delaBellacasa:2011:MCT

[dlB11] Maria Puig de la Bellacasa. Matters of care in technoscience: Assembling neglected things. Social Studies of Science, 41(1): 85–106, February 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312710380301.

delaBellacasa:2015:MTS

[dlB15] Maria Puig de la Bellacasa. Making time for soil: Technoscientific futurity and the pace of care. Social Studies of Science, 45 (5):691–716, October 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312715599851.

deLaet:2021:TAA

[dLDV21] Marianne de Laet, Annelieke Driessen, and Else Vogel. Thinking with attachments: Appreciating a generative analytic. Social Studies of Science, 51(6):799–819, December 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/03063127211048804.

deLaet:2000:ZBP

[dLM00] Marianne de Laet and Annemarie Mol. The Zimbabwe bush pump: Mechanics of a fluid technology. Social Studies of Science, 30(2):225–263, April 1, 2000. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030002002.

deMello:1998:SIC

[dMdF98] José Manoel Carvalho de Mello and Carlos Machado de Freitas. Social interests, contextualizations and uncertainties in risk assessment: The case of methanol as a fuel component in Brazil. Social Studies of Science, 28(3):401–421, June 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631298028003002.

Dean:2008:DAS

[DNTS08] Katrina Dean, Simon Naylor, Simone Turchetti, and Martin Siegert. Data in Antarctic science and politics. Social Studies of Science, 38(4):571–604, August 1, 2008. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312708090693.

Doel:2003:CPE

[Doe03] Ronald E. Doel. Constituting the postwar Earth sciences: The military’s influence on the environmental sciences in the USA after 1945. Social Studies of Science, 33(5):635–666, October 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312703335002.

Doing:2004:LHS

[Doi04] Park Doing. ‘Lab hands’ and the ‘Scarlet O’: Epistemic politics and (scientific) labor. Social Studies of Science, 34(3):299–323, June 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312704043677.

Doing:2011:RET

[Doi11] Park Doing. Review essay: Tacit knowledge: Discovery by or topic for science studies?: Michael Polanyi, The Tacit Dimension (University of Chicago Press, 2009), 128 pp., £11.00 / EUR 24.77 / $16.00. ISBN 978-0-226-67298-4 (pbk). Harry Collins, Tacit and Explicit Knowledge (University of Chicago Press, 2010), xiii + 182 pp., £21.00 / EUR 26.03 / $32.50. ISBN 978-0-226-11380-7 (hbk). Social Studies of Science, 41 (2):301–306, April 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312710397690.

Dolby:1971:SKN

[Dol71] R. G. A. Dolby. Sociology of knowledge in natural science. Social Studies of Science, 1(1):3–21, January 1, 1971. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277100100102.

Dolby:1974:DSC

[Dol74] R. G. A. Dolby. In defence of a social criterion of scientific objectivity. Social Studies of Science, 4(2):187–190, April 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277400400204.

Dolby:1975:WCW

[Dol75] R. G. A. Dolby. What can we usefully learn from the Velikovsky affair? Social Studies of Science, 5(2):165–175, May 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277500500203.

Donnelly:1986:RAS

[Don86] J. F. Donnelly. Representations of applied science: Academics and chemical industry in Late Nineteenth-century England. Social Studies of Science, 16(2):195–234, May 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312786016002001.

Doran:1989:GR

[Dor89a] Chris Doran. Grasping reflexivity. Social Studies of Science, 19 (4):755–759, November 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631289019004015.

Doran:1989:JFR

[Dor89b] Chris Doran. Jumping frames: Reflexivity and recursion in the sociology of science. Social Studies of Science, 19(3): 515–531, August 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631289019003006.

Downey:1988:RCI

[Dow88] Gary L. Downey. Reproducing cultural identity in negotiating nuclear power: The Union of Concerned Scientists and emergency core cooling. Social Studies of Science, 18(2):231–264, May 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631288018002003.

Downer:2007:WCH

[Dow07a] John Downer. When the chick hits the fan: Representativeness and reproducibility in technological tests. Social Studies of Science, 37(1):7–26, February 1, 2007. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706064235.

Downey:2007:PPT

[Dow07b] Greg Downey. Producing pain: Techniques and technologies in no-holds-barred fighting. Social Studies of Science, 37 (2):201–226, April 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706072174.

Denis:2021:MEP

[DP21] Jérôme Denis and David Pontille. Maintenance epistemology and public order: Removing graffiti in Paris. Social Studies of Science, 51(2):233–258, April 1, 2021. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720956720.

Davies:2023:LSR

[DP23] Sarah R. Davies and Bao-Chau Pham. Luck and the ‘situations’ of research. Social Studies of Science, 53(2):287–299, April 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127221125438.

Danesi:2020:TSR

[DPP^{+}20] Giada Danesi, Mélody Pralong, Francesco Panese, Bernard Burnand, and Michèle Grossen. Techno-social reconfigurations in diabetes (self-) care. Social Studies of Science, 50(2):198–220, April 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312720903493.

Dreyfus:1992:RCA

[Dre92] Hubert L. Dreyfus. Response to Collins, artificial experts. Social Studies of Science, 22(4):717–726, November 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631292022004007.

Drori:1993:RBS

[Dro93] Gili S. Drori. The relationship between science, technology and the economy in lesser developed countries. Social Studies of Science, 23(1):201–215, February 1, 1993. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631293023001007.

Danyi:2020:CGT

[DS20] Endre Dányi and Michaela Spencer. Un/common grounds: Tracing politics across worlds. Social Studies of Science, 50 (2):317–334, April 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312720909536.

Dupras:2019:EEL

[DSJ19] Charles Dupras, Katie Michelle Saulnier, and Yann Joly. Epigenetics, ethics, law and society: a multidisciplinary review of descriptive, instrumental, dialectical and reflexive analyses. Social Studies of Science, 49(5):785–810, October 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312719866007.

Price:1971:PPF

[dSP71] Derek J. de Solla Price. Principles for projecting funding of academic science in the 1970s. Social Studies of Science, 1 (1):85–94, January 1, 1971. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277100100106.

Price:1975:RNS

[dSP75] Derek de Solla Price. Review: Nakayama Shigeru, David L. Swain, Yagi Eri (eds.), Science and Society in Modern Japan: Selected Historical Sources (Tokyo: University of Tokyo Press, and Boston: M.I.T. Press), 5700 Yen (£10.00 UK). Social Studies of Science, 5(2):211–212, May 1, 1975. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277500500207.

Decoteau:2015:ANK

[DU15] Claire Laurier Decoteau and Kelly Underman. Adjudicating non-knowledge in the omnibus autism proceedings. Social Studies of Science, 45(4):471–500, August 1, 2015. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312715600278.

Dugan:1987:HSN

[Dug87] Kathy Dugan. History of science in non–Western classrooms: a bridge between cultures. Social Studies of Science, 17(1): 145–161, February 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631287017001007.

Dugdale:1992:GNS

[Dug92] Ann Dugdale. Gender and the new sociology of technology. Social Studies of Science, 22(4):759–762, November 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631292022004012.

Duncan:1974:ISD

[Dun74] Simon S. Duncan. The isolation of scientific discovery: Indifference and resistance to a new idea. Social Studies of Science, 4(2):109–134, April 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277400400201.

Dunlap:1978:SGR

[Dun78] Thomas R. Dunlap. Science as a guide in regulating technology: The case of DDT in the United States. Social Studies of Science, 8(3):265–285, August 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277800800302.

Dunford:1985:PRC

[Dun85] Richard W. Dunford. The problem of relevant collectivities: Solar energy research in Australia. Social Studies of Science, 15 (3):455–474, August 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631285015003003.

Dunlap:1991:OWP

[Dun91] Thomas R. Dunlap. Organization and wildlife preservation: The case of the whooping crane in North America. Social Studies of Science, 21(2):197–221, May 1, 1991. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631291021002002.

Dunlop:2022:EEE

[Dun22] Tessa Dunlop. Energy efficiency: The evolution of a motherhood concept. Social Studies of Science, 52(5):710–732, October 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127221096171.

Durant:2011:MDS

[Dur11] Darrin Durant. Models of democracy in social studies of science. Social Studies of Science, 41(5):691–714, October 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312711414759.

Durant:2020:ILH

[Dur20] Jennie L. Durant. Ignorance loops: How non-knowledge about bee-toxic agrochemicals is iteratively produced. Social Studies of Science, 50(5):751–777, October 1, 2020. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720923390.

deVries:2007:WPS

[dV07] Gerard de Vries. What is political in sub-politics?: How Aristotle might help STS. Social Studies of Science, 37(5): 781–809, October 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706070749. See note

[Lat07].

Duque:2005:CPS

[DYS^{+}05] Ricardo B. Duque, Marcus Ynalvez, R. Sooryamoorthy, Paul Mbatia, Dan-Bright S. Dzorgbo, and Wesley Shrum. Collaboration paradox: Scientific productivity, the Internet, and problems of research in developing areas. Social Studies of Science, 35(5):755–785, October 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312705053048.

Easlea:1973:IHS

[Eas73] Brian Easlea. An introduction to the history and social studies of science: a seminar course for first-year science students. Social Studies of Science, 3(2):185–209, April 1, 1973. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277300300206.

Easlea:1974:WNL

[Eas74] Brian Easlea. Who needs the liberation of nature? Social Studies of Science, 4(1):77–92, January 1, 1974. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277400400105.

Eden:2006:RMC

[Ede06] Lynn Eden. Response to my critics. Social Studies of Science, 36(4):646–656, August 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706064818.

Edge:1991:Ea

[Edg91a] David Edge. Editorial. Social Studies of Science, 21(1):5, February 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631291021001001.

Edge:1991:Eb

[Edg91b] David Edge. Editorial. Social Studies of Science, 21(2):195–196, May 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631291021002001.

Edge:1992:E

[Edg92] David Edge. Editorial. Social Studies of Science, 22(2):211–212, May 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631292022002001.

Edge:1994:Ea

[Edg94a] David Edge. Editorial. Social Studies of Science, 24(1):5–6, February 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631279402400101.

Edge:1994:Eb

[Edg94b] David Edge. Editorial. Social Studies of Science, 24(2):195, May 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631279402400201.

Edge:1995:Ea

[Edg95a] David Edge. Editorial. Social Studies of Science, 25(1):5–8, February 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631295025001001.

Edge:1995:Eb

[Edg95b] David Edge. Editorial. Social Studies of Science, 25(4): 611–612, November 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631295025004001.

Edge:1995:Ec

[Edg95c] David Edge. Editorial. Social Studies of Science, 25(4): 884, November 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631295025004014.

Edge:1995:EN

[Edg95d] David Edge. Editor’s notes. Social Studies of Science, 25(2):400, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631295025002023.

Edge:1996:Ea

[Edg96a] David Edge. Editorial. Social Studies of Science, 26(1):5–6, February 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631296026001001.

Edge:1996:Eb

[Edg96b] David Edge. Editorial. Social Studies of Science, 26(4): 723–732, November 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631296026004001.

Edge:1997:BM

[Edg97a] David Edge. Beam and mote. Social Studies of Science, 27 (2):357–359, April 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631297027002008.

Edge:1997:E

[Edg97b] David Edge. Editorial. Social Studies of Science, 27(1):5–6, February 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631297027001001.

Edge:1997:R

[Edg97c] David Edge. For the record. Social Studies of Science, 27 (3):543–544, June 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631297027003008.

Edge:1999:E

[Edg99a] David Edge. Editorial. Social Studies of Science, 29(1):5–6, February 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631299029001001.

Edge:1999:EP

[Edg99b] David Edge. Editorial postscript. Social Studies of Science, 29 (5):790–799, October 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631299029005007.

Edge:2001:Ea

[Edg01a] David Edge. Editorial. Social Studies of Science, 31(1):5–6, February 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631201031001001.

Edge:2001:Eb

[Edg01b] David Edge. Editorial. Social Studies of Science, 31(2):163, April 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312701031002001.

Edge:2003:CSA

[Edg03] David Edge. Celebration and strategy: The 4S after 25 years, and STS after 9-11. Social Studies of Science, 33(2):161–169, April 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/03063127030332002.

Edmond:2002:LEC

[Edm02] Gary Edmond. Legal engineering: Contested representations of law, science (and non-science) and society. Social Studies of Science, 32(3):371–412, June 1, 2002. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312702032003002.

Edmond:2011:REB

[Edm11] Gary Edmond. Review essay: The building blocks of forensic science and law: Recent work on DNA profiling (and photo comparison). Social Studies of Science, 41(1):127–152, February 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312710379453.

Edwards:1994:HTH

[Edw94] Paul N. Edwards. Hyper text and hypertension: Poststructuralist critical theory, social studies of science and software. Social Studies of Science, 24(2):229–278, May 1, 1994.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631279402400203.

Erikson:2014:TMC

[EE14] Martin G. Erikson and Peter Erlandson. A taxonomy of motives to cite. Social Studies of Science, 44(4):625–637, August 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312714522871.

Elzen:1996:STN

[EES96] Boelie Elzen, Bert Enserink, and Wim A. Smit. Socio-technical networks: How a technology studies approach May help to solve problems related to technical change. Social Studies of Science, 26(1):95–141, February 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631296026001006.

Efmertova:1992:CPJ

[Efm92] Marcela Efmertová. Czech physicist Jaroslav ŠSafránek and his television. Social Studies of Science, 22(2):283–300, May 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631292022002006.

Eglash:1997:AHB

[Egl97] Ron Eglash. The African heritage of Benjamin Banneker. Social Studies of Science, 27(2):307–315, April 1, 1997. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631297027002004.

Edge:1997:TKJ

[EHB^{+}97] David Edge, Rom Harré, Andrew Brown, Barry Barnes, Michael Mulkay, Steve Fuller, Martin Rudwick, Ronald N. Giere, and David Bloor. Thomas S. Kuhn (18 July 1922–17 June 1996). Social Studies of Science, 27(3):483–502, June 1, 1997. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631297027003005.

Edge:2002:E

[EL02] David Edge and Michael Lynch. Editorial. Social Studies of Science, 32(1):5–6, February 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/0306312702032001001.

Elgabsi:2022:EKR

[Elg22] Natan Elgabsi. The ‘ethic of knowledge’ and responsible science: Responses to genetically motivated racism. Social Studies of Science, 52(2):303–323, April 1, 2022. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/03063127211063887.

Elichirigoity:2000:FRE

[Eli00] Fernando Elichirigoity. On failing to reach escape velocity beyond modernity. Social Studies of Science, 30(1):145– 150, February 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631200030001007. See re-

sponse [Fee00].

Elkana:1978:TTT

[Elk78] Yehuda Elkana. Two-tier-thinking: Philosophical realism and historical Relativism. Social Studies of Science, 8(3):309–326, August 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631277800800304.

Elliott:1975:ASA

[Ell75] Clark A. Elliott. The American scientist in Antebellum society: a quantitative view. Social Studies of Science, 5(1): 93–108, February 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277500500107.

Ellis:2011:JJP

[Ell11] Rebecca Ellis. Jizz and the joy of pattern recognition: Virtuosity, discipline and the agency of insight in UK naturalists’ arts of seeing. Social Studies of Science, 41(6):769–790, December 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312711423432.

Evans:2021:STS

[ELR21] Sam Weiss Evans, Matthias Leese, and Dagmar Rychnovská. Science, technology, security: Towards critical collaboration. Social Studies of Science, 51(2):189–213, April 1, 2021. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720953515.

Elzinga:1980:SSS

[Elz80] Aant Elzinga. ‘Science studies’ in Sweden. Social Studies of Science, 10(2):181–214, May 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631278001000204.

Elzen:1986:TUC

[Elz86] Boelie Elzen. Two ultracentrifuges: a comparative study of the social construction of artefacts. Social Studies of Science, 16 (4):621–662, November 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631286016004004.

Elzinga:2003:MDO

[Elz03] Aant Elzinga. In memoriam: David Owen Edge. Social Studies of Science, 33(2):190, April 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127030332008.

Elzinga:2004:BRM

[Elz04] Aant Elzinga. Book review: Making Science and Technology Studies Relevant for Technology Policy — Gains and Losses? Social Studies of Science, 34(6):949–956, December 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704046365.

Edge:1973:EN

[EM73] David Edge and Roy MacLeod. Editorial note. Social Studies of Science, 3(1):1, January 1, 1973. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277300300101.

Edge:1975:E

[EM75] David O. Edge and Roy M. MacLeod. Editorial. Social Studies of Science, 5(1):3–4, February 1, 1975. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277500500101.

Edge:1978:E

[EM78] David Edge and Roy MacLeod. Editorial. Social Studies of Science, 8(3):263, August 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277800800301.

Edge:1979:E

[EM79] David Edge and Roy MacLeod. Editorial. Social Studies of Science, 9(4):401–402, November 1, 1979. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277900900401.

Edge:1982:EN

[EM82] David Edge and Roy MacLeod. Editors’ note. Social Studies of Science, 12(1):i–ii, February 1, 1982. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631282012001001.

Edge:1983:E

[EM83a] David Edge and Roy MacLeod. Editorial. Social Studies of Science, 13(1):i–ii, February 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631283013001002.

Edge:1983:EN

[EM83b] David Edge and Roy MacLeod. o editorial note. Social Studies of Science, 13(4):i, November 1, 1983. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631283013004001.

Edge:1984:EN

[EM84] David Edge and Roy MacLeod. Editorial note. Social Studies of Science, 14(1):i, February 1, 1984. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014001001.

Edge:1985:E

[EM85] David Edge and Roy MacLeod. Editorial. Social Studies of Science, 15(4):591, November 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631285015004001.

Edge:1986:E

[EM86] David Edge and Roy MacLeod. Editorial. Social Studies of Science, 16(1):3–8, February 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631286016001001.

Edge:1987:E

[EM87] David Edge and Roy MacLeod. Editorial. Social Studies of Science, 17(4):579, November 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631287017004001.

Edge:1988:Ea

[EM88a] David Edge and Roy Macleod. Editorial. Social Studies of Science, 18(2):195–196, May 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631288018002001.

Edge:1988:Eb

[EM88b] David Edge and Roy MacLeod. Editorial. Social Studies of Science, 18(4):579, November 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631288018004001.

Edmond:2000:LLL

[EM00] Gary Edmond and David Mercer. Litigation life: Law-science knowledge construction in (Bendectin) mass toxic tort litigation. Social Studies of Science, 30(2):265–316, April 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030002003.

Edmond:2006:ASE

[EM06] Gary Edmond and David Mercer. Anti-social epistemologies. Social Studies of Science, 36(6):843–853, December 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706067900.

Edwards:2011:SFD

[EMB^{+}11] Paul N. Edwards, Matthew S. Mayernik, Archer L. Batcheller, Geoffrey C. Bowker, and Christine L. Borgman. Science friction: Data, metadata, and collaboration. Social Studies of Science, 41(5):667–690, October 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312711413314.

Engelmann:2023:DMC

[EMSL23] Lukas Engelmann, Catherine M. Montgomery, Steve Sturdy, and Cristina Moreno Lozano. Domesticating models: On the contingency of Covid-19 modelling in UK media and policy. Social Studies of Science, 53(1):121–145, February 1, 2023. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

Engel:2020:ADV

[Eng20] Nora Engel. Aligning in the dark: Variable and shifting (user- ) settings in developing point-of-care diagnostics for tuberculosis and HIV. Social Studies of Science, 50(1):50–75, February 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312719900545.

Ensmenger:2012:CDA

[Ens12] Nathan Ensmenger. Is chess the Drosophila of artificial intelligence? A social history of an algorithm. Social Studies of Science, 42(1):5–30, February 1, 2012. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312711424596.

Evans:2007:LPR

[EP07] Robert Evans and Alexandra Plows. Listening without prejudice? Re-discovering the value of the disinterested citizen. Social Studies of Science, 37(6):827–853, December 1, 2007.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312707076602.

Epstein:1997:ADR

[Eps97] Steven Epstein. Activism, drug regulation, and the politics of therapeutic evaluation in the AIDS era: a case study of ddC and the ‘surrogate markers’ debate. Social Studies of Science, 27(5):691–726, October 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631297027005001.

Epstein:1998:HDS

[Eps98] Steven Epstein. History and diagnosis of ‘scientific’ medicine. Social Studies of Science, 28(3):489–495, June 1, 1998. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631298028003006.

Epstein:2008:RRC

[Eps08] Steven Epstein. The rise of ‘recruitmentology’: Clinical research, racial knowledge, and the politics of inclusion and difference. Social Studies of Science, 38(5):801–832, October 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312708091930.

Epstein:2009:RHM

[Eps09] Steven Epstein. Review: Health, materiality, and the politics of office work: Michelle Murphy, Sick Building Syndrome and the Problem of Uncertainty: Environmental Politics, Technoscience, and Women Workers (Duke University Press, 2006), x + 253 pp., £15.99/$22.95 (pbk), £58.00 / $79.95 (hbk). ISBN 0-8223-3671-5 (pbk), 0-8223-3659-6 (hbk). Social Studies of Science, 39(6):953–956, December 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312709345036.

Epstein:2021:CCP

[Eps21] Steven Epstein. Cultivated co-production: Sexual health, human rights, and the revision of the ICD. Social Studies of Science, 51(5):657–682, October 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
03063127211014283.

Eisemon:1978:SBS

[ER78] Thomas O. Eisemon and Yakov M. Rabkin. Science in a bilingual society: The case of two engineering schools in Quebec. Social Studies of Science, 8(2):245–256, May 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277800800205.

Erickson:1971:MIC

[Eri71]  J. Erickson.  The military-industrial complex.  Social

Studies of Science, 1(2):225–233, April 1, 1971. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277100100207.

Erickson:1972:RLA

[Eri72] John Erickson. Radio-location and the air defence problem: The design and development of Soviet radar 1934–40. Social Studies of Science, 2(3):241–263, July 1, 1972. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277200200302.

Eriksson:2004:PSS

[Eri04] Lena Eriksson. The presentation of the scientific self. Social Studies of Science, 34(3):423–426, June 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704043691.

Sugawara:2023:MPT

[eS23] Shin etsu Sugawara. The multistability of predictive technology in nuclear disasters. Social Studies of Science, 53(4): 495–521, August 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127231161609.

Evans:1997:SSF

[Eva97] Robert Evans. Soothsaying or science?: Falsification, uncertainty and social change in macroeconomic modelling. Social Studies of Science, 27(3):395–438, June 1, 1997. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631297027003002.

Evans:1999:CB

[Eva99] Robert Evans. A cause to believe in? Social Studies of Science, 29(1):145–150, February 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631299029001007.

Evans:2010:ICS

[Eva10] James A. Evans. Industry collaboration, scientific sharing, and the dissemination of knowledge. Social Studies of Science, 40 (5):757–791, October 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312710379931.

Evans:2022:SAP

[Eva22] Robert Evans. SAGE advice and political decision-making: ‘following the science’ in times of epistemic uncertainty. Social Studies of Science, 52(1):53–78, February 1, 2022. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/03063127211062586.

Ezrahi:1971:PRA

[Ezr71] Yaron Ezrahi. The political resources of American science. Social Studies of Science, 1(2):117–133, April 1, 1971. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277100100201.

Fangerau:2009:MIJ

[Fan09] Heiner Fangerau. From Mephistopheles to Isaiah: Jacques Loeb, technical biology and war. Social Studies of Science, 39(2):229–256, April 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312708101045.

Farrall:1975:CCS

[Far75] Lyndsay A. Farrall. Controversy and conflict in science: a case study — the English Biometric School and Mendel’s Laws. Social Studies of Science, 5(3):269–301, August 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277500500302.

Faulkner:1992:WSB

[Fau92] Wendy Faulkner. When is small beautiful in biotech? Social Studies of Science, 22(3):586–589, August 1, 1992. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312792022003008.

Faulkner:1998:EJA

[Fau98] Wendy Faulkner. Extraordinary journeys around ordinary technologies in ordinary lives. Social Studies of Science, 28 (3):484–489, June 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631298028003005.

Faulkner:2000:DHG

[Fau00] Wendy Faulkner. Dualisms, hierarchies and gender in engineering. Social Studies of Science, 30(5):759–792, October 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030005005.

Faulkner:2007:NBP

[Fau07] Wendy Faulkner. ‘Nuts and bolts and people’: Gender-troubled engineering identities. Social Studies of Science, 37(3):331–356, June 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312706072175.

Faulkner:2012:LPS

[Fau12] Alex Faulkner. Law’s performativities: Shaping the emergence of regenerative medicine through European Union legislation. Social Studies of Science, 42(5):753–774, October 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712446694.

Frame:1979:IRC

[FC79] J. Davidson Frame and Mark P. Carpenter. International research collaboration. Social Studies of Science, 9(4):481– 497, November 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277900900405.

Friese:2012:TBK

[FC12] Carrie Friese and Adele E. Clarke. Transposing bodies of knowledge and technique: Animal models at work in reproductive sciences. Social Studies of Science, 42(1):31–52, February 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312711429995.

Fujimura:2008:IRG

[FDR08] Joan H. Fujimura, Troy Duster, and Ramya Rajagopalan. Introduction: Race, genetics, and disease: Questions of evidence, matters of consequence. Social Studies of Science, 38 (5):643–656, October 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312708091926.

Feenberg:2000:WRP

[Fee00] Andrew Feenberg. Will the real post-human please stand up!: a response to Elichirigoity. Social Studies of Science, 30(1): 151–157, February 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631200030001008. See

[Eli00].

Fox:2011:WFC

[FFB11] Mary Frank Fox, Carolyn Fonseca, and Jinghui Bao. Work and family conflict in academic science: Patterns and predictors among women and men in research universities. Social Studies of Science, 41(5):715–735, October 1, 2011. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312711417730.

Felt:2008:VVG

[FFMW08] Ulrike Felt, Maximilian Fochler, Astrid Mager, and Peter Winkler. Visions and versions of governing biomedicine: Narratives on power structures, decision-making and public participation in the field of biomedical technology in the Austrian context. Social Studies of Science, 38(2):233–257, April 1, 2008.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312707083489.

French:1973:SNA

[FG73] Richard French and Michael Gross. A survey of North American graduate students in the history of science 1970–71. Social Studies of Science, 3(2):161–171, April 1, 1973. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277300300203.

Fox:1986:FUS

[FG86] Robert Fox and Anna Guagnini. The flexible University: Some historical reflexions on the analysis of education and the modern British economy. Social Studies of Science, 16(3):515–527, August 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631286016003005.

Fishman:2004:MDC

[Fis04] Jennifer R. Fishman. Manufacturing desire: The commodification of female sexual dysfunction. Social Studies of Science, 34(2):187–218, April 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312704043028.

Fisher:2010:BRA

[Fis10] Saul Fisher. Book review: Arriving at norms of scholarly judgment: Michele Lamont, How Professors Think: Inside the Curious World of Academic Judgment (Cambridge, MA: Harvard University Press, 2009), 330 pp., £20.95 / EUR 25.20 / $27.95. ISBN: 978-0-674-03266-8 (hbk). Social Studies of Science, 40 (3):485–491, June 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312710362276.

Fiske:2018:DHT

[Fis18] Amelia Fiske. Dirty hands: The toxic politics of denunciation. Social Studies of Science, 48(3):389–413, June 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312718781505.

Fitzgerald:1986:EAA

[Fit86] Deborah Fitzgerald. Exporting American agriculture: The Rockefeller Foundation in Mexico, 1943–53. Social Studies of Science, 16(3):457–483, August 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631286016003003.

Freudenthal:1988:LFR

[FL88] Gad Freudenthal and Ilana Löwy. Ludwik Fleck’s roles in society: a case study using Joseph Ben-David’s paradigm for a sociology of knowledge. Social Studies of Science, 18(4): 625–651, November 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631288018004003.

Fleck:1994:KER

[Fle94] James Fleck. Knowing engineers?: a response to Forsythe. Social Studies of Science, 24(1):105–113, February 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631279402400106. See [For93, For94].

Fadlon:1997:LSA

[FLE97] Judith Fadlon and Noah Lewin-Epstein. Laughter spreads: Another perspective on boundary crossing in the Benveniste Affair. Social Studies of Science, 27(1):131–141, February 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631297027001007. See comment [Pic97].

Fitzgerald:2014:AEP

[FLK^{+}14] Des Fitzgerald, Melissa M. Littlefield, Kasper J. Knudsen, James Tonks, and Martin J. Dietz. Ambivalence, equivocation and the politics of experimental knowledge: a transdisciplinary neuroscience encounter. Social Studies of Science, 44 (5):701–721, October 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312714531473.

Findlen:2017:BGR

[FM17] Paula Findlen and Hannah Marcus. The breakdown of Galileo’s Roman network: Crisis and community, ca. 1633.

Social Studies of Science, 47(3):326–352, June 1, 2017. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716676657.

Frame:1977:DWS

[FNC77] J. Davidson Frame, Francis Narin, and Mark P. Carpenter. The distribution of world science. Social Studies of Science, 7(4): 501–516, November 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277700700414.

Fuhrman:1986:DAR

[FO86] Ellsworth R. Fuhrman and Kay Oehler. Discourse analysis and reflexivity. Social Studies of Science, 16(2):293–307, May 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312786016002005. See response [Pot87] and re-

ply [FO87].

Fuhrman:1987:RRR

[FO87] Ellsworth R. Fuhrman and Kay Oehler. Reflexivity redux: Reply to Potter. Social Studies of Science, 17(1):177–181, February 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631287017001010. See [Pot87].

Folkers:2022:LCC

[FO22] Andreas Folkers and Sven Opitz. Low-carbon cows: From microbial metabolism to the symbiotic planet. Social Studies of Science, 52(3):330–352, June 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
03063127221077987.

Freeman:1971:GRD

[FOC^{+}71] C. Freeman, C. H. G. Oldham, C. M. Cooper, T. C. Sinclair, and B. G. Achilladelis. The goals of r & d in the 1970s. Social Studies of Science, 1(3–4):357–406, October 1, 1971. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277100100306.

Fochler:2016:BBA

[Foc16] Maximilian Fochler. Beyond and between academia and business: How Austrian biotechnology researchers describe hightech startup companies as spaces of knowledge production. Social Studies of Science, 46(2):259–281, April 1, 2016. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716629831.

Fischer:2020:RHC

[FÖP20] Björn Fischer, Britt Östlund, and Alexander Peine. Of robots and humans: Creating user representations in practice. Social Studies of Science, 50(2):221–244, April 1, 2020. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720905116.

Forge:1989:RTS

[For89] John Forge. A realistic theory of science? Social Studies of Science, 19(1):181–189, February 1, 1989. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019001009.

Forsythe:1993:EKC

[For93] Diana E. Forsythe. Engineering knowledge: The construction of knowledge in artificial intelligence. Social Studies of Science, 23(3):445–477, August 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312793023003002. See

response [Fle94] and reply [For94].

Forsythe:1994:SRC

[For94] Diana E. Forsythe. STS (re)constructs anthropology: a reply to Fleck. Social Studies of Science, 24(1):113–123, February 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631279402400107. See [For93, Fle94].

Fortescue:1999:BRP

[For99] Stephen Fortescue. Book review: Paul R. Josephson, New Atlantis Revisited: Akademgorodok, the Siberian City of Science. Social Studies of Science, 29(1):150–152, February 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631299029001008.

Fourez:1997:STL

[Fou97] Gérard Fourez. Scientific and technological literacy as a social practice. Social Studies of Science, 27(6):903–936, December 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631297027006003.

Fox:1983:PPA

[Fox83] Mary Frank Fox. Publication productivity among scientists: a critical review. Social Studies of Science, 13(2):285–305, May 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631283013002005.

Fox:1994:FSR

[Fox94] Nicholas Fox. Fabricating surgery: a response to Collins. Social Studies of Science, 24(2):347–354, May 1, 1994. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631279402400207.

Fox:2003:MDO

[Fox03] Mary Frank Fox. In memoriam: David Owen Edge. Social Studies of Science, 33(2):190–191, April 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127030332009.

Fox:2005:GFC

[Fox05] Mary Frank Fox. Gender, family characteristics, and publication productivity among scientists. Social Studies of Science, 35(1): 131–150, February 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312705046630.

Freeman:2003:OKP

[FPM03] Chris Freeman, Pari Patel, and Ben Martin. Obituary: Keith Pavitt. Social Studies of Science, 33(3):463–464, June 1, 2003.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127030333010.

Farris:1989:DPC

[FR89] Hilary Farris and Russell Revlin. The discovery process: a counterfactual strategy. Social Studies of Science, 19(3): 497–513, August 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631289019003005. See re-

sponse [Gor91a].

Farris:1991:RDS

[FR91] Hilary Farris and Russell Revlin. Rule discovery strategies: Falsification without disconfirmation (reply to Gorman). Social Studies of Science, 21(3):565–567, August 1, 1991. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631291021003008. See [Gor91a].

Fujimura:2011:DDU

[FR11] Joan H. Fujimura and Ramya Rajagopalan. Different differences: The use of ‘genetic ancestry’ versus race in biomedical human genetic research. Social Studies of Science, 41 (1):5–30, February 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312710379170.

Frankel:1976:COW

[Fra76] Eugene Frankel. Corpuscular optics and the wave theory of light: The science and politics of a revolution in physics. Social Studies of Science, 6(2):141–184, May 1, 1976. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277600600201.

Frame:1979:NER

[Fra79] J. Davidson Frame. National economic resources and the production of research in lesser developed countries. Social Studies of Science, 9(2):233–246, May 1, 1979. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277900900204.

Franklin:1988:TUQ

[Fra88] J. Jeffrey Franklin. Testing and using quantitative methods in science policy contexts: a response to Hicks. Social Studies of Science, 18(2):365–375, May 1, 1988. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631288018002007. See [Hic87, Hic88].

Francoeur:1997:FTD

[Fra97] Eric Francoeur. The forgotten tool: The design and use of molecular models. Social Studies of Science, 27(1):7–40, February 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631297027001002.

Freudenthal:1984:RSK

[Fre84] Gad Freudenthal. The role of shared knowledge in science: The failure of the constructivist programme in the sociology of science. Social Studies of Science, 14(2):285–295, May 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014002010.

Freudenthal:1990:SSF

[Fre90] Gad Freudenthal. Science studies in France: a sociological view. Social Studies of Science, 20(2):353–369, May 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020002007.

Freire:1992:CCP

[Fre92] Olival Freire, Jr. Comment on ‘The Crisis in Physics’. Social Studies of Science, 22(4):739–742, November 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/030631292022004009. See [Cro91] and reply [Cro92].

Fries:1984:ISD

[Fri84] Sylvia D. Fries. The ideology of science during the Nixon years: 1970–76. Social Studies of Science, 14(3):323–341, August 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631284014003001.

Frow:2012:DLS

[Fro12] Emma K. Frow. Drawing a line: Setting guidelines for digital image processing in scientific journal articles. Social Studies of Science, 42(3):369–392, June 1, 2012. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712444303.

Frohlich:2017:ITF

[Fro17] Xaq Frohlich. The informational turn in food politics: The US FDA’s nutrition label as information infrastructure. Social Studies of Science, 47(2):145–171, April 1, 2017. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716671223.

Fortun:1993:SLW

[FS93] M. Fortun and S. S. Schweber. Scientists and the legacy of World War II: The case of operations research (OR). Social Studies of Science, 23(4):595–642, November 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631293023004001.

Fox:2001:CYS

[FS01] Mary Frank Fox and Paula E. Stephan. Careers of young scientists: Preferences, prospects and realities by gender and field. Social Studies of Science, 31(1):109–122, February 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031001006.

Fausto-Sterling:2008:BBR

[FS08] Anne Fausto-Sterling. The bare bones of race. Social Studies of Science, 38(5):657–694, October 1, 2008. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312708091925.

Finn:2023:EGD

[FS23] Megan Finn and Katie Shilton. Ethics governance development: The case of the Menlo Report. Social Studies of Science, 53(3):315–340, June 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127231151708.

Falkenberg:2024:MLC

[FSF24] Ruth Falkenberg, Lisa Sigl, and Maximilian Fochler. From ‘making lists’ to conducting ‘well-rounded’ studies: Epistemic reorientations in soil microbial ecology. Social Studies of Science, 54(1):78–104, February 1, 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127231179700.

Fuerst:1982:RRD

[Fue82] John A. Fuerst. The role of reductionism in the development of molecular biology: Peripheral or central? Social Studies of Science, 12(2):241–278, May 1, 1982. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631282012002003.

Fuerst:1984:DMB

[Fue84] John A. Fuerst. The definition of molecular biology and the definition of policy: The role of the Rockefeller Foundation’s policy for molecular biology. Social Studies of Science, 14(2):225–237, May 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631284014002003. See reply [AA84].

Fujimura:1987:CDA

[Fuj87] Joan H. Fujimura. Constructing ‘do-able’ problems in cancer research: Articulating alignment. Social Studies of Science, 17(2):257–293, May 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631287017002003.

Fukushima:2016:CFB

[Fuk16] Masato Fukushima. Constructing failure in big biology: The socio-technical anatomy of Japan’s protein 3000 project. Social Studies of Science, 46(1):7–33, February 1, 2016. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312715612146.

Fuller:1989:BDV

[Ful89a] Steve Fuller. Back to Descartes? The very idea! Social Studies of Science, 19(2):357–360, May 1, 1989. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019002008.

Fuller:1989:CIV

[Ful89b] Steve Fuller. Of conceptual intersections and verbal collisions: Towards the routing of Slezak. Social Studies of Science, 19(4): 625–638, November 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631289019004004.

Fuller:1990:TSD

[Ful90] Steve Fuller. They shoot dead horses, don’t they?: Philosophical fear and sociological loathing in St. Louis. Social Studies of Science, 20(4):664–681, November 1, 1990. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020004005.

Fuller:1991:SSP

[Ful91] Steve Fuller. Simon says ‘put your foot in your mouth’. Social Studies of Science, 21(1):149–150, February 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631291021001010.

Fuller:1994:BCS

[Ful94a] Steve Fuller. Being civil with scientists: Reply to Wolpert and Weinberg. Social Studies of Science, 24(4):751–757, November 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631279402400408. See [Wol94].

Fuller:1994:BRC

[Ful94b] Steve Fuller. Book review: Can Science Studies be Spoken in a Civil Tongue? Social Studies of Science, 24(1):143–168, February 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631279402400112. See response [Wol94] and

reply [Ful94a].

Fuller:1995:PPR

[Ful95] Steve Fuller. From pox to pax?: Response to Labinger. Social Studies of Science, 25(2):309–314, May 1, 1995. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631295025002006. See [Lab95b].

Fuller:2006:STL

[Ful06] Steve Fuller. A step toward the legalization of science studies. Social Studies of Science, 36(6):827–834, December 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706067898.

Fullwiley:2008:BCR

[Ful08] Duana Fullwiley. The biologistical construction of race: ‘admixture’ technology and the new genetic medicine. Social Studies of Science, 38(5):695–735, October 1, 2008. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312708090796.

Fuller:2012:CKL

[Ful12] Steve Fuller. CSI: Kuhn and Latour. Social Studies of Science, 42(3):429–434, June 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712437228.

Ford:2017:ACE

[FW17] Heather Ford and Judy Wajcman. ‘Anyone can edit’, not everyone does: Wikipedia’s infrastructure and the gender gap. Social Studies of Science, 47(4):511–527, August 1, 2017. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312717692172.

Finkielsztein:2023:SMB

[FW23] Mariusz Finkielsztein and Izabela Wagner. The sense of meaninglessness in bureaucratized science. Social Studies of Science, 53(2):271–286, April 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127221117227.

Galka:2024:BPS

[Gal24] Jonathan M. Galka. Bohemia at the Pacific seabed: Archiving the future of deep-sea mining with the Interoceanmetal Joint Organization. Social Studies of Science, 54(5):678–705, October 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127231226423.

Gaponenko:1995:TRS

[Gap95] Nadezhda Gaponenko. Transformation of the research system in a transitional society: The case of Russia. Social Studies of Science, 25(4):685–703, November 1, 1995. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631295025004005.

Gardiner:1973:RWP

[Gar73a] J. Paul Gardiner. Review: Wintry prospects for Canadian industry and science policy, in a bleak climate. Social Studies of Science, 3(2):181–183, April 1, 1973. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277300300205.

Garvey:1973:MRD

[Gar73b] William D. Garvey. A machine-readable data bank on the communication behaviour of scientists and technologists. Social Studies of Science, 3(1):91–92, January 1, 1973. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277300300109.

Garrety:1997:SWA

[Gar97] Karin Garrety. Social worlds, actor-networks and controversy: The case of cholesterol, dietary fat and heart disease. Social Studies of Science, 27(5):727–773, October 1, 1997. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631297027005002.

Garroutte:1999:GSA

[Gar99] Eva Marie Garroutte. Getting serious about ‘interrogating representation’: an indigenous turn. Social Studies of Science, 29 (6):945–956, December 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631299029006006.

Garfield:2004:UUC

[Gar04] Eugene Garfield. The unintended and unanticipated consequences of Robert K. Merton. Social Studies of Science, 34 (6):845–853, December 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312704042087.

Gardner:2013:HDB

[Gar13] John Gardner. A history of deep brain stimulation: Technological innovation and the role of clinical assessment tools. Social Studies of Science, 43(5):707–728, October 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713483678.

Gascoigne:1984:MME

[Gas84] John Gascoigne. Mathematics and meritocracy: The emergence of the Cambridge Mathematical Tripos. Social Studies of Science, 14(4):547–584, November 1, 1984. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014004003.

Gaster:1990:ASC

[Gas90] Barak Gaster. Assimilation of scientific change: The introduction of molecular genetics into biology textbooks. Social Studies of Science, 20(3):431–454, August 1, 1990. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020003002.

Gascoigne:1992:HDS

[Gas92] Robert Gascoigne. The historical demography of the scientific community, 1450–1900. Social Studies of Science, 22(3): 545–573, August 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312792022003005.

Gascoigne:1995:ECS

[Gas95] John Gascoigne. The Eighteenth-Century scientific community: a prosopographical study. Social Studies of Science, 25 (3):575–581, August 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631295025003007.

Gerber:2009:SCI

[GB09] Theodore P. Gerber and Deborah Yarsike Ball. Scientists in a changed institutional environment: Subjective adaptation and social responsibility norms in Russia. Social Studies of Science, 39(4):529–567, August 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312709103477.

Gaughan:2016:UPG

[GB16] Monica Gaughan and Barry Bozeman. Using the prisms of gender and rank to interpret research collaboration power dynamics. Social Studies of Science, 46(4):536–558, August 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716652249.

Garcia-Deister:2015:PGP

[GDLB15] Vivette García-Deister and Carlos López-Beltrán. País de gordos/país de muertos: Obesity, death and nation in biomedical and forensic genetics in Mexico. Social Studies of Science, 45 (6):797–815, December 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312715608449.

Gillespie:1979:CRA

[GEJ79] Brendan Gillespie, Dave Eva, and Ron Johnston. Carcinogenic risk assessment in the United States and Great Britain: The case of aldrin/dieldrin. Social Studies of Science, 9(3):265–301, August 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277900900301.

Geltzer:2009:RDD

[Gel09] Anna Geltzer. Review: The dilemmas of drug testing. Social Studies of Science, 39(3):481–484, June 1, 2009. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312709103479.

Gerson:1995:BRJ

[Ger95] Elihu M. Gerson. Book review: John Law, Organizing Modernity. Social Studies of Science, 25(2):382–388, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631295025002019.

Gerovitch:2001:MMC

[Ger01] Slava Gerovitch. ‘Mathematical machines’ of the Cold War: Soviet computing, American cybernetics and ideological disputes in the early 1950s. Social Studies of Science, 31(2):253–287, April 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312701031002006.

Geuna:2003:ESG

[Geu03] Aldo Geuna. The economics of science: Good old wine in a new bottle. Social Studies of Science, 33(3):458–461, June 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/03063127030333009.

Glenna:2007:EOS

[GGJ07] Leland L. Glenna, Margaret A. Gollnick, and Stephen S. Jones. Eugenic opportunity structures: Teaching genetic engineering at US land-grant universities since 1911. Social Studies of Science, 37(2):281–296, April 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706066177.

Gieryn:1983:MIS

[GH83] Thomas F. Gieryn and Richard F. Hirsh. Marginality and innovation in science. Social Studies of Science, 13(1):87–106, February 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631283013001005.

Gieryn:1984:MRS

[GH84] Thomas F. Gieryn and Richard F. Hirsh. Marginalia: Reply to Simonton and Handberg. Social Studies of Science, 14 (4):624, November 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631284014004007. See

[Han84, Sim84].

Gross:2000:ACS

[GHR00] Alan G. Gross, Joseph E. Harmon, and Michael S. Reidy. Argument and 17th-century science: a rhetorical analysis with sociological implications. Social Studies of Science, 30(3):371–396, June 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631200030003002.

Gieryn:1982:LWW

[Gie82a] Thomas F. Gieryn. Not-last words: Worn-out dichotomies in the sociology of science (reply). Social Studies of Science, 12 (2):329–335, May 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631282012002009. See

[Gie82b, Kro82].

Gieryn:1982:RCP

[Gie82b] Thomas F. Gieryn. Relativist/constructivist programmes in the sociology of science: Redundance and retreat. Social Studies of Science, 12(2):279–297, May 1, 1982. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/030631282012002004. See [Kro82] and reply [Gie82a].

Giere:1989:CDH

[Gie89] Ronald N. Giere. Computer discovery and human interests. Social Studies of Science, 19(4):638–643, November 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019004005.

Giere:1991:SSH

[Gie91] Ronald N. Giere. Syntax, semantics and human interests. Social Studies of Science, 21(1):150–153, February 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631291021001011.

Giere:1992:CCS

[Gie92] Ronald N. Giere. The cognitive construction of scientific knowledge (response to Pickering). Social Studies of Science, 22(1): 95–107, February 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312792022001006.

Giebelhaus:1996:MKN

[Gie96] August W. Giebelhaus. Melvin Kranzberg (22 November 1917–6 December 1995). Social Studies of Science, 26(2):469–472, May 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631296026002011.

Gieryn:2004:MT

[Gie04] Thomas Gieryn. Merton, teacher. Social Studies of Science, 34 (6):859–861, December 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312704042088.

Giere:2006:BRW

[Gie06a] Ronald Giere. Book review: What Everyone Should Know About STS. Social Studies of Science, 36(3):489–493, June 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706058827.

Gieryn:2006:CTS

[Gie06b] Thomas F. Gieryn. City as truth-spot: Laboratories and field-sites in urban studies. Social Studies of Science, 36(1): 5–38, February 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312705054526.

Gilbert:1976:TRF

[Gil76] G. Nigel Gilbert. The transformation of research findings into scientific knowledge. Social Studies of Science, 6(3–4): 281–306, September 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277600600302.

Gilbert:1977:RP

[Gil77] G. Nigel Gilbert. Referencing as persuasion. Social Studies of Science, 7(1):113–122, February 1, 1977. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700112.

Gillmor:1986:FFK

[Gil86] C. Stewart Gillmor. Federal funding and knowledge growth in ionospheric physics, 1945–81. Social Studies of Science, 16(1): 105–133, February 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631286016001006.

Gillespie:2006:EPE

[Gil06] Tarleton Gillespie. Engineering a principle: ‘end-to-end’ in the design of the Internet. Social Studies of Science, 36(3):427–457, June 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312706056047.

Gingras:1997:NDN

[Gin97] Yves Gingras. The new dialectics of nature. Social Studies of Science, 27(2):317–334, April 1, 1997. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631297027002005. See comments [Pic99].

Gingras:1999:HMR

[Gin99] Yves Gingras. From the heights of metaphysics: A reply to Pickering. Social Studies of Science, 29(2):312–315, April 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631299029002016. See [Pic99].

Gingras:2010:RES

[Gin10] Yves Gingras. Review essay: Sociological reflexivity in action: Pierre Bourdieu, Sketch for a Self-analysis (Chicago, IL: University of Chicago Press, 2008), translated by Richard Nice, x + 118 pp., £12.99 / EUR 18.58 / $20.00 (pbk); £38.00 / $55.00 / EUR 57.20 (hbk). ISBN 0-226-06751-3 (pbk); 0-226-06747-5 (hbk). Social Studies of Science, 40(4):619–631, August 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312710370236.

Gibbons:1972:DOS

[GK72] Michael Gibbons and Philip King. The development of Ovonic switches: a case study of a scientific controversy. Social Studies of Science, 2(4):295–309, October 1, 1972. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277200200401.

Geissler:2016:FSS

[GK16a] P. Wenzel Geissler and Ann H. Kelly. Field station as stage: Re-enacting scientific work and life in Amani, Tanzania. Social Studies of Science, 46(6):912–937, December 1, 2016. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716650045.

Geissler:2016:HSL

[GK16b] P. Wenzel Geissler and Ann H. Kelly. A home for science: The life and times of tropical and polar field stations. Social Studies of Science, 46(6):797–808, December 1, 2016. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716680767.

Groeneveld:1975:AUS

[GKM75] Lyle Groeneveld, Norman Koller, and Nicholas C. Mullins. The advisers of the United States National Science Foundation. Social Studies of Science, 5(3):343–354, August 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277500500305.

Gjodsbol:2024:PCC

[GKSS24] Iben M. Gjødsbøl, Jeanette Bresson Ladegaard Knox, Lea Skovgaard, and Mette N. Svendsen. Population curation: The construction of mutual obligation between individual and state in Danish precision medicine. Social Studies of Science, 54 (6):883–906, December 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127241255971.

Gilbert:2021:SOD

[GL21] Thomas Krendl Gilbert and Andrew Loveridge. Subjectifying objectivity: Delineating tastes in theoretical quantum gravity research. Social Studies of Science, 51(1):73–99, February 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312720949691.

Gieryn:1978:ERS

[GM78] Thomas F. Gieryn and Robert K. Merton. Essay review: The sociological study of scientific specialties: David O. Edge and Michael J. Mulkay, Astronomy Transformed: The Emergence of Radio Astronomy in Britain (New York: Wiley-Interscience, 1976), 473 pp., $30.00 (£19.25). ISBN 0-471-23273-4. Gerard Lemaine, Roy MacLeod, Michael Mulkay and Peter Weingart (eds), Perspectives on the Emergence of Scientific Disciplines (The Hague and Paris: Mouton, and Chicago: Aldine, 1976), 281 pp., $22.50. ISBN 90-279-7743-7 (Mouton); 0-202-30284-9 (Aldine). Social Studies of Science, 8(2):257–261, May 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277800800206.

Gilbert:1982:WSB

[GM82] G. Nigel Gilbert and Michael Mulkay. Warranting scientific belief. Social Studies of Science, 12(3):383–408, August 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631282012003002. See comments [Hal88, PM89,

Hal89].

Giere:2003:DCW

[GM03] Ronald N. Giere and Barton Moffatt. Distributed cognition: Where the cognitive and the social merge. Social Studies of Science, 33(2):301–310, April 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127030332017.

Granja:2023:FDP

[GM23] Rafaela Granja and Helena Machado. Forensic DNA phenotyping and its politics of legitimation and contestation: Views of forensic geneticists in Europe. Social Studies of Science, 53(6):

850–868, December 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/0306312720945033.

Godin:1997:RHT

[God97] Benoît Godin. The rhetoric of a health technology: The microprocessor patient card. Social Studies of Science, 27(6): 865–902, December 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631297027006002.

Godin:1998:WPH

[God98] Benoît Godin. Writing performative history: The new New Atlantis? Social Studies of Science, 28(3):465–483, June 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631298028003004.

Godin:2007:ESG

[God07] Benoît Godin. From eugenics to scientometrics: Galton, Cattell, and men of science. Social Studies of Science, 37(5):691–728, October 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312706075338.

Godin:2015:MIW

[God15] Benoît Godin. Models of innovation: Why models of innovation are models, or what work is being done in calling them models? Social Studies of Science, 45(4):570–596, August 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312715596852.

Goldstone:1979:DEM

[Gol79] Jack A. Goldstone. A deductive explanation of the Matthew Effect in science. Social Studies of Science, 9(3):385–391, August 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277900900306.

Golden:1994:WJP

[Gol94] Gloria Y. Golden. On the way to Jupiter: Psychological dimensions of the Galileo Mission. Social Studies of Science, 24 (3):463–511, August 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631279402400302.

Golan:2004:ESW

[Gol04] Tal Golan. The emergence of the silent witness: The legal and medical reception of X-rays in the USA. Social Studies of Science, 34(4):469–499, August 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704045705.

Gomart:2002:MSE

[Gom02] Emilie Gomart. Methadone: Six effects in search of a substance. Social Studies of Science, 32(1):93–135, February 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312702032001005.

Goodwin:1995:SD

[Goo95] Charles Goodwin. Seeing in depth. Social Studies of Science, 25(2):237–274, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631295025002002.

Goodrum:2010:RSH

[Goo10] Matthew R. Goodrum. Review: The social history of research on an early human fossil: Marianne Sommer, Bones and Ochre: The Curious Afterlife of the Red Lady of Paviland (Cambridge, MA: Harvard University Press, 2007), 398 pp., £31.95 / EUR 36.90 / $42.00. ISBN 0-674-02499-0 (hbk). Social Studies of Science, 40(1):155–158, February 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312709349441.

Goodwin:2018:CCH

[Goo18] Dawn Goodwin. Cultures of caring: Healthcare ‘scandals’, inquiries, and the remaking of accountabilities. Social Studies of Science, 48(1):101–124, February 1, 2018. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312717751051.

Gorman:1977:EIF

[Gor77] Mel Gorman. Electric illumination in the Franco-Prussian war. Social Studies of Science, 7(4):525–529, November 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700416.

Gordon:1982:HSD

[Gor82] Michael D. Gordon. How socially distinctive is cognitive deviance in an emergent science? The case of parapsychology. Social Studies of Science, 12(1):151–165, February 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631282012001012.

Gordon:1984:HAS

[Gor84] Michael D. Gordon. How authors select journals: a test of the reward maximization model of submission behaviour. Social Studies of Science, 14(1):27–43, February 1, 1984. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014001003.

Gorman:1989:AE

[Gor89a]  Michael E. Gorman. Artificial epistemology?  Social

Studies of Science, 19(2):374–380, May 1, 1989. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019002012.

Gorman:1989:BSP

[Gor89b] Michael E. Gorman. Beyond strong programmes: How cognitive approaches can complement SSK. Social Studies of Science, 19 (4):643–653, November 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631289019004006.

Gorman:1990:BRC

[Gor90] Michael E. Gorman. Book review: Charles Bazerman, Shaping Written Knowledge: The Genre and Activity of the Experimental Article in Science. Social Studies of Science, 20(1):

169–172, February 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631290020001008.

Gorman:1991:CSS

[Gor91a] Michael E. Gorman. Counterfactual simulations of science: a response to Farris and Revlin. Social Studies of Science, 21 (3):561–564, August 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631291021003007. See

[FR89, FR91].

Gorman:1991:TPS

[Gor91b] Michael E. Gorman. Towards a psychology of science. Social Studies of Science, 21(2):369–374, May 1, 1991. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631291021002008.

Gorman:1991:WSS

[Gor91c] Michael E. Gorman. What Simon said. Social Studies of Science, 21(1):153–154, February 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631291021001012.

Gorman:1992:UTM

[Gor92] Michael E. Gorman. Using technology to model technoscience. Social Studies of Science, 22(2):407–411, May 1, 1992. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631292022002013.

Gorman:1997:MWC

[Gor97] Michael E. Gorman. Mind in the world: Cognition and practice in the invention of the telephone. Social Studies of Science, 27 (4):583–624, August 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631297027004002.

Gorman:2002:LET

[Gor02] Michael E. Gorman. Levels of expertise and trading zones: a framework for multidisciplinary collaboration. Social Studies of Science, 32(5–6):933–938, December 1, 2002. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631270203200511.

Gorman:2006:SEK

[Gor06] Michael E. Gorman. STS, ethics, and knowledge transfer in the courtroom: Personal experiences. Social Studies of Science, 36 (6):861–866, December 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706067902.

Gorman:2008:OMJ

[Gor08] Michael E. Gorman. Obituary: Michael J. Mahoney. Social Studies of Science, 38(1):159–160, February 1, 2008. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706088564.

Gorman:2010:ODC

[Gor10] Michael E. Gorman. Obituary: David Charles Gooding (1947– 2009). Social Studies of Science, 40(2):341–343, April 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312710364086.

Gottweis:1995:GPG

[Got95] Herbert Gottweis. German politics of genetic engineering and its deconstruction. Social Studies of Science, 25(2):195–235, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631295025002001.

Gale:1997:STP

[GP97] George Gale and Cassandra L. Pinnick. Stalking theoretical physicists: an ethnography flounders: a response to Merz and Knorr-Cetina. Social Studies of Science, 27(1):113–123, February 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631297027001005. See reply [KCM97].

Geissler:2009:ACA

[GP09] P. Wenzel Geissler and Ruth J. Prince. Active compounds and atoms of society: Plants, bodies, minds and cultures in the work of Kenyan ethnobotanical knowledge. Social Studies of Science, 39(4):599–634, August 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312709104075.

Griggs:1986:SST

[GR86] Richard A. Griggs and Sarah E. Ransdell. Scientists and the selection task. Social Studies of Science, 16(2):319–330, May 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312786016002007.

Goedeke:2008:OAS

[GR08] T. L. Goedeke and S. Rikoon. Otters as actors: Scientific controversy, dynamism of networks, and the implications of power in ecological restoration. Social Studies of Science, 38 (1):111–132, February 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312707077363.

Graham:1975:FSR

[Gra75] Loren R. Graham. The formation of Soviet research institutes: a combination of revolutionary innovation and international borrowing. Social Studies of Science, 5(3):303–329, August 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277500500303.

Graham:1985:SPR

[Gra85] Loren R. Graham. The socio-political roots of Boris Hessen: Soviet Marxism and the history of science. Social Studies of Science, 15(4):705–722, November 1, 1985. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631285015004005.

Grabiner:1998:SDC

[Gra98] Judith V. Grabiner. ‘Some disputes of consequence’: Maclaurin among the molasses barrels. Social Studies of Science, 28(1): 139–168, February 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631298028001005.

Graber:2007:ODD

[Gra07] Frédéric Graber. Obvious decisions: Decision-making among French Ponts-et-Chaussées engineers around 1800. Social Studies of Science, 37(6):935–960, December 1, 2007. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312707078013.

Grant:2016:SRB

[Gra16] Jenna M. Grant. From subjects to relations: Bioethics and the articulation of postcolonial politics in the Cambodia preexposure prophylaxis trial. Social Studies of Science, 46(2): 236–258, April 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312716632617.

Greene:2004:ADE

[Gre04] Jeremy A. Greene. Attention to ‘details’: Etiquette and the pharmaceutical salesman in postwar American. Social Studies of Science, 34(2):271–292, April 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704043029.

Greiffenhagen:2010:RES

[Gre10] Christian Greiffenhagen. Review essay: a sociology of formal logic?: Claude Rosental, Weaving Self-Evidence: a Sociology of Logic (Princeton, NJ: Princeton University Press, 2008), 296 pp. £59.00 / EUR 78.99 / $85.00; ISBN 978-0-691-13741-4 (cloth); £16.95 / EUR 19.99 / $24.95; ISBN 978-0-691-13940-1 (pbk). Translated from French by Catherine Porter. Originally published in French as La trame de l’évidence (Paris: Presses Universitaires de France, 2003). Social Studies of Science, 40 (3):471–480, June 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312709357229.

Green:2016:KGB

[Gre16a] Adam Isaiah Green. Keeping gay and bisexual men safe: The arena of HIV prevention science and praxis. Social Studies of Science, 46(2):210–235, April 1, 2016. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312715624982.

Greenhalgh:2016:NSC

[Gre16b] Susan Greenhalgh. Neoliberal science, Chinese style: Making and managing the ‘obesity epidemic’. Social Studies of Science, 46(4):485–510, August 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312716655501.

Greene:2022:LIB

[Gre22] Daniel Greene. Landlords of the Internet: Big data and big real estate. Social Studies of Science, 52(6):904–927, December 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

Greiffenhagen:2024:CCM

[Gre24] Christian Greiffenhagen. Checking correctness in mathematical peer review. Social Studies of Science, 54(2):184–209, April 1, 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127231200274.

Griffiths:2018:SSS

[Gri18] David Andrew Griffiths. Shifting syndromes: Sex chromosome variations and intersex classifications. Social Studies of Science, 48(1):125–148, February 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312718757081.

Gingras:1986:CC

[GS86] Yves Gingras and Silvan S. Schweber. Constraints on construction. Social Studies of Science, 16(2):372–383, May 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312786016002011.

Gupta:2002:WAS

[GS02] Namrata Gupta and Arun K. Sharma. Women academic scientists in India. Social Studies of Science, 32(5–6):901–915, December 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631270203200505.

Greiffenhagen:2011:DML

[GS11] Christian Greiffenhagen and Wes Sharrock. Does mathematics look certain in the front, but fallible in the back? Social Studies of Science, 41(6):839–866, December 1, 2011. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312711424789.

Glenhaber:2025:PVP

[GS25] Mehitabel Glenhaber and Hamsini Sridharan. Precog visions: Predicting the future with the minority report sociotechnical imaginary. Social Studies of Science, 55(1):37–61, February 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127241270991.

Griffith:1974:SSL

[GSSD74] Belver C. Griffith, Henry G. Small, Judith A. Stonehill, and Sandra Dey. The structure of scientific literatures II: Toward a macro- and microstructure for science. Social Studies of Science, 4(4):339–365, October 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277400400402.

Garvey:1972:CPS

[GT72] William D. Garvey and Kazuo Tomita. Continuity of productivity by scientists in the years 1968–71. Social Studies of Science, 2(4):379–383, October 1, 1972. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277200200406.

Golub:1977:MMC

[GT77] Robert Golub and Joe Townsend. Malthus, multinationals and the Club of Rome. Social Studies of Science, 7(2):201–222, May 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277700700204.

Gingras:1993:CTP

[GT93] Yves Gingras and Michel Trépanier. Constructing a tokamak: Political, economic and technical factors as constraints and resources. Social Studies of Science, 23(1):5–36, February 1, 1993.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631293023001001.

Ghamari-Tabrizi:2000:SUG

[GT00] Sharon Ghamari-Tabrizi. Simulating the unthinkable: Gaming future war in the 1950s and 1960s. Social Studies of Science, 30(2):163–223, April 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631200030002001.

Gu:2023:CPN

[Gu23] Chao Gu. The co-production of normal science: a social history of high-temperature superconductivity research in China (1987–2008). Social Studies of Science, 53(1):81–101, February 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

Guice:1998:CSL

[Gui98] Jon Guice. Controversy and the state: Lord ARPA and intelligent computing. Social Studies of Science, 28(1):103–138, February 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631298028001004.

Gummett:1990:ISR

[Gum90] Philip Gummett. Issues for STS raised by defence science and technology policy. Social Studies of Science, 20(3):541–558, August 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631290020003007.

Guma:2020:IUI

[Gum20] Prince K. Guma. Incompleteness of urban infrastructures in transition: Scenarios from the mobile age in Nairobi. Social Studies of Science, 50(5):728–750, October 1, 2020. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720927088.

Gunnarsdottir:2005:SJP

[Gun05] Kristrún Gunnarsdóttir. Scientific journal publications: On the role of electronic preprint exchange in the distribution of scientific literature. Social Studies of Science, 35(4):549–579, August 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312705052358.

Guston:1999:SBB

[Gus99] David H. Guston. Stabilizing the boundary between US politics and science: The rôle of the Office of Technology Transfer as a boundary organization. Social Studies of Science, 29 (1):87–111, February 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631299029001004.

Gusterson:2006:IAC

[Gus06] Hugh Gusterson. Inventing annihilation: Comments on Lynn Eden’s Whole World on Fire: Organizations, Knowledge, & Nuclear Weapons Devastation. Social Studies of Science, 36 (4):634–640, August 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706064813.

Guston:2014:UAG

[Gus14] David H. Guston. Understanding ‘anticipatory governance’. Social Studies of Science, 44(2):218–242, April 1, 2014. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713508669.

Gilbert:1974:ERQ

[GW74] G. Nigel Gilbert and Steve Woolgar. Essay review: The quantitative study of science: an examination of the literature. Social Studies of Science, 4(3):279–294, July 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277400400305.

Habgood:1997:DER

[Hab97] John Habgood. Debating embryo research. Social Studies of Science, 27(3):537–542, June 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631297027003007.

Haase:2023:DSD

[HAB^{+}23] Christoffer Bjerre Haase, Rola Ajjawi, Margaret Bearman, John Brandt Brodersen, Torsten Risor, and Klaus Hoeyer. Data as symptom: Doctors’ responses to patient-provided data in general practice. Social Studies of Science, 53(4):522–544, August 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
abs/10.1177/03063127231164345.

Hacking:1984:WR

[Hac84] Ian Hacking. Wittgenstein rules. Social Studies of Science, 14 (3):469–476, August 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631284014003008.

Hacker:1992:GPF

[Hac92] Barton C. Hacker. The Gemini paraglider: a failure of scheduled innovation, 1961–64. Social Studies of Science, 22(2):387–406, May 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631292022002012.

Hackett:2005:ETI

[Hac05a] Edward J. Hackett. Essential tensions: Identity, control, and risk in research. Social Studies of Science, 35(5):787–826, October 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312705056045.

Hackett:2005:ISG

[Hac05b] Edward J. Hackett. Introduction to the special guest-edited issue on scientific collaboration. Social Studies of Science, 35 (5):667–671, October 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312705057569.

Hagendijk:1994:BRT

[Hag94] Rob Hagendijk. Book review: Towards a Sociology of Science. Social Studies of Science, 24(1):135–139, February 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631279402400110.

Hagendijk:1999:ASP

[Hag99] Rob Hagendijk. An agenda for STS: Porter on trust and quantification in science, politics and society. Social Studies of Science, 29(4):629–637, August 1, 1999. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029004007.

Haines:2019:CCE

[Hai19] Monamie Bhadra Haines. Contested credibility economies of nuclear power in India. Social Studies of Science, 49(1): 29–51, February 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312719827114.

Hakken:1995:CRS

[Hak95] David Hakken. The cultural reconstruction of science: a response to Labinger. Social Studies of Science, 25(2):317–320, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631295025002008. See [Lab95b].

Halfpenny:1988:TTW

[Hal88] Peter Halfpenny. Talking of talking, writing of writing: Some reflections on Gilbert and Mulkay’s discourse analysis. Social Studies of Science, 18(1):169–182, February 1, 1988. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631288018001007. See [GM82, PM89, Hal89].

Halfpenny:1989:RPM

[Hal89] Peter Halfpenny. Reply to Potter and McKinlay. Social Studies of Science, 19(1):145–152, February 1, 1989. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019001006. See [GM82, Hal88].

Halfon:1998:CTC

[Hal98] Saul Halfon. Collecting, testing and convincing: Forensic DNA experts in the courts. Social Studies of Science, 28(5– 6):801–828, October 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631298028005006.

Halfon:2006:DCI

[Hal06] Saul Halfon. The disunity of consensus: International population policy coordination as socio-technical practice. Social Studies of Science, 36(5):783–807, October 1, 2006. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706059745.

Halverson:2019:SLP

[Hal19] Colin Michael Egenberger Halverson. Standards and legacies: Pragmatic constraints on a uniform gene nomenclature. Social Studies of Science, 49(3):432–455, June 1, 2019. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312719850335.

Hammond:1982:ENH

[Ham82] Michael Hammond. The expulsion of the Neanderthals from human ancestry: Marcellin Boule and the social context of scientific research. Social Studies of Science, 12(1):1–036, February 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631282012001002.

Hamlin:1986:SME

[Ham86] Christopher Hamlin. Scientific method and expert witnessing: Victorian perspectives on a modern problem. Social Studies of Science, 16(3):485–513, August 1, 1986. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631286016003004.

Hamlin:1992:RTS

[Ham92] Christopher Hamlin. Reflexivity in technology studies: Toward a technology of technology (and science)? Social Studies of Science, 22(3):511–544, August 1, 1992. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312792022003004.

Hamann:2019:MPA

[Ham19] Julian Hamann. The making of professors: Assessment and recognition in academic recruitment. Social Studies of Science, 49(6):919–941, December 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312719880017.

Handberg:1984:RGH

[Han84] Roger Handberg. Response to Gieryn and Hirsh. Social Studies of Science, 14(4):622–624, November 1, 1984. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014004006. See reply [GH84].

Handberg:1986:PWS

[Han86] Roger Handberg. Practising Western science inside the West: Psychological and institutional parallels between Western and nonwestern academic cultures. Social Studies of Science, 16 (3):529–534, August 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631286016003006.

Harwood:1976:RIC

[Har76] Jonathan Harwood. The race–intelligence controversy: a sociological approach I — professional factors. Social Studies of Science, 6(3–4):369–394, September 1, 1976. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277600600305.

Harwood:1977:RIC

[Har77] Jonathan Harwood. The race-intelligence controversy: a sociological approach II — ‘external’ factors. Social Studies of Science, 7(1):1–30, February 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631277700700108.

Hardin:1980:RD

[Har80] C. L. Hardin. Rationality and disconfirmation. Social Studies of Science, 10(4):509–514, November 1, 1980. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278001000407.

Hardin:1981:TTP

[Har81a] C. L. Hardin. Table-turning, parapsychology and fraud. Social Studies of Science, 11(2):249–255, May 1, 1981. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278101100203. See reply [Pin81a].

Harvey:1981:PEK

[Har81b] Bill Harvey. Plausibility and the evaluation of knowledge: a case-study of experimental quantum mechanics. Social Studies of Science, 11(1):95–130, February 1, 1981. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278101100105.

Hart:1983:RHS

[Har83a] David Hart. Review: Hart on Sharma: Dhirenda Sharma, India’s Nuclear Estate (New Delhi: Lancers Publishers, 1983), 195 pp., Rs. 115. Social Studies of Science, 13(4):627–629, November 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631283013004008.

Harwood:1983:IH

[Har83b] Jonathan Harwood. The IQ in history. Social Studies of Science, 13(3):465–477, August 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631283013003006.

Harwood:1986:LFS

[Har86] Jonathan Harwood. Ludwik Fleck and the sociology of knowledge. Social Studies of Science, 16(1):173–187, February 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631286016001009.

Harwood:1992:SS

[Har92] Jonathan Harwood. Styles of science. Social Studies of Science, 22(1):205–208, February 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312792022001012.

Haard:1994:TPL

[Hår94a] Mikael Hård. Technology as practice: Local and global closure processes in Diesel-engine design. Social Studies of Science, 24 (3):549–585, August 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631279402400304.

Hardy:1994:BRL

[Har94b] Susan Hardy. Book review: The Laboratory Revolution in Medicine. Social Studies of Science, 24(4):759–762, November 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631279402400409.

Hartland:1996:ABP

[Har96] Joanne Hartland. Automating blood pressure measurements: The division of labour and the transformation of method. Social Studies of Science, 26(1):71–94, February 1, 1996. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631296026001005.

Hartley:2001:WAG

[Har01] James Hartley. Where is the address: Give readers and writers a chance by providing postal details. Social Studies of Science, 31(4):627–628, August 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631201031004006.

Hartley:2002:HAI

[Har02a] James Hartley. Here’s the address: Improving information in abstracting journals and databases. Social Studies of Science, 32(4):637–638, August 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312702032004007.

Hartley:2002:CTS

[Har02b] James Hartley. On choosing typographic settings for reference lists. Social Studies of Science, 32(5–6):917–932, December 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631270203200512;   https://www.jstor.org/
stable/3183059.

Harper:2003:RBL

[Har03] Kristine C. Harper. Research from the boundary layer: Civilian leadership, military funding and the development of numerical weather prediction (1946–55). Social Studies of Science, 33 (5):667–696, October 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312703335003.

Haraway:2023:PBD

[Har23] Donna Haraway. Present to Bruno, from Donna. Social Studies of Science, 53(2):165–168, April 1, 2023. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231157395.

Hasson:2012:MAS

[Has12] Katie Ann Hasson. Making appropriation ‘stick’: Stabilizing politics in an ‘inherently feminist’ tool. Social Studies of Science, 42(5):638–661, October 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712440750.

Hawthorn:1973:ERI

[Haw73] Geoffrey Hawthorn. Essay reviews: The interest of scientific knowledge. Social Studies of Science, 3(1):78–87, January 1, 1973. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277300300107.

Hayden:2007:TGB

[Hay07] Cori Hayden. Taking as giving: Bioscience, exchange, and the politics of benefit-sharing. Social Studies of Science, 37 (5):729–758, October 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312707078012.

Henriksen:2023:ECA

[HB23] Anne Henriksen and Lasse Blond. Executive-centered AI? Designing predictive systems for the public sector. Social Studies of Science, 53(5):738–760, October 1, 2023. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231163756.

Hoeyer:2019:DAP

[HBP19] Klaus Hoeyer, Susanne Bauer, and Martyn Pickersgill. Datafication and accountability in public health: Introduction to a special issue. Social Studies of Science, 49(4):459–475, August 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312719860202.

Hess:1998:DFN

[HDS^{+}98] David Hess, Gary Downey, Lucy Suchman, David Hakken, and Leigh Star. Diana E. Forsythe (11 November 1947–14 August 1997). Social Studies of Science, 28(1):175–182, February 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631298028001007.

Healy:2004:SII

[Hea04] David Healy. Shaping the intimate: Influences on the experience of everyday nerves. Social Studies of Science, 34(2):219–245, April 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312704042620.

Healy:2007:DDW

[Hea07] Stephen Healy. Deadly dingoes: ‘wild’ or simply requiring ‘due process’? Social Studies of Science, 37(3):443–471, June 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312706070746.

Hecht:1996:RPT

[Hec96] Gabrielle Hecht. Rebels and pioneers: Technocratic ideologies and social identities in the French nuclear workplace, 1955– 69. Social Studies of Science, 26(3):483–530, August 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631296026003001.

Hecht:2002:RTN

[Hec02] Gabrielle Hecht. Rupture-talk in the nuclear age: Conjugating colonial power in Africa. Social Studies of Science, 32(5–6):

691–727, December 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631270203200504.

Hedgecoe:2001:SNE

[Hed01] Adam Hedgecoe. Schizophrenia and the narrative of enlightened geneticization. Social Studies of Science, 31(6):875–911, December 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631201031006004.

Hedgecoe:2006:PAS

[Hed06] Adam Hedgecoe. Pharmacogenetics as alien science: Alzheimer’s disease, core sets and expectations. Social Studies of Science, 36(5):723–752, October 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706059746.

Hedfors:2008:MSLb

[Hed08a] Eva Hedfors. Medical science in the light of the Holocaust. Social Studies of Science, 38(6):945–950, December 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/0306312708098610. See [Hed08a, ABB^{+}08, Hed08b].

Hedfors:2008:MSLa

[Hed08b] Eva Hedfors. Medical science in the light of the Holocaust: Departing from a post-war paper by Ludwik Fleck. Social Studies of Science, 38(2):259–283, April 1, 2008. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/0306312707082953. See comment [ABB^{+}08] and reply [Hed08a].

Hedgecoe:2010:BRS

[Hed10] Adam Hedgecoe. Bioethics and the reinforcement of sociotechnical expectations. Social Studies of Science, 40(2):163–186, April 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312709349781.

Hedgecoe:2012:TRO

[Hed12] Adam M. Hedgecoe. Trust and regulatory organisations: The role of local knowledge and facework in research ethics review. Social Studies of Science, 42(5):662–683, October 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712446364.

Hedgecoe:2014:DSD

[Hed14] Adam Hedgecoe. A deviation from standard design? Clinical trials, research ethics committees, and the regulatory co-construction of organizational deviance. Social Studies of Science, 44(1):59–81, February 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312713506141.

Heffner:1979:ARS

[Hef79] Alan G. Heffner. Authorship recognition of subordinates in collaborative research. Social Studies of Science, 9(3):377–384, August 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277900900305.

Heggie:2016:HCS

[Heg16] Vanessa Heggie. Higher and colder: The success and failure of boundaries in high altitude and Antarctic research stations. Social Studies of Science, 46(6):809–832, December 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716636249.

Heintz:2003:WPP

[Hei03] Bettina Heintz. When is a proof a proof? Social Studies of Science, 33(6):929–943, December 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312703336005.

Helmreich:1998:RRR

[Hel98] Stefan Helmreich. Recombination, rationality, reductionism and romantic reactions: Culture, computers, and the genetic algorithm. Social Studies of Science, 28(1):39–71, February 1, 1998.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631298028001002.

Henderson:1990:SSC

[Hen90] David K. Henderson. On the sociology of science and the continuing importance of epistemologically couched accounts. Social Studies of Science, 20(1):113–148, February 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020001004.

Henke:2000:MPS

[Hen00] Christopher R. Henke. Making a place for science: The field trial. Social Studies of Science, 30(4):483–511, August 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030004001.

Henke:2007:SNR

[Hen07] Christopher R. Henke. Situation normal? Repairing a risky ecology. Social Studies of Science, 37(1):135–142, February 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312706069436.

Hennessy:2018:PNL

[Hen18] Elizabeth Hennessy. The politics of a natural laboratory: Claiming territory and governing life in the Galápagos Islands. Social Studies of Science, 48(4):483–506, August 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312718788179.

Hercock:2003:MSC

[Her03a] Marion Hercock. Masters and servants: The contrasting roles of scientists in island management. Social Studies of Science, 33 (1):117–136, February 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312703033001293.

Hermanowicz:2003:SS

[Her03b] Joseph C. Hermanowicz. Scientists and satisfaction. Social Studies of Science, 33(1):45–73, February 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312703033001177.

Hermanowicz:2007:AOS

[Her07] Joseph C. Hermanowicz. Argument and outline for the sociology of scientific (and other) careers. Social Studies of Science, 37 (4):625–646, August 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706075337.

Hesse:1986:CCS

[Hes86] Mary Hesse. Changing concepts and stable order. Social Studies of Science, 16(4):714–726, November 1, 1986. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631286016004008.

Hess:1987:RHS

[Hes87] David Hess. Religion, heterodox science and Brazilian culture. Social Studies of Science, 17(3):465–477, August 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631287017003004.

Hessenbruch:2000:CWX

[Hes00] Arne Hessenbruch. Calibration and work in the X-ray economy, 1896–1928. Social Studies of Science, 30(3):397–420, June 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631200030003003.

Hollin:2017:DEB

[HFGP17] Gregory Hollin, Isla Forsyth, Eva Giraud, and Tracey Potts. (Dis)entangling Barad: Materialisms and ethics. Social Studies of Science, 47(6):918–941, December 1, 2017. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312717728344.

Hargens:1988:NCM

[HG88] Lowell Hargens and Thomas F. Gieryn. Nicholas C. Mullins (1939–88). Social Studies of Science, 18(3):572–573, August 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631288018003013.

Hargreaves:1983:SMS

[HH83] Judith Hargreaves and Tony Hargreaves. Some models of school science in British curriculum projects, and their implications for STS teaching at the secondary level. Social Studies of Science, 13(4):569–604, November 1, 1983. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631283013004005.

Helosvuori:2022:WCK

[HH22] Elina Helosvuori and Riikka Homanen. When craft kicks back: Embryo culture as knowledge production in the context of the transnational fertility industry. Social Studies of Science, 52 (3):425–446, June 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/03063127221083869.

Harsh:2019:SSA

[HHW^{+}19] Matthew Harsh, Kerry Holden, Jameson Wetmore, G. Pascal Zachary, and Ravtosh Bal. Situating science in Africa: The dynamics of computing research in Nairobi and Kampala. Social Studies of Science, 49(1):52–76, February 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312719829595.

Horst:2010:NER

[HI10] Maja Horst and Alan Irwin. Nations at ease with radical knowledge: On consensus, consensusing and false consensusness. Social Studies of Science, 40(1):105–126, February 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312709341500.

Hicks:1987:LCC

[Hic87] Diana Hicks. Limitations of co-citation analysis as a tool for science policy. Social Studies of Science, 17(2):295–316, May 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/

10.1177/030631287017002004. See response [Fra88] and reply [Hic88].

Hicks:1988:LML

[Hic88] Diana Hicks. Limitations and more limitations of co-citation Analysis/ bibliometric modelling: a reply to Franklin. Social Studies of Science, 18(2):375–384, May 1, 1988. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631288018002008. See [Fra88].

Higgitt:2021:NFT

[Hig21] Ryan Higgitt. Neanderthal and the fossilization of the Third World. Social Studies of Science, 51(3):439–462, June 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720983460.

Hill:1974:QIS

[Hil74] Stephen C. Hill. Questioning the influence of a ‘social system of science’: a study of Australian scientists. Social Studies of Science, 4(2):135–163, April 1, 1974. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277400400202.

Hill:1986:NKC

[Hil86] Christopher Hill. ‘A new kind of clergy’: Ideology and the experimental method. Social Studies of Science, 16(4):726– 735, November 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631286016004009.

Hilgartner:1990:DVP

[Hil90] Stephen Hilgartner. The dominant view of popularization: Conceptual problems, political uses. Social Studies of Science, 20 (3):519–539, August 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631290020003006.

Hilgartner:2001:EPU

[Hil01] Stephen Hilgartner. Election 2000 and the production of the unknowable. Social Studies of Science, 31(3):439–441, June 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631201031003010.

Hilgartner:2004:CSS

[Hil04] Stephen Hilgartner. The credibility of science on stage. Social Studies of Science, 34(3):443–452, June 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704043694.

Hilgartner:2007:OCA

[Hil07] Stephen Hilgartner. Overflow and containment in the aftermath of disaster. Social Studies of Science, 37(1):153–158, February 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312706069439.

Hine:2006:DSI

[Hin06] Christine Hine. Databases as scientific instruments and their role in the ordering of scientific work. Social Studies of Science, 36(2):269–298, April 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706054047.

Hinterberger:2013:CPC

[Hin13] Amy Hinterberger. Curating postcolonial critique. Social Studies of Science, 43(4):619–627, August 1, 2013. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713478678.

Hindmarsh:2014:HAA

[Hin14] Richard Hindmarsh. Hot air ablowin! ‘media-speak’, social conflict, and the Australian ‘decoupled’ wind farm controversy. Social Studies of Science, 44(2):194–217, April 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713504239.

Hirschauer:1991:MBS

[Hir91] Stefan Hirschauer. The manufacture of bodies in surgery. Social Studies of Science, 21(2):279–319, May 1, 1991. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/030631291021002005. See comment [Col94a] and reply [Hir94].

Hirschauer:1994:TMI

[Hir94] Stefan Hirschauer. Towards a methodology of investigations into the strangeness of one’s own culture: a response to Collins. Social Studies of Science, 24(2):335–346, May 1, 1994. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631279402400206. See [Hir91, Col94a].

Hirschauer:2010:EJP

[Hir10] Stefan Hirschauer. Editorial judgments: a praxeology of ‘voting’ in peer review. Social Studies of Science, 40(1):71–103, February 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312709335405.

Hoeyer:2013:TEP

[HJ13] Klaus L. Hoeyer and Anja Mb. Jensen. Transgressive ethics: Professional work ethics as a perspective on ‘aggressive organ harvesting’. Social Studies of Science, 43(4):598–618, August 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312712460341.

Hetherington:2023:BFM

[HJ23] Kregg Hetherington and Elie Jalbert. The Big Flush of Montreal: On affective maintenance and infrastructural events. Social Studies of Science, 53(1):102–120, February 1, 2023. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

Hedgecoe:2023:RSR

[HJC23] Adam Hedgecoe, Kathleen Job, and Angus Clarke. Reflexive standardization and the resolution of uncertainty in the genomics clinic. Social Studies of Science, 53(3):358–378, June 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127231154863.

Hunter:2010:PRP

[HL10] Laura A. Hunter and Erin Leahey. Parenting and research productivity: New evidence and methods. Social Studies of Science, 40(3):433–451, June 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312709358472.

Hillman:2019:SMU

[HL19] Alexandra Hillman and Joanna Latimer. Somaticization, the making and unmaking of minded persons and the fabrication of dementia. Social Studies of Science, 49(2):208–226, April 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312719834069.

Hird:2014:MWM

[HLRK14] Myra J. Hird, Scott Lougheed, R. Kerry Rowe, and Cassandra Kuyvenhoven. Making waste management public (or falling back to sleep). Social Studies of Science, 44(3):441–465, June 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312713518835.

Hedgecoe:2003:DDW

[HM03] Adam Hedgecoe and Paul Martin. The drugs don’t work: Expectations and the shaping of pharmacogenetics. Social Studies of Science, 33(3):327–364, June 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127030333002.

Healy:2014:SS

[HMA14] David Healy, Derelie Mangin, and Kalman Applbaum. The shipwreck of the singular. Social Studies of Science, 44(4): 518–523, August 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312714536270. See re-

ply [NK14b].

Hargens:1980:RAS

[HMH80] Lowell L. Hargens, Nicholas C. Mullins, and Pamela K. Hecht. Research areas and stratification processes in science. Social Studies of Science, 10(1):55–74, February 1, 1980. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278001000103.

Hoddeson:1983:EKJ

[Hod83] Lillian Hoddeson. Establishing KEK in Japan and Fermilab in the US: Internationalism, nationalism and high energy accelerators. Social Studies of Science, 13(1):1–48, February 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631283013001003.

Hoeppe:2014:WDT

[Hoe14] Götz Hoeppe. Working data together: The accountability and reflexivity of digital astronomical practice. Social Studies of Science, 44(2):243–270, April 1, 2014. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713509705.

Hoeppe:2019:MCP

[Hoe19a] Götz Hoeppe. Medium, calculation, play: On digital images in scientific practice. Social Studies of Science, 49(5):758–784, October 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312719871356.

Hoeyer:2019:DPR

[Hoe19b] Klaus Hoeyer. Data as promise: Reconfiguring Danish public health through personalized medicine. Social Studies of Science, 49(4):531–555, August 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312719858697.

Hoffmann:1979:CST

[Hof79] Erik P. Hoffmann. Contemporary Soviet theories of scientific, technological and social change. Social Studies of Science, 9 (1):101–113, February 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277900900106.

Hoffman:2015:TST

[Hof15] Steve G. Hoffman. Thinking science with thinking machines: The multiple realities of basic and applied knowledge in a research border zone. Social Studies of Science, 45(2):242–269, April 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312714564912.

Hogle:2009:POS

[Hog09] Linda F. Hogle. Pragmatic objectivity and the standardization of engineered tissues. Social Studies of Science, 39(5): 717–742, October 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312709103478.

Hogle:2019:AAC

[Hog19] Linda F. Hogle. Accounting for accountable care: Value-based population health management. Social Studies of Science, 49 (4):556–582, August 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312719840429.

Holloway:1973:ERS

[Hol73] David Holloway. Essay reviews: Science in Russian and Soviet society. Social Studies of Science, 3(1):61–72, January 1, 1973. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277300300105.

Holloway:1974:ISC

[Hol74] David Holloway. Innovation in science-the case of cybernetics in the Soviet Union. Social Studies of Science, 4(4):299–337, October 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277400400401.

Holloway:1981:ENA

[Hol81a] David Holloway. Entering the nuclear arms race: The Soviet decision to build the atomic bomb, 1939–45. Social Studies of Science, 11(2):159–197, May 1, 1981. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL http://www.

jstor.org/stable/284865;   https://journals.sagepub.
com/doi/pdf/10.1177/030631278101100201.

Holloway:1981:ERP

[Hol81b] David Holloway. Essay review: The politics of Soviet science and technology: Kendall E. Bailes, Technology and Society under Lenin and Stalin: Origins of the Soviet Technical Intelligentsia, 1917–1941 (Princeton, NJ: Princeton University Press, 1978), xiii + 472 pp., $34.30. ISBN 0-691-05265-4 (paperback — 0-691-10063-2). Nicholas Lampert, The Technical Intelligentsia and the Soviet State: a Study of Soviet Managers and Technicians, 1928–1935 (London: Macmillan, 1979), x + 191 pp., £12.00. ISBN 0-333-23759-5. Robert Lewis, Science and Industrialisation in the USSR: Industrial Research and Development, 1917–1940 (London: Macmillan, 1979), xiv + 211 pp., £12.00. ISBN 0-333-23758-7. Frederic J. Fleron, Jr (ed.), Technology and Communist Culture: The Socio-Cultural Impact of Technology under Socialism (New York and London: Praeger, 1977), 518 pp., £26.00. ISBN 0-03-021821-7. (Praeger Special Studies in International Politics and. Social Studies of Science, 11(2):259–274, May 1, 1981. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631278101100205.

Hollin:2022:CWC

[Hol22] Gregory Hollin. Consider the woodpecker: The contested morethan-human ethics of biomimetic technology and traumatic brain injury. Social Studies of Science, 52(2):149–173, April 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/03063127211052513.

Home:1990:BRR

[Hom90] R. W. Home. Book review: Rudolf Stichweh, Zur Enstehung des modernen Systems wissenschaftlicher Disziplinen: Physik in Deutschland, 1740–1890. Social Studies of Science, 20(4): 761–763, November 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631290020004010.

Hommels:2003:SC

[Hom03] Anique Hommels. STS and the city. Social Studies of Science, 33 (6):945–950, December 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312703336006.

Hong:2008:DSF

[Hon08] Wei Hong. Domination in a scientific field: Capital struggle in a Chinese isotope lab. Social Studies of Science, 38 (4):543–570, August 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706092456.

Hood:2017:SVE

[Hoo17] Katherine Hood. The science of value: Economic expertise and the valuation of human life in US federal regulatory agencies. Social Studies of Science, 47(4):441–465, August 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312717693465.

Hopman:2023:FFO

[Hop23] Roos Hopman. The face as folded object: Race and the problems with ‘progress’ in forensic DNA phenotyping. Social Studies of Science, 53(6):869–890, December 1, 2023. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127211035562.

Horning:2004:EPR

[Hor04] Susan Schmidt Horning. Engineering the performance: Recording engineers, tacit knowledge and the art of controlling sound. Social Studies of Science, 34(5):703–731, October 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704047536.

Horst:2025:AST

[Hor25a] Maja Horst. The art, science and technology studies movement: an essay review. Social Studies of Science, 55(1):131–150, February 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127241270917.

Horton:2025:ZVE

[Hor25b] Lynn Horton. ‘Zoonati’ vs. ‘epistemic trespassers’: Science identity in contentious online advocacy campaigns on the origins of SARS-CoV-2. Social Studies of Science, 55(3):423–443, June 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127241294028.

Hounshell:2001:RCW

[Hou01] David A. Hounshell. Rethinking the Cold War; rethinking science and technology in the Cold War; rethinking the social study of science and technology. Social Studies of Science, 31 (2):289–297, April 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312701031002007.

Hicks:1991:SSK

[HP91] Diana Hicks and Jonathan Potter. Sociology of scientific knowledge: a reflexive citation analysis or science disciplines and disciplining science. Social Studies of Science, 21(3):459–501, August 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631291021003003.

Heinsler:1987:CAC

[HR87] Janet M. Heinsler and Rachel A. Rosenfeld. Charting academic careers: Does data source make a difference? Social Studies of Science, 17(1):135–144, February 1, 1987. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631287017001006.

Hesselmann:2021:CIL

[HR21] Felicitas Hesselmann and Martin Reinhart. Cycles of invisibility: The limits of transparency in dealing with scientific misconduct. Social Studies of Science, 51(3):414–438, June 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720975201.

Hirsch:2022:SLD

[HRI22] Shana Lee Hirsch, David Ribes, and Sarah Inman. Sedimentary legacy and the disturbing recurrence of the human in long-term ecological research. Social Studies of Science, 52 (4):561–580, August 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127221101171.

Hartley:2002:SSP

[HSP02] James Hartley, Eric Sotto, and James Pennebaker. Style and substance in psychology: Are influential articles more readable than less influential ones? Social Studies of Science, 32 (2):321–334, April 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312702032002005.

Hsu:2013:MGZ

[Hsu13] Hansen Hsu. Machine gambling in the ‘zone’: Natasha Dow Schüll’s Addiction by Design: Natasha Dow Schüll, Addiction By Design: Machine Gambling in Las Vegas (Princeton, NJ: Princeton University Press, 2012), 442 pp., ISBN 978-0-691-12755-2. Social Studies of Science, 43(6):952–956, December 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312713507496.

Haard:2020:TZC

[HTB20] Mikael Hård and Mai Lin Tjoa-Bonatz. Trading zones in a colony: Transcultural techniques at missionary stations in the Dutch East Indies, 1860–1940. Social Studies of Science, 50(6): 932–955, December 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312720925913.

Hu:2025:PTR

[Hu25] Cameron Hu. Postcolonial technoscience revisited. Social Studies of Science, 55(4):613–630, August 2025. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127241303778.

Hudson:1973:FAS

[Hud73] Liam Hudson. Fertility in the arts and sciences. Social Studies of Science, 3(3):305–310, July 1, 1973. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277300300304.

Hufbauer:1986:FFS

[Huf86] Karl Hufbauer. Federal funding and sudden infant death research, 1945–80. Social Studies of Science, 16(1):61–78, February 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631286016001004.

Hughes:1980:WFR

[Hug80] Phillip S. Hughes. Wartime fission research in Japan. Social Studies of Science, 10(3):345–349, August 1, 1980. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278001000304.

Hughes:1986:SWT

[Hug86] Thomas P. Hughes. The seamless Web: Technology, science, etcetera, etcetera. Social Studies of Science, 16(2):281–292, May 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312786016002004.

Hughes:2006:CLE

[Hug06] Thomas P. Hughes. Comments on Lynn Eden’s Whole World on Fire: Organizations, Knowledge, & Nuclear Weapons Devastation. Social Studies of Science, 36(4):641–645, August 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312706064814. See [Ano06c, Ans06].

Hughes:2024:HMT

[Hug24] Stephen Hughes. Hearts and minds: The technopolitical role of affect in sociotechnical imaginaries. Social Studies of Science, 54(6):907–930, December 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127241257489.

Hult:1992:IPB

[Hul92] Jan Hult. The Itera plastic bicycle. Social Studies of Science, 22(2):373–385, May 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631292022002011.

Hunter:1982:RRS

[Hun82] Michael Hunter. Reconstructing restoration science: Problems and pitfalls in institutional history. Social Studies of Science, 12(3):451–466, August 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631282012003006.

Hunter:2024:LIB

[Hun24] Hannah Hunter. Listening for the ivory-billed woodpecker: Sonic geography and the making of extinction knowledge. Social Studies of Science, 54(3):325–351, June 1, 2024. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231214501.

Hutchinson:1971:GLI

[Hut71] Eric Hutchinson. Government laboratories and the influence of organized scientists. Social Studies of Science, 1(3–4): 331–356, October 1, 1971. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277100100305.

Hutchison:1991:ICL

[Hut91] Keith Hutchison. Individualism, causal location, and the eclipse of scholastic philosophy. Social Studies of Science, 21 (2):321–350, May 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631291021002006.

Huutoniemi:2012:CCD

[Huu12] Katri Huutoniemi. Communicating and compromising on disciplinary expertise in the peer review of research proposals. Social Studies of Science, 42(6):897–921, December 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712458478.

Hart:1993:SEM

[HV93] David M. Hart and David G. Victor. Scientific elites and the making of US policy for climate change research, 1957– 74. Social Studies of Science, 23(4):643–680, November 1, 1993.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631293023004002.

Heath:2008:CIE

[HvL08] Christian Heath and Dirk vom Lehn. Configuring ‘interactivity’: Enhancing engagement in science centres and museums. Social Studies of Science, 38(1):63–91, February 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312707084152.

Hyysalo:2006:RUP

[Hyy06] Sampsa Hyysalo. Representations of use and practice-bound imaginaries in automating the safety of the elderly. Social Studies of Science, 36(4):599–626, August 1, 2006. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706058426.

Hyysalo:2009:RHP

[Hyy09] Sampsa Hyysalo. Review: How packaged software conquers the organization: Neil Pollock and Robin Williams, Software and Organizations: The Biography of the Packaged Enterprise System or How SAP Conquered the World (London: Routledge, 2008), 348 pp., £65.00 / EUR 87.99 / $135.00. ISBN 978-0-415-40397-9 (hbk); ISBN 978-0-203-89194-0 (ebk). Social Studies of Science, 39(4):635–642, August 1, 2009. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312709335847.

Ialenti:2021:DBO

[Ial21] Vincent Ialenti. Drum breach: Operational temporalities, error politics and WIPP’s kitty litter nuclear waste accident. Social Studies of Science, 51(3):364–391, June 1, 2021. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720986609.

Irwin:2002:R

[IG02] Alan Irwin and Jeanne Guillemin. Reviews. Social Studies of Science, 32(4):615–630, August 1, 2002. CODEN SSSCDH.

ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/0306312702032004005.

Iida:2015:CIC

[Iid15] Kaori Iida. A controversial idea as a cultural resource: The Lysenko controversy and discussions of genetics as a ‘democratic’ science in postwar Japan. Social Studies of Science, 45 (4):546–569, August 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312715596460.

Irwin:2013:GBP

[IJJ13] Alan Irwin, Torben Elgaard Jensen, and Kevin E. Jones. The good, the bad and the perfect: Criticizing engagement practice. Social Studies of Science, 43(1):118–135, February 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712462461.

Irvine:1983:INT

[IM83a] John Irvine and Ben Martin. The Isaac Newton Telescope. Social Studies of Science, 13(2):321–322, May 1, 1983. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631283013002007.

Irvine:1983:ABR

[IM83b] John Irvine and Ben R. Martin. Assessing basic research: The case of the Isaac Newton Telescope. Social Studies of Science, 13(1):49–86, February 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631283013001004.

Irvine:1985:BRE

[IM85] John Irvine and Ben R. Martin. Basic research in the East and West: a comparison of the scientific performance of high-energy physics accelerators. Social Studies of Science, 15(2):293–341, May 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/
doi/pdf/10.1177/030631285015002004. See critique [Bud85,

KP85, MvR85].

Inhaber:1977:SEG

[Inh77a] Herbert Inhaber. Scientists and economic growth. Social Studies of Science, 7(4):517–524, November 1, 1977. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700415.

Inhaber:1977:WSP

[Inh77b] Herbert Inhaber. Where scientists publish. Social Studies of Science, 7(3):388–394, August 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631277700700308.

Inkster:1985:SEC

[Ink85] Ian Inkster. Scientific enterprise and the colonial ‘model’: Observations on Australian experience in historical context. Social Studies of Science, 15(4):677–704, November 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631285015004004.

Inkster:1988:CES

[Ink88] Ian Inkster. Cultural enterprise: Science, steam intellect and social class in Rochdale circa 1833–1900. Social Studies of Science, 18(2):291–330, May 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631288018002005.

Inkster:1991:MAL

[Ink91] Ian Inkster. Made in America but lost to Japan: Science, technology and economic performance in the two capitalist superpowers. Social Studies of Science, 21(1):157–178, February 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631291021001014.

Ramirez-i-Olle:2018:CSS

[iO18] Meritxell Ramírez i Ollé. ‘Civil skepticism’ and the social construction of knowledge: a case in dendroclimatology. Social Studies of Science, 48(6):821–845, December 1, 2018. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312718763119.

Inhaber:1976:QRN

[IP76] H. Inhaber and K. Przednowek. Quality of research and the Nobel Prizes. Social Studies of Science, 6(1):33–50, February 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277600600102.

Irwin:1994:BRS

[Irw94] Alan Irwin. Book review: Science and its Publics: Continuity and Change in the Risk Society. Social Studies of Science, 24 (1):168–184, February 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631279402400113.

Irwin:2006:PTC

[Irw06] Alan Irwin. The politics of talk: Coming to terms with the ‘new’ scientific governance. Social Studies of Science, 36(2):299–320, April 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312706053350.

Jablecka:1995:CMF

[Jab95] Julita Jablecka. Changes in the management and finance of the research system in Poland: a survey of the opinions of Grant applicants. Social Studies of Science, 25(4):727–753, November 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631295025004007.

Jacob:1975:NEC

[Jac75] J. R. Jacob. The New England Company, the Royal Society and the Indians. Social Studies of Science, 5(4):450–455, November 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277500500405.

Jacob:1978:BAR

[Jac78] J. R. Jacob. Boyle’s atomism and the restoration assault on pagan naturalism. Social Studies of Science, 8(2):211–233, May 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277800800203.

Jackson:1994:SBG

[Jac94] Myles W. Jackson. A spectrum of belief: Goethe’s ‘Republic’ versus Newtonian ‘despotism’. Social Studies of Science, 24(4): 673–701, November 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631279402400403. See

comment [BL94].

Jacob:2019:URP

[Jac19] Marie-Andrée Jacob. Under repair: a publication ethics and research record in the making. Social Studies of Science, 49 (1):77–101, February 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312718824663.

Jahoda:1972:RFM

[Jah72] Marie Jahoda. Review: Foucault, Michel, The Order of Things: an Archaeology of the Human Sciences. London: Tavistock Publications Ltd., 1970. p. 387, £3.75. Social Studies of Science, 2 (1):99–103, January 1, 1972. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277200200104.

Jankovic:2004:SMM

[Jan04] Vladimir Jankovic. Science migrations: Mesoscale weather prediction from Belgrade to Washington, 1970–2000. Social Studies of Science, 34(1):45–75, February 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704040490.

Jankovic:2018:SPS

[Jan18] Vladimir Jankovic. ‘The sun without a permit’: Serbian solar politics, informational risk cascades, and the Great Disappearing Act of August 1999. Social Studies of Science, 48 (4):589–614, August 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312718790812.

Jasanoff:1987:CBP

[Jas87] Sheila S. Jasanoff. Contested boundaries in policy-relevant science. Social Studies of Science, 17(2):195–230, May 1, 1987.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631287017002001.

Jasanoff:1995:CWV

[Jas95] Sheila Jasanoff. Cooperation for what?: a view from the Sociological/ cultural study of science policy. Social Studies of Science, 25(2):314–317, May 1, 1995. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631295025002007.

Jasanoff:1996:BER

[Jas96] Sheila Jasanoff. Beyond epistemology: Relativism and engagement in the politics of science. Social Studies of Science, 26 (2):393–418, May 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631296026002008.

Jasanoff:1997:PKP

[Jas97] Sheila Jasanoff. Public knowledge, private fears. Social Studies of Science, 27(2):350–355, April 1, 1997. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631297027002007.

Jasanoff:1998:EEW

[Jas98] Sheila Jasanoff. The eye of everyman: Witnessing DNA in the Simpson trial. Social Studies of Science, 28(5–6):713–740, October 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631298028005003.

Jasanoff:2000:RPC

[Jas00] Sheila Jasanoff. Reconstructing the past, constructing the present: Can science studies and the history of science live happily ever after? Social Studies of Science, 30(4):621–631, August 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631200030004005.

Jasanoff:2001:EME

[Jas01] Sheila Jasanoff. Election 2000: Mechanical error or system failure. Social Studies of Science, 31(3):461–467, June 1, 2001.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031003016.

Jasanoff:2002:SSV

[Jas02] Sheila Jasanoff. Science and the statistical victim: Modernizing knowledge in breast implant litigation. Social Studies of Science, 32(1):37–69, February 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312702032001003.

Jasanoff:2003:BWS

[Jas03a] Sheila Jasanoff. Breaking the waves in science studies: Comment on H. M. Collins and Robert Evans, ‘The Third Wave of Science Studies’. Social Studies of Science, 33(3):389–400, June 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/03063127030333004. See [CE02].

Jasanoff:2003:MDO

[Jas03b] Sheila Jasanoff. In memoriam: David Owen Edge. Social Studies of Science, 33(2):177–179, April 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127030332004.

Jasanoff:2012:GS

[Jas12] Sheila Jasanoff. Genealogies of STS. Social Studies of Science, 42(3):435–441, June 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712440174.

Jaton:2017:WGA

[Jat17] Florian Jaton. We get the algorithms of our ground truths: Designing referential databases in digital image processing. Social Studies of Science, 47(6):811–840, December 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312717730428.

Jaton:2023:GAE

[Jat23] Florian Jaton. Groundwork for AI: Enforcing a benchmark for neoantigen prediction in personalized cancer immunotherapy. Social Studies of Science, 53(5):787–810, October 1, 2023.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231192857.

Jauho:2015:MSL

[Jau15] Mikko Jauho. The mutual shaping of life insurance and medicine in Finland. Social Studies of Science, 45(4):501–524, August 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312715599850.

Jamison:1975:STP

[JE75] Andrew Jamison and Aant Elzinga. Science: Its theory and practice, an introduction for doctoral students. Social Studies of Science, 5(2):223–234, May 1, 1975. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277500500209.

Jeffrey:2003:SWO

[Jef03] Paul Jeffrey. Smoothing the waters: Observations on the process of cross-disciplinary research collaboration. Social Studies of Science, 33(4):539–562, August 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312703334003.

Jensen:2005:EPH

[Jen05] Casper Bruun Jensen. An experiment in performative history: Electronic patient records as a future-generating device. Social Studies of Science, 35(2):241–267, April 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312705047737.

Jensen:2007:RBT

[Jen07] Casper Bruun Jensen. Review: Beyond the two cultures with scandalous knowledge: Relativism and constructivism revisited. Social Studies of Science, 37(4):647–654, August 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706071791.

Jensen:2020:FMM

[Jen20] Casper Bruun Jensen. A flood of models: Mekong ecologies of comparison. Social Studies of Science, 50(1):76–93, February 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312719871616.

Jensen:2023:EIS

[Jen23] Casper Bruun Jensen. Exercises in irreduction: Some Latourian favourites. Social Studies of Science, 53(2):183–187, April 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127231156649.

Jeon:2019:IPA

[Jeo19] June Jeon. Invisibilizing politics: Accepting and legitimating ignorance in environmental sciences. Social Studies of Science, 49 (6):839–862, December 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312719872823.

Jacob:1972:PMB

[JL72] M. C. Jacob and W. A. Lockwood. Political millenarianism and Burnet’s sacred theory. Social Studies of Science, 2(3):265–279, July 1, 1972. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277200200303.

Jordan:1998:DSR

[JL98] Kathleen Jordan and Michael Lynch. The dissemination, standardization and routinization of a molecular biological technique. Social Studies of Science, 28(5–6):773–800, October 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631298028005005.

Jean:2018:EFD

[JL18] Jason Jean and Yixi Lu. Evolution as a fact? A discourse analysis. Social Studies of Science, 48(4):615–632, August 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312718785773.

Jones:1973:AST

[JLM73] C. L. Jones, G. M. Littlejohn, and A. F. McPherson. Aversion to school teaching as a factor in the choice of science subjects. Social Studies of Science, 3(2):173–179, April 1, 1973. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277300300204.

Jones:1973:FIC

[JM73] C. L. Jones and A. F. McPherson. Fertile imaginations and contrary findings: a comment on subject specialization and sexuality. Social Studies of Science, 3(4):389–391, October 1, 1973. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277300300403.

Jensen:2001:MCP

[JM01] Casper Bruun Jensen and Randi Markussen. Mårup Church and the politics of hybridization: On complexities of choice. Social Studies of Science, 31(6):795–819, December 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031006001.

Joerges:1994:ELE

[Joe94] Bernward Joerges. Expertise lost: an early case of technology assessment. Social Studies of Science, 24(1):96–104, February 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631279402400105.

Joerges:1999:DPA

[Joe99a]  Bernward Joerges. Do politics have artefacts?  So-

cial Studies of Science, 29(3):411–431, June 1, 1999. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029003004.

Joerges:1999:SCB

[Joe99b] Bernward Joerges. Scams cannot be busted: Reply to Woolgar & Cooper. Social Studies of Science, 29(3):450–457, June 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631299029003006. See [WC99].

Johnston:1984:CTI

[Joh84] Ron Johnston. Controlling technology: an issue for the social studies of science. Social Studies of Science, 14(1):97– 113, February 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631284014001008. See re-

sponse [Col85a] and reply [Joh85].

Johnston:1985:SCT

[Joh85] Ron Johnston. The social character of technology (reply to Collingridge). Social Studies of Science, 15(2):381–383, May 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631285015002007. See [Joh84, Col85a].

Johnson:1989:SSR

[Joh89] Michael Johnson. Social science research. Social Studies of Science, 19(4):759–762, November 1, 1989. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019004016.

Johnson:2007:SSS

[Joh07] Ericka Johnson. Surgical simulators and simulated surgeons: Reconstituting medical practice and practitioners in simulations. Social Studies of Science, 37(4):585–608, August 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706072179.

Johnson:2013:AMO

[Joh13] Andi Johnson. The athlete as model organism: The everyday practice of the science of human performance. Social Studies of Science, 43(6):878–904, December 1, 2013. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713490448.

Johnson:2018:BLB

[Joh18] Mark R. Johnson. ‘The biggest legal battle in UK casino history’: Processes and politics of ‘cheating’ in sociotechnical networks. Social Studies of Science, 48(2):304–327, April 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312718771212.

Jones:2009:ESR

[Jon09] Mark Peter Jones. Entrepreneurial science: The rules of the game. Social Studies of Science, 39(6):821–851, December 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312709104434.

Jonsson:2016:BTH

[Jön16] Erik Jönsson. Benevolent technotopias and hitherto unimaginable meats: Tracing the promises of in vitro meat. Social Studies of Science, 46(5):725–748, October 1, 2016. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716658561.

Jones:2017:RHL

[Jon17a] Meg Leta Jones. The right to a human in the loop: Political constructions of computer automation and personhood. Social Studies of Science, 47(2):216–239, April 1, 2017. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312717699716.

Jones:2017:WMR

[Jon17b]  Raya A. Jones. What makes a robot ‘social’?  Social

Studies of Science, 47(4):556–579, August 1, 2017. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312717704722.

Jong:2023:PRU

[Jon23] Lisette Jong. On the persistence of race: Unique skulls and average tissue depths in the practice of forensic craniofacial depiction. Social Studies of Science, 53(6):891–915, December 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127221112073.

Joyce:2005:AIM

[Joy05] Kelly Joyce. Appealing images: Magnetic resonance imaging and the production of authoritative knowledge. Social Studies of Science, 35(3):437–462, June 1, 2005. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312705050180.

Jacobi:1989:SIP

[JS89] Daniel Jacobi and Bernard Schiele. Scientific imagery and popularized imagery: Differences and similarities in the photographic portraits of scientists. Social Studies of Science, 19(4): 731–753, November 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631289019004014.

Jasanoff:2017:NFB

[JS17] Sheila Jasanoff and Hilton R. Simmet. No funeral bells: Public reason in a ‘post-truth’ age. Social Studies of Science, 47 (5):751–770, October 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312717731936.

Jones:2018:MHL

[JS18] David S. Jones and Kavita Sivaramakrishnan. Making heartlung machines work in India: Imports, indigenous innovation and the challenge of replicating cardiac surgery in Bombay, 1952–1962. Social Studies of Science, 48(4):507–539, August 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312718794034.

Jaton:2023:EAS

[JS23] Florian Jaton and Philippe Sormani. Enabling ‘AI’? The situated production of commensurabilities. Social Studies of Science, 53(5):625–634, October 1, 2023. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231194591.

Jeske:2025:MMH

[JSLS25] Melanie Jeske, Aliya Saperstein, Sandra Soo-Jin Lee, and Janet K. Shim. Marginalized measures: The harmonization of diversity in precision medicine research. Social Studies of Science, 55(2):178–208, April 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127241288498.

Jungnickel:2023:SSR

[Jun23] Kat Jungnickel. Speculative sewing: Researching, reconstructing, and re-imagining wearable technoscience. Social Studies of Science, 53(1):146–162, February 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

Jerak-Zuiderent:2012:CUM

[JZ12] Sonja Jerak-Zuiderent. Certain uncertainties: Modes of patient safety in healthcare. Social Studies of Science, 42(5): 732–752, October 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712448122.

Kahn:2008:ERD

[Kah08] Jonathan Kahn. Exploiting race in drug development: BiDil’s interim model of pharmacogenomics. Social Studies of Science, 38(5):737–758, October 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312708091928.

Kaiser:2006:WMI

[Kai06] David Kaiser. Whose mass is it anyway? Particle cosmology and the objects of theory. Social Studies of Science, 36 (4):533–564, August 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706059457.

Kantor:1978:CR

[Kan78a] Wallace Kantor. Closing response. Social Studies of Science, 8(3):359, August 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277800800310.

Kantor:1978:PEE

[Kan78b] Wallace Kantor. ‘Pseudo-effects’ in experimental physics: a study in mistaken identity. Social Studies of Science, 8(3):

355–358, August 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277800800308. See re-

ply [Mag78].

Kantor:2021:PAA

[Kan21] Brian James Kantor. Performing ‘the authoritative account’: How the BBC’s Horizon produces epistemic authority. Social Studies of Science, 51(3):463–483, June 1, 2021. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720947181.

Kang:2022:BIF

[Kan22] Edward B. Kang. Biometric imaginaries: Formatting voice, body, identity to data. Social Studies of Science, 52(4): 581–602, August 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127221079599.

Kerr:2019:PRS

[KBRB19] Anne Kerr, Tineke Broer, Emily Ross, and Sarah Cunningham Burley. Polygenic risk-stratified screening for cancer: Responsibilization in public health genomics. Social Studies of Science, 49(4):605–626, August 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312719858404.

Kaltenbrunner:2022:CPP

[KBvLA22] Wolfgang Kaltenbrunner, Kean Birch, Thed van Leeuwen, and Maria Amuchastegui. Changing publication practices and the typification of the journal article in science and technology studies. Social Studies of Science, 52(5):758–782, October 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127221110623.

Knorr-Cetina:1982:CPS

[KC82a] Karin D. Knorr-Cetina. The constructivist programme in the sociology of science: Retreats or advances? Social Studies of Science, 12(2):320–324, May 1, 1982. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631282012002007.

Knorr-Cetina:1982:RWN

[KC82b] Karin D. Knorr-Cetina. Relativism — what now? Social Studies of Science, 12(1):133–136, February 1, 1982. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631282012001007.

Knorr-Cetina:1982:SCT

[KC82c] Karin D. Knorr-Cetina. Scientific communities or transepistemic arenas of research? A critique of quasi-economic models of science. Social Studies of Science, 12(1):101–130, February 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631282012001005.

Knorr-Cetina:1985:GW

[KC85] Karin Knorr-Cetina. Germ warfare. Social Studies of Science, 15(3):577–586, August 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631285015003011.

Knorr-Cetina:1993:SCS

[KC93] Karin Knorr-Cetina. Strong constructivism — from a sociologist’s point of view: a personal addendum to Sismondo’s paper. Social Studies of Science, 23(3):555–563, August 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312793023003005. See [Sis93b] and response

[Sis93a].

Knorr-Cetina:1995:HSC

[KC95] Karin Knorr-Cetina. How superorganisms change: Consensus formation and the social ontology of high-energy physics experiments. Social Studies of Science, 25(1):119–147, February 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631295025001006.

Kerr:2007:SSP

[KCBT07] Anne Kerr, Sarah Cunningham-Burley, and Richard Tutton. Shifting subject positions: Experts and lay people in public dialogue. Social Studies of Science, 37(3):385–411, June 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706068492.

Knorr-Cetina:1997:FFH

[KCM97] Karin Knorr-Cetina and Martina Merz. Floundering or frolicking — how does ethnography fare in theoretical physics? (and what sort of ethnography?): a reply to Gale and Pinnick. Social Studies of Science, 27(1):123–131, February 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631297027001006. See [GP97].

Kaltenbrunner:2019:FGI

[KdR19] Wolfgang Kaltenbrunner and Sarah de Rijcke. Filling in the gaps: The interpretation of curricula vitae in peer review. Social Studies of Science, 49(6):863–883, December 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312719864164.

Kearney:1973:ERM

[Kea73] H. F. Kearney. Essay reviews: Merton revisited. Social Studies of Science, 3(1):72–78, January 1, 1973. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277300300106.

Keith:1995:RL

[Kei95] William Keith. Response to Labinger. Social Studies of Science, 25(2):321–324, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631295025002009. See

[Lab95b].

Keller:1989:JWD

[Kel89] Evelyn Fox Keller. Just what is so difficult about the concept of gender as a social category? (response to Richards and Schuster). Social Studies of Science, 19(4):721–724, November 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631289019004012. See reply [RS89b].

Kennefick:2000:SCT

[Ken00] Daniel Kennefick. Star crushing: Theoretical practice and the theoreticians’ regress. Social Studies of Science, 30(1): 5–40, February 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631200030001001.

Kent:2013:IBU

[Ken13] Michael Kent. The importance of being Uros: Indigenous identity politics in the genomic age. Social Studies of Science, 43 (4):534–556, August 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712468520.

Kenney:2015:CAA

[Ken15] Martha Kenney. Counting, accounting, and accountability: Helen Verran’s relational empiricism. Social Studies of Science, 45 (5):749–771, October 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312715607413.

Kennedy:2017:MMC

[Ken17] Devin Kennedy. The machine in the market: Computers and the infrastructure of price at the New York Stock Exchange, 1965–1975. Social Studies of Science, 47(6):888–917, December 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312717739367.

Kerr:2000:RCG

[Ker00] Anne Kerr. (re)constructing genetic disease: The clinical continuum between cystic fibrosis and male infertility. Social Studies of Science, 30(6):847–894, December 1, 2000. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030006002.

Kelly:2012:PPI

[KG12] Kimberly Kelly and Linda Grant. Penalties and premiums: The impact of gender, marriage, and parenthood on faculty salaries in science, engineering and mathematics (SEM) and non-SEM fields. Social Studies of Science, 42(6):869–896, December 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312712457111.

Kent:2015:BGN

[KGDLB^{+}15] Michael Kent, Vivette García-Deister, Carlos López-Beltrán, Ricardo Ventura Santos, Ernesto Schwartz-Marín, and Peter Wade. Building the genomic nation: ‘Homo Brasilis’ and the ‘Genoma Mexicano’ in comparative cultural perspective. Social Studies of Science, 45(6):839–861, December 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312715611262.

Kasperowski:2018:ECO

[KH18] Dick Kasperowski and Thomas Hillman. The epistemic culture in an online citizen science project: Programs, antiprograms and epistemic subjects. Social Studies of Science, 48(4):564–588, August 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312718778806.

Kalender:2019:CWC

[KH19] Ute Kalender and Christine Holmberg. Courtesy work: Care practices for quality assurance in a cohort study. Social Studies of Science, 49(4):583–604, August 1, 2019. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312719863139.

Kasperowski:2022:MPT

[KH22] Dick Kasperowski and Niclas Hagen. Making particularity travel: Trust and citizen science data in Swedish environmental governance. Social Studies of Science, 52(3):447–462, June 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/03063127221085241.

Klenk:2010:ESC

[KHM10] Nicole L. Klenk, Gordon M. Hickey, and James Ian MacLellan. Evaluating the social capital accrued in large research networks: The case of the sustainable forest management network (1995– 2009). Social Studies of Science, 40(6):931–960, December 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312710374130.

Kidd:1988:SEJ

[Kid88] J. S. Kidd. Scholarly excess and journalistic restraint in the popular treatment of cannibalism. Social Studies of Science, 18 (4):749–754, November 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631288018004007.

Kaiser:2004:STT

[KIH04] David Kaiser, Kenji Ito, and Karl Hall. Spreading the tools of theory: Feynman diagrams in the USA, Japan, and the Soviet Union. Social Studies of Science, 34(6):879–922, December 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL http://www.jstor.org/stable/4144350;
 https://journals.sagepub.com/doi/pdf/10.1177/0306312704046628.

Kimmelman:1983:ABA

[Kim83] Barbara A. Kimmelman. The American Breeders’ Association: Genetics and eugenics in an agricultural context, 1903–

13. Social Studies of Science, 13(2):163–204, May 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631283013002001.

Kim:1991:RJP

[Kim91] Kyung-Man Kim. On the reception of Johannsen’s pure line theory: Toward a sociology of scientific validity. Social Studies of Science, 21(4):649–679, November 1, 1991. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631291021004002.

Kim:1994:BRN

[Kim94] Kyung-Man Kim. Book review: Natural versus Normative Rationality: Reassessing the Strong Programme in the Sociology of Knowledge: David Bloor, Knowledge and Social Imagery, second edition, (London & Chicago, IL: The University of Chicago Press, 1991), xi + 203 pp., $32.00, $13.95/£11.25 pbk. ISBN 0-226-06096-9 (-06097-7 pbk). Social Studies of Science, 24 (2):391–403, May 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631279402400210.

Kim:2007:AME

[Kim07] Jongyoung Kim. Alternative medicine’s encounter with laboratory science: The scientific construction of Korean medicine in a global age. Social Studies of Science, 37(6):855–880, December 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312707076600.

Kim:2014:AGG

[Kim14] Mi Gyung Kim. Archeology, genealogy, and geography of experimental philosophy. Social Studies of Science, 44(1): 150–162, February 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312713507329.

Kim:2016:SPC

[Kim16] Eun-Sung Kim. The sensory power of cameras and noise meters for protest surveillance in South Korea. Social Studies of Science, 46(3):396–416, June 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312716648403.

King:2002:SDC

[Kin02] Nicholas B. King. Security, disease, commerce: Ideologies of postcolonial global health. Social Studies of Science, 32(5–6): 763–789, December 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631270203200507.

Kirejczyk:1999:PCH

[Kir99] Marta Kirejczyk. Parliamentary cultures and human embryos: The Dutch and British debates compared. Social Studies of Science, 29(6):889–912, December 1, 1999. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029006004.

Kirby:2003:SCF

[Kir03] David A. Kirby. Science consultants, fictional films, and scientific practice. Social Studies of Science, 33(2):231–268, April 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/03063127030332015.

Kirby:2010:FND

[Kir10] David Kirby. The future is now: Diegetic prototypes and the role of popular films in generating real-world technological development. Social Studies of Science, 40(1):41–70, February 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312709338325.

Kirkland:2012:CBA

[Kir12] Anna Kirkland. Credibility battles in the autism litigation. Social Studies of Science, 42(2):237–261, April 1, 2012. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312711435832.

Kistner:2013:WFW

[Kis13] Kelly Kistner. ‘A word factory was wanted’: Organizational objectivity in the making of the Oxford English Dictionary. Social Studies of Science, 43(6):801–828, December 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713496603.

Kinchy:2003:OCD

[KK03] Abby J. Kinchy and Daniel Lee Kleinman. Organizing credibility: Discursive and organizational orthodoxy on the borders of ecology and politics. Social Studies of Science, 33(6): 869–896, December 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312703336003.

Kameda:2021:IDT

[KKLL21] Koichi Kameda, Ann H. Kelly, Javier Lezaun, and Ilana Löwy. Imperfect diagnosis: The truncated legacies of Zika testing. Social Studies of Science, 51(5):683–706, October 1, 2021.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/03063127211035492.

Kohli-Laven:2011:CCT

[KLBKC11] Nina Kohli-Laven, Pascale Bourret, Peter Keating, and Alberto Cambrosio. Cancer clinical trials in the era of genomic signatures: Biomedical innovation, clinical utility, and regulatoryscientific hybrids. Social Studies of Science, 41(4):487–513, August 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312711398741.

Kleinman:1991:CPS

[Kle91] Daniel Lee Kleinman. Conceptualizing the politics of science: a response to Cambrosio, Limoges and Pronovost. Social Studies of Science, 21(4):769–774, November 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/030631291021004007. See [CLP90] and reply [CLP91].

Klintman:2002:GMG

[Kli02] Mikael Klintman. The genetically modified (GM) food labelling controversy: Ideological and epistemic crossovers. Social Studies of Science, 32(1):71–91, February 1, 2002. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312702032001004.

Kline:2009:WCC

[Kli09] Ronald Kline. Where are the cyborgs in cybernetics? Social Studies of Science, 39(3):331–362, June 1, 2009. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312708101046.

Kenny:2019:SPF

[KLW19] Catherine Kenny, Max Liboiron, and Sara Ann Wylie. Seeing power with a flashlight: DIY thermal sensing technology in the classroom. Social Studies of Science, 49(1):3–28, February 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312718823282.

Kern:1983:SUP

[KMH83] Leslie H. Kern, Herbert L. Mirels, and Virgil G. Hinshaw. Scientists’ understanding of propositional logic: an experimental investigation. Social Studies of Science, 13(1):131–146, February 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631283013001007.

Knaapen:2013:BEB

[Kna13] Loes Knaapen. Being ‘evidence-based’ in the absence of evidence: The management of non-evidence in guideline development. Social Studies of Science, 43(5):681–706, October 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713483679.

Kochan:2010:CES

[Koc10a] Jeff Kochan. Contrastive explanation and the ‘strong programme’ in the sociology of scientific knowledge. Social Studies of Science, 40(1):127–144, February 1, 2010. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312709104780.

Kochan:2010:LH

[Koc10b] Jeff Kochan. Latour’s Heidegger. Social Studies of Science, 40 (4):579–598, August 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312709360263.

Koertge:1999:ZSA

[Koe99] Noretta Koertge. The zero-sum assumption and the symmetry thesis. Social Studies of Science, 29(5):777–784, October 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631299029005005. See reply [Col99a].

Komasova:2022:AST

[Kom22] Sarah Komasová. Airport security as translation through division and movement. Social Studies of Science, 52(1):35–52, February 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/full/10.1177/03063127211037733.

Kopp:1979:OAS

[Kop79] Carolyn Kopp. The origins of the American scientific debate over fallout hazards. Social Studies of Science, 9(4): 403–422, November 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277900900402.

Kowal:2013:ODI

[Kow13] Emma Kowal. Orphan DNA: Indigenous samples, ethical biovalue and postcolonial science. Social Studies of Science, 43 (4):577–597, August 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712471739.

Krige:1985:CIM

[KP85] John Krige and Dominique Pestre. A critique of Irvine and Martin’s methodology for evaluating big science. Social Studies of Science, 15(3):525–539, August 1, 1985. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631285015003006. See [IM85, MI85].

Krohn:2000:ZTC

[KR00] Wolfgang Krohn and Diederick Raven. The ‘Zilsel Thesis’ in the context of Edgar Zilsel’s research programme. Social Studies of Science, 30(6):925–933, December 1, 2000. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030006005.

Kaplan:2011:BET

[KR11] Sarah Kaplan and Joanna Radin. Bounding an emerging technology: Para-scientific media and the Drexler–Smalley debate about nanotechnology. Social Studies of Science, 41(4): 457–485, August 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312711402722.

Krauze:1972:SIS

[Kra72] Tadeusz K. Krauze. Social and intellectual structures of sciencea mathematical analysis. Social Studies of Science, 2(4):

369–378, October 1, 1972. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277200200405.

Kranakis:1989:SDE

[Kra89] Eda Kranakis. Social determinants of engineering practice: a comparative view of France and America in the Nineteenth Century. Social Studies of Science, 19(1):5–70, February 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019001001.

Kreilkamp:1971:HRW

[Kre71] Karl Kreilkamp. Hindsight and the real world of science policy. Social Studies of Science, 1(1):43–66, January 1, 1971. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277100100104.

Kreilkamp:1973:TTS

[Kre73] Karl Kreilkamp. Towards a theory of science policy. Social Studies of Science, 3(1):3–29, January 1, 1973. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277300300102.

Krige:1978:PEA

[Kri78] John Krige. Popper’s epistemology and the autonomy of science. Social Studies of Science, 8(3):287–307, August 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277800800303.

Krieger:1987:ESP

[Kri87] Martin H. Krieger. The elementary structures of particles. Social Studies of Science, 17(4):749–752, November 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631287017004007.

Krohn:1982:GRC

[Kro82] Roger Krohn. On Gieryn on the ‘relativist/constructivist’ programme in the sociology of science: Naïveté and reaction. Social Studies of Science, 12(2):325–328, May 1, 1982.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/030631282012002008. See [Gie82b] and reply [Gie82a].

Krolokke:2019:LCK

[Kro19] Charlotte Kroløkke. Life in the cryo-kennel: The ‘exceptional’ life of frozen pet DNA. Social Studies of Science, 49(2):162–179, April 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312719837610.

Kowal:2013:IBP

[KRR13] Emma Kowal, Joanna Radin, and Jenny Reardon. Indigenous body parts, mutating temporalities, and the half-lives of postcolonial technoscience. Social Studies of Science, 43 (4):465–483, August 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312713490843.

Kruse:2013:BAF

[Kru13] Corinna Kruse. The Bayesian approach to forensic evidence: Evaluating, communicating, and distributing responsibility. Social Studies of Science, 43(5):657–680, October 1, 2013. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712472572.

Kachaunov:1979:SSS

[KS79] Stefan Kachaunov and Kostadinka Simeonova. Social studies of science in Bulgaria. Social Studies of Science, 9(1):91–99, February 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277900900105.

Koester:1982:TSP

[KSW82] David Koester, Daniel Sullivan, and D. Hywel White. Theory selection in particle physics: a quantitative case study of the evolution of weak-electromagnetic unification theory. Social Studies of Science, 12(1):73–100, February 1, 1982. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631282012001004.

Kuchinskaya:2013:TIF

[Kuc13] Olga Kuchinskaya. Twice invisible: Formal representations of radiation danger. Social Studies of Science, 43(1):78–96, February 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312712465356.

Kuklick:1994:BRD

[Kuk94] Henrika Kuklick. Book review: The Definition of a Profession: The Authority of Metaphor in the History of Intelligence Testing, 1980–1930. Social Studies of Science, 24(1):185–188, February 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631279402400114.

Kumaki:2024:IME

[Kum24] Hiroko Kumaki. Intra-mediary expertise: Trans-science and expert understanding of the public. Social Studies of Science, 54 (4):512–535, August 1, 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127241229076.

Kusch:2004:RMC

[Kus04a] Martin Kusch. Reply to my critics. Social Studies of Science, 34(4):615–620, August 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312704046600. See [Blo04,

Sha04].

Kusch:2004:RSS

[Kus04b] Martin Kusch. Rule-scepticism and the sociology of scientific knowledge: The Bloor–Lynch debate revisited. Social Studies of Science, 34(4):571–591, August 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/0306312704044168. See replies [Blo04, Sha04].

Kusiak:2010:IRC

[Kus10] Pauline Kusiak. Instrumentalized rationality, cross-cultural mediators, and civil epistemologies of late colonialism. Social Studies of Science, 40(6):871–902, December 1, 2010. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312710372436.

Koch:2021:MSI

[KVA21] Tomás Koch, Raf Vanderstraeten, and Ricardo Ayala. Making science international: Chilean journals and communities in the world of science. Social Studies of Science, 51(1): 121–138, February 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312720949709.

Kent:2015:GAR

[KW15] Michael Kent and Peter Wade. Genetics against race: Science, politics and affirmative action in Brazil. Social Studies of Science, 45(6):816–838, December 1, 2015. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312715610217.

Kwa:1987:RNM

[Kwa87] Chunglin Kwa. Representations of nature mediating between ecology and science policy: The case of the international biological programme. Social Studies of Science, 17(3):413–442, August 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631287017003002.

Kwa:2005:LEG

[Kwa05] Chunglin Kwa. Local ecologies and global science: Discourses and strategies of the International Geosphere–Biosphere Programme. Social Studies of Science, 35(6):923–950, December 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312705052100.

Kyvik:1990:MSP

[Kyv90] Svein Kyvik. Motherhood and scientific productivity. Social Studies of Science, 20(1):149–160, February 1, 1990. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020001005.

Loh:2019:HFR

[LA19] Shi Lin Loh and Sulfikar Amir. Healing Fukushima: Radiation hazards and disaster medicine in post-3.11 Japan. Social Studies of Science, 49(3):333–354, June 1, 2019. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312719854540.

Labinger:1995:PDE

[Lab95a] Jay A. Labinger. Out of the Petri dish endlessly rocking: Reply to my responders. Social Studies of Science, 25(2):341–348, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631295025002014. See [Lab95b].

Labinger:1995:SCV

[Lab95b] Jay A. Labinger. Science as culture: a view from the Petri dish. Social Studies of Science, 25(2):285–306, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631295025002004. See comments [Col95, Ful95,

Hak95, Kei95, Lyn95a, Mar95, Sto95] and reply [Lab95a].

Labinger:2004:LE

[Lab04] Jay A. Labinger. Logic and the Editor. Social Studies of Science, 34(1):91–92, February 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312704041341.

Lachenal:2016:HPE

[Lac16] Guillaume Lachenal. At home in the postcolony: Ecology, empire and domesticity at the Lamto field station, Ivory Coast. Social Studies of Science, 46(6):877–893, December 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716649800.

Lagesen:2007:SNS

[Lag07] Vivian Anette Lagesen. The strength of numbers: Strategies to include women into computer science. Social Studies of Science, 37(1):67–92, February 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706063788.

Lagesen:2012:RGA

[Lag12] Vivian Anette Lagesen. Reassembling gender: Actor-network theory (ANT) and the making of the technology in gender. Social Studies of Science, 42(3):442–448, June 1, 2012. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712437078.

Lahsen:2005:SSU

[Lah05] Myanna Lahsen. Seductive simulations? Uncertainty distribution around climate models. Social Studies of Science, 35(6): 895–922, December 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312705053049.

Lahn:2021:CCC

[Lah21] Bård Lahn. Changing climate change: The carbon budget and the modifying-work of the IPCC. Social Studies of Science, 51 (1):3–27, February 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312720941933.

Lakoff:2004:AGA

[Lak04] Andrew Lakoff. The anxieties of globalization: Antidepressant sales and economic crisis in Argentina. Social Studies of Science, 34(2):247–269, April 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312704042624.

Lakoff:2017:FAM

[Lak17] Andrew Lakoff. A fragile assemblage: Mutant bird flu and the limits of risk assessment. Social Studies of Science, 47 (3):376–397, June 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312716666420.

Lambert:2006:FFK

[Lam06] Kevin Lambert. Fuller’s folly, Kuhnian paradigms, and intelligent design. Social Studies of Science, 36(6):835–842, December 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312706067899.

Lam:2010:ITT

[Lam10a] Alice Lam. From ‘Ivory Tower traditionalists’ to ‘entrepreneurial scientists’?: Academic scientists in fuzzy university– industry boundaries. Social Studies of Science, 40(2):307–340, April 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312709349963.

Lampland:2010:FNF

[Lam10b] Martha Lampland. False numbers as formalizing practices. Social Studies of Science, 40(3):377–404, June 1, 2010. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312709359963.

Lankford:1981:AAN

[Lan81] John Lankford. Amateurs and astrophysics: a neglected aspect in the development of a scientific specialty. Social Studies of Science, 11(3):275–303, August 1, 1981. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278101100301.

Landstrom:2000:OPS

[Lan00] Catharina Landström. The ontological politics of staying true to complexity. Social Studies of Science, 30(3):475–480, June 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631200030003007.

Landstrom:2001:ARC

[Lan01a] Catharina Landström. The Australian rabbit calicivirus disease program: a story about technoscience and culture. Social Studies of Science, 31(6):912–949, December 1, 2001. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031006005.

Langfeldt:2001:DMC

[Lan01b] Liv Langfeldt. The decision-making constraints and processes of grant peer review, and their effects on the review outcome. Social Studies of Science, 31(6):820–841, December 1, 2001.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031006002.

Langlitz:2009:PPB

[Lan09] Nicolas Langlitz. Pharmacovigilance and post-black market surveillance. Social Studies of Science, 39(3):395–420, June 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312708101977.

Lanzarotta:2020:ERB

[Lan20] Tess Lanzarotta. Ethics in retrospect: Biomedical research, colonial violence, and Iñupiat sovereignty in the Alaskan Arctic. Social Studies of Science, 50(5):778–801, October 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720943678.

Lappe:2016:MBE

[Lap16] Martine Lappé. The maternal body as environment in autism science. Social Studies of Science, 46(5):675–700, October 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312716659372.

Latour:1986:WLP

[Lat86] Bruno Latour. Will the last person to leave the social studies of science please turn on the tape-recorder? Social Studies of Science, 16(3):541–548, August 1, 1986. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631286016003008.

Latour:1988:RAE

[Lat88] Bruno Latour. A relativistic account of Einstein’s relativity. Social Studies of Science, 18(1):3–44, February 1, 1988. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL http://www.jstor.org/stable/285375;   https://
journals.sagepub.com/doi/pdf/10.1177/030631288018001001.

Latour:1992:RCG

[Lat92] Bruno Latour. A reply to Carlson and Gorman. Social Studies of Science, 22(1):91–95, February 1, 1992. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312792022001005. See [LMT92, CG92].

Latour:2003:MDO

[Lat03] Bruno Latour. In memoriam: David Owen Edge. Social Studies of Science, 33(2):191–192, April 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127030332010.

Latour:2007:TAP

[Lat07] Bruno Latour. Turning around politics: a note on Gerard de Vries’ paper. Social Studies of Science, 37(5):811–820, October 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312707081222. See [dV07].

Latour:2008:REN

[Lat08] Bruno Latour. Review essay: The netz-works of Greek deductions. Social Studies of Science, 38(3):441–459, June 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312707087973.

Latour:2010:CP

[Lat10] Bruno Latour. Coming out as a philosopher. Social Studies of Science, 40(4):599–608, August 1, 2010. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312710367697.

Latour:2013:BIB

[Lat13] Bruno Latour. Biography of an inquiry: On a book about modes of existence. Social Studies of Science, 43(2):287–301, April 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312712470751.

Laudan:1982:NCB

[Lau82] Larry Laudan. A note on Collins’s blend of Relativism and empiricism. Social Studies of Science, 12(1):131–132, February 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631282012001006. See [Col81b, Ste87].

Laurent:1984:SSP

[Lau84] John Laurent. Science, society and politics in Late Nineteenth-Century England: a further look at Mechanics’ Institutes. Social Studies of Science, 14(4):585–619, November 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014004004.

Laudan:1987:DIC

[Lau87] Rachel Laudan. Drifting interests and colliding continents: a response to Stewart. Social Studies of Science, 17(2):317–321, May 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631287017002005. See [Ste86] and reply

[Ste87].

Laurent:1990:BRR

[Lau90] John Laurent. Book review: Robert J. Richards, Darwin and the Emergence of Evolutionary Theories of Mind and Behaviour. Social Studies of Science, 20(1):161–165, February 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020001006.

Laursen:2013:CPA

[Lau13] Ditte Laursen. Co-participation among school children around a computer-based exhibit. Social Studies of Science, 43(1): 97–117, February 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712455114.

Laurent:2016:PEM

[Lau16] Brice Laurent. Political experiments that matter: Ordering democracy from experimental sites. Social Studies of Science, 46(5):773–794, October 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312716668587.

Laurent:2025:CDS

[Lau25] Brice Laurent. Can democracy save children’s lives? Addressing the constitutional problem of expertise. Social Studies of Science, 55(2):288–294, April 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127241310461.

Law:1973:DSS

[Law73] John Law. The development of specialties in science: the case of X-ray protein crystallography. Social Studies of Science, 3 (3):275–303, July 1, 1973. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277300300303.

Law:1977:PFA

[Law77] John Law. Prophecy failed (for the actors)!: a note on ‘Recovering Relativity’. Social Studies of Science, 7(3):367–372, August 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277700700305. See [CC76].

Law:1980:FIS

[Law80] John Law. Fragmentation and investment in sedimentology. Social Studies of Science, 10(1):1–22, February 1, 1980. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278001000101.

Law:1981:BES

[Law81] John Law. On benthic ecology, sociological determinism, and other matters. Social Studies of Science, 11(3):398–401, August 1, 1981. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631278101100306.

Lawrence:1984:MNS

[Law84] Christopher Lawrence. Making the nervous system. Social Studies of Science, 14(1):153–158, February 1, 1984. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014001011.

Law:1987:TTS

[Law87] John Law. Technology transfer and social engineering. Social Studies of Science, 17(3):564–569, August 1, 1987. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631287017003008.

Lawrence:1997:SSM

[Law97] Christopher Lawrence. Science, skill and medical art. Social Studies of Science, 27(1):171–179, February 1, 1997. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631297027001010.

Lawson:2011:SEC

[Law11] Sean Lawson. Surfing on the edge of chaos: Nonlinear science and the emergence of a doctrine of preventive war in the US. Social Studies of Science, 41(4):563–584, August 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312711402866.

Lawless:2013:LTD

[Law13] Christopher J. Lawless. The low template DNA profiling controversy: Biolegality and boundary work among forensic scientists. Social Studies of Science, 43(2):191–214, April 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712465665.

Lawless:2020:AAS

[Law20a] Christopher Lawless. Assembling airspace: The Single European Sky and contested transnationalities of European air traffic management. Social Studies of Science, 50(4):680–704, August 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312720920704.

Lawrence:2020:HGT

[Law20b] Christopher Lawrence. Heralds of global transparency: Remote sensing, nuclear fuel-cycle facilities, and the modularity of imagination. Social Studies of Science, 50(4):508–541, August 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312719879769.

Lawrence:2025:GAS

[Law25] Christopher Lawrence. Gathering around a satellite image: Visual media cycles of the nuclear nonproliferation complex. Social Studies of Science, 55(1):3–36, February 2025. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127241274793.

Lee:2005:IRC

[LB05] Sooho Lee and Barry Bozeman. The impact of research collaboration on scientific productivity. Social Studies of Science, 35 (5):673–702, October 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312705052359.

Lisle:2019:MLB

[LB19] Debbie Lisle and Mike Bourne. The many lives of border automation: Turbulence, coordination and care. Social Studies of Science, 49(5):682–706, October 1, 2019. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312719870868.

Lomnitz:1999:BAT

[LC99] Larissa Adler Lomnitz and Laura Cházaro. Basic, applied and technological research: Computer science and applied mathematics at the National Autonomous University of Mexico. Social Studies of Science, 29(1):113–134, February 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029001005.

Lynch:2005:STS

[LC05] Michael Lynch and Simon Cole. Science and technology studies on trial: Dilemmas of expertise. Social Studies of Science, 35(2):269–311, April 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312705048715.

Leahey:2013:SSA

[LC13] Erin Leahey and Cindy L. Cain. Straight from the source: Accounting for scientific success. Social Studies of Science, 43(6): 927–951, December 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312713484820.

Ledger:1982:KR

[LcaM82] Christine Ledger and colleagues at Murdoch. Keith Roby 1941–81. Social Studies of Science, 12(2):–, May 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631282012002010.

LeDantec:2013:IFP

[LD13] Christopher A. Le Dantec and Carl DiSalvo. Infrastructuring and the formation of publics in participatory design. Social Studies of Science, 43(2):241–264, April 1, 2013. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712471581.

Lariviere:2016:CDL

[LDM^{+}16] Vincent Larivière, Nadine Desrochers, Benoît Macaluso, Philippe Mongeon, Adèle Paul-Hus, and Cassidy R. Sugimoto. Contributorship and division of labor in knowledge production. Social Studies of Science, 46(3):417–435, June 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716650046.

Lave:2010:PSR

[LDR10] Rebecca Lave, Martin Doyle, and Morgan Robertson. Privatizing stream restoration in the US. Social Studies of Science, 40 (5):677–703, October 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312710379671.

Leem:2016:APB

[Lee16] So Yeon Leem. The anxious production of beauty: Unruly bodies, surgical anxiety and invisible care. Social Studies of Science, 46(1):34–55, February 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312715615971.

Lehman:2018:SRS

[Leh18] Jessica Lehman. From ships to robots: The social relations of sensing the world ocean. Social Studies of Science, 48(1): 57–79, February 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312717743579.

Leiss:1975:IS

[Lei75] William Leiss. Ideology and science. Social Studies of Science, 5(2):193–200, May 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277500500205.

Leibler:2014:DES

[Lei14] Anat Leibler. Disciplining ethnicity: Social sorting intersects with political demography in Israel’s pre-state period. Social Studies of Science, 44(2):271–292, April 1, 2014. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713509309.

Lenoir:1979:QFS

[Len79] Timothy Lenoir. Quantitative foundations for the sociology of science: On linking blockmodeling with co-citation analysis. Social Studies of Science, 9(4):455–480, November 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277900900404.

Lentacker:2016:SED

[Len16] Antoine Lentacker. The symbolic economy of drugs. Social Studies of Science, 46(1):140–156, February 1, 2016. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312715624983.

Leonardi:2010:RLM

[Leo10] Paul M. Leonardi. From road to lab to math: The coevolution of technological, regulatory, and organizational innovations for automotive crash testing. Social Studies of Science, 40(2):243–274, April 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312709346079.

Leonelli:2012:WHE

[Leo12] Sabina Leonelli. When humans are the exception: Crossspecies databases at the interface of biological and clinical research. Social Studies of Science, 42(2):214–236, April 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312711436265.

Leroux:2018:WBH

[Ler18] Darryl Leroux. ‘We’ve been here for 2,000 years’: White settlers, Native American DNA and the phenomenon of indigenization. Social Studies of Science, 48(1):80–100, February 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312717751863.

Levitan:1979:SST

[Lev79] Karen B. Levitan. Scientific societies and their journals: Biomedical scientists assess the relationship. Social Studies of Science, 9(3):393–400, August 1, 1979. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277900900307.

Levin:1984:SSS

[Lev84] Alexey Levin. Soviet science studies: a dissident view. Social Studies of Science, 14(3):451–467, August 1, 1984. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014003007.

Levidow:2001:PUR

[Lev01] Les Levidow. Precautionary uncertainty: Regulating GM crops in Europe. Social Studies of Science, 31(6):842–874, December 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631201031006003.

Levin:2014:MSE

[Lev14] Nadine Levin. Multivariate statistics and the enactment of metabolic complexity. Social Studies of Science, 44(4):555–578, August 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/full/10.1177/0306312714524845.

Levy:2022:ATH

[Lev22] Moran Levy. Adequate trials: How the search for a cure shaped leukemia diagnosis. Social Studies of Science, 52(6):878–903, December 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

Lewis:1972:SAR

[Lew72] Robert A. Lewis. Some aspects of the research and development effort of the Soviet Union, 1924–35. Social Studies of Science, 2(2):153–179, April 1, 1972. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277200200202.

Lewis:1980:RCD

[Lew80] Gwendolyn L. Lewis. The relationship of conceptual development to consensus: an exploratory analysis of three subfields. Social Studies of Science, 10(3):285–308, August 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278001000302.

Lewenstein:1995:FFC

[Lew95] Bruce V. Lewenstein. From fax to facts: Communication in the cold fusion saga. Social Studies of Science, 25(3):403–436, August 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631295025003001.

Lewenstein:2001:EM

[Lew01] Bruce V. Lewenstein. Expertise in the media. Social Studies of Science, 31(3):441–444, June 1, 2001. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031003011.

Leydesdorff:1991:SEN

[Ley91] Loet Leydesdorff. In search of epistemic networks. Social Studies of Science, 21(1):75–110, February 1, 1991. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631291021001005.

Lezaun:2006:CNO

[Lez06] Javier Lezaun. Creating a new object of government: Making genetically modified organisms traceable. Social Studies of Science, 36(4):499–531, August 1, 2006. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706059461.

Lezaun:2010:EII

[Lez10] Javier Lezaun. Eloquence and incommensurability: an investigation into the grammar of irreconcilable differences. Social Studies of Science, 40(3):349–375, June 1, 2010. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312709358119.

Lichtenberg:1989:ISD

[Lic89] Frank R. Lichtenberg. The impact of the strategic defense initiative on US civilian R&D investment and industrial competitiveness. Social Studies of Science, 19(2):265–282, May 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019002003.

Liebenau:1987:BSP

[Lie87] Jonathan Liebenau. The British success with penicillin. Social Studies of Science, 17(1):69–86, February 1, 1987. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631287017001003.

Lindsey:1978:CQR

[Lin78] Duncan Lindsey. The corrected quality ratio: a composite index of scientific contribution to knowledge. Social Studies of Science, 8(3):349–354, August 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277800800307.

Lindsey:1980:PCM

[Lin80] Duncan Lindsey. Production and citation measures in the sociology of science: The problem of multiple authorship. Social Studies of Science, 10(2):145–162, May 1, 1980. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278001000202.

Lindee:2016:SSH

[Lin16] Susan Lindee. Survivors and scientists: Hiroshima, Fukushima, and the radiation effects research foundation, 1975–2014. Social Studies of Science, 46(2):184–209, April 1, 2016. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716632933.

Lipsey:1974:PPP

[Lip74] Mark W. Lipsey. Psychology: Preparadigmatic, postparadigmatic, or misparadigmatic? Social Studies of Science, 4(4): 406–410, October 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277400400405.

Lipp:2023:CRH

[Lip23] Benjamin Lipp. Caring for robots: How care comes to matter in human-machine interfacing. Social Studies of Science, 53 (5):660–685, October 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127221081446.

Livingston:1999:CP

[Liv99] Eric Livingston. Cultures of proving. Social Studies of Science, 29(6):867–888, December 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631299029006003.

Livingston:2006:CP

[Liv06] Eric Livingston. The context of proving. Social Studies of Science, 36(1):39–68, February 1, 2006. CODEN SSSCDH.

ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/0306312705053055.

Lynch:1998:CIS

[LJ98] Michael Lynch and Sheila Jasanoff. Contested identities: Science, law and forensic practice. Social Studies of Science, 28 (5–6):675–686, October 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631298028005001.

Lei:1999:CNA

[lL99] Sean Hsiang lin Lei. From Changshan to a new anti-malarial drug: Re-networking Chinese drugs and excluding Chinese doctors. Social Studies of Science, 29(3):323–358, June 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029003001.

Law:2013:SFN

[LL13] John Law and Marianne Elisabeth Lien. Slippery: Field notes in empirical ontology. Social Studies of Science, 43(3):363–378, June 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312712456947.

LaPorte:1975:PAT

[LM75] Todd La Porte and Daniel Metlay. Public attitudes toward present and future technologies: Satisfactions and apprehensions. Social Studies of Science, 5(4):373–398, November 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277500500401.

Leahy:1980:RFP

[LM80] Peter J. Leahy and Allan Mazur. The rise and fall of public opposition in specific social movements. Social Studies of Science, 10(3):259–284, August 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631278001000301.

Lightman:1989:CCB

[LM89] Alan P. Lightman and Jon D. Miller. Contemporary cosmological beliefs. Social Studies of Science, 19(1):127–136, February 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631289019001004.

Long:1980:PJA

[LMA80] J. Scott Long, Robert McGinnis, and Paul D. Allison. The problem of junior-authored papers in constructing citation counts. Social Studies of Science, 10(2):127–143, May 1, 1980. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278001000201.

Lave:2010:ISN

[LMR10] Rebecca Lave, Philip Mirowski, and Samuel Randalls. Introduction: STS and neoliberal science. Social Studies of Science, 40(5):659–675, October 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312710378549.

Latour:1992:NST

[LMT92] Bruno Latour, Philippe Mauguin, and Geneviève Teil. A note on socio-technical graphs. Social Studies of Science, 22(1): 33–57, February 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312792022001002. See

response [CG92] and reply [Lat92].

Lajoie-OMalley:2023:CER

[LOBB23] Alana Lajoie-O’Malley, Kelly Bronson, and Gwendolyn Blue. ‘Consent’ as epistemic recognition: Indigenous knowledges, Canadian impact assessment, and the colonial liberal democratic order. Social Studies of Science, 53(4):545–571, August 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127231177311.

Lohan:2000:CTF

[Loh00] Maria Lohan. Constructive tensions in feminist technology studies. Social Studies of Science, 30(6):895–916, December 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030006003.

Long:1971:GSC

[Lon71] T. Dixon Long. The government of science: a comparative approach. Social Studies of Science, 1(3–4):263–286, October 1, 1971. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277100100302.

Lopez:2022:LTL

[Lóp22] Iván Chaar López. Latina/o/e technoscience: Labor, race, and gender in cybernetics and computing. Social Studies of Science, 52(6):829–852, December 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

Lorimer:2008:CCA

[Lor08] Jamie Lorimer. Counting corncrakes: The affective science of the UK corncrake census. Social Studies of Science, 38 (3):377–405, June 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312707084396.

Lorimer:2017:ASG

[Lor17] Jamie Lorimer. The anthropo-scene: a guide for the perplexed. Social Studies of Science, 47(1):117–142, February 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716671039.

Low:1989:BFS

[Low89] Morris Fraser Low. The butterfly and the Frigate: Social studies of science in Japan. Social Studies of Science, 19(2):313–342, May 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631289019002006.

Littlefield:2011:RET

[LP11] Melissa M. Littlefield and Anne Pollock. Review essay: Troubling with ‘the ethics of the thing’ in Culturing Life: How Cells Became Technologies and The Immortal Life of Henrietta Lacks. Social Studies of Science, 41(4):609–618, August 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312711408485.

Lincoln:2012:MES

[LPKL12] Anne E. Lincoln, Stephanie Pincus, Janet Bandows Koster, and Phoebe S. Leboy. The Matilda Effect in science: Awards and prizes in the US, 1990s and 2000s. Social Studies of Science, 42(2):307–320, April 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312711435830.

Lee:2001:HDD

[LR01] Stuart Lee and Wolff-Michael Roth. How ditch and drain become a healthy creek: Re-presentations, translations and agency during the re/design of a watershed. Social Studies of Science, 31(3):315–356, June 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631201031003001.

Leahey:2008:RSC

[LR08] Erin Leahey and Ryan C. Reikowsky. Research specialization and collaboration patterns in sociology. Social Studies of Science, 38(3):425–440, June 1, 2008. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312707086190.

Levidow:2020:SIL

[LR20] Les Levidow and Sujatha Raman. Sociotechnical imaginaries of low-carbon waste-energy futures: UK techno-market fixes displacing public accountability. Social Studies of Science, 50 (4):609–641, August 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312720905084.

Lomnitz:1987:PRP

[LRC87] Larissa Adler Lomnitz, Martha W. Rees, and Leon Cameo. Publication and referencing patterns in a Mexican research institute. Social Studies of Science, 17(1):115–133, February 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631287017001005.

Lafuente:2004:USS

[LS04] Antonio Lafuente and Tiago Saraiva. The urban scale of science and the enlargement of Madrid (1851–1936). Social Studies of Science, 34(4):531–569, August 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704042638.

Leuenberger:2010:PMC

[LS10] Christine Leuenberger and Izhak Schnell. The politics of maps: Constructing national territories in Israel. Social Studies of Science, 40(6):803–842, December 1, 2010. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312710370377.

Laube:2020:CIE

[LSS20] Stefan Laube, Jan Schank, and Thomas Scheffer. Constitutive invisibility: Exploring the work of staff advisers in political position-making. Social Studies of Science, 50(2):292–316, April 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312720911715.

Levin:1995:PPR

[LSW95] Sharon G. Levin, Paula E. Stephan, and Mary Beth Walker. Planck’s principle revisited: a note. Social Studies of Science, 25(2):275–283, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631295025002003.

Landstrom:2024:SED

[LSW24] Catharina Landström, Eric Sarmiento, and Sarah J. Whatmore. Stakeholder engagement does not guarantee impact: a co-productionist perspective on model-based drought research. Social Studies of Science, 54(2):210–230, April 1, 2024. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231199220.

Lewis:1992:APS

[LT92] W. David Lewis and William F. Trimble. The airmail pickup system of all American aviation: a failed innovation? Social Studies of Science, 22(2):301–315, May 1, 1992. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631292022002007.

Lujan:2012:PT

[LT12] José Luis Luján and Oliver Todt. Precaution: a taxonomy. Social Studies of Science, 42(1):143–157, February 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312711431836.

Liboiron:2018:TPA

[LTC18] Max Liboiron, Manuel Tironi, and Nerea Calvillo. Toxic politics: Acting in a permanently polluted world. Social Studies of Science, 48(3):331–349, June 1, 2018. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312718783087.

Lubrano:1981:NIP

[Lub81] Linda L. Lubrano. National and international politics in US-USSR scientific cooperation. Social Studies of Science, 11(4): 451–480, November 1, 1981. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631278101100403.

Lucier:2001:OL

[Luc01] Paul Lucier. An object lesson. Social Studies of Science, 31 (3):444–446, June 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631201031003012.

Lukes:1975:RBB

[Luk75] Steven Lukes. Review: Barry Barnes, Scientific Knowledge and Sociological Theory (London and Boston: Routledge and Kegan Paul, 1974), x + 192 pp. £3.95 (hardcover), £1.95 (paperback). Social Studies of Science, 5(4):501–505, November 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277500500409.

Lowe:1975:TSS

[LW75] Philip D. Lowe and Michael Worboys. The teaching of social studies of science and technology in British polytechnics. Social Studies of Science, 5(2):177–192, May 1, 1975.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277500500204.

Law:1982:PFT

[LW82] John Law and R. J. Williams. Putting facts together: a study of scientific persuasion. Social Studies of Science, 12(4): 535–558, November 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631282012004003.

Lawless:2010:HIH

[LW10] Christopher J. Lawless and Robin Williams. Helping with inquiries or helping with profits? The trials and tribulations of a technology of forensic reasoning. Social Studies of Science, 40 (5):731–755, October 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312710378787.

Lynch:1982:TWC

[Lyn82] Michael E. Lynch. Technical work and critical inquiry: Investigations in a scientific laboratory. Social Studies of Science, 12 (4):499–533, November 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631282012004002.

Lynch:1985:DMF

[Lyn85] Michael Lynch. Discipline and the material form of images: an analysis of scientific visibility. Social Studies of Science, 15 (1):37–66, February 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631285015001002.

Lynch:1988:STA

[Lyn88] Michael E. Lynch. Sacrifice and the transformation of the animal body into a scientific object: Laboratory culture and ritual practice in the neurosciences. Social Studies of Science, 18(2):265–289, May 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631288018002004.

Lynch:1994:CHW

[Lyn94a] Michael Lynch. Collins, Hirschauer and Winch: Ethnography, exoticism, surgery, antisepsis and dehorsification. Social Studies of Science, 24(2):354–369, May 1, 1994. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631279402400208.

Lynch:1994:ISS

[Lyn94b] William T. Lynch. Ideology and the sociology of scientific knowledge. Social Studies of Science, 24(2):197–227, May 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631279402400202.

Lynch:1995:CSC

[Lyn95a] Michael Lynch. Collaboration and scandal: a comment on Labinger. Social Studies of Science, 25(2):324–329, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631295025002010. See [Lab95b].

Lynch:1995:IA

[Lyn95b] Michael Lynch. The idylls of the academy. Social Studies of Science, 25(3):582–600, August 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631295025003008.

Lynch:1998:DPU

[Lyn98a] Michael Lynch. The discursive production of uncertainty: The OJ Simpson ‘Dream Team’ and the sociology of knowledge machine. Social Studies of Science, 28(5–6):829–868, October 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631298028005007.

Lynch:1998:EYE

[Lyn98b] Michael Lynch. Everything you ever wanted to know about technicians, and more. Social Studies of Science, 28(1):186– 190, February 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631298028001009.

Lynch:2001:NP

[Lyn01a] Michael Lynch. Noise and politics. Social Studies of Science, 31(3):446–454, June 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631201031003013.

Lynch:2001:PBB

[Lyn01b] Michael Lynch. Pandora’s ballot box: Comments on the 2000 US Presidential Election. Social Studies of Science, 31 (3):417–419, June 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631201031003004.

Lynch:2002:E

[Lyn02] Michael Lynch. Editorial. Social Studies of Science, 32(5–6): 641–642, December 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631270203200513.

Lynch:2003:Ea

[Lyn03a] Michael Lynch. Editorial. Social Studies of Science, 33 (1):5–6, February 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312703033001011.

Lynch:2003:Eb

[Lyn03b] Michael Lynch. Editorial. Social Studies of Science, 33(2): 155–159, April 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/03063127030332001.

Lynch:2003:Ec

[Lyn03c] Michael Lynch. Editorial. Social Studies of Science, 33(3):325, June 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/03063127030333001.

Lynch:2003:Ed

[Lyn03d] Michael Lynch. Editorial. Social Studies of Science, 33(6): 829–830, December 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312703336001.

Lynch:2004:Ea

[Lyn04a] Michael Lynch. Editorial. Social Studies of Science, 34 (1):5–6, February 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312704043153.

Lynch:2004:Eb

[Lyn04b] Michael Lynch. Editorial. Social Studies of Science, 34(4): 467–468, August 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312704048182.

Lynch:2004:Ec

[Lyn04c] Michael Lynch. Editorial. Social Studies of Science, 34(6): 827–828, December 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312704049799.

Lynch:2004:GWO

[Lyn04d] Michael Lynch. Ghost writing and other matters. Social Studies of Science, 34(2):147–148, April 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704044964.

Lynch:2004:SPC

[Lyn04e] Michael Lynch. Scientism and philosophism: Comment on ‘Kinder, Gentler Science Wars’ by Gabriel Stolzenberg. Social Studies of Science, 34(1):93–98, February 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704041340. See [Sto04a].

Lynch:2005:E

[Lyn05] Michael Lynch. Editorial. Social Studies of Science, 35(2): 171–172, April 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312705054445.

Lynch:2006:EM

[Lyn06a] Michael Lynch. Expert metascientists. Social Studies of Science, 36(6):867–868, December 1, 2006. CODEN SSSCDH.

ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/0306312706067903.

Lynch:2006:RF

[Lyn06b] Michael Lynch. From ruse to farce. Social Studies of Science, 36 (6):819–826, December 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706067897.

Lynch:2009:Ea

[Lyn09a] Michael Lynch. Editorial. Social Studies of Science, 39 (1):5–6, February 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312708100841.

Lynch:2009:Eb

[Lyn09b] Michael Lynch. Editorial. Social Studies of Science, 39(6): 819–820, December 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312709350402.

Lynch:2010:Ea

[Lyn10a] Michael Lynch. Editorial. Social Studies of Science, 40 (1):5–6, February 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312710361277.

Lynch:2010:Eb

[Lyn10b] Michael Lynch. Editorial. Social Studies of Science, 40(3):347, June 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312710370923.

Lynch:2011:AHS

[Lyn11a] Michael Lynch. Ad hoc special section on ethnomethodological studies of science, mathematics, and technical activity: Introduction. Social Studies of Science, 41(6):835–837, December 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312711427369.

Lynch:2011:E

[Lyn11b] Michael Lynch. Editorial. Social Studies of Science, 41(6): 767–768, December 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312711427368.

Lynch:2011:HGO

[Lyn11c] Michael Lynch. Harold Garfinkel (29 October 1917–21 April 2011): a remembrance and reminder. Social Studies of Science, 41(6):927–942, December 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312711423434.

Lynch:2011:SEA

[Lyn11d] Michael Lynch. Still emerging after all these years. Social Studies of Science, 41(1):3–4, February 1, 2011. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312710396569.

Lynch:2012:C

[Lyn12a] Michael Lynch. Changes. Social Studies of Science, 42(2): 167–169, April 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712437521.

Lynch:2012:SER

[Lyn12b] Michael Lynch. Self-exemplifying revolutions? Notes on Kuhn and Latour. Social Studies of Science, 42(3):449–455, June 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312712439120. See [Col12].

Lynch:2013:OIP

[Lyn13] Michael Lynch. Ontography: Investigating the production of things, deflating ontology. Social Studies of Science, 43 (3):444–462, June 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312713475925.

Lynch:2014:GHE

[Lyn14] Michael Lynch. Genius, Hawking, and expertise. Social Studies of Science, 44(5):793–798, October 1, 2014. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312714550228.

Lynch:2017:SSP

[Lyn17] Michael Lynch. STS, symmetry and post-truth. Social Studies of Science, 47(4):593–599, August 1, 2017. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312717720308.

Lyons:2018:CWC

[Lyo18] Kristina Lyons. Chemical warfare in Colombia, evidentiary ecologies and senti-actuando practices of justice. Social Studies of Science, 48(3):414–437, June 1, 2018. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312718765375.

Ma:2019:BRS

[Ma19] Dali Ma. Boundary repair: Science and enterprise at the Chinese Academy of Sciences. Social Studies of Science, 49(3):381–402, June 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312719846560.

MacKenzie:1976:EB

[Mac76] Donald MacKenzie. Eugenics in Britain. Social Studies of Science, 6(3–4):499–532, September 1, 1976. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277600600310.

MacKenzie:1978:PWS

[Mac78a] Donald MacKenzie. A PAREX workshop on the sociology of mathematical knowledge and the social history of mathematics. Social Studies of Science, 8(1):141–142, February 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277800800106.

MacKenzie:1978:STS

[Mac78b] Donald MacKenzie. Statistical theory and social interests: a case-study. Social Studies of Science, 8(1):35–83, February 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277800800102.

MacLeod:1979:CPS

[Mac79] Kay MacLeod. The contradictions of professionalism: Scientists, trade unionism and the First World War. Social Studies of Science, 9(1):1–32, February 1, 1979. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277900900101.

MacKenzie:1981:IPH

[Mac81] Donald MacKenzie. Interests, positivism and history. Social Studies of Science, 11(4):498–504, November 1, 1981. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278101100405.

Mackay:1984:DJS

[Mac84] Alan Mackay. Derek John de Solla Price: an appreciation. Social Studies of Science, 14(2):315–320, May 1, 1984. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014002013.

MacKenzie:1986:STS

[Mac86] Donald MacKenzie. Science and technology studies and the question of the military. Social Studies of Science, 16(2):361– 371, May 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/0306312786016002010.

MacLeod:1990:MC

[Mac90] Roy MacLeod. Moas in the cathedral. Social Studies of Science, 20(1):165–169, February 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631290020001007.

MacKenzie:1993:NAC

[Mac93] Donald MacKenzie. Negotiating arithmetic, constructing proof: The sociology of mathematics and information technology. Social Studies of Science, 23(1):37–65, February 1, 1993. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://ieeemilestones.ethw.org/w/images/
1/10/Negotiating_Arithmetic_MacKenzie.pdf;   https:/
/journals.sagepub.com/doi/pdf/10.1177/030631293023001002.

Summary of the process of IEEE 754, several years after its initial adoption.

MacKenzie:1999:SWP

[Mac99a] Donald MacKenzie. The science wars and the past’s quiet voices. Social Studies of Science, 29(2):199–213, April 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029002002.

MacKenzie:1999:SKS

[Mac99b] Donald MacKenzie. Slaying the kraken: The sociohistory of a mathematical proof. Social Studies of Science, 29(1):7–60, February 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631299029001002. See correction

[Mac00].

MacKenzie:1999:ZSA

[Mac99c] Donald MacKenzie. The zero-sum assumption: Reply to Sullivan. Social Studies of Science, 29(2):223–234, April 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029002004. See [Sul99].

MacKenzie:2000:SKC

[Mac00] Donald MacKenzie. Slaying the kraken: a correction. Social Studies of Science, 30(2):317–318, April 1, 2000. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030002004. See [Mac99b].

MacFarlane:2003:UYM

[Mac03a] Allison MacFarlane. Underlying Yucca Mountain: The interplay of geology and policy in nuclear waste disposal. Social Studies of Science, 33(5):783–807, October 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312703335006.

Mackenzie:2003:TCS

[Mac03b] Adrian Mackenzie. These things called systems: Collective imaginings and infrastructural software. Social Studies of Science, 33(3):365–387, June 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/03063127030333003.

MacKenzie:2003:EWB

[Mac03c] Donald MacKenzie. An equation and its worlds: Bricolage, exemplars, disunity and performativity in financial economics. Social Studies of Science, 33(6):831–868, December 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312703336002.

MacLeod:2003:MDO

[Mac03d] Roy MacLeod. In memoriam: David Owen Edge. Social Studies of Science, 33(2):181–183, April 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127030332005.

MacKenzie:2017:MPE

[Mac17] Donald MacKenzie. A material political economy: Automated trading desk and price prediction in high-frequency trading. Social Studies of Science, 47(2):172–194, April 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716676900.

Maerker:2007:RWD

[Mae07] Anna Maerker. Review: Why do they look like That? threedimensional models in science. de Chadarevian, Soraya & Nick Hopwood (eds) (2004) Models: The Third Dimension of Science (Stanford: Stanford University Press), 488 pp, $70.00 / £42.95 / EUR 58.90 (hbk), $25.95 / £18.50 / EUR 22.00 (pbk), ISBN 0-8047-3971-4 (hbk), 0-8047-3972-2 (pbk). Social Studies of Science, 37(6):961–965, December 1, 2007. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312707087013.

Magyar:1975:TRP

[Mag75] George Magyar. Typology of research in physics. Social Studies of Science, 5(1):79–85, February 1, 1975. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277500500105.

Magyar:1977:PEE

[Mag77] George Magyar. ‘pseudo-effects’ in experimental physics: Some notes for case-studies. Social Studies of Science, 7(2):241–256, May 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277700700206.

Magyar:1978:RK

[Mag78] George Magyar. Reply to Kantor. Social Studies of Science, 8 (3):358–359, August 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277800800309. See

[Kan78b].

Magnus:2007:DCT

[Mag07] P. D. Magnus. Distributed cognition and the task of science. Social Studies of Science, 37(2):297–310, April 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706072177.

Mager:2017:SEI

[Mag17] Astrid Mager. Search engine imaginary: Visions and values in the co-production of search technology and Europe. Social Studies of Science, 47(2):240–262, April 1, 2017. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716671433.

Mahoney:1979:RPP

[Mah79] Michael J. Mahoney. Review paper: Psychology of the scientist: an evaluative review. Social Studies of Science, 9(3):

349–375, August 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277900900304.

Mahoney:1980:RAC

[Mah80] Michael J. Mahoney. Rationality and authority: On the confusion of justification and permission. Social Studies of Science, 10 (4):515–518, November 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631278001000408.

Mahony:2014:PSS

[Mah14] Martin Mahony. The predictive state: Science, territory and the future of the Indian climate. Social Studies of Science, 44 (1):109–133, February 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312713501407.

Mallard:1998:CSS

[Mal98] Alexandre Mallard. Compare, standardize and settle agreement: On some usual metrological problems. Social Studies of Science, 28(4):571–601, August 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631298028004003.

Manier:1987:EFI

[Man87] Edward Manier. ‘external factors’ and ‘ideology’ in the earliest drafts of Darwin’s theory. Social Studies of Science, 17(4): 581–609, November 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631287017004002.

Mann:1993:IDS

[Man93] Gregory Mann. Institutional dynamics of scientific change: Ben-David’s legacy. Social Studies of Science, 23(4):757–763, November 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631293023004006.

Mansfield:2021:DSC

[Man21] Becky Mansfield. Deregulatory science: Chemical risk analysis in Trump’s EPA. Social Studies of Science, 51(1):28–50, February 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312720970284.

Marshall:1977:MMM

[Mar77] John C. Marshall. Minds, machines and metaphors. Social Studies of Science, 7(4):475–488, November 1, 1977. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700412.

Martin:1978:SUG

[Mar78] Brian Martin. The selective usefulness of game theory. Social Studies of Science, 8(1):85–110, February 1, 1978. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277800800103.

Martin:1988:AFC

[Mar88] Brian Martin. Analyzing the fluoridation controversy: Resources and structures. Social Studies of Science, 18(2):331–363, May 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631288018002006.

Marks:1995:OVR

[Mar95] Harry M. Marks. Other voices: a response to Labinger. Social Studies of Science, 25(2):329–334, May 1, 1995. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631295025002011. See [Lab95b].

Martin:1996:SCA

[Mar96a] Brian Martin. Social construction of an ‘attack on science’. Social Studies of Science, 26(1):161–173, February 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631296026001008.

Martin:1996:SNS

[Mar96b] Brian Martin. Sticking a needle into science: The case of polio vaccines and the origin of AIDS. Social Studies of Science, 26(2):245–276, May 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631296026002003.

Martin:1997:STN

[Mar97] Brian Martin. Science, technology and nonviolent action: The case for a utopian dimension in the social analysis of science and technology. Social Studies of Science, 27(3):439–463, June 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631297027003003.

Martin:2004:CBC

[Mar04] Aryn Martin. Can’t any body count?: Counting as an epistemic theme in the history of human chromosomes. Social Studies of Science, 34(6):923–948, December 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704046843.

Marie:2005:MMF

[Mar05] Jenny Marie. Making a mouse-focused history? The story of C. C. Little and the Jackson Laboratory. Social Studies of Science, 35(6):955–958, December 1, 2005. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312705057230.

Marres:2007:IDM

[Mar07] Noortje Marres. The issues deserve more credit: Pragmatist contributions to the study of public involvement in controversy. Social Studies of Science, 37(5):759–780, October 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312706077367.

Marie:2008:SLM

[Mar08a] Jenny Marie. For science, love and money: The social worlds of poultry and rabbit breeding in Britain, 1900–1940. Social Studies of Science, 38(6):919–936, December 1, 2008. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312708098608.

Martello:2008:AIP

[Mar08b] Marybeth Long Martello. Arctic indigenous peoples as representations and representatives of climate change. Social Studies of Science, 38(3):351–376, June 1, 2008. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312707083665.

Martucci:2010:NEM

[Mar10] Jessica Martucci. Negotiating exclusion: MSM, identity, and blood policy in the age of AIDS. Social Studies of Science, 40(2):215–241, April 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312709346579.

Marres:2013:WPO

[Mar13] Noortje Marres. Why political ontology must be experimentalized: On eco-show homes as devices of participation. Social Studies of Science, 43(3):417–443, June 1, 2013. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712475255.

Margocsy:2017:LHB

[Mar17] Dániel Margócsy. A long history of breakdowns: a historiographical review. Social Studies of Science, 47(3):307–325, June 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312717706559.

Marshall:2020:SSI

[Mar20] Owen Marshall. Shibboleths in the studio: Informal demarcation practices among audio engineers. Social Studies of Science, 50(6):881–900, December 1, 2020. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720922314.

Marres:2023:SMB

[Mar23] Noortje Marres. Seven moments with Bruno Latour. Social Studies of Science, 53(2):188–193, April 1, 2023. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231159852.

Marathe:2024:TVT

[Mar24] Megh Marathe. Therapeutic value in the time of digital brainwaves. Social Studies of Science, 54(6):931–954, December 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127241241032.

Masco:2010:BWP

[Mas10] Joseph Masco. Bad weather: On planetary crisis. Social Studies of Science, 40(1):7–40, February 1, 2010. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312709341598.

Matsumoto:2005:UCR

[Mat05] Miwao Matsumoto. The uncertain but crucial relationship between a ‘new energy’ technology and global environmental problems: The complex case of the ‘Sunshine’ Project. Social Studies of Science, 35(4):623–651, August 1, 2005. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312705052107.

Matusall:2013:HTF

[Mat13] Svenja Matusall. Hormone troubles: Feminist analyses of contemporary neurosciences. Social Studies of Science, 43(5): 780–791, October 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312713488843.

Mathews:2014:SAF

[Mat14] Andrew S. Mathews. Scandals, audits, and fictions: Linking climate change to Mexican forests. Social Studies of Science, 44 (1):82–108, February 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312713490330.

Maxigas:2017:HAT

[Max17] Maxigas. Hackers against technology: Critique and recuperation in technological cycles. Social Studies of Science, 47(6): 841–860, December 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312717736387.

Mayernik:2019:MAA

[May19] Matthew S. Mayernik. Metadata accounts: Achieving data and evidence in scientific research. Social Studies of Science, 49 (5):732–757, October 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312719863494.

Mazur:1977:PCS

[Maz77] Allan Mazur. Public confidence in science. Social Studies of Science, 7(1):123–125, February 1, 1977. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700113.

Mazur:1981:LSO

[MB81] Allan Mazur and Elma Boyko. Large-scale ocean research projects: What makes them succeed or fail? Social Studies of Science, 11(4):425–449, November 1, 1981. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278101100402.

Michael:1994:ECS

[MB94] Mike Michael and Lynda Birke. Enrolling the core set: The case of the animal experimentation controversy. Social Studies of Science, 24(1):81–95, February 1, 1994. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631279402400104.

Min:2022:SFO

[MB22] Bo Hee Min and Christian Borch. Systemic failures and organizational risk management in algorithmic trading: Normal accidents and high reliability in financial markets. Social Studies of Science, 52(2):277–302, April 1, 2022. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/03063127211048515.

Marres:2025:MEA

[MB25a] Noortje Marres and Matías Valderrama Barragán. Making expert advice public in a time of emergency: Independent SAGE and the contestation of science during the Covid pandemic in the UK. Social Studies of Science, 55(4):512–541, August 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127241309071.

Miroshnichenko:2025:CDA

[MB25b] Alia Miroshnichenko and Kean Birch. Constructing digital assets through blockchain technologies? Unpacking the techno-economic configuration of non-fungible tokens. Social Studies of Science, 55(4):631–651, August 2025. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127241286447.

Malone:2000:SNJ

[MBB00] Ruth E. Malone, Elizabeth Boyd, and Lisa A. Bero. Science in the news: Journalists’ constructions of passive smoking as a social problem. Social Studies of Science, 30(5):713–735, October 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631200030005003.

Murray:2024:CCR

[MBDR24] Andy Murray, Dennis Browe, Katherine Weatherford Darling, and Jenny Reardon. Cells and the city: The rise and fall of urban biopolitics in San Francisco, 1970–2020. Social Studies of Science, 54(6):805–835, December 2024. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127241261376.

Meadows:1971:BSG

[MC71] A. J. Meadows and J. G. O’ Connor. Bibliographical statistics as a guide to growth points in science. Social Studies of Science, 1(1):95–99, January 1, 1971. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277100100107.

Mazur:1978:CLN

[MC78] Allan Mazur and Beverlie Conant. Controversy over a local nuclear waste repository. Social Studies of Science, 8(2):235–243, May 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277800800204.

Mitroff:1979:PRN

[MC79] Ian I. Mitroff and Daryl E. Chubin. Peer review at the NSF: a dialectical policy analysis. Social Studies of Science, 9(2):199–232, May 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277900900203.

Mitchell:1997:ITP

[MC97] Lisa M. Mitchell and Alberto Cambrosio. The invisible topography of power: Electromagnetic fields, bodies and the environment. Social Studies of Science, 27(2):221–271, April 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631297027002002.

Mcharek:2023:CAR

[M’c23] Amade M’charek. Curious about race: Generous methods and modes of knowing in practice. Social Studies of Science, 53(6): 826–849, December 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127231201178.

Mackay:2000:RUU

[MCBDT00] Hugh Mackay, Chris Carne, Paul Beynon-Davies, and Doug Tudhope. Reconfiguring the user: Using rapid application development. Social Studies of Science, 30(5):737–757, October 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631200030005004.

McCulloch:1988:TVI

[McC88] Gary McCulloch. A technocratic vision: The ideology of school science reform in Britain in the 1950s. Social Studies of Science, 18(4):703–724, November 1, 1988. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631288018004005.

McCutcheon:1992:STS

[McC92] Robert T. McCutcheon. Science, technology and the state in the provision of low-income accommodation: The case of industrialized house-building, 1955–77. Social Studies of Science, 22(2):353–371, May 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631292022002010.

McCray:2000:LTM

[McC00] W. Patrick McCray. Large telescopes and the moral economy of recent astronomy. Social Studies of Science, 30(5):685–711, October 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631200030005002.

McCormick:2007:DSM

[McC07] Sabrina McCormick. Democratizing science movements: a new framework for mobilization and contestation. Social Studies of Science, 37(4):609–623, August 1, 2007. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312707076598.

McDonagh:1976:ACP

[McD76] Eileen L. McDonagh. Attitude changes and paradigm shifts: Social psychological foundations of the Kuhnian thesis. Social Studies of Science, 6(1):51–76, February 1, 1976. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277600600103.

McDonell:1997:SEK

[McD97] Gavan McDonell. Scientific and everyday knowledge: Trust and the politics of environmental initiatives. Social Studies of Science, 27(6):819–863, December 1, 1997. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631297027006001.

McHenry:2009:GMC

[McH09] Leemon McHenry. Ghosts in the machine: Comment on Sismondo. Social Studies of Science, 39(6):943–947, December 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/

10.1177/0306312709345358. See [Sis09a] and reply [Sis09b].

McKinney:1999:PHB

[McK99] William J. McKinney. Partial houses built on common ground: Reply to Pinch. Social Studies of Science, 29(2):240–246, April 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631299029002006. See [Pin99b].

McLevey:2015:UPR

[McL15] John McLevey. Understanding policy research in liminal spaces: Think tank responses to diverging principles of legitimacy. Social Studies of Science, 45(2):270–293, April 1, 2015. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312715575054.

Mclellan:2021:ITC

[Mcl21] Timothy Mclellan. Impact, theory of change, and the horizons of scientific practice. Social Studies of Science, 51(1): 100–120, February 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312720950830.

Medina-Domenech:2009:STN

[MD09] Rosa Medina-Doménech. Scientific technologies of national identity as colonial legacies: Extracting the Spanish nation from Equatorial Guinea. Social Studies of Science, 39(1): 81–112, February 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312708097658.

Mulinari:2020:WCP

[MD20] Shai Mulinari and Courtney Davis. The will of Congress? Permissive regulation and the strategic use of labeling for the anti-influenza drug Relenza. Social Studies of Science, 50(1): 145–169, February 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312719890015.

Mehrtens:1980:WSH

[Meh80] Herbert Mehrtens. Workshop on the Social History of Mathematics (West Berlin, 5–8 July 1979). Social Studies of Science, 10(1):121–125, February 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631278001000107.

Mellor:2003:BFF

[Mel03] Felicity Mellor. Between fact and fiction: Demarcating science from non-science in popular physics books. Social Studies of Science, 33(4):509–538, August 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312703334002.

Mellor:2007:CWA

[Mel07] Felicity Mellor. Colliding worlds: Asteroid research and the legitimization of war in space. Social Studies of Science, 37 (4):499–531, August 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706075336.

Mellstrom:2009:IGR

[Mel09] Ulf Mellström. The intersection of gender, race and cultural boundaries, or why is computer science in Malaysia dominated by women? Social Studies of Science, 39(6):885–907, December 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312709334636.

Mercer:1996:RSC

[Mer96] David Mercer. The revealing science of cod. Social Studies of Science, 26(1):182–186, February 1, 1996. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631296026001010.

Mermin:1998:APR

[Mer98a] N. David Mermin. Abandoning preconceptions: Reply to Bloor and Barnes. Social Studies of Science, 28(4):641–647, August 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631298028004007.

Mermin:1998:SSP

[Mer98b] N. David Mermin. The science of science: a physicist reads Barnes, Bloor and Henry. Social Studies of Science, 28(4):603– 623, August 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631298028004004. See book Scien-

tific Knowledge, by Barry Barnes, David Bloor and John Henry, and replies [Bar98, Blo98a].

Mercer:2002:SMD

[Mer02] David Mercer. Scientific method discourses in the construction of ‘EMF Science’: Interests, resources and rhetoric in submissions to a public inquiry. Social Studies of Science, 32 (2):205–233, April 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312702032002002.

Merchant:2022:EMD

[Mer22] Emily Klancher Merchant. Environmental Malthusianism and demography. Social Studies of Science, 52(4):536–560, August 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127221104929.

Messeri:1988:ADR

[Mes88] Peter Messeri. Age differences in the reception of new scientific theories: The case of plate tectonics theory. Social Studies of Science, 18(1):91–112, February 1, 1988. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631288018001004.

Messeri:2010:PPC

[Mes10] Lisa R. Messeri. The problem with Pluto: Conflicting cosmologies and the classification of planets. Social Studies of Science, 40(2):187–214, April 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312709347809.

Messeri:2017:ETI

[Mes17] Lisa Messeri. Extra-terra incognita: Martian maps in the digital age. Social Studies of Science, 47(1):75–94, February 1, 2017.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716656820.

Metcalfe:1988:TD

[Met88] Stan Metcalfe. Technology and development. Social Studies of Science, 18(1):183–187, February 1, 1988. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631288018001008.

Metlay:2006:RRS

[Met06] Grischa Metlay. Reconsidering renormalization: Stability and change in 20th-Century views on university patents. Social Studies of Science, 36(4):565–597, August 1, 2006. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706058581.

Metcalf:2025:CMM

[Met25] Kathryne Metcalf. Categorical misalignment: Making autism(s) in big data biobanking. Social Studies of Science, 55(2):209–237, April 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
abs/10.1177/03063127241288223.

Meynell:1974:DBR

[Mey74] Hugo Meynell. David Bloor’s Reductio Ad Absurdum of ‘objective’ truth. Social Studies of Science, 4(2):190–193, April 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277400400205.

Meynell:1977:LSK

[Mey77] Hugo Meynell. On the limits of the sociology of knowledge. Social Studies of Science, 7(4):489–500, November 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700413.

Meyer:2022:TRM

[Mey22] Morgan Meyer. Taking responsibility, making irresponsibility: Controversies in human gene editing. Social Studies of Science, 52(1):127–143, February 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/03063127211025631.

Mosoni-Fried:1995:IRH

[MF95] Judith Mosoni-Fried. Industrial research in Hungary: a victim of structural change. Social Studies of Science, 25(4): 777–803, November 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631295025004009.

Mulkay:1982:JAS

[MG82a] Michael Mulkay and G. Nigel Gilbert. Joking apart: Some recommendations concerning the analysis of scientific culture. Social Studies of Science, 12(4):585–613, November 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631282012004005.

Mulkay:1982:WUQ

[MG82b] Michael Mulkay and G. Nigel Gilbert. What is the ultimate question? Some remarks in defence of the analysis of scientific discourse. Social Studies of Science, 12(2):309–319, May 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631282012002006.

Mackay:1992:ESS

[MG92] Hughie Mackay and Gareth Gillespie. Extending the social shaping of technology approach: Ideology and appropriation. Social Studies of Science, 22(4):685–716, November 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631292022004006.

Martin:1985:EER

[MI85] Ben R. Martin and John Irvine. Evaluating the evaluators: a reply to our critics. Social Studies of Science, 15(3):558–575, August 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/
doi/pdf/10.1177/030631285015003010. See [Bud85, KP85,

MvR85].

Mialet:1999:DAB

[Mia99] Hélène Mialet. Do angels have bodies? Two stories about subjectivity in science: The cases of William X and Mister H. Social Studies of Science, 29(4):551–581, August 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029004004.

Mialet:2003:RWP

[Mia03] Hélène Mialet. The ‘righteous wrath’ of Pierre Bourdieu. Social Studies of Science, 33(4):613–621, August 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312703334006.

Mialet:2012:WWS

[Mia12] Hélène Mialet. Where would STS be without Latour? What would be missing? Social Studies of Science, 42(3):456–461, June 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312712440311.

Michael:2020:LFA

[Mic20] Mike Michael. London’s fatbergs and affective infrastructuring. Social Studies of Science, 50(3):377–397, June 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720917754.

Miettinen:1998:OCN

[Mie98] Reijo Miettinen. Object construction and networks in research work: The case of research on cellulose-degrading enzymes. Social Studies of Science, 28(3):423–463, June 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631298028003003.

Milch:1976:IPU

[Mil76] Jerome Milch. Inverted pyramids: The use and misuse of aviation forecasting. Social Studies of Science, 6(1):5–31, February 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277600600101.

Millstone:1978:FSK

[Mil78] Erik Millstone. A framework for the sociology of knowledge. Social Studies of Science, 8(1):111–125, February 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277800800104.

Mills:1981:AFS

[Mil81] Eric L. Mills. Another fragmentation of sedimentology. Social Studies of Science, 11(3):395–398, August 1, 1981. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278101100305.

Miller:1984:SHB

[Mil84] David Philip Miller. The social history of British science: After the harvest? Social Studies of Science, 14(1):115–135, February 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631284014001009.

Miller:1992:VR

[Mil92] David Philip Miller. Values redivivus? Social Studies of Science, 22(2):419–427, May 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631292022002015.

Miller:2001:MDC

[Mil01] Clark A. Miller. Making democracy count. Social Studies of Science, 31(3):454–458, June 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631201031003014.

Miller:2003:DPM

[Mil03] Fiona Alice Miller. Dermatoglyphics and the persistence of ‘Mongolism’: Networks of technology, disease and discipline. Social Studies of Science, 33(1):75–94, February 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312703033001311.

Miller:2004:ICE

[Mil04a] Clark A. Miller. Interrogating the civic epistemology of American democracy: Stability and instability in the 2000 US Presidential Election. Social Studies of Science, 34(4):501–530, August 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312704045661.

Miller:2004:AFH

[Mil04b] David Philip Miller. Absolutely fabulous history of science? Social Studies of Science, 34(3):453–459, June 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704042094.

Miller:2009:WDI

[Mil09] Boaz Miller. What does it mean that PRIMES is in P?: Popularization and distortion revisited. Social Studies of Science, 39(2):257–288, April 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312708101131.

Milne:2012:PPB

[Mil12] Richard Milne. Pharmaceutical prospects: Biopharming and the geography of technological expectations. Social Studies of Science, 42(2):290–306, April 1, 2012. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312711436266.

Mirsky:1972:SSU

[Mir72] E. M. Mirsky. Science studies in the USSR (history, problems, prospects). Social Studies of Science, 2(3):281–294, July 1, 1972. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277200200304.

Mirskaya:1995:RAS

[Mir95] Elena Z. Mirskaya. Russian academic science today: Its societal standing and the situation within the scientific community. Social Studies of Science, 25(4):705–725, November 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631295025004006.

Mirowski:1999:CAE

[Mir99] Philip Mirowski. Cyborg agonistes: Economics meets operations research in mid-century. Social Studies of Science, 29 (5):685–718, October 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631299029005002.

Mirowski:2007:REJ

[Mir07] Philip Mirowski. Review essay: Johnny’s in the basement, mixin’ up the medicine: Review of Angell, Avorn, and Daemmrich on the modern pharmaceutical predicament: Marcia Angell (2004) The Truth About the Drug Companies: How They Deceive Us and What to Do About It (New York: Random House), US$24.95. ISBN 0-375-50846-5. Jerry Avorn (2004) Powerful Medicines: The Benefits, Risks and Costs of Prescription Drugs (New York: Knopf), US$27.50. ISBN 0-375 41483-5. Arthur Daemmrich (2004) Pharmacopolitics: Drug Regulation in the US and Germany (Chapel Hill, NC: University of North Carolina Press), US$34.95. ISBN 0-8078-2844-0. Social Studies of Science, 37(2):311–327, April 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706068494.

Mirowski:2018:FOS

[Mir18] Philip Mirowski. The future(s) of open science. Social Studies of Science, 48(2):171–203, April 1, 2018. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312718772086.

Missner:1985:WEB

[Mis85] Marshall Missner. Why Einstein became famous in America. Social Studies of Science, 15(2):267–291, May 1, 1985. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL http://www.jstor.org/stable/285389;   https://
journals.sagepub.com/doi/pdf/10.1177/030631285015002003.

Mitroff:1977:SSS

[MJM77] Ian I. Mitroff, Theodore Jacob, and Eileen Trauth Moore. On the shoulders of the spouses of scientists. Social Studies of Science, 7(3):303–327, August 1, 1977. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700303.

Maguire:1984:SIA

[MK84] Carmel Maguire and Robin Kench. Sources of ideas for applied University research, and their effect on the application of findings in Australian industry. Social Studies of Science, 14 (3):371–397, August 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631284014003003.

McCarthy:2010:RN

[MK10] Elise McCarthy and Christopher Kelty. Responsibility and nanotechnology. Social Studies of Science, 40(3):405–432, June 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312709351762.

Merz:1997:DTS

[MKC97] Martina Merz and Karin Knorr-Cetina. Deconstruction in a ‘thinking’ science: Theoretical physicists at work. Social Studies of Science, 27(1):73–111, February 1, 1997. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631297027001004.

Mazur:1993:SGA

[ML93] Allan Mazur and Jinling Lee. Sounding the global alarm: Environmental issues in the US national news. Social Studies of Science, 23(4):681–720, November 1, 1993. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631293023004003.

Mol:1994:RNF

[ML94] Annemarie Mol and John Law. Regions, networks and fluids: Anaemia and social topology. Social Studies of Science, 24(4): 641–671, November 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631279402400402.

Murphy:2006:RSE

[MLC06] Joseph Murphy, Les Levidow, and Susan Carr. Regulatory standards for environmental risks: Understanding the US– European Union conflict over genetically modified crops. Social Studies of Science, 36(1):133–160, February 1, 2006. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312705054591.

Moravcsik:1975:SRF

[MM75] Michael J. Moravcsik and Poovanalingam Murugesan. Some results on the function and quality of citations. Social Studies of Science, 5(1):86–92, February 1, 1975. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277500500106.

Mulligan:1981:RRS

[MM81] Lotte Mulligan and Glenn Mulligan. Reconstructing restoration science: Styles of leadership and social composition of the early Royal Society. Social Studies of Science, 11(3):327–364, August 1, 1981. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631278101100303.

MacRoberts:1982:REL

[MM82] Michael H. MacRoberts and Barbara R. MacRoberts. A re-evaluation of Lotka’s Law of Scientific Productivity. Social Studies of Science, 12(3):443–450, August 1, 1982. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631282012003005.

MacRoberts:1984:NRA

[MM84a] Michael H. MacRoberts and Barbara R. MacRoberts. The negational reference: or the art of dissembling. Social Studies of Science, 14(1):91–94, February 1, 1984. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014001006.

McCrea:1984:ERC

[MM84b] Frances B. McCrea and Gerald E. Markle. The estrogen replacement controversy in the USA and UK: Different answers to the same question? Social Studies of Science, 14(1):1–26, February 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631284014001002.

MacRoberts:1986:QMC

[MM86] Michael H. MacRoberts and Barbara R. MacRoberts. Quantitative measures of communication in science: a study of the formal level. Social Studies of Science, 16(1):151–172, February 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631286016001008. See [Sti87] and reply [MM87].

MacRoberts:1987:MFU

[MM87] Michael H. MacRoberts and Barbara R. MacRoberts. Measurement in the face of universal uncertainty: a reply to Stigler. Social Studies of Science, 17(2):334–336, May 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631287017002008. See [MM86].

Mol:1996:NFP

[MM96] Annemarie Mol and Jessica Mesman. Neonatal food and the politics of theory: Some questions of method. Social Studies of Science, 26(2):419–444, May 1, 1996. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631296026002009.

Mort:1998:HTR

[MM98] Maggie Mort and Mike Michael. Human and technological ‘redundancy’: Phantom intermediaries in a nuclear submarine industry. Social Studies of Science, 28(3):355–400, June 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631298028003001.

Moreira:2009:ROA

[MMB09] Tiago Moreira, Carl May, and John Bond. Regulatory objectivity in action: Mild cognitive impairment and the collective production of uncertainty. Social Studies of Science, 39 (5):665–690, October 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312709103481.

Martin:2015:PCT

[MMV15] Aryn Martin, Natasha Myers, and Ana Viseu. The politics of care in technoscience. Social Studies of Science, 45(5): 625–641, October 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312715602073.

Milyaeva:2023:LAA

[MN23] Sveta Milyaeva and Daniel Neyland. Let’s agree to agree: The situational academic quality of the UK REF as consensual public knowledge. Social Studies of Science, 53(3):427–448, June 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127231152915.

Montgomery:2009:SGE

[MO09] Kathleen Montgomery and Amalya L. Oliver. Shifts in guidelines for ethical scientific conduct: How public and private organizations create and change norms of research integrity1. Social Studies of Science, 39(1):137–155, February 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312708097659.

Mody:2000:NWF

[Mod00] Cyrus C. M. Mody. ‘A new way of flying’: Difference, rhetoric and the autogiro in interwar aviation. Social Studies of Science, 30(4):513–543, August 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631200030004002.

Mody:2001:LDN

[Mod01] Cyrus C. M. Mody. A little dirt never hurt anyone: Knowledge-making and contamination in materials science. Social Studies of Science, 31(1):7–36, February 1, 2001. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031001002.

Mody:2012:FTL

[Mod12] Cyrus C. M. Mody. Faster-than-light reading. Social Studies of Science, 42(1):159–164, February 1, 2012. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312711430259.

Mol:2013:MYP

[Mol13] Annemarie Mol. Mind your plate! The ontonorms of Dutch dieting. Social Studies of Science, 43(3):379–396, June 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712456948.

Mollers:2016:SCT

[Möl16] Norma Möllers. Shifting in and out of context: Technoscientific drama as technology of the self. Social Studies of Science, 46(3):351–373, June 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312716638951.

Montgomery:2012:MPP

[Mon12] Catherine M. Montgomery. Making prevention public: The co-production of gender and technology in HIV prevention research. Social Studies of Science, 42(6):922–944, December 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312712457707.

Moore:2010:RBP

[Moo10] Alfred Moore. Review: Beyond participation: Opening up political theory in STS: Mark Brown, Science in Democracy: Expertise, Institutions and Representation (Cambridge, MA: MIT Press, 2009), 368 pp., £20.95 / $28.00 / EUR 23.99 (pbk), £41.95 / $56.00 / EUR 48.99 (hbk). ISBN 978-0-262-51304-3 (pbk), 978-0-262-01324-6 (hbk). Social Studies of Science, 40 (5):793–799, October 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312710383070.

Morrell:1973:PMV

[Mor73] J. B. Morrell. The patronage of mid-Victorian science in the University of Edinburgh. Social Studies of Science, 3(4):

353–388, October 1, 1973. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277300300402.

Moravcsik:1982:ERD

[Mor82] Michael J. Moravcsik. The effectiveness of research in developing countries. Social Studies of Science, 12(1):144–147, February 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631282012001010. See reply [Cla82].

Moravcsik:1986:AMF

[Mor86a] Michael J. Moravcsik. Assessing the methodology for finding a methodology for assessment. Social Studies of Science, 16 (3):534–539, August 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631286016003007.

Morrell:1986:BB

[Mor86b] Jack Morrell. Brains of Britain. Social Studies of Science, 16 (4):735–745, November 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631286016004010.

Moravcsik:1988:CCC

[Mor88a] Michael J. Moravcsik. Citation context classification of a citation classic concerning citation context classification. Social Studies of Science, 18(3):515–521, August 1, 1988. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631288018003005.

Morus:1988:SSE

[Mor88b] Iwan Rhys Morus. The sociology of sparks: an episode in the history and meaning of electricity. Social Studies of Science, 18 (3):387–417, August 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631288018003001.

Mort:1994:BRW

[Mor94] Maggie Mort. Book review: What About the Workers? Social Studies of Science, 24(3):596–606, August 1, 1994. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631279402400306.

Moreira:2000:TDO

[Mor00] Tiago Moreira. Translation, difference and ontological fluidity: Cerebral angiography and neurosurgical practice (1926– 45). Social Studies of Science, 30(3):421–446, June 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030003004.

Moreira:2006:HCB

[Mor06] Tiago Moreira. Heterogeneity and coordination of blood pressure in neurosurgery. Social Studies of Science, 36(1):69–97, February 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/0306312705053051.

Moreira:2019:DFP

[Mor19] Tiago Moreira. Devicing future populations: Problematizing the relationship between quantity and quality of life. Social Studies of Science, 49(1):118–137, February 1, 2019. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312719829841.

Moseley:1977:TFS

[Mos77] Russell Moseley. Tadpoles and frogs: Some aspects of the professionalization of British physics, 1870–1939. Social Studies of Science, 7(4):423–446, November 1, 1977. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700410.

Moseley:1979:ERA

[Mos79] Russell Moseley. Essay review: From avocation to job: The changing nature of scientific practice: David Knight, The Nature of Science. The History of Science in Western Culture since 1600 (London: Andre Deutsch, 1976), 215 pp., £4.95. ISBN 0-223-96814-5. Robert Charles Post, Physics, Patents and Politics. A Biography of Charles Grafton Page (New York: Science History Publications, 1976), 227 pp., $15.95. ISBN 0-88202-046-3. Colin A. Russell, with Noel G. Coley and Gerrylynn K.

Roberts, Chemists by Profession. The Origins and Rise of the Royal Institute of Chemistry (Milton Keynes, Bucks.: The Open University Press, 1977), x + 342 pp., £9.50. ISBN 0-335-00041-X. Morris Berman, Social Change and Scientific Organization. The Royal Institution, 1799–1844 (London: Heinemann, 1978), xxv + 224 pp., £11.00. ISBN 0-435-54060-2. Daniel J. Kevles, The Physicists. The History of a Scientific Community in Modern America (New York: A. Social Studies of Science, 9(4): 511–522, November 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277900900411.

Moss:2016:DSM

[Mos16] Timothy Moss. Discarded surrogates, modified traditions, welcome complements: The chequered careers of alternative technologies in Berlin’s infrastructure systems. Social Studies of Science, 46(4):559–582, August 1, 2016. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716657205.

McKinlay:1987:MDI

[MP87] Andrew McKinlay and Jonathan Potter. Model discourse: Interpretative repertoires in scientists’ conference talk. Social Studies of Science, 17(3):443–463, August 1, 1987. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631287017003003.

Mulkay:1987:CMD

[MPA87] Michael Mulkay, Trevor Pinch, and Malcolm Ashmore. Colonizing the mind: Dilemmas in the application of social science. Social Studies of Science, 17(2):231–256, May 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631287017002002.

Molella:1973:TIM

[MR73] Arthur P. Molella and Nathan Reingold. Theorists and ingenious mechanics: Joseph Henry defines science. Social Studies of Science, 3(4):323–351, October 1, 1973. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277300300401.

Monteiro:2017:SCK

[MR17] Marko Monteiro and Raoni Rajão. Scientists as citizens and knowers in the detection of deforestation in the Amazon. Social Studies of Science, 47(4):466–484, August 1, 2017. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716679746.

Mazur:2001:BAM

[MRL01] Allan Mazur, Stanley Rothman, and S. Robert Lichter. Biases about man-made cancer among researchers. Social Studies of Science, 31(5):771–778, October 1, 2001. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031005004.

MacKenzie:1988:SNWa

[MS88a] Donald MacKenzie and Graham Spinardi. The shaping of nuclear weapon system technology: US fleet ballistic missile guidance and navigation: I: From Polaris to Poseidon. Social Studies of Science, 18(3):419–463, August 1, 1988. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631288018003002.

MacKenzie:1988:SNWb

[MS88b] Donald MacKenzie and Graham Spinardi. The shaping of nuclear weapon system technology: US fleet ballistic missile guidance and navigation: II: ‘going for broke’ — the path to Trident II. Social Studies of Science, 18(4):581–624, November 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631288018004002.

Maynard:2000:TSS

[MS00] Douglas W. Maynard and Nora Cate Schaeffer. Toward a sociology of social scientific knowledge: Survey research and ethnomethodology’s asymmetric alternates. Social Studies of Science, 30(3):323–370, June 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631200030003001.

MacKenzie:2014:DBA

[MS14a] Donald MacKenzie and Taylor Spears. ‘A device for being able to book P&L’: The organizational embedding of the Gaussian copula. Social Studies of Science, 44(3):418–440, June 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713517158.

MacKenzie:2014:FKW

[MS14b] Donald MacKenzie and Taylor Spears. ‘The formula that killed Wall Street’: The Gaussian copula and modelling practices in investment banking. Social Studies of Science, 44(3):393–417, June 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312713517157.

Molldrem:2024:HPC

[MS24] Stephen Molldrem and Anthony K. J. Smith. Health policy counterpublics: Enacting collective resistances to US molecular HIV surveillance and cluster detection and response programs. Social Studies of Science, 54(3):451–477, June 1, 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231211933.

Mukerji:2007:GFS

[Muk07a] Chandra Mukerji. The Great Forestry Survey of 1669–1671: The use of archives for political reform. Social Studies of Science, 37(2):227–253, April 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706063785.

Mukerji:2007:SPC

[Muk07b] Chandra Mukerji. Stewardship politics and the control of wild weather: Levees, seawalls, and state building in 17th-Century France. Social Studies of Science, 37(1):127–133, February 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312706069435.

Mulkay:1971:SSS

[Mul71] Michael Mulkay. Some suggestions for sociological research. Social Studies of Science, 1(2):207–213, April 1, 1971. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277100100205.

Mullins:1972:SEA

[Mul72] Nicholas C. Mullins. The structure of an elite: the advisory structure of the U.S. Public Health Service. Social Studies of Science, 2(1):3–29, January 1, 1972. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277200200101.

Mullins:1973:DSS

[Mul73] Nicholas C. Mullins. The development of specialties in social science: The case of ethnomethodology. Social Studies of Science, 3(3):245–273, July 1, 1973. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277300300302.

Mulkay:1974:CDM

[Mul74] Michael Mulkay. Conceptual displacement and migration in science: a prefatory paper. Social Studies of Science, 4(3):205–234, July 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277400400301.

Mulkay:1976:MRS

[Mul76] Michael Mulkay. The mediating role of the scientific elite. Social Studies of Science, 6(3–4):445–470, September 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277600600308.

Mulkay:1979:KUI

[Mul79] Michael Mulkay. Knowledge and utility: Implications for the sociology of knowledge. Social Studies of Science, 9(1): 63–80, February 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277900900103.

Mulkay:1984:STB

[Mul84] Michael Mulkay. The scientist talks back: a one-act play, with a moral, about replication in science and reflexivity in sociology. Social Studies of Science, 14(2):265–282, May 1, 1984.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014002008.

Mulkay:1993:RHF

[Mul93] Michael Mulkay. Rhetorics of hope and fear in the great embryo debate. Social Studies of Science, 23(4):721–742, November 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631293023004004.

Mulkay:1994:TPE

[Mul94] Michael Mulkay. The triumph of the pre-embryo: Interpretations of the human embryo in parliamentary debate over embryo research. Social Studies of Science, 24(4):611–639, November 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631279402400401.

Mulkay:1995:GER

[Mul95a] Michael Mulkay. Galileo and the embryos: Religion and science in parliamentary debate over research on human embryos. Social Studies of Science, 25(3):499–532, August 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631295025003004.

Mulkay:1995:PAR

[Mul95b] Michael Mulkay. Parliamentary ambivalence in relation to embryo research. Social Studies of Science, 25(1):149–163, February 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631295025001007.

Muller:1995:CBB

[Mül95c] Karel Müller. Changes on the ‘borderlines’ between research and industry following economic transformation in the Czech Republic. Social Studies of Science, 25(4):805–828, November 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631295025004010.

Mulcahy:1998:DUU

[Mul98] Monica Dianne Mulcahy. Designing the user/using the design: The shifting relations of a curriculum technology change. Social Studies of Science, 28(1):5–37, February 1, 1998. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631298028001001.

Muller:2018:DCA

[Mül18] Sophie Merit Müller. Distributed corporeality: Anatomy, knowledge and the technological reconfiguration of bodies in ballet. Social Studies of Science, 48(6):869–890, December 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312718811636.

Murray:2013:TBD

[Mur13] Cameron M. Murray. Tracing biomedical dispositifs: Play, rigor, and cooperation in medical oncology. Social Studies of Science, 43(5):792–797, October 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312713487024.

Murphy:2015:UCT

[Mur15] Michelle Murphy. Unsettling care: Troubling transnational itineraries of care in feminist health practices. Social Studies of Science, 45(5):717–737, October 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312715589136.

Mirowski:2005:CRO

[MV05] Philip Mirowski and Robert Van Horn. The contract research organization and the commercialization of scientific research. Social Studies of Science, 35(4):503–548, August 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312705052103.

Mcharek:2023:PFT

[MvO23] Amade M’charek and Irene van Oorschot. The politics of face and the trouble with race: Exploring relations at the interface between the individual and the collective in forensic practice.

Social Studies of Science, 53(6):813–825, December 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231211817.

Moed:1985:CRI

[MvR85] H. F. Moed and A. F. J. van Raan. Critical remarks on Irvine and Martin’s methodology for evaluating scientific performance. Social Studies of Science, 15(3):539–547, August 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631285015003007. See [IM85, MI85].

Myers:1985:TKC

[Mye85] Greg Myers. Texts as knowledge claims: The social construction of two biology articles. Social Studies of Science, 15(4): 593–630, November 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631285015004002.

Myers:1989:PCD

[Mye89] Greg Myers. Postscript: Conversation over dinner, on the usefulness of paradox in controversies. Social Studies of Science, 19 (4):668–670, November 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631289019004009. See

[Woo89].

Myers:1990:SSS

[Mye90] Greg Myers. Sociology of science without the sociology. Social Studies of Science, 20(3):559–563, August 1, 1990. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020003008.

Myers:1991:PCL

[Mye91] Greg Myers. Politeness and certainty: The language of collaboration in an Al project. Social Studies of Science, 21 (1):37–73, February 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631291021001004.

Myers:1995:DIW

[Mye95] Greg Myers. From discovery to invention: The writing and rewriting of two patents. Social Studies of Science, 25(1): 57–105, February 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631295025001004.

Myers:1998:FTT

[Mye98] Greg Myers. Facts, thinking and talk. Social Studies of Science, 28(4):666–672, August 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631298028004010.

Myers:2008:MEB

[Mye08] Natasha Myers. Molecular embodiments and the body-work of modeling in protein crystallography. Social Studies of Science, 38(2):163–199, April 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312707082969.

Nadel:1980:MCA

[Nad80] Edward Nadel. Multivariate citation analysis and the changing cognitive organization in a specialty of physics. Social Studies of Science, 10(4):449–473, November 1, 1980. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278001000402.

Nadel:1983:CCC

[Nad83] Edward Nadel. Commitment and co-citation: an indicator of incommensurability in patterns of formal communication. Social Studies of Science, 13(2):255–283, May 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631283013002004.

Nadel:1984:PPR

[Nad84] Edward Nadel. Programmes and paradigms: a reply to Zandvoort. Social Studies of Science, 14(3):446–450, August 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014003006. See [Zan84].

Nafus:2024:UAD

[Naf24] Dawn Nafus. Unclearing the air: Data’s unexpected limitations for environmental advocacy. Social Studies of Science, 54(2):163–183, April 1, 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127231201169.

Nakajima:2012:KSJ

[Nak12] Hideto Nakajima. Kuhn’s structure in Japan. Social Studies of Science, 42(3):462–466, June 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312712437619.

Nandy:1972:DCS

[Nan72] Ashis Nandy. Defiance and conformity in science: the identity of Jagadis Chandra Bose. Social Studies of Science, 2 (1):31–85, January 1, 1972. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277200200102.

Navon:2011:GDH

[Nav11] Daniel Navon. Genomic designation: How genetics can delineate new, phenotypically diffuse medical categories. Social Studies of Science, 41(2):203–226, April 1, 2011. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312710391923.

Nelkin:1971:SEC

[Nel71] Dorothy Nelkin. Scientists in an environmental controversy. Social Studies of Science, 1(3–4):245–261, October 1, 1971. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277100100301.

Nelkin:1975:PIT

[Nel75] Dorothy Nelkin. The political impact of technical expertise. Social Studies of Science, 5(1):35–54, February 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277500500103.

Nelkin:1977:SPR

[Nel77] Dorothy Nelkin. Scientists and professional responsibility: The experience of American ecologists. Social Studies of Science, 7 (1):75–95, February 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277700700110.

Nelson:2008:BSG

[Nel08] Alondra Nelson. Bio science: Genetic genealogy testing and the pursuit of African ancestry. Social Studies of Science, 38 (5):759–783, October 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312708091929.

Nelson:2013:MMH

[Nel13] Nicole C. Nelson. Modeling mouse, human, and discipline: Epistemic scaffolds in animal behavior genetics. Social Studies of Science, 43(1):3–29, February 1, 2013. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712463815.

Nelson:2022:CNA

[Nel22] Nicole Nelson. Call for new Associate Editors. Social Studies of Science, 52(6):809–811, December 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

Nesvetailov:1995:CCP

[Nes95] Gennady Nesvetailov. Changing centre-periphery relations in the former Soviet republics: The case of Belarus. Social Studies of Science, 25(4):853–871, November 1, 1995. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631295025004012.

Neve:1980:ERN

[Nev80] Michael Neve. Essay review: The naturalization of science: Barry Barnes and Steven Shapin (eds), Natural Order: Historical Studies of Scientific Culture (Beverly Hills, Calif. and London: Sage Publications, 1979), 255 pp., £10.00, $18.50 (£5.00, $9.95 pbk.). ISBN 0-8039-0958-6 (0-8039-0959-4 pbk.). Roy Wallis (ed.), On the Margins of Science: The Social Construction of Rejected Knowledge (Keele, Staffs.: University of Keele, Sociological Review Monograph No. 27, 1979), 337 pp., £4.95, $12.00. ISBN 0-904425-06-1 pbk. Ludmilla J. Jordanova and Roy S. Porter (eds), Images of the Earth: Essays in the History of the Environmental Sciences (Chalfont St Giles, Bucks.: The British Society for the History of Science Monographs, 1979), xx + 282 pp., £5.95 (£6.55 outside UK). ISBN-0-906450-00-4 pbk. Social Studies of Science, 10(3):375–391, August 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278001000306.

Narin:1983:HCS

[NFC83] Francis Narin, J. Davidson Frame, and Mark P. Carpenter. Highly cited Soviet papers: an exploratory investigation. Social Studies of Science, 13(2):307–319, May 1, 1983. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631283013002006.

Nelkin:1974:ICT

[NHM74] Dorothy Nelkin, Cary Hershey, and Dennis Mueller. The impact and control of technological change. Social Studies of Science, 4(1):97–103, January 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277400400107.

Nickles:1984:RFC

[Nic84] Thomas Nickles. A revolution that failed: Collins and Pinch on the paranormal. Social Studies of Science, 14(2):297–308, May 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631284014002011.

Nickles:1990:HTS

[Nic90] Thomas Nickles. How to talk with sociologists (or philosophers). Social Studies of Science, 20(4):633–638, November 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631290020004002.

Nicolson:1996:MIH

[Nic96] Malcolm Nicolson. Medical innovations: Historiography, heterogeneity and ‘the mangle of practice’. Social Studies of Science, 26(4):863–874, November 1, 1996. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631296026004007.

Nicolson:1997:IHV

[Nic97] Malcolm Nicolson. Interdisciplinary history: Visions of empire, dreams of youth. Social Studies of Science, 27(5):805–814, October 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631297027005004.

Nissani:1995:POI

[Nis95] Moti Nissani. The plight of the obscure innovator in science: a few reflections on Campanario’s note. Social Studies of Science, 25(1):165–183, February 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631295025001008. See

[Cam93].

Nishizaka:2011:EOR

[Nis11] Aug Nishizaka. The embodied organization of a real-time fetus: The visible and the invisible in prenatal ultrasound examinations. Social Studies of Science, 41(3):309–336, June 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312710386842.

Nik-Khah:2014:NPS

[NK14a] Edward Nik-Khah. Neoliberal pharmaceutical science and the Chicago school of economics. Social Studies of Science, 44 (4):489–517, August 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312714520864.

Nik-Khah:2014:PPR

[NK14b] Edward Nik-Khah. ‘Power to the people’: a reply to Healy, Mangin, and Applbaum. Social Studies of Science, 44(4):524– 530, August 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/full/10.1177/0306312714541495. See [HMA14].

Noll:1986:DC

[NKB86] Roger Noll, Carlos Kruytbosch, and Stuart Blume. Discussants’ comments. Social Studies of Science, 16(1):135–150, February 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631286016001007.

Nielsen:2018:TMP

[NL18] Karen Dam Nielsen and Henriette Langstrup. Tactics of material participation: How patients shape their engagement through e-health. Social Studies of Science, 48(2):259–282, April 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312718769156.

Neale:2020:FBS

[NM20] Timothy Neale and Daniel May. Fuzzy boundaries: Simulation and expertise in bushfire prediction. Social Studies of Science, 50(6):837–859, December 1, 2020. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720906869.

Nolan:1990:PVI

[Nol90]  Richard Nolan. Progress: The very idea!  Social Stud-

ies of Science, 20(1):175–184, February 1, 1990. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020001010.

Nola:1992:OHI

[Nol92] Robert Nola. Ordinary human inference as refutation of the strong programme. Social Studies of Science, 22(1):107–129, February 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/0306312792022001007.

Norton:1978:KPS

[Nor78] Bernard J. Norton. Karl Pearson and statistics: The social origins of scientific innovation. Social Studies of Science, 8 (1):3–34, February 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277800800101.

Norton:2017:FMP

[Nor17] Aaron T. Norton. Foreskin and the molecular politics of risk. Social Studies of Science, 47(5):655–680, October 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312717707360.

Nowotny:1987:BSM

[Now87] Helga Nowotny. The beginnings of scientific modernity. Social Studies of Science, 17(4):753–759, November 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631287017004008.

Nelson:2023:FRB

[NTW^{+}23] Nicole C. Nelson, Stefan Timmermans, Andrew Warwick, Zdeněk Konopásek, Russell E. Vance, and Wen-Hua Kuo. On first reading Bruno Latour. Social Studies of Science, 53 (2):174–179, April 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127231158086.

Nurok:2003:EME

[Nur03] Michael Nurok. Elements of the medical emergency’s epistemological alignment: 18–20th-Century perspectives. Social Studies of Science, 33(4):563–579, August 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312703334004.

OConnell:1993:MCU

[O’C93] Joseph O’Connell. Metrology: The creation of universality by the circulation of particulars. Social Studies of Science, 23(1): 129–173, February 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631293023001005.

Oikkonen:2017:ATT

[Oik17] Venla Oikkonen. Affect, technoscience and textual analysis: Interrogating the affective dynamics of the Zika epidemic through media texts. Social Studies of Science, 47(5):681–702, October 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312717723760.

Olson:2018:TAW

[OL18] Philip Olson and Christine Labuski. ‘There’s always a [white] man in the loop’: The gendered and racialized politics of civilian drones. Social Studies of Science, 48(4):540–563, August 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312718792619.

Olazaran:1996:SSO

[Ola96] Mikel Olazaran. A sociological study of the official history of the perceptrons controversy. Social Studies of Science, 26 (3):611–659, August 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631296026003005.

Olby:1984:SCW

[Olb84] Robert Olby. The sheriff and the cowboys: or Weaver’s support of Astbury and Pauling. Social Studies of Science, 14 (2):244–247, May 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631284014002005. See re-

ply [AA84].

Oldroyd:1988:BRP

[Old88] David Oldroyd. Book review: Peter J. Bowler, Evolution: The History of an Idea. Social Studies of Science, 18(3):562–564, August 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631288018003011.

Oldroyd:1990:SHS

[Old90a] David Oldroyd. Social and historical studies of science in the classroom? Social Studies of Science, 20(4):747–756, November 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631290020004008.

Oldroyd:1990:PPD

[Old90b] David R. Oldroyd. Picking at/on Pickering: The deconstruction of the social construction of scientific knowledge. Social Studies of Science, 20(4):638–657, November 1, 1990. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020004003.

Oldroyd:1993:HG

[Old93] David R. Oldroyd. Honouring a Goodman. Social Studies of Science, 23(3):583–590, August 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/0306312793023003008.

Oliver:2025:CMI

[Oli25] Catherine Oliver. Chicken metabolism, immobilization, and post-industrial production. Social Studies of Science, 55(1): 85–108, February 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127241247022.

OConnor:1976:SPB

[OM76] Jean G. O’Connor and A. J. Meadows. Specialization and professionalization in British geology. Social Studies of Science, 6 (1):77–89, February 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277600600104.

Onuigbo:1984:URP

[Onu84] Wilson I. B. Onuigbo. The utilization of request-a-print. Social Studies of Science, 14(1):94–96, February 1, 1984. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014001007.

OReilly:2012:RDP

[OOO12] Jessica O’Reilly, Naomi Oreskes, and Michael Oppenheimer. The rapid disintegration of projections: The West Antarctic Ice Sheet and the Intergovernmental Panel on Climate Change. Social Studies of Science, 42(5):709–731, October 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712448130.

Oppenheim:1979:CNP

[Opp79] Charles Oppenheim. Could the 1978 Nobel Prizewinner in Chemistry have been predicted? Social Studies of Science, 9 (4):507–508, November 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277900900407.

ORiordan:1983:RBW

[O’R83] Timothy O’Riordan. Review: Brain Wynne, Rationality and Ritual: The Windscale Inquiry and Nuclear Decisions in Britain, (Chalfont St Giles, Bucks.: The British Society for the History of Science, 1982). BSHS Monographs, No. 3. x + 222 pp. ISBN 0-906450-02-0. Obtainable from the Administrator, BSHS, Halfpenny Furze, Mill Lane, Chalfont St Giles, Bucks. HP8 4NR, England, UK: £4.90, $9.50, for BSHS members: £6.50, $13.50, for non-members. (Prices incl. p.&p.). Social Studies of Science, 13(4):621–626, November 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631283013004007.

Orange:1971:BAA

[Ora71] A. D. Orange. The British Association for the Advancement of Science: the provincial background. Social Studies of Science, 1 (3–4):315–329, October 1, 1971. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277100100304.

Oreskes:2003:CMU

[Ore03] Naomi Oreskes. A context of motivation: US Navy oceanographic research and the discovery of sea-floor hydrothermal vents. Social Studies of Science, 33(5):697–742, October 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312703335004.

Orlans:1973:DRA

[Orl73] Harold Orlans. ‘D & R’ allocations in the United States. Social Studies of Science, 3(2):119–159, April 1, 1973. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277300300202.

Oromaner:1977:CSL

[Oro77a] Mark Oromaner. The career of sociological literature: a diachronous study. Social Studies of Science, 7(1):126–132, February 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277700700114.

Oromaner:1977:PAR

[Oro77b] Mark Oromaner. Professional age and the reception of sociological publications: a test of the Zuckerman–Merton hypothesis. Social Studies of Science, 7(3):381–388, August 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700307.

OMahony:2005:BLP

[OS05a] Patrick O’Mahony and Mike Steffen Schäfer. The ‘book of life’ in the press: Comparing German and Irish media discourse on human genome research. Social Studies of Science, 35(1): 99–130, February 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312705046542.

Owen-Smith:2005:DDS

[OS05b] Jason Owen-Smith. Dockets, deals, and sagas: Commensuration and the rationalization of experience in university licensing. Social Studies of Science, 35(1):69–97, February 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312705047738.

Ozden-Schilling:2024:TNS

[ÖS24] Tom Özden-Schilling. Trust in numbers: Serious numbers and speculative fictions in rare earth elements exploration. Social Studies of Science, 54(2):281–304, April 1, 2024. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231205044.

Osborne:1993:MPM

[Osb93] Thomas Osborne. Mobilizing psychoanalysis: Michael Balint and the general practitioners. Social Studies of Science, 23(1): 175–200, February 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631293023001006.

Oudshoorn:1990:MSH

[Oud90] Nelly Oudshoorn. On the making of sex hormones: Research materials and the production of knowledge. Social Studies of Science, 20(1):5–33, February 1, 1990. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020001001.

Oudshoorn:1997:PCP

[Oud97] Nelly Oudshoorn. From population control politics to chemicals: The WHO as an intermediary organization in contraceptive development. Social Studies of Science, 27(1):41–72, February 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631297027001003.

Oudshoorn:2012:HPM

[Oud12] Nelly Oudshoorn. How places matter: Telecare technologies and the changing spatial dimensions of healthcare. Social Studies of Science, 42(1):121–142, February 1, 2012. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312711431817.

Oudshoorn:2015:SCS

[Oud15] Nelly Oudshoorn. Sustaining cyborgs: Sensing and tuning agencies of pacemakers and implantable cardioverter defibrillators. Social Studies of Science, 45(1):56–76, February 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312714557377.

Outhwaite:1976:RRB

[Out76] William Outhwaite. Reviews: Roy Bhaskar, A Realist Theory of Science (Leeds: Leeds Books, 1975), 260 pp., £5.95 hardback, £2.25 paperback. ISBN 0-85952-013-7 (cloth) 0-85952-014-5 (paperback). Social Studies of Science, 6(1):123–127, February 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277600600108.

Plotkin-Amrami:2015:MNT

[PAB15] Galia Plotkin-Amrami and José Brunner. Making up ‘national trauma’ in Israel: From collective identity to collective vulnerability. Social Studies of Science, 45(4):525–545, August 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312715589846.

Palermo:1971:SRT

[Pal71] David S. Palermo. Is a scientific revolution taking place in psychology? Social Studies of Science, 1(2):135–155, April 1, 1971. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277100100202.

Palladino:1990:ETP

[Pal90] Paolo Palladino. Ecological theory and pest control practice: a study of the institutional and conceptual dimensions of a scientific debate. Social Studies of Science, 20(2):255–281, May 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631290020002003.

Palladino:1994:BRR

[Pal94] Paolo Palladino. Book review: The Right Tools for the Job: At Work in Twentieth-Century Life Sciences. Social Studies of Science, 24(2):404–409, May 1, 1994. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631279402400211.

Palladino:2002:BKP

[Pal02] Paolo Palladino. Between knowledge and practice: On medical professionals, patients, and the making of the genetics of cancer. Social Studies of Science, 32(1):137–165, February 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312702032001006.

Pancaldi:1980:HSS

[Pan80] Giuliano Pancaldi. The history and social studies of science in Italy. Social Studies of Science, 10(3):351–374, August 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631278001000305.

Panofsky:2011:GSD

[Pan11] Aaron Panofsky. Generating sociability to drive science: Patient advocacy organizations and genetics research. Social Studies of Science, 41(1):31–57, February 1, 2011. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312710385852.

Papadopoulos:2011:AOT

[Pap11] Dimitris Papadopoulos. Alter-ontologies: Towards a constituent politics in technoscience. Social Studies of Science, 41(2):177–201, April 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312710385853.

Pape:2021:CPM

[Pap21] Madeleine Pape. Co-production, multiplied: Enactments of sex as a biological variable in US biomedicine. Social Studies of Science, 51(3):339–363, June 1, 2021. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720985939.

Parssinen:1988:JMD

[Par88] Terry M. Parssinen. James McClenon, deviant science: The case of parapsychology. Social Studies of Science, 18(3):560–562, August 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631288018003010.

Parthasarathy:2005:AGM

[Par05] Shobita Parthasarathy. Architectures of genetic medicine: Comparing genetic testing for breast cancer in the USA and the UK. Social Studies of Science, 35(1):5–40, February 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312705047172.

Parthasarathy:2022:HSP

[Par22] Shobita Parthasarathy. How sanitary pads came to save the world: Knowing inclusive innovation through science and the marketplace. Social Studies of Science, 52(5):637–663, October 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127221122457.

Parthasarathy:2025:BPP

[Par25] Shobita Parthasarathy. From the bench to public policy: Enhancing public trust in science. Social Studies of Science, 55(2):295–302, April 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127241310587.

Pattison:1983:SIM

[Pat83] Michael Pattison. Scientists, inventors and the military in Britain, 1915–19: The munitions inventions department. Social Studies of Science, 13(4):521–568, November 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631283013004004.

Pinch:1984:SCF

[PB84] Trevor J. Pinch and Wiebe E. Bijker. The social construction of facts and artefacts: or how the sociology of science and the sociology of technology might benefit each other. Social Studies of Science, 14(3):399–441, August 1, 1984. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/030631284014003004. See response [Rus86] and reply [PB86].

Pinch:1986:SRN

[PB86] Trevor Pinch and Wiebe Bijker. Science, Relativism and the new sociology of technology: Reply to Russell. Social Studies of Science, 16(2):347–360, May 1, 1986. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312786016002009. See [PB84, Rus86].

Pinch:2004:SSN

[PB04] Trevor Pinch and Karin Bijsterveld. Sound studies: New technologies and music. Social Studies of Science, 34(5):635–648, October 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/0306312704047615.

Pullen-Blasnik:2024:YAM

[PBEW24] Hannah Pullen-Blasnik, Gil Eyal, and Amy Weissenbach. ‘Is your accuser me, or is it the software?’ ambiguity and contested expertise in probabilistic DNA profiling. Social Studies of Science, 54(1):30–58, February 1, 2024. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231186646.

Popa:2025:PCD

[PBSK25] Eugen Octav Popa, Vincent Blok, Cornelius Schubert, and Georgios Katsoukis. Path creation as a discursive process: a study of discussion starters in the field of solar fuels. Social Studies of Science, 55(1):62–84, February 2025. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127241271024.

Pinch:1984:PSP

[PC84] T. J. Pinch and H. M. Collins. Private science and public knowledge: The committee for the scientific investigation of the claims of the paranormal and its use of the literature. Social Studies of Science, 14(4):521–546, November 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014004002.

Parker:2012:BAT

[PC12] John Parker and Beatrice Crona. On being all things to all people: Boundary organizations and the contemporary research university. Social Studies of Science, 42(2):262–289, April 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312711435833.

Palmer:1998:PGW

[PCL98] Anthony Palmer, Harry Collins, and Michael Lynch. Peter G. Winch (14 January 1926–27 April 1997). Social Studies of Science, 28(1):169–175, February 1, 1998. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631298028001006.

Panofsky:2019:GAT

[PD19] Aaron Panofsky and Joan Donovan. Genetic ancestry testing among white nationalists: From identity repair to citizen science. Social Studies of Science, 49(5):653–681, October 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312719861434.

Pels:1996:PS

[Pel96] Dick Pels. The politics of symmetry. Social Studies of Science, 26(2):277–304, May 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631296026002004.

Pelaez:1999:SPC

[Pel99] Eloína Peláez. The stored-program computer: Two conceptions. Social Studies of Science, 29(3):359–389, June 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029003002.

Pelizza:2021:ITA

[Pel21] Annalisa Pelizza. Identification as translation: The art of choosing the right spokespersons at the securitized border. Social Studies of Science, 51(4):487–511, August 1, 2021. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720983932.

Perkins:1989:MNS

[Per89] John Perkins. Murderous Nazi science. Social Studies of Science, 19(2):368–370, May 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631289019002010.

Perlman:2004:GEM

[Per04] Marc Perlman. Golden ears and meter readers: The contest for epistemic authority in audiophilia. Social Studies of Science, 34 (5):783–807, October 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312704047613.

Petersen:2000:MTA

[Pet00] Arthur C. Petersen. Models as technological artefacts. Social Studies of Science, 30(5):793–799, October 1, 2000. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030005006.

Pettit:2015:SMH

[Pet15] Michael Pettit. Subject matter: Human behavior, psychological expertise, and therapeutic lives. Social Studies of Science, 45 (1):146–158, February 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312714546366.

Peterson:2017:DFM

[Pet17] David Peterson. The depth of fields: Managing focus in the epistemic subcultures of mind and brain science. Social Studies of Science, 47(1):53–74, February 1, 2017. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716663047.

Pfetsch:1971:RPW

[Pfe71] Frank Pfetsch. Reviews: Peter Weingart, Die Amerikanische Wissenschaftslobby. Dusseldorf: Bertelsmann Universitätsverlag, 1970. 253 pp. DM 28. Social Studies of Science, 1(2):235–237, April 1, 1971. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277100100208.

Pfetsch:1979:FDG

[Pfe79] Frank R. Pfetsch. The ‘finalization’ debate in Germany: Some comments and explanations. Social Studies of Science, 9(1): 115–124, February 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277900900107.

Petty:2011:ERH

[PH11] JuLeigh Petty and Carol A. Heimer. Extending the rails: How research reshapes clinics. Social Studies of Science, 41 (3):337–360, June 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312710396402.

Paxson:2014:PPM

[PH14] Heather Paxson and Stefan Helmreich. The perils and promises of microbial abundance: Novel natures and model ecosystems, from artisanal cheese to alien seas. Social Studies of Science, 44(2):165–193, April 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312713505003.

Phillips:2016:TMS

[Phi16] Christopher J. Phillips. The taste machine: Sense, subjectivity, and statistics in the California wine world. Social Studies of Science, 46(3):461–481, June 1, 2016. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716651504.

Pesch:2019:FFP

[PI19] Udo Pesch and Georgy Ishmaev. Fictions and frictions: Promises, transaction costs and the innovation of network technologies. Social Studies of Science, 49(2):264–277, April 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312719838339.

Pickering:1980:EAC

[Pic80a] Andy Pickering. Exemplars and analogies: a comment on Crane’s study of Kuhnian paradigms in high energy physics. Social Studies of Science, 10(4):497–502, November 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/030631278001000404. See [Cra80a, Cra80b, Pic80b].

Pickering:1980:RC

[Pic80b] Andy Pickering. Reply to Crane. Social Studies of Science, 10 (4):507–508, November 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631278001000406. See

[Cra80b].

Pickering:1981:CCC

[Pic81] Andrew Pickering. Constraints on controversy: The case of the magnetic monopole. Social Studies of Science, 11(1):

63–93, February 1, 1981. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631278101100104.

Pickering:1990:KPM

[Pic90] Andy Pickering. Knowledge, practice and mere construction. Social Studies of Science, 20(4):682–729, November 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020004006. See [RB90b].

Pickering:1991:PNB

[Pic91] Andy Pickering. Philosophy naturalized a bit. Social Studies of Science, 21(3):575–584, August 1, 1991. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631291021003011.

Picart:1994:SCF

[Pic94] Caroline Joan S. Picart. Scientific controversy as farce: The Benveniste–Maddox counter trials. Social Studies of Science, 24(1):7–37, February 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631279402400102.

Picart:1997:BBR

[Pic97] Caroline Joan Picart. Blurring boundaries: a reply to Fadlon and Lewin-Epstein. Social Studies of Science, 27(1):142–146, February 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631297027001008. See [FLE97].

Pickering:1999:LBL

[Pic99] Andy Pickering. In the land of the blind …:: Thoughts on Gingras. Social Studies of Science, 29(2):307–311, April 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/030631299029002015. See [Gin97] and reply [Gin99].

Pickering:2002:CMA

[Pic02] Andrew Pickering. Cybernetics and the mangle: Ashby, Beer and Pask. Social Studies of Science, 32(3):413–437, June 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312702032003003.

Pickering:2012:WSK

[Pic12] Andrew Pickering. The world since Kuhn. Social Studies of Science, 42(3):467–473, June 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312712440503.

Pickersgill:2013:HPB

[Pic13] Martyn Pickersgill. How personality became treatable: The mutual constitution of clinical knowledge and mental health law. Social Studies of Science, 43(1):30–53, February 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712457722.

Pickersgill:2019:AAP

[Pic19] Martyn Pickersgill. Access, accountability, and the proliferation of psychological therapy: On the introduction of the IAPT initiative and the transformation of mental healthcare. Social Studies of Science, 49(4):627–650, August 1, 2019. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312719834070.

Pihl:2025:BTH

[Pih25] Eva Vibeke Kofoed Pihl. Bettie’s travels: How pigs enable new connections between human health innovations and industrial agricultural pork production in Denmark. Social Studies of Science, 55(1):109–130, February 2025. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127241268772.

Pinch:1979:NEP

[Pin79] Trevor J. Pinch. Normal explanations of the paranormal: The demarcation problem and fraud in parapsychology. Social Studies of Science, 9(3):329–348, August 1, 1979. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277900900303.

Pinch:1981:RH

[Pin81a] Trevor Pinch. Reply to Hardin. Social Studies of Science, 11(2):255–257, May 1, 1981. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631278101100204. See

[Har81a].

Pinch:1981:SSP

[Pin81b] Trevor J. Pinch. The Sun-set: The presentation of certainty in scientific life. Social Studies of Science, 11(1):131–158, February 1, 1981. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631278101100106.

Pinch:1985:TAS

[Pin85] Trevor Pinch. Towards an analysis of scientific observation: The externality and evidential significance of observational reports in physics. Social Studies of Science, 15(1):3–36, February 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631285015001001.

Pinch:1986:SV

[Pin86] Trevor Pinch. Strata various. Social Studies of Science, 16(4): 705–713, November 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631286016004007.

Pinch:1990:DRB

[Pin90] Trevor Pinch. Deconstructing Roth and Barrett. Social Studies of Science, 20(4):658–663, November 1, 1990. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020004004.

Pinch:1992:OBB

[Pin92] Trevor J. Pinch. Opening black boxes: Science, technology and society. Social Studies of Science, 22(3):487–510, August 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312792022003003.

Pinch:1993:GS

[Pin93] Trevor Pinch. Generations of SSK. Social Studies of Science, 23(2):363–373, May 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631293023002006.

Pinch:1995:PDS

[Pin95] Trevor J. Pinch. In and out of the Petri dish: Science and S&TS. Social Studies of Science, 25(2):334–337, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631295025002012.

Pinch:1997:KCR

[Pin97] T. J. Pinch. Kuhn — the conservative and radical interpretations: Are some Mertonians ‘Kuhnians’ and some Kuhnians ‘Mertonians’? Social Studies of Science, 27(3):465–482, June 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631297027003004.

Pinch:1999:FRM

[Pin99a] Trevor Pinch. Final response to McKinney. Social Studies of Science, 29(2):246–247, April 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631299029002007.

See [Pin99b].

Pinch:1999:HHR

[Pin99b] Trevor Pinch. Half a house: A response to McKinney. Social Studies of Science, 29(2):235–240, April 1, 1999. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029002005. See [Pin99a, McK99].

Pinnick:1999:CSS

[Pin99c] Cassandra L. Pinnick. Caught in a sandy shoal of the shallow: Reply to Shapin and Schaffer. Social Studies of Science, 29(2):253–257, April 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631299029002009. See

[SS99b, SS99a].

Pinch:2003:MDO

[Pin03] Trevor Pinch. In memoriam: David Owen Edge. Social Studies of Science, 33(2):192–193, April 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127030332011.

Pinch:2011:REK

[Pin11] Trevor Pinch. Review essay: Karen Barad, Quantum mechanics, and the paradox of mutual exclusivity. Social Studies of Science, 41(3):431–441, June 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312711400657.

Pinto:2015:TAN

[Pin15] Manuela Fernández Pinto. Tensions in agnotology: Normativity in the studies of commercially driven ignorance. Social Studies of Science, 45(2):294–315, April 1, 2015. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312714565491.

Pinzur:2021:IOM

[Pin21] David Pinzur. Infrastructure, ontology and meaning: The endogenous development of economic ideas. Social Studies of Science, 51(6):914–937, December 1, 2021. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/03063127211011524.

Pfotenhauer:2017:PDI

[PJ17] Sebastian Pfotenhauer and Sheila Jasanoff. Panacea or diagnosis? Imaginaries of innovation and the ‘MIT model’ in three political cultures. Social Studies of Science, 47(6):783–810, December 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312717706110.

Prainsack:2009:DBB

[PK09] Barbara Prainsack and Martin Kitzberger. DNA behind bars: Other ways of knowing forensic DNA technologies. Social Studies of Science, 39(1):51–79, February 1, 2009. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312708097289.

Palmer:1988:OUF

[PKS88] David D. Palmer, Lauriston R. King, and W. Wayne Shannon. Oceanographers and the US Federal patron: Perceptions of agency–university relations. Social Studies of Science, 18 (1):73–90, February 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631288018001003.

Plajas:2023:IRV

[Plá23] Ildikó Zonga Plájás. InterFaces: On the relationality of vision, face and race in practices of identification. A multimodal intervention. Social Studies of Science, 53(6):938–953, December 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127231151237.

Plemons:2014:DSD

[Ple14] Eric D. Plemons. Description of sex difference as prescription for sex change: On the origins of facial feminization surgery. Social Studies of Science, 44(5):657–679, October 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312714531349.

Pfotenhauer:2022:PS

[PLPS22] Sebastian Pfotenhauer, Brice Laurent, Kyriaki Papageorgiou, and Jack Stilgoe. The politics of scaling. Social Studies of Science, 52(1):3–34, February 1, 2022. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/03063127211048945.

Petersen:1979:PSL

[PM79] James C. Petersen and Gerald E. Markle. Politics and science in the Laetrile controversy. Social Studies of Science, 9(2):139–166, May 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277900900201.

Potter:1989:DPR

[PM89] Jonathan Potter and Andy McKinlay. Discourse — philosophy — reflexivity: Comment on Halfpenny. Social Studies of Science, 19(1):137–145, February 1, 1989. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019001005. See [GM82, Hal88, Hal89].

Pickering:1987:CRQ

[PN87] Andy Pickering and Edward Nadel. Charm revisited: a quantitative analysis of the HEP literature. Social Studies of Science, 17(1):87–113, February 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631287017001004.

Poros:2004:WRK

[PN04] Maritsa V. Poros and Elizabeth Needham. Writings of Robert K. Merton. Social Studies of Science, 34(6):863–878, December 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312704048252.

Pollock:2014:PPK

[Pol14] Anne Pollock. Places of pharmaceutical knowledge-making: Global health, postcolonial science, and hope in South African drug discovery. Social Studies of Science, 44(6):848–873, December 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312714543285.

Polleri:2020:PPU

[Pol20] Maxime Polleri. Post-political uncertainties: Governing nuclear controversies in post-Fukushima Japan. Social Studies of Science, 50(4):567–588, August 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312719889405.

Popa:2021:ISH

[Pop21] Ioana Popa. Internationalized science and human rights activism during the late Cold War: The French Committee of Mathematicians. Social Studies of Science, 51(6):871–894, December 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/full/10.1177/03063127211033976.

Porter:1977:CAQ

[Por77] Alan L. Porter. Citation analysis: Queries and caveats. Social Studies of Science, 7(2):257–267, May 1, 1977. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700207.

Porter:1982:SP

[Por82] Roy Porter. Shutting people up. Social Studies of Science, 12 (3):467–476, August 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631282012003007.

Porter:1992:I

[Por92a] Theodore M. Porter. Introduction. Social Studies of Science, 22(4):595, November 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631292022004001.

Porter:1992:QAI

[Por92b] Theodore M. Porter. Quantification and the accounting ideal in science. Social Studies of Science, 22(4):633–651, November 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631292022004004.

Porter:1996:AMV

[Por96] Theodore M. Porter. Accounting made visible. Social Studies of Science, 26(3):712–715, August 1, 1996. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631296026003008.

Porcello:2004:SSL

[Por04] Thomas Porcello. Speaking of sound: Language and the professionalization of sound-recording engineers. Social Studies of Science, 34(5):733–758, October 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704047328.

Porter:2023:RTB

[Por23] Catherine Porter. Reflections on translating Bruno Latour. Social Studies of Science, 53(2):180–182, April 1, 2023. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231156650.

Potter:1987:DAT

[Pot87] Jonathan Potter. Discourse analysis and the turn of the reflexive screw: a response to Fuhrman and Oehler. Social Studies of Science, 17(1):171–177, February 1, 1987. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/030631287017001009. See [FO86] and reply [FO87].

Pottage:2011:LMS

[Pot11] Alain Pottage. Law machines: Scale models, forensic materiality and the making of modern patent law. Social Studies of Science, 41(5):621–643, October 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312711408484.

Powell:2001:WSA

[Pow01] Christopher Powell. What’s social about social construction?: How to bridge a political schism by reaffirming an ontological divide. Social Studies of Science, 31(2):299–307, April 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312701031002008.

Peterson:2021:SCS

[PP21] David Peterson and Aaron Panofsky. Self-correction in science: The diagnostic and integrative motives for replication. Social Studies of Science, 51(4):583–605, August 1, 2021. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/03063127211005551.

Passanti:2022:MET

[PP22] Cecilia Passanti and Marie-Emmanuelle Pommerolle. The (un)making of electoral transparency through technology: The 2017 Kenyan presidential election controversy. Social Studies of Science, 52(6):928–953, December 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

Philipps:2024:IRL

[PP24] Axel Philipps and Laura Paruschke. Inside regular lab meetings: The social construction of a research team and ideas in optical physics. Social Studies of Science, 54(2):257–280, April 1, 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127231188132.

Pinel:2020:CDV

[PPM20] Clémence Pinel, Barbara Prainsack, and Christopher Mckevitt. Caring for data: Value creation in a data-intensive research laboratory. Social Studies of Science, 50(2):175–197, April 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720906567.

Prasad:2005:SCO

[Pra05] Amit Prasad. Scientific culture in the ‘other’ theater of ‘modern science’: an analysis of the culture of magnetic resonance imaging research in India. Social Studies of Science, 35(3):463–489, June 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312705050831.

Prainsack:2006:NLR

[Pra06] Barbara Prainsack. ‘Negotiating life’: The regulation of human cloning and embryonic stem cell research in Israel. Social Studies of Science, 36(2):173–205, April 1, 2006. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706053348.

Prasad:2007:AAI

[Pra07] Amit Prasad. The (amorphous) anatomy of an invention: The case of magnetic resonance imaging (MRI). Social Studies of Science, 37(4):533–560, August 1, 2007. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706075334.

Prasad:2009:RMO

[Pra09] Amit Prasad. Review: Memory and its ‘other’: Geoffrey C. Bowker, Memory Practices in the Sciences (Cambridge, MA: MIT Press, 2005), 261 pp., £11.95 / $17.95 / EUR 13.99 (pbk), £22.95 / $90.00 / EUR 34.95 (hbk). ISBN 0-262-52489-9 (pbk), 0-262-02589-2 (hbk). Social Studies of Science, 39(1): 157–161, February 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312708097326.

Presser:1980:CQR

[Pre80] Stanley Presser. Collaboration and the quality of research. Social Studies of Science, 10(1):95–101, February 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278001000105.

Prentice:2005:ASS

[Pre05] Rachel Prentice. The anatomy of a surgical simulation: The mutual articulation of bodies in and through the machine. Social Studies of Science, 35(6):837–866, December 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312705053351.

Preda:2006:STA

[Pre06] Alex Preda. Socio-technical agency in financial markets: The case of the stock ticker. Social Studies of Science, 36(5): 753–782, October 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706059543.

Preda:2008:RES

[Pre08] Alex Preda. Review: Engineering statehood. Social Studies of Science, 38(2):313–318, April 1, 2008. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127080380020602.

Prickett:1971:RGD

[Pri71] Stephen Prickett. Reviews: Geoffrey Durrant, Wordsworth and the Great System. Cambridge: Cambridge University Press, 1970. £1.75. A. J. Meadows, The High Firmament. Leicester: Leicester University Press, 1969. £2.10. Social Studies of Science, 1(2):237–241, April 1, 1971. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277100100209.

Pritchard:2012:HHF

[Pri12] Sara B. Pritchard. From hydroimperialism to hydrocapitalism: ‘French’ hydraulics in France, North Africa, and beyond. Social Studies of Science, 42(4):591–615, August 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712443018.

Pyenson:1977:EPG

[PS77] Lewis Pyenson and Douglas Skopp. Educating physicists in Germany circa 1900. Social Studies of Science, 7(3):329–366, August 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277700700304.

Pinel:2024:DDT

[PS24] Clémence Pinel and Mette N. Svendsen. Domesticating data: Traveling and value-making in the data economy. Social Studies of Science, 54(3):429–450, June 1, 2024. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231212506.

Pearman:1972:PIS

[PT72] J. R. Pearman and Winnifred M. Thomet. Professional implications for social work in a cybernetic state. Social Studies of Science, 2(2):181–190, April 1, 1972. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277200200203.

Panikkar:2018:LMK

[PT18] Bindu Panikkar and Jonathan Tollefson. Land as material, knowledge and relationships: Resource extraction and subsistence imaginaries in Bristol Bay, Alaska. Social Studies of Science, 48(5):715–739, October 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312718803453.

Putz:2021:MEV

[Püt21] Ole Pütz. Managing exactness and vagueness in computer science work: Programming and self-repair in meetings. Social Studies of Science, 51(6):938–961, December 1, 2021. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/03063127211010972.

Packer:1995:IBP

[PW95] Kathryn Packer and Andrew Webster. Inventing boundaries: The prior art of the social world. Social Studies of Science, 25 (1):107–117, February 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631295025001005.

Pollock:2010:BEH

[PW10] Neil Pollock and Robin Williams. The business of expectations: How promissory organizations shape technology and innovation. Social Studies of Science, 40(4):525–548, August 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312710362275.

Philipps:2019:UIC

[PW19] Axel Philipps and Leonie Weißenborn. Unconventional ideas conventionally arranged: a study of grant proposals for exceptional research. Social Studies of Science, 49(6):884–897, December 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/full/10.1177/0306312719857156.

Pollock:2007:GSP

[PWD07] Neil Pollock, Robin Williams, and Luciana D’Adderio. Global software and its provenance: Generification work in the production of organizational software packages. Social Studies of Science, 37(2):254–280, April 1, 2007. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706066022.

Pycior:1993:RBC

[Pyc93] Helena M. Pycior. Reaping the benefits of collaboration while avoiding its pitfalls: Marie Curie’s rise to scientific prominence. Social Studies of Science, 23(2):301–323, May 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631293023002003.

Richards:1996:MSP

[RA96] Evelleen Richards and Malcolm Ashmore. More sauce please! The politics of SSK: Neutrality, commitment and beyond. Social Studies of Science, 26(2):219–228, May 1, 1996. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631296026002001.

Rabinowitz:2022:IJA

[RA22] Or Rabinowitz and Yehonatan Abramson. Imagining a ‘Jewish atom bomb’, constructing a scientific diaspora. Social Studies of Science, 52(2):253–276, April 1, 2022. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/03063127221077313.

Raab:1997:SS

[Raa97] Charles D. Raab. Surveying surveillance. Social Studies of Science, 27(6):951–956, December 1, 1997. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631297027006005.

Rabkin:1976:RNI

[Rab76] Yakov M. Rabkin. Reviews: Naukometricheskie issledovania v khimii (Scientometric Studies in Chemistry), Moskva: Izd. Moskovskogo Universiteta, 1974. 136 pp. 33kop. Social Studies of Science, 6(1):128–132, February 1, 1976. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277600600109.

Radder:1986:ETI

[Rad86] Hans Radder. Experiment, technology and the intrinsic connection between knowledge and power. Social Studies of Science, 16 (4):663–683, November 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631286016004005.

Radder:1992:NRC

[Rad92] Hans Radder. Normative reflexions on constructivist approaches to science and technology. Social Studies of Science, 22(1): 141–173, February 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312792022001009.

Radder:1998:PS

[Rad98a] Hans Radder. The politics of STS. Social Studies of Science, 28(2):325–331, April 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631298028002005. See

[Sin98].

Radder:1998:STP

[Rad98b] Hans Radder. Second thoughts on the politics of STS: a response to the replies by Singleton and Wynne. Social Studies of Science, 28(2):344–348, April 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631298028002008.

Radin:2013:LLC

[Rad13] Joanna Radin. Latent life: Concepts and practices of human tissue preservation in the international biological program. Social Studies of Science, 43(4):484–508, August 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713476131.

Rahaman:2011:BNP

[Rah11] Mizanur Rahaman. Biotechnology, neoliberal politics of life and the spirit of biocapital. Social Studies of Science, 41(5): 759–763, October 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312711413824.

Rall:2006:HDB

[Ral06] Denise N. Rall. The ‘house that Dick built’: Constructing the team that built the bomb. Social Studies of Science, 36(6):

943–957, December 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706062676.

Ramsden:2002:CPS

[Ram02] Edmund Ramsden. Carving up population science: Eugenics, demography and the controversy over the ‘biological law’ of population growth. Social Studies of Science, 32(5–6): 857–899, December 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631270203200509.

Ramsden:2009:CSE

[Ram09] Edmund Ramsden. Confronting the stigma of eugenics: Genetics, demography and the problems of population. Social Studies of Science, 39(6):853–884, December 1, 2009. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312709335406.

Randalls:2010:WPW

[Ran10] Samuel Randalls. Weather profits: Weather derivatives and the commercialization of meteorology. Social Studies of Science, 40 (5):705–730, October 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312710378448.

Rankin:2017:ZPN

[Ran17] William Rankin. Zombie projects, negative networks, and multigenerational science: The temporality of the international map of the world. Social Studies of Science, 47(3):353–375, June 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312716658138.

Rappert:2001:DRA

[Rap01] Brian Rappert. The distribution and resolution of the ambiguities of technology, or why bobby can’t spray. Social Studies of Science, 31(4):557–591, August 1, 2001. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031004004.

Rappert:2005:PWC

[Rap05] Brian Rappert. Prohibitions, weapons and controversy: Managing the problems of ordering. Social Studies of Science, 35 (2):211–240, April 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312705046629.

Rasmussen:1998:ES

[Ras98] Nicolas Rasmussen. Down-to-Earth science. Social Studies of Science, 28(1):183–186, February 1, 1998. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631298028001008.

Rasmussen:2003:PP

[Ras03] Nicolas Rasmussen. A prescription for psychiatry. Social Studies of Science, 33(3):453–458, June 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127030333008.

Rasmussen:2004:MED

[Ras04] Nicolas Rasmussen. The moral economy of the drug companymedical scientist collaboration in interwar America. Social Studies of Science, 34(2):161–185, April 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704042623.

Rasmussen:2020:MMM

[Ras20] Erik Børve Rasmussen. Making and managing medical anomalies: Exploring the classification of ‘medically unexplained symptoms’. Social Studies of Science, 50(6):901–931, December 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312720940405.

Ravetz:1974:R

[Rav74] J. R. Ravetz. Review. Social Studies of Science, 4(3):295–297, July 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277400400306.

Ravetz:1999:LTS

[Rav99] Jerry Ravetz. Less than the sum of its parts: Or, studying periods, not problems. Social Studies of Science, 29(2):302–306, April 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631299029002014.

Rayner:1986:MRH

[Ray86] Steve Rayner. Management of radiation hazards in hospitals: Plural rationalities in a single institution. Social Studies of Science, 16(4):573–591, November 1, 1986. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631286016004002.

Rip:1975:SSR

[RB75] Arie Rip and Egbert Boeker. Scientists and social responsibility in the Netherlands. Social Studies of Science, 5(4): 457–484, November 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277500500406.

Roth:1990:DQ

[RB90a] Paul Roth and Robert Barrett. Deconstructing quarks. Social Studies of Science, 20(4):579–632, November 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020004001.

Roth:1990:RAS

[RB90b] Paul Roth and Robert Barrett. Reply: Aspects of sociological explanation. Social Studies of Science, 20(4):729–746, November 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631290020004007. See [Pic90].

Roth:1999:DLT

[RB99] Wolff-Michael Roth and G. Michael Bowen. Digitizing lizards: The topology of ‘vision’ in ecological fieldwork. Social Studies of Science, 29(5):719–764, October 1, 1999. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029005003.

Roth:2001:CSF

[RB01] Wolff-Michael Roth and G. Michael Bowen. ‘Creative solutions’ and ‘fibbing results’: Enculturation in field ecology. Social Studies of Science, 31(4):533–556, August 1, 2001. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031004003.

Rabeharisoa:2009:SWE

[RB09] Vololona Rabeharisoa and Pascale Bourret. Staging and weighting evidence in biomedicine: Comparing clinical practices in cancer genetics and psychiatric genetics. Social Studies of Science, 39(5):691–715, October 1, 2009. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312709103501.

Reinecke:2022:MAM

[RB22] David Reinecke and Jordan Bimm. The maintenance of ambiguity in Martian exobiology. Social Studies of Science, 52 (2):199–226, April 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/03063127221077207.

Rappert:2015:CCE

[RC15] Brian Rappert and Catelijne Coopmans. On conveying and not conveying expertise. Social Studies of Science, 45(4): 611–619, August 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312715595297.

Rosemann:2016:PIR

[RC16] Achim Rosemann and Nattaka Chaisinthop. The pluralization of the international: Resistance and alter-standardization in regenerative stem cell medicine. Social Studies of Science, 46(1): 112–139, February 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312715619783.

Richter:2018:NSS

[RCB18] Lauren Richter, Alissa Cordner, and Phil Brown. Non-stick science: Sixty years of research and (in)action on fluorinated compounds. Social Studies of Science, 48(5):691–714, October 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312718799960.

Restivo:2008:OBB

[RD08] Sal Restivo and Rachel Dowty. Obituary: Bernard Barber and Mary Douglas. Social Studies of Science, 38(4):635–640, August 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312708095712.

Roth:2003:FPP

[RDB03] Andrew L. Roth, Joshua Dunsby, and Lisa A. Bero. Framing processes in public commentary on US Federal tobacco control regulation. Social Studies of Science, 33(1):7–44, February 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312703033001038.

Reardon:2001:HGD

[Rea01] Jenny Reardon. The Human Genome Diversity Project: a case study in coproduction. Social Studies of Science, 31(3):357–388, June 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631201031003002.

Reddy:1975:ATV

[Red75] Amulya Kumar N. Reddy. Alternative technology: a viewpoint from India. Social Studies of Science, 5(3):331–342, August 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277500500304.

Redfield:2002:HLE

[Red02] Peter Redfield. The half-life of empire in outer space. Social Studies of Science, 32(5–6):791–825, December 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631270203200508.

Redfield:2016:FTB

[Red16] Peter Redfield. Fluid technologies: The bush pump, the LifeStraw(R) and microworlds of humanitarian design. Social Studies of Science, 46(2):159–183, April 1, 2016. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312715620061.

Rees:2001:PIO

[Ree01] Amanda Rees. Practising infanticide, observing narrative: Controversial texts in a field science. Social Studies of Science, 31 (4):507–531, August 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631201031004002.

Rees:2006:BRG

[Ree06] Teresa Rees. Book review: Gendered Opportunities in Biotech. Social Studies of Science, 36(1):164–168, February 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312705056408.

Rees:2007:RFP

[Ree07] Amanda Rees. Reflections on the field: Primatology, popular science and the politics of personhood. Social Studies of Science, 37(6):881–907, December 1, 2007. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312707077368.

Reingold:1980:TPL

[Rei80] Nathan Reingold. Through paradigm-land to a normal history of science. Social Studies of Science, 10(4):475–496, November 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631278001000403.

Reikowsky:2006:BRP

[Rei06] Ryan C. Reikowsky. Book review: Personalized medicine?: Expectations, hype, and the truth about pharmacogenetics: Adam Hedgecoe, The Politics of Personalised Medicine: Pharmacogenetics in the Clinic (Cambridge: Cambridge University Press, 2004)., 216 pp., £39.30 / $75.00 / EUR 64.90 (hbk), £17.99 / $32.99 / EUR 28.50 (pbk). ISBN: 0-521-84177-1 (hbk), 0-521-60265-3. Social Studies of Science, 36(1):161–163, February 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312705057334.

Reichmann:2013:EPH

[Rei13] Werner Reichmann. Epistemic participation: How to produce knowledge about the economic future. Social Studies of Science, 43(6):852–877, December 1, 2013. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713498641.

Reinecke:2021:WFF

[Rei21] David Reinecke. When funding fails: Planetary exploration at NASA in an era of austerity, 1967–1976. Social Studies of Science, 51(5):750–779, October 1, 2021. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/03063127211021913.

Reichmann:2023:MRI

[Rei23] Stefan Reichmann. Mobile researchers, immobile data: Managing data (producers). Social Studies of Science, 53(3):341–357, June 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
abs/10.1177/03063127231156862.

Rekdal:2014:AUL

[Rek14] Ole Bjørn Rekdal. Academic urban legends. Social Studies of Science, 44(4):638–654, August 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312714535679.

Rella:2024:CMT

[Rel24] Ludovico Rella. Close to the metal: Towards a material political economy of the epistemology of computation. Social Studies of Science, 54(1):3–29, February 1, 2024. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231185095.

Rabkin:1979:CVA

[RELHR79] Yakov M. Rabkin, Thomas O. Eisemon, Jean-Jacques Lafitte-Houssat, and Eva McLean Rathgeber. Citation visibility of Africa’s science. Social Studies of Science, 9(4):499–506, November 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277900900406.

Remington:1988:BBS

[Rem88] John A. Remington. Beyond big science in America: The binding of inquiry. Social Studies of Science, 18(1):45–72, February 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631288018001002.

Restivo:1978:PPM

[Res78] Sal P. Restivo. Parallels and paradoxes in modern physics and Eastern mysticism: I — a critical reconnaissance. Social Studies of Science, 8(2):143–181, May 1, 1978. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277800800201.

Restivo:1982:PPM

[Res82] Sal Restivo. Parallels and paradoxes in modern physics and Eastern mysticism: II — a sociological perspective on parallelism. Social Studies of Science, 12(1):37–071, February 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631282012001003.

Restivo:1996:JND

[Res96] Sal Restivo. Joseph Needham (9 December 1900–24 March 1995). Social Studies of Science, 26(1):7–8, February 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631296026001002.

Reubi:2012:HCR

[Reu12] David Reubi. The human capacity to reflect and decide: Bioethics and the reconfiguration of the research subject in the British biomedical sciences. Social Studies of Science, 42 (3):348–368, June 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712439457.

Reyes-Galindo:2014:LSP

[RG14] Luis Reyes-Galindo. Linking the subcultures of physics: Virtual empiricism and the bonding role of trust. Social Studies of Science, 44(5):736–757, October 1, 2014. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312714539058.

Reyes-Galindo:2016:AHB

[RG16] Luis Reyes-Galindo. Automating the horae: Boundary-work in the age of computers. Social Studies of Science, 46(4): 586–606, August 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312716642317.

Reyes-Galindo:2023:VVP

[RG23] Luis Reyes-Galindo. Values and vendettas: Populist science governance in Mexico. Social Studies of Science, 53(2):213–241, April 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
abs/10.1177/03063127221140020.

Roll-Hansen:1983:DSG

[RH83] Nils Roll-Hansen. The death of spontaneous generation and the birth of the gene: Two case studies of Relativism. Social Studies of Science, 13(4):481–519, November 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631283013004003.

Roll-Hansen:1985:ESP

[RH85] Nils Roll-Hansen. Empirical studies or philosophy? (reply to Barnes and Collins). Social Studies of Science, 15(1):178–180, February 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631285015001009. See [Bar85a,

Col85d].

Raina:1990:RTT

[RH90] Dhruv Raina and S. Irfan Habib. Ramchandra’s treatise through ‘the haze of the golden sunset’: an aborted pedagogy. Social Studies of Science, 20(3):455–472, August 1, 1990.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020003003.

Raina:1996:MLM

[RH96] Dhruv Raina and S. Irfan Habib. The moral legitimation of modern science: Bhadralok reflections on theories of evolution. Social Studies of Science, 26(1):9–42, February 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631296026001003.

Ribes:2019:LD

[RHSB19] David Ribes, Andrew S. Hoffman, Steven C. Slota, and Geoffrey C. Bowker. The logic of domains. Social Studies of Science, 49(3):281–309, June 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312719849709.

Rystedt:2011:RRN

[RIA^{+}11] Hans Rystedt, Jonas Ivarsson, Sara Asplund, Åse Allansdotter Johnsson, and Magnus Båth. Rediscovering radiology: New technologies and remedial action at the worksite. Social Studies of Science, 41(6):867–891, December 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312711423433.

Ribeiro:2007:LBA

[Rib07] Rodrigo Ribeiro. The language barrier as an aid to communication. Social Studies of Science, 37(4):561–584, August 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312706070738.

Richards:1988:PTE

[Ric88] Evelleen Richards. The politics of therapeutic evaluation: The vitamin c and cancer controversy. Social Studies of Science, 18 (4):653–701, November 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631288018004004.

Richards:1996:BM

[Ric96] Evelleen Richards. (un)boxing the monster. Social Studies of Science, 26(2):323–356, May 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631296026002006.

Richardson:2004:RKM

[Ric04] Alan Richardson. Robert K. Merton and philosophy of science. Social Studies of Science, 34(6):855–858, December 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312704042086.

Rier:2003:GLP

[Rie03] David A. Rier. Gender, lifecourse and publication decisions in toxic-exposure epidemiology: ‘now!’ versus ‘wait a minute!’. Social Studies of Science, 33(2):269–300, April 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127030332016.

Rieppel:2015:PDM

[Rie15] Lukas Rieppel. Prospecting for dinosaurs on the mining frontier: The value of information in America’s Gilded Age. Social Studies of Science, 45(2):161–186, April 1, 2015. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312715570650.

Rifkin:1972:CAA

[Rif72] Susan B. Rifkin. On ‘contradictions’ among academics: (a commentary on a workshop). Social Studies of Science, 2(4): 395–398, October 1, 1972. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277200200408.

Riis:2008:SBB

[Rii08] Søren Riis. The symmetry between Bruno Latour and Martin Heidegger: The technique of turning a police officer into a speed bump. Social Studies of Science, 38(2):285–301, April 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312707081379.

Rilling:1986:SGD

[Ril86] Rainer Rilling. The structure of the Gesellschaft Deutscher Chemiker (Society of German Chemists). Social Studies of Science, 16(2):235–260, May 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/0306312786016002002.

Rip:1979:SSE

[Rip79] Arie Rip. Science, society and education (Amsterdam, 14–17 August 1978). Social Studies of Science, 9(1):135–138, February 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277900900109.

Rip:1991:RUW

[Rip91] Arie Rip. Review: United we fall: Michiel Schwarz and Michael Thompson, Divided We Stand. Redefining Politics, Technology and Social Choice (Hemel Hempstead, Herts. & New York: Harvester Wheatsheaf, 1990), 176 pp., £35.00, £12.95 pbk. ISBN 0-7450-0416-4 (0-7450-0787-2 pbk). Social Studies of Science, 21 (1):179–186, February 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631291021001015.

Rip:2003:CET

[Rip03] Arie Rip. Constructing expertise: In a third wave of science studies? Social Studies of Science, 33(3):419–434, June 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127030333006.

Ritson:2016:CAR

[Rit16] Sophie Ritson. ‘Crackpots’ and ‘active researchers’: The controversy over links between arXiv and the scientific blogosphere. Social Studies of Science, 46(4):607–628, August 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716647508.

Robbins:1976:RCO

[RJ76] David Robbins and Ron Johnston. The role of cognitive and occupational differentiation in scientific controversies. Social Studies of Science, 6(3–4):349–368, September 1, 1976.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277600600304.

Ribeiro:2016:VPC

[RL16] Rodrigo Ribeiro and Francisco Pa Lima. The value of practice: a critique of interactional expertise. Social Studies of Science, 46(2):282–311, April 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312715615970.

Raffn:2021:PNB

[RL21] Jakob Raffn and Frederik Lassen. Politics of Nature: The board game. Social Studies of Science, 51(1):139–164, February 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312720983907.

Robin:1996:EAN

[Rob96] Libby Robin. Educating the activist: Natural and unnatural visions. Social Studies of Science, 26(4):857–862, November 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631296026004006.

Robins:1999:RPP

[Rob99] Rosemary Robins. Review: Public and popular representations of ‘Frankenscience’: Jon Turney, Frankenstein’s Footsteps: Science, Genetics, and Popular Culture (New Haven, CT: Yale University Press, 1998), ix + 276 pp., $30.00. ISBN 0-300-07417-4. Social Studies of Science, 29(2):295–301, April 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029002013.

Robins:2002:RRG

[Rob02] Rosemary Robins. The realness of risk: Gene technology in Germany. Social Studies of Science, 32(1):7–35, February 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312702032001002.

Robinson:2016:BPT

[Rob16] Joan H. Robinson. Bringing the pregnancy test home from the hospital. Social Studies of Science, 46(5):649–674, October 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312716664599.

Roche:1974:SVI

[Roc74] Marcel Roche. Science in Venezuela: Implications of the scientific census of 1970/71. Social Studies of Science, 4(4): 397–405, October 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277400400404.

Rocca:1981:SPO

[Roc81] Gordon L. Rocca. ‘A second party in our midst’: The history of the Soviet Scientific Forecasting Association. Social Studies of Science, 11(2):199–247, May 1, 1981. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278101100202.

Ronayne:1978:SRS

[Ron78] Jarlath Ronayne. Scientific research, science policy, and social studies of science and technology in Australia. Social Studies of Science, 8(3):361–384, August 1, 1978. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277800800311.

Rosenberg:1977:RRS

[Ros77] Charles E. Rosenberg. Rationalization and reality in the shaping of American agricultural research, 1875–1914. Social Studies of Science, 7(4):401–422, November 1, 1977. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700409.

Rosenbloom:1981:PAS

[Ros81] Joshua Rosenbloom. The politics of the American SST programme: Origin, opposition and termination. Social Studies of Science, 11(4):403–423, November 1, 1981. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278101100401.

Rosen:1993:SCM

[Ros93a] Paul Rosen. The social construction of mountain bikes: Technology and postmodernity in the cycle industry. Social Studies of Science, 23(3):479–513, August 1, 1993. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312793023003003.

Rossiter:1993:MME

[Ros93b] Margaret W. Rossiter. The Matthew Matilda Effect in science. Social Studies of Science, 23(2):325–341, May 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631293023002004.

Rosen:1996:AAA

[Ros96a] Paul Rosen. Of artifacts, analysis and alliteration: Theory and politics in constructivist technology studies. Social Studies of Science, 26(3):705–711, August 1, 1996. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631296026003007.

Rosenwein:1996:DCN

[Ros96b] Robert Rosenwein. Donald T. Campbell (20 November 1916–5 May 1996). Social Studies of Science, 26(3):717–720, August 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631296026003009.

Rosental:2010:BRS

[Ros10] Claude Rosental. Book review: Social studies of evaluation: Michele Lamont, How Professors Think: Inside the Curious World of Academic Judgment (Cambridge, MA: Harvard University Press, 2009), 330 pp., £20.95 / EUR 25.20 / $27.95. ISBN: 978-0-674-03266-8 (hbk). Social Studies of Science, 40 (3):481–484, June 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312709357645.

Ross:2011:PUU

[Ros11] Philippe Ross. Problematizing the user in user-centered production: a new media lab meets its audiences. Social Studies of Science, 41(2):251–270, April 1, 2011. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312710385851.

Rothenberg:1981:OCP

[Rot81] Marc Rothenberg. Organization and control: Professionals and amateurs in American astronomy, 1899–1918. Social Studies of Science, 11(3):305–325, August 1, 1981. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278101100302.

Roth:2005:MCW

[Rot05] Wolff-Michael Roth. Making classifications (at) work: Ordering practices in science. Social Studies of Science, 35(4): 581–621, August 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312705052102.

Rottner:2019:WBM

[Rot19] Renee Rottner. Working at the boundary: Making space for innovation in a NASA megaproject. Social Studies of Science, 49(3):403–431, June 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312719851557.

Roumbanis:2017:AJU

[Rou17] Lambros Roumbanis. Academic judgments under uncertainty: a study of collective anchoring effects in Swedish research council panel groups. Social Studies of Science, 47(1):95–116, February 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312716659789.

Rowland:2005:STS

[Row05] Nicholas J. Rowland. Science and technology studies saves Planet Earth via Latour. Social Studies of Science, 35(6):

951–954, December 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312705056049.

Rowland:2012:RGG

[Row12] Nicholas J. Rowland. Review of Govind Gopakumar, Transforming Urban Water Supply in India: The Role of Reform and Partnerships in Globalization. Social Studies of Science, 42(4):628–631, August 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312711435417.

Ribes:2015:OOC

[RP15a] David Ribes and Jessica Beth Polk. Organizing for ontological change: The kernel of an AIDS research infrastructure. Social Studies of Science, 45(2):214–241, April 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312714558136.

Rowland:2015:ISS

[RP15b] Nicholas J. Rowland and Jan-Hendrik Passoth. Infrastructure and the state in science and technology studies. Social Studies of Science, 45(1):137–145, February 1, 2015. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312714537566.

Roberts:2016:SNB

[RP16] Peder Roberts and Eric Paglia. Science as national belonging: The construction of Svalbard as a Norwegian space. Social Studies of Science, 46(6):894–911, December 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716639153.

Roderick:1974:SSS

[RS74] Gordon Roderick and Michael Stephens. Scientific studies and scientific manpower in the English civic universities 1870– 1914. Social Studies of Science, 4(1):41–63, January 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277400400103.

Richards:1989:FMM

[RS89a] Evelleen Richards and John Schuster. The feminine method as myth and accounting resource: a challenge to gender studies and social studies of science. Social Studies of Science, 19(4): 697–720, November 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631289019004011.

Richards:1989:WSC

[RS89b] Evelleen Richards and John Schuster. So what’s not a social category? or you can’t have it both ways: (reply to Keller). Social Studies of Science, 19(4):725–729, November 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019004013. See [Kel89].

Roth:2011:CBR

[RS11] Wendy D. Roth and Gerhard Sonnert. The costs and benefits of ‘red tape’: Anti-bureaucratic structure and gender inequity in a science research organization. Social Studies of Science, 41(3):385–409, June 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312710391494.

Ryghaug:2018:CEC

[RSH18] Marianne Ryghaug, Tomas Moe Skjølsvold, and Sara Heidenreich. Creating energy citizenship through material participation. Social Studies of Science, 48(2):283–303, April 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312718770286.

Rudd:1977:EAO

[Rud77] Ernest Rudd. The effect of alphabetical order of author listing on the careers of scientists. Social Studies of Science, 7 (2):268–269, May 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277700700208.

Rudwick:1982:RSS

[Rud82] Martin J. S. Rudwick. Review of sociology of the sciences yearbook, volumes IV and v. Social Studies of Science, 12(4):

627–632, November 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631282012004007.

Rudwick:1996:GTT

[Rud96] Martin Rudwick. Geological travel and theoretical innovation: The role of ‘liminal’ experience. Social Studies of Science, 26 (1):143–159, February 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631296026001007.

Ruff:1979:CAS

[Ruf79] Imre Ruff. Citation analysis of a scientific career: a case study. Social Studies of Science, 9(1):81–90, February 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277900900104.

Russell:1983:AG

[Rus83] Denise Russell. Anything goes. Social Studies of Science, 13 (3):437–464, August 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631283013003005.

Russell:1986:SCA

[Rus86] Stewart Russell. The social construction of artefacts: a response to Pinch and Bijker. Social Studies of Science, 16(2):331–346, May 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312786016002008. See [PB84, PB86].

Russell:1994:BRH

[Rus94] Stewart Russell. Book review: Heating Networks. Social Studies of Science, 24(3):587–595, August 1, 1994. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631279402400305.

Rommetveit:2022:PET

[RvD22] Kjetil Rommetveit and Niels van Dijk. Privacy engineering and the techno-regulatory imaginary. Social Studies of Science, 52 (6):853–877, December 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

Scheffler:2022:SPC

[SA22] Robin W. Scheffler and Natalie B. Aviles. State planning, cancer vaccine infrastructure, and the origins of the oncogene theory. Social Studies of Science, 52(2):174–198, April 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/03063127211070304.

Sabanovic:2014:IJR

[Sab14] Selma Sabanović. Inventing Japan’s ‘robotics culture’: The repeated assembly of science, technology, and culture in social robotics. Social Studies of Science, 44(3):342–367, June 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713509704.

Shrum:2020:WAE

[SAC^{+}20] Wesley Shrum, John Aggrey, Andre Campos, Janaina Pamplona da Costa, Jan Joseph, Pablo Kreimer, Rhiannon Kroeger, Leandro Rodriguez Medina, Paige Miller, Antony Palackal, Ana Pandal de la Peza, and Abou Traore. Who’s afraid of Ebola? Epidemic fires and locative fears in the Information Age. Social Studies of Science, 50(5):707–727, October 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720927781.

Sadowski:2024:TLI

[Sad24] Jathan Sadowski. Total life insurance: Logics of anticipatory control and actuarial governance in insurance technology. Social Studies of Science, 54(2):231–256, April 1, 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231186437.

Sagasti:1973:UST

[Sag73] Francisco R. Sagasti. Underdevelopment, science and technology: the point of view of the underdeveloped countries. Social Studies of Science, 3(1):47–59, January 1, 1973. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277300300104.

Stephens:2011:DDD

[SAG11] Neil Stephens, Paul Atkinson, and Peter Glasner. Documenting the doable and doing the documented: Bridging strategies at the UK stem cell bank. Social Studies of Science, 41(6): 791–813, December 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312711423306.

Sahota:2012:GHN

[Sah12] Puneet Chawla Sahota. Genetic histories: Native Americans’ accounts of being at risk for diabetes. Social Studies of Science, 42(6):821–842, December 1, 2012. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712454044.

Saha:2025:RDW

[Sah25] Madhumita Saha. Right data, wrong data: Statistical sampling and the making of modern agriculture in India. Social Studies of Science, 55(4):591–612, August 2025. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127241307947.

Salomon:1971:IS

[Sal71] Jean-Jacques Salomon. The Internationale of science. Social Studies of Science, 1(1):23–42, January 1, 1971. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277100100103.

Salomon:1976:MNK

[Sal76] Jean-Jacques Salomon. In memoriam: Norman Kaplan 1923– 1975. Social Studies of Science, 6(2):279–280, May 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277600600207.

Sanderson:1972:RFB

[San72] Michael Sanderson. Research and the firm in British industry, 1919–39. Social Studies of Science, 2(2):107–151, April 1, 1972. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277200200201.

Sandbach:1978:RFL

[San78] Francis Sandbach. The rise and fall of the limits to growth debate. Social Studies of Science, 8(4):495–520, November 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277800800404.

Sariola:2023:MCE

[Sar23] Salla Sariola. Mistrust: Community engagement in global health research in coastal Kenya. Social Studies of Science, 53(3):449–471, June 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127231162082.

Sato:2007:SEC

[Sat07] Yasushi Sato. Systems engineering and contractual individualism: Linking engineering processes to macro social values. Social Studies of Science, 37(6):909–934, December 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312707076601.

Sattlegger:2021:NAP

[Sat21] Lukas Sattlegger. Negotiating attachments to plastic. Social Studies of Science, 51(6):820–845, December 1, 2021. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/03063127211027950.

Saulson:2004:US

[Sau04] Peter R. Saulson. Understanding ‘social’. Social Studies of Science, 34(1):99–101, February 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704041339.

Sayes:2014:ANT

[Say14] Edwin Sayes. Actor-network theory and methodology: Just what does it mean to say that nonhumans have agency? Social Studies of Science, 44(1):134–149, February 1, 2014. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713511867.

Shapin:1977:SNC

[SB77] Steven Shapin and Barry Barnes. Science, nature and control: Interpreting Mechanics’ Institutes. Social Studies of Science, 7 (1):31–74, February 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277700700109.

Spackman:2018:SPT

[SB18] Christy Spackman and Gary A. Burlingame. Sensory politics: The tug-of-war between potability and palatability in municipal water production. Social Studies of Science, 48(3):350–371, June 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312718778358.

Sovacool:2020:ISE

[SBH^{+}20] Benjamin K. Sovacool, Noam Bergman, Debbie Hopkins, Kirsten Eh Jenkins, Sabine Hielscher, Andreas Goldthau, and Brent Brossmann. Imagining sustainable energy and mobility transitions: Valence, temporality, and radicalism in 38 visions of a low-carbon future. Social Studies of Science, 50 (4):642–679, August 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312720915283.

Sundqvist:2015:FSS

[SBHY15] Göran Sundqvist, Ingemar Bohlin, Erlend At Hermansen, and Steven Yearley. Formalization and separation: a systematic basis for interpreting approaches to summarizing science for climate policy. Social Studies of Science, 45(3):416–440, June 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312715583737.

Smith:2000:SGH

[SBS^{+}00] Laurence D. Smith, Lisa A. Best, D. Alan Stubbs, John Johnston, and Andrea Bastiani Archibald. Scientific graphs and the hierarchy of the sciences: A Latourian survey of inscription practices. Social Studies of Science, 30(1):73–94, February 1, 2000.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030001003.

Stark:2018:IFS

[SC18] Laura Stark and Nancy D. Campbell. The ineffable: a framework for the study of methods through the case of midcentury mind-brain sciences. Social Studies of Science, 48(6): 789–820, December 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312718816807.

Shrum:2001:TCP

[SCG01] Wesley Shrum, Ivan Chompalov, and Joel Genuth. Trust, conflict and performance in scientific collaborations. Social Studies of Science, 31(5):681–730, October 1, 2001. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031005002.

Schultz:1972:KAW

[Sch72a] Theodore W. Schultz. Knowledge, agriculture and welfare. Social Studies of Science, 2(4):361–368, October 1, 1972. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277200200404.

Schurr:1972:ERA

[Sch72b] George M. Schurr. Essay review: Array and disarray on the medico–ethical front. Social Studies of Science, 2(2):191–202, April 1, 1972. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277200200204.

Schaffer:1984:MC

[Sch84] Simon Schaffer. Making certain. Social Studies of Science, 14 (1):137–152, February 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631284014001010.

Schaffer:1986:SDE

[Sch86] Simon Schaffer. Scientific discoveries and the end of natural philosophy. Social Studies of Science, 16(3):387–420, August 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631286016003001.

Schwartzman:1987:PS

[Sch87] Simon Schwartzman. Peripheral science. Social Studies of Science, 17(3):569–573, August 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631287017003009.

Schaffer:1988:SP

[Sch88a] Simon Schaffer. Science and Puritanism. Social Studies of Science, 18(3):551–556, August 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631288018003008.

Schumacher:1988:OFR

[Sch88b] John A. Schumacher. The observer’s frame of reference in natural and social science: a response to Latour. Social Studies of Science, 18(3):523–531, August 1, 1988. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631288018003006.

Schulze:1990:RSI

[Sch90] Annedore Schulze. On the rise of scientific innovations and their acceptance in research groups: a socio-psychological study. Social Studies of Science, 20(1):35–64, February 1, 1990. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020001002.

Schulman:1994:BRM

[Sch94] Paul R. Schulman. Book review: The Management of Scale: Big Organizations, Big Decisions, Big Mistakes. Social Studies of Science, 24(1):188–191, February 1, 1994. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631279402400115.

Schimank:1995:TRS

[Sch95] Uwe Schimank. Transformation of research systems in Central and Eastern Europe: a coincidence of opportunities and trouble. Social Studies of Science, 25(4):633–653, November 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631295025004003.

Schaffer:2002:RSP

[Sch02] Simon Schaffer. Roy Sydney Porter. Social Studies of Science, 32(3):477–486, June 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312702032003006.

Schmid:2006:CTT

[Sch06] Sonja D. Schmid. Celebrating tomorrow today: The peaceful atom on display in the Soviet Union. Social Studies of Science, 36(3):331–365, June 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706055534.

Schrader:2010:RPP

[Sch10] Astrid Schrader. Responding to Pfiesteria piscicida (the fish killer): Phantomatic ontologies, indeterminacy, and responsibility in toxic microbiology. Social Studies of Science, 40 (2):275–306, April 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312709344902.

Schiolin:2012:FVC

[Sch12] Kasper Schiølin. Follow the verbs! A contribution to the study of the Heidegger–Latour connection. Social Studies of Science, 42(5):775–786, October 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712448441.

Schleifer:2013:CCT

[Sch13] David Schleifer. Categories count: Trans fat labeling as a technique of corporate governance. Social Studies of Science, 43 (1):54–77, February 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712461573.

Schrader:2015:AIT

[Sch15] Astrid Schrader. Abyssal intimacies and temporalities of care: How (not) to care about deformed leaf bugs in the aftermath of Chernobyl. Social Studies of Science, 45(5):665–690, October 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312715603249.

Schinkel:2016:MCC

[Sch16] Willem Schinkel. Making climates comparable: Comparison in paleoclimatology. Social Studies of Science, 46(3):374–395, June 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312716633537.

Schiolin:2020:RDF

[Sch20] Kasper Schiølin. Revolutionary dreams: Future essentialism and the sociotechnical imaginary of the fourth industrial revolution in Denmark. Social Studies of Science, 50(4):542–566, August 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312719867768.

Scott:1991:LCL

[Sco91] Pam Scott. Levers and counterweights: a laboratory that failed to raise the world. Social Studies of Science, 21(1):7–35, February 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631291021001003.

Scott:1992:EST

[Sco92] James K. Scott. Exploring socio-technical analysis: Monsieur Latour is not joking! Social Studies of Science, 22 (1):59–80, February 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312792022001003.

Shim:2014:HHS

[SDL^{+}14] Janet K. Shim, Katherine Weatherford Darling, Martine D. Lappe, L. Katherine Thomson, Sandra Soo-Jin Lee, Robert A. Hiatt, and Sara L. Ackerman. Homogeneity and heterogeneity as situational properties: Producing — and moving beyond?

— race in post-genomic science. Social Studies of Science, 44 (4):579–599, August 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312714531522.

Samimian-Darash:2025:UIT

[SDSR25] Limor Samimian-Darash, Amit Sheniak, and Nir Rotem. Unboxing the imaginary: Typology of future imagination techniques in high-tech development. Social Studies of Science, 55(4):542–564, August 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127241300625.

Secord:1985:NHD

[Sec85] James A. Secord. Natural history in depth. Social Studies of Science, 15(1):181–200, February 1, 1985. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631285015001010.

Segal:2018:SDR

[Seg18] Judy Z. Segal. Sex, drugs, and rhetoric: The case of flibanserin for ‘female sexual dysfunction’. Social Studies of Science, 48 (4):459–482, August 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312718778802.

Sleeboom-Faulkner:2019:RBC

[SF19] Margaret Sleeboom-Faulkner. Regulatory brokerage: Competitive advantage and regulation in the field of regenerative medicine. Social Studies of Science, 49(3):355–380, June 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312719850628.

Sleeboom-Faulkner:2012:GSC

[SFH12] Margaret Sleeboom-Faulkner and Seyoung Hwang. Governance of stem cell research: Public participation and decisionmaking in China, Japan, South Korea and Taiwan. Social Studies of Science, 42(5):684–708, October 1, 2012. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712450939.

Sleeboom-Faulkner:2011:ESC

[SFP11] Margaret Sleeboom-Faulkner and Prasanna Kumar Patra. Experimental stem cell therapy: Biohierarchies and bionetworking in Japan and India. Social Studies of Science, 41(5):

645–666, October 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312711409792.

Schroeder-Gudehus:1973:CTL

[SG73] Brigitte Schroeder-Gudehus. Challenge to transnational loyalties: International scientific organizations after the First World War. Social Studies of Science, 3(2):93–118, April 1, 1973. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277300300201.

Schroeder-Gudehus:1974:ICS

[SG74a] Brigitte Schroeder-Gudehus. International Commission for Science Policy Studies: New Delhi meeting. Social Studies of Science, 4(1):105–107, January 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631277400400108.

Small:1974:SSL

[SG74b] Henry Small and Belver C. Griffith. The structure of scientific literatures I: Identifying and graphing specialties. Social Studies of Science, 4(1):17–40, January 1, 1974. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277400400102.

Star:1989:IET

[SG89] Susan Leigh Star and James R. Griesemer. Institutional ecology, ‘translations’ and boundary objects: Amateurs and professionals in Berkeley’s Museum of Vertebrate Zoology, 1907– 39. Social Studies of Science, 19(3):387–420, August 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019003001.

Sweet:2021:BE

[SG21] Paige L. Sweet and Danielle Giffort. The bad expert. Social Studies of Science, 51(3):313–338, June 1, 2021. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720970282.

Stokes:1989:CSS

[SH89] T. D. Stokes and J. A. Hartley. Coauthorship, social structure and influence within specialties. Social Studies of Science, 19 (1):101–125, February 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631289019001003.

Sims:2012:RCR

[SH12] Benjamin Sims and Christopher R. Henke. Repairing credibility: Repositioning nuclear weapons knowledge after the Cold War. Social Studies of Science, 42(3):324–347, June 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712437778.

Sovacool:2017:OTT

[SH17] Benjamin K. Sovacool and David J. Hess. Ordering theories: Typologies and conceptual frameworks for sociotechnical change. Social Studies of Science, 47(5):703–750, October 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312717709363.

Swallow:2019:FAA

[SH19] Julia Swallow and Alexandra Hillman. Fear and anxiety: Affects, emotions and care practices in the memory clinic. Social Studies of Science, 49(2):227–244, April 1, 2019. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312718820965.

Shapin:1972:PPS

[Sha72] Steven A. Shapin. The Pottery Philosophical Society, 1819– 1835: an examination of the cultural uses of provincial science. Social Studies of Science, 2(4):311–336, October 1, 1972. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277200200402.

Shapin:1980:CSH

[Sha80] Steven Shapin. A course in the social history of science. Social Studies of Science, 10(2):231–258, May 1, 1980. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278001000206.

Shapin:1984:PCR

[Sha84] Steven Shapin. Pump and circumstance: Robert Boyle’s literary technology. Social Studies of Science, 14(4):481–520, November 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631284014004001.

Shapin:1988:FSA

[Sha88] Steven Shapin. Following scientists around. Social Studies of Science, 18(3):533–550, August 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631288018003007.

Shapin:1997:ST

[Sha97] Steven Shapin. Signs of the times. Social Studies of Science, 27(2):335–349, April 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631297027002006.

Shapin:2001:PEH

[Sha01] Steven Shapin. Proverbial economies: How an understanding of some linguistic and social features of common sense can throw light on more prestigious bodies of knowledge, science for example. Social Studies of Science, 31(5):731–769, October 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631201031005003.

Sharrock:2004:NCA

[Sha04] Wes Sharrock. No case to answer: a response to Martin Kusch’s ‘Rule-Scepticism and the Sociology of Scientific Knowledge’. Social Studies of Science, 34(4):603–614, August 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/0306312704046366. See [Kus04b] and reply [Kus04a].

Shapin:2012:SS

[Sha12] Steven Shapin. The sciences of subjectivity. Social Studies of Science, 42(2):170–184, April 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312711435375.

Shapin:2016:TSM

[Sha16] Steven Shapin. A taste of science: Making the subjective objective in the California wine world. Social Studies of Science, 46(3):436–460, June 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312716651346.

Shafiee:2020:CBA

[Sha20] Katayoun Shafiee. Cost-benefit analysis at the floodgates: Governing democratic futures through the reassembly of Iran’s waterways. Social Studies of Science, 50(1):94–120, February 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312719891564.

Shermer:2002:VSS

[She02] Michael B. Shermer. The view of science: Stephen Jay Gould as historian of science and scientific historian, popular scientist and scientific popularizer. Social Studies of Science, 32 (4):489–524, August 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312702032004001.

Shepherd:2006:VSP

[She06] Chris J. Shepherd. From In Vitro to In Situ: On the precarious extension of agricultural science in the indigenous ‘Third World’. Social Studies of Science, 36(3):399–426, June 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706055272.

Sheldrake:2020:ERS

[She20] Merlin Sheldrake. The ‘enigma’ of Richard Schultes, Amazonian hallucinogenic plants, and the limits of ethnobotany. Social Studies of Science, 50(3):345–376, June 1, 2020. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720920362.

Shinn:1999:PFE

[Shi99] Terry Shinn. Pillars of French engineering. Social Studies of Science, 29(1):135–144, February 1, 1999. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029001006.

Shinn:2002:THN

[Shi02] Terry Shinn. The triple helix and new production of knowledge: Prepackaged thinking on science and technology. Social Studies of Science, 32(4):599–614, August 1, 2002. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312702032004004.

Shipley:2003:WCM

[Shi03] Brian C. Shipley. The world that changed the map. Social Studies of Science, 33(5):821–825, October 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312703335008.

Shim:2005:CRA

[Shi05] Janet K. Shim. Constructing ‘race’ across the science-lay divide: Racial formation in the epidemiology and experience of cardiovascular disease. Social Studies of Science, 35(3):405–436, June 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312705052105.

Shostak:2005:ETC

[Sho05] Sara Shostak. The emergence of toxicogenomics: a case study of molecularization. Social Studies of Science, 35(3):367–403, June 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312705049882.

Shrum:1984:SST

[Shr84] Wesley Shrum. Scientific specialties and technical systems. Social Studies of Science, 14(1):63–90, February 1, 1984. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014001005.

Shrum:2000:SSD

[Shr00] Wesley Shrum. Science and story in development: The emergence of non-governmental organizations in agricultural research. Social Studies of Science, 30(1):95–124, February 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631200030001004.

Shrum:2003:MDO

[Shr03] Wes Shrum. In memoriam: David Owen Edge. Social Studies of Science, 33(2):193–194, April 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127030332012.

Shrum:2004:SCP

[Shr04] Wesley Shrum. Science and the committee process: The presentation of the scientific self. Social Studies of Science, 34 (3):427–432, June 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312704043692.

Shrum:2005:RIH

[Shr05] Wesley Shrum. Reagency of the Internet, or, how i became a guest for science. Social Studies of Science, 35(5):723–754, October 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312705052106.

Shrum:2007:HSW

[Shr07] Wesley Shrum. Hurricane stories, from within. Social Studies of Science, 37(1):97–102, February 1, 2007. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706069430.

Shrum:2014:WCF

[Shr14] Wesley Shrum. What caused the flood? Controversy and closure in the Hurricane Katrina disaster. Social Studies of Science, 44 (1):3–33, February 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312713498654.

Silverman:1974:NLP

[Sil74] S. M. Silverman. Notes and letters: Parental loss and scientists. Social Studies of Science, 4(3):259–264, July 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277400400303.

Siles:2011:OFW

[Sil11] Ignacio Siles. From online filter to web format: Articulating materiality and meaning in the early history of blogs. Social Studies of Science, 41(5):737–758, October 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312711420190.

Simonton:1978:IDS

[Sim78] Dean Keith Simonton. Independent discovery in science and technology: a closer look at the Poisson distribution. Social Studies of Science, 8(4):521–532, November 1, 1978. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277800800405.

Simonton:1979:ESS

[Sim79] Dean Keith Simonton. To the Editors, social studies of science Davis, 13 February 1979 dear sirs. Social Studies of Science, 9 (4):509–510, November 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277900900409.

Simonton:1984:MEA

[Sim84] Dean Keith Simonton. Is the marginality effect all that marginal? Social Studies of Science, 14(4):621–622, November 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631284014004005. See reply [GH84].

Simon:1991:CSC

[Sim91] Herbert A. Simon. Comments on the Symposium on ‘Computer Discovery and the Sociology of Scientific Knowledge’. Social Studies of Science, 21(1):143–148, February 1, 1991.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631291021001008.

Simeonova:1995:RDS

[Sim95] Kostadinka Simeonova. Radical and defensive strategies in the democratization of the Bulgarian Academy of Sciences. Social Studies of Science, 25(4):755–775, November 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631295025004008.

Simon:1999:USM

[Sim99a] Bart Simon. Undead science: Making sense of cold fusion after the (arti)fact. Social Studies of Science, 29(1):61–85, February 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631299029001003.

Sims:1999:CPT

[Sim99b] Benjamin Sims. Concrete practices: Testing in an earthquakeengineering laboratory. Social Studies of Science, 29(4):483–518, August 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631299029004002.

Sims:2005:SSM

[Sim05] Benjamin Sims. Safe science: Material and social order in laboratory work. Social Studies of Science, 35(3):333–366, June 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312705052362.

Sims:2007:DAH

[Sim07a] Benjamin Sims. ‘The day after the hurricane’: Infrastructure, order, and the New Orleans Police Department’s response to Hurricane Katrina. Social Studies of Science, 37(1): 111–118, February 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706069432.

Sims:2007:TFA

[Sim07b] Benjamin Sims. Things fall apart: Disaster, infrastructure, and risk. Social Studies of Science, 37(1):93–95, February 1, 2007.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706069429.

Simakova:2010:RTP

[Sim10] Elena Simakova. RFID ‘theatre of the proof’: Product launch and technology demonstration as corporate practices. Social Studies of Science, 40(4):549–576, August 1, 2010. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312710365587.

Simmet:2025:MCP

[Sim25] Hilton R. Simmet. Making citizens, procedures, and outcomes: Theorizing politics in a co-productionist idiom. Social Studies of Science, 55(2):153–177, April 2025. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127241269804.

Sinding:1996:LGC

[Sin96a] Christiane Sinding. Literary genres and the construction of knowledge in biology: Semantic shifts and scientific change. Social Studies of Science, 26(1):43–70, February 1, 1996. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631296026001004.

Singleton:1996:FSS

[Sin96b] Vicky Singleton. Feminism, sociology of scientific knowledge and postmodernism: Politics, theory and me. Social Studies of Science, 26(2):445–468, May 1, 1996. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631296026002010.

Singleton:1998:PIS

[Sin98] Vicky Singleton. The politic(ian)s of SSK: A reply to Radder. Social Studies of Science, 28(2):332–338, April 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631298028002006. See [Rad98a, Rad98b].

Sismondo:1993:RKC

[Sis93a] Sergio Sismondo. Response to Knorr-Cetina. Social Studies of Science, 23(3):563–569, August 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/0306312793023003006.

See [Sis93b, KC93, Sis95].

Sismondo:1993:SSC

[Sis93b] Sergio Sismondo. Some social constructions. Social Studies of Science, 23(3):515–553, August 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/0306312793023003004.

See [Tay95, KC93, Sis93a, Sis95].

Sismondo:1995:RT

[Sis95] Sergio Sismondo. Reply to Taylor. Social Studies of Science, 25(2):359–362, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631295025002016. See

[Sis93b, Tay95].

Sismondo:2004:PM

[Sis04] Sergio Sismondo. Pharmaceutical maneuvers. Social Studies of Science, 34(2):149–159, April 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/0306312704042575.

Sismondo:2009:GMP

[Sis09a] Sergio Sismondo. Ghosts in the machine: Publication planning in the medical sciences. Social Studies of Science, 39(2):171–198, April 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312708101047. See comments [McH09].

Sismondo:2009:GMR

[Sis09b] Sergio Sismondo. Ghosts in the machine: Reply to McHenry (2009). Social Studies of Science, 39(6):949–952, December 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312709345359. See [McH09].

Sismondo:2010:REI

[Sis10] Sergio Sismondo. Review essay: Inclusion by numbers: New biomedical subjects and biopolitical citizens: Jeremy A. Greene, Prescribing by Numbers: Drugs and the Definition of Disease (Johns Hopkins University Press, 2007), 318 pp., £14.00 / EUR 23.99 / $25.00. ISBN 0-801-8910-0 (pbk); ISBN 0-8018-8477-2 (hbk). Steven Epstein, Inclusion: The Politics of Difference in Medical Research (University of Chicago Press, 2007), 413 pp., £13.00 / EUR 18.99 / $19.00. ISBN 0-226-21310-2 (pbk); ISBN 0-226-21309-9 (hbk). Social Studies of Science, 40 (4):633–640, August 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312710363512.

Sismondo:2012:CPI

[Sis12a] Sergio Sismondo. 2012 changes, part II. Social Studies of Science, 42(5):635–637, October 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312712456946.

Sismondo:2012:FYS

[Sis12b] Sergio Sismondo. Fifty years of the structure of scientific revolutions, twenty-five of science in action. Social Studies of Science, 42(3):415–419, June 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712439255.

Sismondo:2013:E

[Sis13] Sergio Sismondo. Editorial. Social Studies of Science, 43 (2):161–162, April 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312713482058.

Sismondo:2015:OTT

[Sis15] Sergio Sismondo. Ontological turns, turnoffs and roundabouts. Social Studies of Science, 45(3):441–448, June 1, 2015. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312715574681.

Sismondo:2016:NVS

[Sis16a] Sergio Sismondo. New venues in STS. Social Studies of Science, 46(1):3–6, February 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312715625854.

Sismondo:2016:SAI

[Sis16b] Sergio Sismondo. Sorting on arXiv: Introduction to an ad hoc section. Social Studies of Science, 46(4):583–585, August 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312716661429.

Sismondo:2017:CWN

[Sis17a] Sergio Sismondo. Casting a wider net: a reply to Collins, Evans and Weinel. Social Studies of Science, 47(4):587–592, August 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312717721410. See [CEW17].

Sismondo:2017:PT

[Sis17b] Sergio Sismondo. Post-truth? Social Studies of Science, 47 (1):3–6, February 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312717692076.

Sismondo:2017:T

[Sis17c] Sergio Sismondo. Transmissions. Social Studies of Science, 47(5):603, October 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312717732190.

Sismondo:2019:ALC

[Sis19] Sergio Sismondo. Academic lives and cultures. Social Studies of Science, 49(6):813–816, December 1, 2019. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312719886138.

Sismondo:2020:C

[Sis20a] Sergio Sismondo. COVID-19. Social Studies of Science, 50 (2):173–174, April 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312720918403.

Sismondo:2020:SIA

[Sis20b] Sergio Sismondo. Sociotechnical imaginaries: an accidental themed issue. Social Studies of Science, 50(4):505–507, August 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312720944753.

Sismondo:2022:MBL

[Sis22a] Sergio Sismondo. In memoriam: Bruno Latour (22 June 1947–9 October 2022). Social Studies of Science, 52(6):954–956, December 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

Sismondo:2022:NEC

[Sis22b] Sergio Sismondo. A new Editor-in-Chief. Social Studies of Science, 52(3):327–329, June 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
03063127221095117.

Sismondo:2025:VDP

[Sis25] Sergio Sismondo. Virtual diversity: a proposal and discussion. Social Studies of Science, 55(2):260–261, April 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127251334495.

Shenhav:1991:CII

[SK91] Yehouda A. Shenhav and David H. Kamens. The ‘costs’ of institutional isomorphism: Science in non–Western countries. Social Studies of Science, 21(3):527–545, August 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631291021003005.

Svendsen:2008:USE

[SK08] Mette N. Svendsen and Lene Koch. Unpacking the ‘spare embryo’: Facilitating stem cell research in a moral landscape. Social Studies of Science, 38(1):93–110, February 1, 2008.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312707082502.

Suryanarayanan:2013:CEC

[SK13] Sainath Suryanarayanan and Daniel Lee Kleinman. Be(e)coming experts: The controversy over insecticides in the honey bee colony collapse disorder. Social Studies of Science, 43(2):215– 240, April 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/full/10.1177/0306312712466186.

Skinner:2006:RFB

[Ski06] David Skinner. Racialized futures: Biologism and the changing politics of identity. Social Studies of Science, 36(3):459–488, June 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312706054859.

Skolnikoff:1973:ICS

[Sko73] Eugene B. Skolnikoff. International Commission for Science Policy Studies. Social Studies of Science, 3(1):89–90, January 1, 1973. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277300300108.

Skrydstrup:2016:SSC

[Skr16] Martin Skrydstrup. Of spheres and squares: Can Sloterdijk help us rethink the architecture of climate science? Social Studies of Science, 46(6):854–876, December 1, 2016. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716647214.

Stephan:1991:ISP

[SL91] Paula E. Stephan and Sharon G. Levin. Inequality in scientific performance: Adjustment for attribution and journal impact. Social Studies of Science, 21(2):351–368, May 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631291021002007.

Singh:2025:PES

[SL25] Ranjit Singh and Michael Lynch. Proverbial economies of STS. Social Studies of Science, 55(3):327–349, June 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127241294038.

Slayton:2007:DCB

[Sla07] Rebecca Slayton. Discursive choices: Boycotting Star Wars between science and politics. Social Studies of Science, 37 (1):27–66, February 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706063787.

Slayton:2012:IC

[Sla12] Rebecca Slayton. An instrumental concept. Social Studies of Science, 42(5):787–792, October 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312712450773.

Slezak:1989:CCC

[Sle89a] Peter Slezak. Computers, contents and causes: Replies to my respondents. Social Studies of Science, 19(4):671–695, November 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631289019004010.

Slezak:1989:SDC

[Sle89b] Peter Slezak. Scientific discovery by computer as empirical refutation of the strong programme. Social Studies of Science, 19 (4):563–600, November 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631289019004001.

Slezak:1991:HSS

[Sle91] Peter Slezak. How strong is the ‘strong programme’? Social Studies of Science, 21(1):154–156, February 1, 1991. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631291021001013.

Slezak:1992:AE

[Sle92] Peter Slezak. Artificial experts. Social Studies of Science, 22 (1):175–201, February 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312792022001010. See

[Col92a].

Schubert:1993:CCS

[SM93a] András Schubert and Hajnalka Maczelka. Cognitive changes in scientometrics during the 1980s, as reflected by the reference patterns of its core journal. Social Studies of Science, 23 (3):571–581, August 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312793023003007.

Singleton:1993:ANA

[SM93b] Vicky Singleton and Mike Michael. Actor-networks and ambivalence: General practitioners in the UK cervical screening programme. Social Studies of Science, 23(2):227–264, May 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631293023002001.

Santesmases:1997:SOS

[SM97] María Jesús Santesmases and Emilio Muñoz. Scientific organizations in Spain (1950–1970): Social isolation and international legitimation of biochemists and molecular biologists on the periphery. Social Studies of Science, 27(2):187–219, April 1, 1997. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631297027002001.

Slayton:2025:CUP

[SM25] Rebecca Slayton and Lilly Muller. Coordinating uncertainty in the political economy of cyber threat intelligence. Social Studies of Science, 55(4):481–511, August 2025. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127241299132.

Small:1977:CCM

[Sma77] Henry G. Small. A co-citation model of a scientific specialty: a longitudinal study of collagen research. Social Studies of Science, 7(2):139–166, May 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277700700202.

Small:1978:CDC

[Sma78] Henry G. Small. Cited documents as concept symbols. Social Studies of Science, 8(3):327–340, August 1, 1978. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277800800305.

Smallman:2020:NDS

[Sma20] Melanie Smallman. ‘Nothing to do with the science’: How an elite sociotechnical imaginary cements policy resistance to public perspectives on science and technology through the machinery of government. Social Studies of Science, 50(4):589–608, August 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312719879768.

Smith:1983:INT

[Smi83] F. Graham Smith. The Isaac Newton Telescope. Social Studies of Science, 13(1):161–162, February 1, 1983. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631283013001009.

Smith:1985:ECA

[Smi85] Roger Smith. Expertise and causal attribution in deciding between crime and mental disorder. Social Studies of Science, 15 (1):67–98, February 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631285015001003.

Smith:2009:TUF

[Smi09] Wally Smith. Theatre of use: a frame analysis of information technology demonstrations. Social Studies of Science, 39 (3):449–480, June 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312708101978.

Smith:2014:ASD

[Smi14] Frank L. Smith III. Advancing science diplomacy: Indonesia and the US Naval Medical Research Unit. Social Studies of Science, 44(6):825–847, December 1, 2014. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312714535864.

Smith:2015:TSM

[Smi15] Wally Smith. Technologies of stage magic: Simulation and dissimulation. Social Studies of Science, 45(3):319–343, June 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312715577461.

Smith:2017:MMD

[Smi17] Lindsay A. Smith. The missing, the martyred and the disappeared: Global networks, technical intensification and the end of human rights genetics. Social Studies of Science, 47(3):398–416, June 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312716678489.

Schwartz-Marin:2015:EVI

[SMW15] Ernesto Schwartz-Marín and Peter Wade. Explaining the visible and the invisible: Public knowledge of genetics, ancestry, physical appearance and race in Colombia. Social Studies of Science, 45(6):886–906, December 1, 2015. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312715621182.

Schwartz-Marin:2015:CFG

[SMWCSC15] Ernesto Schwartz-Marín, Peter Wade, Arely Cruz-Santiago, and Roosbelinda Cárdenas. Colombian forensic genetics as a form of public science: The role of race, nation and common sense in the stabilization of DNA populations. Social Studies of Science, 45 (6):862–885, December 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312715574158.

Smyth:2001:CUS

[Smy01] Mary M. Smyth. Certainty and uncertainty sciences: Marking the boundaries of psychology in introductory textbooks. Social Studies of Science, 31(3):389–416, June 1, 2001. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031003003.

Smith:1989:GSP

[SN89] David Smith and Malcolm Nicolson. The ‘Glasgow School’ of Paton, Findlay and Cathcart: Conservative thought in chemical physiology, nutrition and public health. Social Studies of Science, 19(2):195–238, May 1, 1989. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019002001.

Sneddon:2012:SDC

[Sne12] Chris Sneddon. The ‘sinew of development’: Cold War geopolitics, technical expertise, and water resource development in Southeast Asia, 1954–1975. Social Studies of Science, 42(4): 564–590, August 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712445835.

Sapir:2017:ALM

[SO17] Adi Sapir and Amalya L. Oliver. From academic laboratory to the market: Disclosed and undisclosed narratives of commercialization. Social Studies of Science, 47(1):33–52, February 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716667647.

Sokolov:2023:AIW

[Sok23] Mikhail Sokolov. The art of ignoring others’ work among academics: a guessing game model of scholarly information search. Social Studies of Science, 53(2):300–312, April 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127221119808.

Solomon:1985:LES

[Sol85] Joan Solomon. Learning and evaluation: a study of school children’s views on the social uses of energy. Social Studies of Science, 15(2):343–371, May 1, 1985. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631285015002005.

Solomon:1992:SLA

[Sol92] Susan Gross Solomon. The Soviet legalization of abortion in German medical discourse: a study of the use of selective perceptions in cross-cultural scientific relations. Social Studies of Science, 22(3):455–485, August 1, 1992. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312792022003002.

Solovey:2001:PCE

[Sol01a] Mark Solovey. Project Camelot and the 1960s epistemological revolution: Rethinking the politics–patronage–social science nexus. Social Studies of Science, 31(2):171–206, April 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312701031002003.

Solovey:2001:SSD

[Sol01b] Mark Solovey. Science and the state during the Cold War: Blurred boundaries and a contested legacy. Social Studies of Science, 31(2):165–170, April 1, 2001. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312701031002002.

Sommer:2006:MMW

[Som06a] Marianne Sommer. Mirror, mirror on the wall: Neanderthal as image and ‘distortion’ in Early 20th-Century French science and press. Social Studies of Science, 36(2):207–240, April 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312706054527.

Sommerlund:2006:CMM

[Som06b] Julie Sommerlund. Classifying microorganisms: The multiplicity of classifications and research practices in molecular microbial ecology. Social Studies of Science, 36(6):909–928, December 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312706063029.

Sonnert:1995:WMG

[Son95] Gerhard Sonnert. What makes a good scientist?: Determinants of peer evaluation among biologists. Social Studies of Science, 25(1):35–55, February 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631295025001003.

Sonnert:2009:PWI

[Son09] Gerhard Sonnert. Parents who influence their children to become scientists: Effects of gender and parental education. Social Studies of Science, 39(6):927–941, December 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312709335843.

Sorensen:1992:TFT

[Sør92] Knut H. Sørensen. Towards a feminized technology? Gendered values in the construction of technology. Social Studies of Science, 22(1):5–31, February 1, 1992. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312792022001001.

Sorensen:1993:AEK

[Sør93] Knut H. Sørensen. An anatomy of engineering knowledge. Social Studies of Science, 23(2):373–377, May 1, 1993. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631293023002007.

Sormani:2023:IAE

[Sor23] Philippe Sormani. Interfacing AlphaGo: Embodied play, object agency, and algorithmic drama. Social Studies of Science, 53 (5):686–711, October 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127231191284.

Sovacool:2010:IOC

[Sov10] Benjamin K. Sovacool. The importance of open and closed styles of energy research. Social Studies of Science, 40(6): 903–930, December 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312710373842.

Spencer:2024:PIS

[SP24] Matt Spencer and Daniele Pizio. The de-perimeterisation of information security: The Jericho Forum, zero trust, and narrativity. Social Studies of Science, 54(5):655–677, October 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231221107.

Spackman:2020:SSM

[Spa20] Christy Spackman. In smell’s shadow: Materials and politics at the edge of perception. Social Studies of Science, 50(3):418–439, June 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312720918946.

Sonnberger:2024:WET

[SPBG24] Marco Sonnberger, Maria Pfeiffer, Alena Bleicher, and Matthias Gross. Wake effects and temperature plumes: Coping with non-knowledge in the expansion of wind and geothermal energy. Social Studies of Science, 54(6):859–882, December 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127241246551.

Spinardi:1987:TTC

[Spi87] Graham Spinardi. Trident: Tracing the course of nuclear weapons technology. Social Studies of Science, 17(2):371–381, May 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631287017002011.

Spinardi:1997:ABN

[Spi97] Graham Spinardi. Aldermaston and British nuclear weapons development: Testing the ‘Zuckerman Thesis’. Social Studies of Science, 27(4):547–582, August 1, 1997. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631297027004001.

Spiegel-Rosing:1972:JAI

[SR72] I. S. Spiegel-Rösing. Journal authors as an indicator of scientific manpower; a methodological study using data for the two Germanies and Europe. Social Studies of Science, 2(4):337–359, October 1, 1972. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277200200403.

Spiegel-Rosing:1973:SPS

[SR73] I. S. Spiegel-Rosing. Science policy studies in a political context: the conceptual and institutional development of science policy studies in the German Democratic Republic. Social Studies of Science, 3(4):393–413, October 1, 1973. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277300300404.

Spiegel-Rosing:1977:SSB

[SR77] Ina Spiegel-Rosing. Science studies: Bibliometric and content analysis. Social Studies of Science, 7(1):97–113, February 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277700700111.

Spiegel-Rosing:1976:ICS

[SRM76] Ina Spiegel-Rösing and Roy MacLeod. The international council for science policy studies. Social Studies of Science, 6(1): 133–136, February 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277600600110.

Stout:1991:GBR

[SS91] Dale Stout and Sue Stuart. E. G. Boring’s review of Brigham’s A Study of American Intelligence: a case-study in the politics of reviews. Social Studies of Science, 21(1):133–142, February 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631291021001007.

Soderqvist:1994:PSM

[SS94] Thomas Söderqvist and Arthur M. Silverstein. Participation in scientific meetings: a new prosopographical approach to the disciplinary history of science — the case of immunology, 1951– 72. Social Studies of Science, 24(3):513–548, August 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631279402400303.

Shapin:1999:BHR

[SS99a] Steven Shapin and Simon Schaffer. On bad history: Reply to Pinnick. Social Studies of Science, 29(2):257–259, April 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029002010. See [Pin99c].

Shapin:1999:RP

[SS99b] Steven Shapin and Simon Schaffer. Response to Pinnick. Social Studies of Science, 29(2):249–253, April 1, 1999. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029002008. See [Pin99c].

Salter:2017:CNK

[SS17a] Brian Salter and Charlotte Salter. Controlling new knowledge: Genomic science, governance and the politics of bioinformatics. Social Studies of Science, 47(2):263–287, April 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716681210.

Strang:2017:JFM

[SS17b] David Strang and Kyle Siler. From ‘just the facts’ to ‘more theory and methods, please’: The evolution of the research article in Administrative Science Quarterly, 1956–2008. Social Studies of Science, 47(4):528–555, August 1, 2017. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312717694512.

Scott-Smith:2018:STP

[SS18] Tom Scott-Smith. Sticky technologies: Plumpy’nut(R), emergency feeding and the viscosity of humanitarian design. Social Studies of Science, 48(1):3–24, February 1, 2018. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312717747418.

Smith:2024:GBP

[SSB24] Robert D. J. Smith, Stefan Schäfer, and Michael J. Bernstein. Governing beyond the project: Refocusing innovation governance in emerging science and technology funding. Social Studies of Science, 54(3):377–404, June 1, 2024. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231205043.

Schuster:1997:BTG

[ST97] John A. Schuster and Alan B. H. Taylor. Blind trust: The gentlemanly origins of experimental science. Social Studies of Science, 27(3):503–536, June 1, 1997. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631297027003006.

Smith:2016:ELE

[ST16] Jessica M. Smith and Abraham Sd Tidwell. The everyday lives of energy transitions: Contested sociotechnical imaginaries in the American West. Social Studies of Science, 46(3):327–350, June 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312716644534.

Street:2024:EDM

[ST24] Alice Street and Emma Michelle Taylor. Equivocal diagnostics: Making a ‘good’ point-of-care test for elimination in global health. Social Studies of Science, 54(6):836–858, December 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127241246727.

Star:1983:SSW

[Sta83] Susan Leigh Star. Simplification in scientific work: an example from neuroscience research. Social Studies of Science, 13 (2):205–228, May 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631283013002002.

Star:1985:SWU

[Sta85] Susan Leigh Star. Scientific work and uncertainty. Social Studies of Science, 15(3):391–427, August 1, 1985. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631285015003001.

Staley:2002:II

[Sta02] Richard Staley. Interstitial instruments. Social Studies of Science, 32(3):469–476, June 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/0306312702032003005.

Stark:2018:APS

[Sta18] Luke Stark. Algorithmic psychometrics and the scalable subject. Social Studies of Science, 48(2):204–231, April 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312718772094.

Stepan:1978:IBS

[Ste78a] Nancy Stepan. The interplay between socio-economic factors and medical science: Yellow fever research, Cuba and the United States. Social Studies of Science, 8(4):397–423, November 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277800800401.

Stern:1978:AAM

[Ste78b] Nancy Stern. Age and achievement in mathematics: a casestudy in the sociology of science. Social Studies of Science, 8 (1):127–140, February 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277800800105.

Stewart:1986:DCC

[Ste86] John A. Stewart. Drifting continents and colliding interests: a quantitative application of the interests perspective. Social Studies of Science, 16(2):261–279, May 1, 1986. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/0306312786016002003. See [Lau87] and reply [Ste87].

Stewart:1987:DCI

[Ste87] John A. Stewart. Drifting or colliding interests? A reply to Laudan with some new results. Social Studies of Science, 17 (2):321–331, May 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631287017002006. See

[Ste86, Lau87].

Stewart:1990:BRJ

[Ste90] Larry Stewart. Book review: John Gascoigne, Cambridge in the Age of the Enlightenment: Science, Religion and Politics from the Restoration to the French Revolution. Social Studies of Science, 20(4):764–767, November 1, 1990. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020004011.

Steiner:1999:CST

[Ste99] Carol J. Steiner. Constructive science and technology studies: On the path to being? Social Studies of Science, 29 (4):583–616, August 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631299029004005.

Stengel:2006:BM

[Ste06] Katrina Stengel. Bodies that matter. Social Studies of Science, 36(5):809–813, October 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706056409.

Steingart:2012:GTG

[Ste12] Alma Steingart. A group theory of group theory: Collaborative mathematics and the ‘uninvention’ of a 1000-page proof. Social Studies of Science, 42(2):185–213, April 1, 2012. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712436547.

Shadish:1995:AJA

[STGG95] William R. Shadish, Donna Tolliver, Maria Gray, and Sunil K. Sen Gupta. Author judgements about works they cite: Three studies from psychology journals. Social Studies of Science, 25 (3):477–498, August 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631295025003003.

Stigler:1987:PMF

[Sti87] Stephen M. Stigler. Precise measurement in the face of error: a comment on MacRoberts and MacRoberts. Social Studies of Science, 17(2):332–334, May 1, 1987. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/030631287017002007. See [MM86] and reply [MM87].

Stilgoe:2018:MLS

[Sti18] Jack Stilgoe. Machine learning, social learning and the governance of self-driving cars. Social Studies of Science, 48(1): 25–56, February 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312717741687.

Smart:2008:SRE

[STM^{+}08] Andrew Smart, Richard Tutton, Paul Martin, George T. H. Ellison, and Richard Ashcroft. The standardization of race and ethnicity in biomedical science editorials and UK biobanks. Social Studies of Science, 38(3):407–423, June 1, 2008. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312707083759.

Stokes:1982:DHW

[Sto82] T. D. Stokes. The double helix and the warped zipper — an exemplary tale. Social Studies of Science, 12(2):207–240, May 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631282012002002.

Stockdale:1995:STA

[Sto95] Alan Stockdale. ‘Stop talking about science!’: A response to Labinger. Social Studies of Science, 25(2):337–341, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631295025002013. See [Lab95b].

Stolzenberg:2004:KGS

[Sto04a] Gabriel Stolzenberg. Kinder, gentler science wars. Social Studies of Science, 34(1):77–89, February 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/0306312704041336. See comment [Lyn04e] and replies [BS04b, Sto04b].

Stolzenberg:2004:RR

[Sto04b] Gabriel Stolzenberg. Replies to the replies. Social Studies of Science, 34(1):115–132, February 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704041343. See [BS04b, Sto04a].

Stolzenberg:2008:REV

[Sto08] Gabriel Stolzenberg. Review essay: a very bad argument: Paul Boghossian, Fear of Knowledge: Against Relativism and Constructivism (Oxford: Clarendon Press, 2005), 135 pp., £15.99 / EUR 25.30 / $24.95. ISBN 0-19-928718-X. Social Studies of Science, 38(6):951–957, December 1, 2008. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312708098611.

Stranges:1992:FDW

[Str92] Anthony N. Stranges. Farrington Daniels and the Wisconsin process for nitrogen fixation. Social Studies of Science, 22(2):317–337, May 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631292022002008.

Street:2011:AKM

[Str11] Alice Street. Artefacts of not-knowing: The medical record, the diagnosis and the production of uncertainty in Papua New Guinean biomedicine. Social Studies of Science, 41(6): 815–834, December 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312711419974.

Street:2016:HHI

[Str16] Alice Street. The hospital and the hospital: Infrastructure, human tissue, labour and the scientific production of relational value. Social Studies of Science, 46(6):938–960, December 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312716638953.

Sturdy:2007:KCB

[Stu07] Steve Sturdy. Knowing cases: Biomedicine in Edinburgh, 1887– 1920. Social Studies of Science, 37(5):659–689, October 1, 2007.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312707076597.

Subramaniam:2025:SL

[Sub25] Banu Subramaniam. Silence of the labs. Social Studies of Science, 55(2):303–307, April 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127251314452.

Suchman:2006:WCC

[Suc06] Lucy Suchman. Wajcman confronts cyberfeminism. Social Studies of Science, 36(2):321–327, April 1, 2006. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706058828.

Suchman:2023:IOA

[Suc23] Lucy Suchman. Imaginaries of omniscience: Automating intelligence in the US Department of Defense. Social Studies of Science, 53(5):761–786, October 1, 2023. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127221104938.

Sudai:2022:WNM

[Sud22] Maayan Sudai. ‘A woman and now a man’: The legitimation of sex-assignment surgery in the United States (1849– 1886). Social Studies of Science, 52(1):79–105, February 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/03063127211056891.

Sullivan:1999:RM

[Sul99] Philip Sullivan. Response to MacKenzie. Social Studies of Science, 29(2):215–223, April 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631299029002003.

See reply [Mac99c].

Sundberg:2007:PBO

[Sun07] Mikaela Sundberg. Parameterizations as boundary objects on the climate arena. Social Studies of Science, 37(3):473–488, June 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312706075330.

Sundberg:2011:DCC

[Sun11] Mikaela Sundberg. The dynamics of coordinated comparisons: How simulationists in astrophysics, oceanography and meteorology create standards for results. Social Studies of Science, 41 (1):107–125, February 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312710385743.

Supper:2014:SFC

[Sup14] Alexandra Supper. Sublime frequencies: The construction of sublime listening experiences in the sonification of scientific data. Social Studies of Science, 44(1):34–58, February 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713496875.

Sutton:1984:OAP

[Sut84] John R. Sutton. Organizational autonomy and professional norms in science: a case study of the Lawrence Livermore Laboratory. Social Studies of Science, 14(2):197–224, May 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014002002.

Suzuki:2021:ICM

[Suz21] Wakana Suzuki. Improvising care: Managing experimental animals at a Japanese laboratory. Social Studies of Science, 51 (5):729–749, October 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/03063127211010223.

Saikkonen:2022:TPE

[SV22] Sampsa Saikkonen and Esa Väliverronen. The trickle-down of political and economic control: On the organizational suppression of environmental scientists in government science. Social Studies of Science, 52(4):603–617, August 1, 2022. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127221093397.

Sideri:2025:TPC

[SvD25] Katerina Sideri and Niels van Dijk. The techno-politics of computing the mind: Opening the black box of digital psychiatry. Social Studies of Science, 55(3):382–403, June 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127241273067.

Sargent:2021:SDC

[SVS21] Adam Sargent, Alexandra H. Vinson, and Reed Stevens. Sensing defects: Collaborative seeing in engineering work. Social Studies of Science, 51(4):564–582, August 1, 2021. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312721991919.

Swales:1989:EBE

[Swa89]  John M. Swales. Are Editors being exploited?  Social

Studies of Science, 19(1):152–153, February 1, 1989. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019001007.

Swanson:2007:BCL

[Swa07] Kara Swanson. Biotech in court: a legal lesson on the unity of science. Social Studies of Science, 37(3):357–384, June 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312706071793.

Swallow:2024:EBA

[Swa24] Julia Swallow. Enrolling the body as active agent in cancer treatment: Tracing immunotherapy metaphors and materialities. Social Studies of Science, 54(2):305–321, April 1, 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231199217.

Sullivan:1977:CCA

[SWB77a] Daniel Sullivan, D. Hywel White, and Edward J. Barboni. Cocitation analyses of science: an evaluation. Social Studies of Science, 7(2):223–240, May 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277700700205.

Sullivan:1977:SSI

[SWB77b] Daniel Sullivan, D. Hywel White, and Edward J. Barboni. The state of a science: Indicators in the specialty of weak interactions. Social Studies of Science, 7(2):167–200, May 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277700700203.

Schull:2011:SBN

[SZ11] Natasha Dow Schüll and Caitlin Zaloom. The shortsighted brain: Neuroeconomics and the governance of choice in time. Social Studies of Science, 41(4):515–538, August 1, 2011. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312710397689.

Szostak:1992:TCR

[Szo92] Rick Szostak. Technology and culture revisited. Social Studies of Science, 22(2):412–418, May 1, 1992. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631292022002014.

Taipale:2019:JST

[Tai19] Jaakko Taipale. Judges’ socio-technical review of contested expertise. Social Studies of Science, 49(3):310–332, June 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312719854538.

Takahashi:2024:IES

[Tak24] Makoto Takahashi. The improvised expert: Staging authority at an OECD Nuclear Energy Agency workshop in Fukushima. Social Studies of Science, 54(4):481–511, August 1, 2024. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127241231822.

TallBear:2013:GAI

[Tal13] Kim TallBear. Genomic articulations of indigeneity. Social Studies of Science, 43(4):509–533, August 1, 2013. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713483893.

Tancredi:2008:RPC

[Tan08] Laurence R. Tancredi. Review: Political cultures of science. Social Studies of Science, 38(3):470–475, June 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127080380030702.

Tan:2021:DSO

[Tan21a] Catherine D. Tan. Defending ‘snake oil’: The preservation of contentious knowledge and practices. Social Studies of Science, 51(4):538–563, August 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312721992543.

Tantchou:2021:MDP

[Tan21b] Josiane Carine Tantchou. Medical disposals and problem solving: About high blood pressure in Morocco. Social Studies of Science, 51(1):51–72, February 1, 2021. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720946487.

Tatarewicz:1986:FFP

[Tat86] Joseph N. Tatarewicz. Federal funding and planetary astronomy, 1950–75: a case study. Social Studies of Science, 16(1): 79–103, February 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631286016001005.

Taylor:1995:CCP

[Tay95] Peter Taylor. Co-construction and process: a response to Sismondo’s classification of constructivisms. Social Studies of Science, 25(2):348–359, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631295025002015.

See [Sis93b] and reply [Sis95].

Timmermans:1997:SAA

[TB97] Stefan Timmermans and Marc Berg. Standardization in action: Achieving local universality through medical protocols.

Social Studies of Science, 27(2):273–305, April 1, 1997. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631297027002003.

Tweney:1982:RDF

[TDM82] Ryan D. Tweney, Michael E. Doherty, and Clifford R. Mynatt. Rationality and disconfirmation: Further evidence. Social Studies of Science, 12(3):435–441, August 1, 1982. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631282012003004.

Teich:1986:ESE

[TG86] Albert H. Teich and Barry D. Gold. Education in science, engineering and public policy: a stocktaking. Social Studies of Science, 16(4):685–704, November 1, 1986. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631286016004006.

Turnhout:2024:ROG

[TH24] Sander Turnhout and Willem Halffman. Readjusting observational grids in dragonfly field guides. Social Studies of Science, 54(1):105–132, February 1, 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127231183011.

Thagard:1989:WCR

[Tha89] Paul Thagard. Welcome to the cognitive revolution. Social Studies of Science, 19(4):653–657, November 1, 1989. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019004007.

Theberge:2004:NSH

[Thé04] Paul Théberge. The network studio: Historical and technological paths to a new ideal in music making. Social Studies of Science, 34(5):759–781, October 1, 2004. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312704047173.

Thevenot:2009:PSI

[Thé09] Laurent Thévenot. Postscript to the special issue: Governing life by standards: a view from engagements. Social Studies of Science, 39(5):793–813, October 1, 2009. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312709338767.

Thorpe:1999:ASA

[Tho99] Charles R. Thorpe. Atomic spaces and American culture. Social Studies of Science, 29(4):617–627, August 1, 1999. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029004006.

Thorpe:2002:DES

[Tho02] Charles Thorpe. Disciplining experts: Scientific authority and liberal democracy in the Oppenheimer case. Social Studies of Science, 32(4):525–562, August 1, 2002. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312702032004002.

Thorpe:2025:VDV

[Tho25] Charles Thorpe. Virtual diversity and the value-ladenness of science. Social Studies of Science, 55(2):308–315, April 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127251314860.

Timmermans:1994:SS

[Tim94] Stefan Timmermans. Science saga. Social Studies of Science, 24 (1):133–134, February 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631279402400109.

Timmermans:1996:SLS

[Tim96] Stefan Timmermans. Saving lives or saving multiple identities?: The double dynamic of resuscitation scripts. Social Studies of Science, 26(4):767–797, November 1, 1996. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631296026004003.

Timmermans:2003:BTB

[Tim03] Stefan Timmermans. A black technician and blue babies. Social Studies of Science, 33(2):197–229, April 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127030332014.

Timmermans:2015:TST

[Tim15] Stefan Timmermans. Trust in standards: Transitioning clinical exome sequencing from bench to bedside. Social Studies of Science, 45(1):77–99, February 1, 2015. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312714559323.

Tironi:2018:HII

[Tir18] Manuel Tironi. Hypo-interventions: Intimate activism in toxic environments. Social Studies of Science, 48(3):438–455, June 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312718784779.

Tibbetts:1985:DPM

[TJ85] Paul Tibbetts and Patricia Johnson. The discourse and praxis models in recent reconstructions of scientific knowledge generation. Social Studies of Science, 15(4):739–749, November 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631285015004007.

Tjhin:2024:RNN

[Tjh24] Daniel Aditya Tjhin. The return of nature? Negotiating the ‘renaturation’ of the Isar as an envirotechnical landscape. Social Studies of Science, 54(4):557–574, August 1, 2024. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231217577.

Timmermans:2000:RTS

[TL00] Stefan Timmermans and Valerie Leiter. The redemption of thalidomide: Standardizing the risk of birth defects. Social Studies of Science, 30(1):41–71, February 1, 2000. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030001002.

Teixeira:2011:ITD

[TL11] Luiz Antonio Teixeira and Ilana Löwy. Imperfect tools for a difficult job: Colposcopy, ‘colpocytology’ and screening for cervical cancer in Brazil. Social Studies of Science, 41(4):585–608, August 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312711408380.

Tempini:2018:CDR

[TL18] Niccolò Tempini and Sabina Leonelli. Concealment and discovery: The role of information security in biomedical data reuse. Social Studies of Science, 48(5):663–690, October 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312718804875.

Todd:1987:TTC

[Tod87] Edmund N. Todd. A tale of three cities: Electrification and the structure of choice in the Ruhr, 1886–1900. Social Studies of Science, 17(3):387–412, August 1, 1987. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631287017003001.

Todd:1992:EPW

[Tod92] Edmund N. Todd. Electric ploughs in Wilhelmine Germany: Failure of an agricultural system. Social Studies of Science, 22(2):263–281, May 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631292022002005.

Tolwinski:2019:FCI

[Tol19] Kasia Tolwinski. Fraught claims at the intersection of biology and sociality: Managing controversy in the neuroscience of poverty and adversity. Social Studies of Science, 49(2):141–161, April 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312719839149.

Tomes:1987:AMN

[Tom87] Nancy Tomes. The anatomy of madness: New directions in the history of psychiatry. Social Studies of Science, 17(2):358–371, May 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631287017002010.

Toren:1983:ATW

[Tor83] Nina Toren. Attitudes towards work: a comparison of Soviet and American immigrant scientists in Israel. Social Studies of Science, 13(2):229–253, May 1, 1983. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631283013002003.

Torrens:1992:SFS

[Tor92] Hugh S. Torrens. A study of ‘failure’ with a ‘successful innovation’: Joseph Day and the two-stroke internal-combustion engine. Social Studies of Science, 22(2):245–262, May 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631292022002004.

Toren:1994:PSI

[Tor94] Nina Toren. Professional-support and intellectual-influence networks of Russian immigrant scientists in Israel. Social Studies of Science, 24(4):725–743, November 1, 1994. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631279402400405.

Toumey:1991:MCS

[Tou91] Christopher P. Toumey. Modern Creationism and scientific authority. Social Studies of Science, 21(4):681–699, November 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631291021004003.

Toumey:2000:RCM

[Tou00] Christopher P. Toumey. Rationalization, Creationism and the mechanics of contradiction. Social Studies of Science, 30(1):

141–145, February 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631200030001006.

Tousignant:2013:BTM

[Tou13] Noemi Tousignant. Broken tempos: Of means and memory in a Senegalese university laboratory. Social Studies of Science, 43 (5):729–753, October 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312713482187.

Travis:1981:RRA

[Tra81] G. D. L. Travis. Replicating replication? Aspects of the social construction of learning in planarian worms. Social Studies of Science, 11(1):11–32, February 1, 1981. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631278101100102.

Tresch:2013:ATA

[Tre13] John Tresch. Another turn after ANT: an interview with Bruno Latour. Social Studies of Science, 43(2):302–313, April 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712468362.

Thorpe:2000:WWJ

[TS00] Charles Thorpe and Steven Shapin. Who was J. Robert Oppenheimer? Charisma and complex organization: Charisma and complex organization. Social Studies of Science, 30 (4):545–590, August 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL http://www.

jstor.org/stable/285781;   https://journals.sagepub.
com/doi/pdf/10.1177/030631200030004003.

Tennant:2021:AAV

[TS21] Chris Tennant and Jack Stilgoe. The attachments of ‘autonomous’ vehicles. Social Studies of Science, 51(6):846–870, December 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/full/10.1177/03063127211038752.

Tanweer:2024:ADS

[TS24] Anissa Tanweer and James Steinhoff. Academic data science: Transdisciplinary and extradisciplinary visions. Social Studies of Science, 54(1):133–160, February 1, 2024. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127231184443.

Tsai:2024:IDM

[Tsa24] Yu-Yueh Tsai. Indigenous DNA as a metaphor: Nation-building and scientific debates on the rediscovery of Taiwanese ancestry. Social Studies of Science, 54(5):777–802, October 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127241226869.

Tarkkala:2018:SSN

[TT18] Heta Tarkkala and Aaro Tupasela. Shortcut to success? Negotiating genetic uniqueness in global biomedicine. Social Studies of Science, 48(5):740–761, October 1, 2018. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312718801165.

Tuan:1971:ERE

[Tua71] Yi-Fu Tuan. Essay reviews: Environmental attitudes. Social Studies of Science, 1(2):215–224, April 1, 1971. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277100100206.

Turkeli:1973:DTE

[Tür73] Arif Türkeli. The doctoral training environment and postdoctorate productivity among Turkish physicists. Social Studies of Science, 3(3):311–318, July 1, 1973. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277300300305.

Turkle:1982:SCS

[Tur82] Sherry Turkle. The subjective computer: a study in the psychology of personal computation. Social Studies of Science, 12(2):173–205, May 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631282012002001.

Turner:1987:PPC

[Tur87] R. Steven Turner. Paradigms and productivity: The case of physiological optics, 1840–94. Social Studies of Science, 17 (1):35–68, February 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631287017001002.

Turnbull:1989:PMV

[Tur89a] David Turnbull. The push for a malaria vaccine. Social Studies of Science, 19(2):283–300, May 1, 1989. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/030631289019002004. See comments [Dal90] and reply [Tur90].

Turner:1989:DP

[Tur89b] Stephen Turner. Depoliticizing power. Social Studies of Science, 19(3):533–560, August 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631289019003007.

Turner:1989:TIU

[Tur89c] Stephen Turner. Towards an integrated understanding of science. Social Studies of Science, 19(2):370–374, May 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019002011.

Turnbull:1990:TTK

[Tur90] David Turnbull. Transporting technoscientific knowledge/ practice: The case of a malaria vaccine (reply to Daly). Social Studies of Science, 20(2):379–384, May 1, 1990. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.

1177/030631290020002009. See [Tur89a, Dal90] and reply [Tur91].

Turnbull:1991:LKA

[Tur91] David Turnbull. Local knowledge and ‘absolute standards’: a reply to Daly. Social Studies of Science, 21(3):571–573, August 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631291021003010. See [Dal91, Tur90].

Turnbull:1995:RTO

[Tur95a] David Turnbull. Rendering turbulence orderly. Social Studies of Science, 25(1):9–33, February 1, 1995. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631295025001002.

Turner:1995:ESJ

[Tur95b] Stephen Turner. Edward Shils (1 July 1910–23 January 1995). Social Studies of Science, 25(2):397–399, May 1, 1995. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631295025002022.

Turner:2001:TRT

[Tur01a] R. Steven Turner. On telling regulatory tales: rBST comes to Canada. Social Studies of Science, 31(4):475–506, August 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631201031004001.

Turner:2001:RML

[Tur01b] Stephen Turner. The representation machine … breaks. Social Studies of Science, 31(3):458–461, June 1, 2001. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031003015.

Turner:2001:WPE

[Tur01c] Stephen Turner. What is the problem with experts? Social Studies of Science, 31(1):123–149, February 1, 2001. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031001007.

Turner:2003:TSW

[Tur03] Stephen Turner. The third science war. Social Studies of Science, 33(4):581–611, August 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/0306312703334005.

Turner:2012:PD

[Tur12a] Stephen Turner. Polanyi defanged. Social Studies of Science, 42 (6):945–953, December 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312712458479.

Turner:2012:WHK

[Tur12b] Stephen Turner. Whatever happened to knowledge? Social Studies of Science, 42(3):474–480, June 1, 2012. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712436555.

Tutton:2011:PPR

[Tut11] Richard Tutton. Promising pessimism: Reading the futures to be avoided in biotech. Social Studies of Science, 41(3):411–429, June 1, 2011. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312710397398.

Tuunainen:2005:CHF

[Tuu05] Juha Tuunainen. Contesting a hybrid firm at a traditional University. Social Studies of Science, 35(2):173–210, April 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312705047825.

Tweedale:1993:BRI

[Twe93] Geoffery Tweedale. Book review: Ian Inkster, Science and Techonology in History: an Approach to Industrial Development. Social Studies of Science, 23(2):378–381, May 1, 1993. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631293023002008.

Twine:2005:BRB

[Twi05] Richard Twine. Book review: Biotechnology and Human Meaning. Social Studies of Science, 35(1):165–168, February 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312705046364.

Tweney:1985:CSR

[TY85] Ryan D. Tweney and Stephen A. Yachanin. Can scientists rationally assess conditional inferences? Social Studies of Science, 15 (1):155–173, February 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631285015001006.

Underhill:2025:WWD

[UB25] Vivian Underhill and Karen Barad. What work does ‘contamination’ do? An agential realist account of oil wastewater and radium in groundwater. Social Studies of Science, 55(3):350–381, June 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
abs/10.1177/03063127241281708.

Urban:1982:MGS

[Urb82] Dieter Urban. Mobility and the growth of science. Social Studies of Science, 12(3):409–433, August 1, 1982. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631282012003003.

Ureta:2014:NTC

[Ure14] Sebastián Ureta. Normalizing Transantiago: On the challenges (and limits) of repairing infrastructures. Social Studies of Science, 44(3):368–392, June 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312714523855.

Ureta:2025:BSR

[Ure25] Sebastián Ureta. Beyond samplism: Rethinking the field in exposure science. Social Studies of Science, 55(3):404–422, June 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127241288581.

Uruena:2022:AMP

[Uru22] Sergio Urueña. Anticipation and modal power: Opening up and closing down the momentum of sociotechnical systems. Social Studies of Science, 52(5):783–805, October 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/03063127221111469.

Unander:2020:RLH

[US20] Trine E. Unander and Knut H. Sørensen. Rhizomic learning: How environmental non-governmental organizations (ENGOs) acquire and assemble knowledge. Social Studies of Science, 50 (5):821–833, October 1, 2020. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312720908343.

Voss:2016:IPP

[VA16] Jan-Peter Voß and Nina Amelung. Innovating public participation methods: Technoscientization and reflexive engagement. Social Studies of Science, 46(5):749–772, October 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716641350.

Vallgaarda:2010:FHR

[Val10] Signild Vallgårda. Is the focus on health-related behaviours a new phenomenon? Social Studies of Science, 40(4):609–614, August 1, 2010. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312710366067. See reply [Arm10].

Vanderpool:1983:MSL

[Van83] Christopher K. Vanderpool. Marine science and the law of the sea. Social Studies of Science, 13(1):107–129, February 1, 1983. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631283013001006.

VanRaan:1989:MMM

[Van89] Anthony Van Raan. Michael Moravcsik (1928–89): In memoriam. Social Studies of Science, 19(4):763–764, November 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631289019004017.

VandePoel:2008:BEW

[Van08] Ibo Van de Poel. The bugs eat the waste: What else is there to know? Changing professional hegemony in the design of sewage treatment plants. Social Studies of Science, 38(4):605–634, August 1, 2008. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312707088149.

Vardy:2017:RPT

[Var17] Mark Vardy. Reading for precarious times. Social Studies of Science, 47(5):771–779, October 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312717704313.

Vardy:2020:RAV

[Var20] Mark Vardy. Relational agility: Visualizing near-real-time Arctic sea ice data as a proxy for climate change. Social Studies of Science, 50(5):802–820, October 1, 2020. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312720906532.

Vasileva:2015:STC

[Vas15] Bistra Vasileva. Stuck with/in a ‘turn’: Can we metaphorize better in science and technology studies? Social Studies of Science, 45(3):454–461, June 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL https://journals.sagepub.com/doi/full/10.1177/
0306312715576018.

Vaughan:1999:ROP

[Vau99] Diane Vaughan. The Rôle of the organization in the production of techno-scientific knowledge. Social Studies of Science, 29(6): 913–943, December 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631299029006005.

Vaughan:2000:EE

[Vau00] Diane Vaughan. An ethnographic excursion. Social Studies of Science, 30(6):951–956, December 1, 2000. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030006007.

Vincenti:2003:BCR

[VB03] Walter G. Vincenti and David Bloor. Boundaries, contingencies and rigor: Thoughts on mathematics prompted by a case study in transonic aerodynamics. Social Studies of Science, 33 (4):469–507, August 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312703334001.

vanderHeide:2020:MMR

[vdH20] Arjen van der Heide. Model migration and rough edges: British actuaries and the ontologies of modelling. Social Studies of Science, 50(1):121–144, February 1, 2020. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312719893465.

vanderSluijs:1998:ADS

[vdSvESW98] Jeroen van der Sluijs, Josée van Eijndhoven, Simon Shackley, and Brian Wynne. Anchoring devices in science for policy: The case of consensus around climate sensitivity. Social Studies of Science, 28(2):291–323, April 1, 1998. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631298028002004.

vanEijndhoven:1991:CEA

[vEG91] José van Eijndhoven and Peter Groenewegen. The construction of expert advice on health risks. Social Studies of Science, 21(2):257–278, May 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631291021002004.

Velho:1990:SIP

[Vel90] Léa Velho. Sources of influence on problem choice in Brazilian university agricultural science. Social Studies of Science, 20 (3):503–517, August 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631290020003005.

Vergragt:1988:SSI

[Ver88] Philip J. Vergragt. The social shaping of industrial innovations. Social Studies of Science, 18(3):483–513, August 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631288018003004.

Verran:2002:PMS

[Ver02] Helen Verran. A postcolonial moment in science studies: Alternative firing regimes of environmental scientists and Aboriginal landowners. Social Studies of Science, 32(5–6):729–762, December 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631270203200506.

Vertesi:2008:MGL

[Ver08] Janet Vertesi. Mind the gap: The London Underground map and users’ representations of urban space. Social Studies of Science, 38(1):7–33, February 1, 2008. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312707084153.

Vertesi:2012:SLR

[Ver12] Janet Vertesi. Seeing like a Rover: Visualization, embodiment, and interaction on the Mars exploration Rover mission. Social Studies of Science, 42(3):393–414, June 1, 2012. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712444645.

Vessuri:1987:SSS

[Ves87] Hebe M. C. Vessuri. The social study of science in Latin America. Social Studies of Science, 17(3):519–554, August 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631287017003006.

Vessuri:1996:BRJ

[Ves96] Hebe Vessuri. Book review: Jacqueline Fortes and Larissa Lomnitz, Becoming a Scientist in Mexico: The Challenge of Creating a Scientific Community in an Underdeveloped Country. Social Studies of Science, 26(1):186–191, February 1, 1996.

CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631296026001011.

Villanueva-Felez:2015:NRC

[VFWC15] Africa Villanueva-Felez, Richard Woolley, and Carolina Cañibano. Nanotechnology researchers’ collaboration relationships: a gender analysis of access to scientific information. Social Studies of Science, 45(1):100–129, February 1, 2015. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312714552347.

vanHeur:2013:TOS

[vHLW13] Bas van Heur, Loet Leydesdorff, and Sally Wyatt. Turning to ontology in STS? Turning to STS through ‘ontology’. Social Studies of Science, 43(3):341–362, June 1, 2013. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712458144.

vanHilvoorde:2007:FKG

[vHVdW07] Ivo van Hilvoorde, Rein Vos, and Guido de Wert. Flopping, klapping and gene doping: Dichotomies between ‘natural’ and ‘artificial’ in elite sport. Social Studies of Science, 37(2):173–200, April 1, 2007. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706063784.

Vedel:2017:WWG

[VI17] Jane Bjørn Vedel and Alan Irwin. ‘This is what we got, what would you like?’: Aligning and unaligning academic–industry relations. Social Studies of Science, 47(3):417–438, June 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312716689346.

Vicedo:1995:WTC

[Vic95] Marga Vicedo. What is that thing called Mendelian genetics? Social Studies of Science, 25(2):370–382, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631295025002018.

Vincenti:1991:SSI

[Vin91] Walter G. Vincenti. The scope for social impact in engineering outcomes: a diagrammatic aid to analysis. Social Studies of Science, 21(4):761–767, November 1, 1991. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631291021004006.

Vincenti:1995:TST

[Vin95] Walter G. Vincenti. The technical shaping of technology: Realworld constraints and technical logic in Edison’s electrical lighting system. Social Studies of Science, 25(3):553–574, August 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631295025003006.

Viner:1999:PSL

[Vin99] Russell Viner. Putting stress in life: Hans Selye and the making of stress theory. Social Studies of Science, 29(3):391–410, June 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631299029003003.

Viseu:2015:CNB

[Vis15] Ana Viseu. Caring for nanotechnology? Being an integrated social scientist. Social Studies of Science, 45(5):642–664, October 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0306312715598666.

Velho:1984:PCP

[VK84] Lea Velho and John Krige. Publication and citation practices of Brazilian agricultural scientists. Social Studies of Science, 14 (1):45–62, February 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631284014001004.

vanKeuren:2001:CWS

[vK01] David K. van Keuren. Cold War science in black and white: US intelligence gathering and its scientific cover at the Naval Research Laboratory, 1948–62. Social Studies of Science, 31 (2):207–229, April 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312701031002004.

vanLente:1998:RMT

[vLR98] Harro van Lente and Arie Rip. The rise of membrane technology: From rhetorics to social reality. Social Studies of Science, 28(2):221–254, April 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631298028002002.

Vogel:2006:BPW

[Vog06] Kathleen Vogel. Bioweapons proliferation: Where science studies and public policy collide. Social Studies of Science, 36(5): 659–690, October 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312706059460.

Voskuhl:2004:HMC

[Vos04] Adelheid Voskuhl. Humans, machines, and conversations: an ethnographic study of the making of automatic speech recognition technologies. Social Studies of Science, 34(3):393–421, June 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312704043576.

Velho:1998:DMP

[VP98] Léa Velho and Osvaldo Pessoa, Jr. The decision-making process in the construction of the Synchrotron Light National Laboratory in Brazil. Social Studies of Science, 28(2):195–219, April 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631298028002001.

Voss:2022:PDI

[VSS22] Jan-Peter Voß, Jannik Schritt, and Volkan Sayman. Politics at a distance: Infrastructuring knowledge flows for democratic innovation. Social Studies of Science, 52(1):106–126, February 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/03063127211033990.

Vezyridis:2021:IDA

[VT21] Paraskevas Vezyridis and Stephen Timmons. E-infrastructures and the divergent assetization of public health data: Expectations, uncertainties, and asymmetries. Social Studies of Science, 51(4):606–627, August 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312721989818.

vanWichelen:2023:ABM

[vW23] Sonja van Wichelen. After biosovereignty: The material transfer agreement as technology of relations. Social Studies of Science, 53(4):599–621, August 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127231177455.

Wadmann:2014:PIC

[Wad14] Sarah Wadmann. Physician-industry collaboration: Conflicts of interest and the imputation of motive. Social Studies of Science, 44(4):531–554, August 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312714525678.

Wagner:2009:MMI

[Wag09] Roy Wagner. Mathematical marriages: Intercourse between mathematics and semiotic choice. Social Studies of Science, 39(2):289–308, April 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312708099443.

Wagner:2013:MUU

[Wag13] Sarah Wagner. The making and unmaking of an unknown soldier. Social Studies of Science, 43(5):631–656, October 1, 2013. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713484646.

Wajcman:1986:TCP

[Waj86] Judy Wajcman. Technological choice and the politics of production. Social Studies of Science, 16(4):746–753, November 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631286016004011.

Wajcman:2000:RGT

[Waj00] Judy Wajcman. Reflections on gender and technology studies: In what state is the art? Social Studies of Science, 30 (3):447–464, June 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631200030003005.

Waksman:2004:CNT

[Wak04] Steve Waksman. California noise: Tinkering with hardcore and heavy metal in Southern California. Social Studies of Science, 34(5):675–702, October 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312704047614.

Walentynowicz:1975:SSP

[Wal75] Bohdan Walentynowicz. The science of science in Poland: Present state and prospects of development. Social Studies of Science, 5(2):213–222, May 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://

journals.sagepub.com/doi/pdf/10.1177/030631277500500208.

Wallis:1982:R

[Wal82] Roy Wallis. Review. Social Studies of Science, 12(3):477–480, August 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631282012003008.

Waldby:2000:FBI

[Wal00] Catherine Waldby. Fragmented bodies, incoherent medicine. Social Studies of Science, 30(3):465–475, June 1, 2000. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030003006.

Waldby:2001:CUH

[Wal01] Catherine Waldby. Code unknown: Histories of the gene. Social Studies of Science, 31(5):779–791, October 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631201031005005.

Warren:1971:SRT

[War71] Neil Warren. Is a scientific revolution taking place in psychology? — doubts and reservations. Social Studies of Science, 1 (3–4):407–413, October 1, 1971. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277100100307.

Warren:1974:NSN

[War74] Neil Warren. Normal science and the normal standards of scholarly debate. Social Studies of Science, 4(2):195–197, April 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277400400207.

Watson:1990:ISF

[Wat90] Helen Watson. Investigating the social foundations of mathematics: Natural number in culturally diverse forms of life. Social Studies of Science, 20(2):283–312, May 1, 1990. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020002004.

Waterton:2002:FFC

[Wat02] Claire Waterton. From field to fantasy: Classifying nature, constructing Europe. Social Studies of Science, 32(2):177–204, April 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312702032002001.

Watson:2014:LPC

[Wat14] Matthew C. Watson. Listening in the Pakal controversy: a matter of care in Ancient Maya studies. Social Studies of Science, 44 (6):930–954, December 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312714543964.

Wazeck:2013:MPS

[Waz13] Milena Wazeck. Marginalization processes in science: The controversy about the theory of relativity in the 1920s. Social Studies of Science, 43(2):163–190, April 1, 2013. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312712469855.

Wield:1978:STD

[WB78] David Wield and Carol Barker. Science, technology and development: Part of a course in Development Studies for first and second year engineering and medical students at the University of Dar Es Salaam, Tanzania. Social Studies of Science, 8(3):385–395, August 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277800800312.

Walsh:1996:CNS

[WB96] John P. Walsh and Todd Bayma. Computer networks and scientific work. Social Studies of Science, 26(3):661–703, August 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631296026003006.

Woolgar:1999:DAA

[WC99] Steve Woolgar and Geoff Cooper. Do artefacts have ambivalence: Moses’ bridges, winner’s bridges and other urban legends in S&TS. Social Studies of Science, 29(3):433–449, June 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631299029003005. See reply [Joe99b].

Williams:2016:EEE

[WC16] James W. Williams and Nikolai M. Cook. Econometrics as evidence? Examining the ‘causal’ connections between financial speculation and commodities prices. Social Studies of Science, 46(5):701–724, October 1, 2016. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312716658980.

Wynn:2018:PPD

[WC18] Alison T. Wynn and Shelley J. Correll. Puncturing the pipeline: Do technology companies alienate women in recruiting sessions? Social Studies of Science, 48(1):149–164, February 1, 2018. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312718756766.

Wagner-Dobler:1997:SOS

[WD97] Roland Wagner-Döbler. Self-organization of scientific specialization and diversification: a quantitative case study. Social Studies of Science, 27(1):147–170, February 1, 1997. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631297027001009.

Waidzunas:2015:MAO

[WE15] Tom Waidzunas and Steven Epstein. ‘For men arousal is orientation’: Bodily truthing, technosexual scripts, and the materialization of sexualities through the phallometric test. Social Studies of Science, 45(2):187–213, April 1, 2015. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312714562103.

Weingart:1973:SCS

[Wei73] Peter Weingart. Symposium On ‘Conditions of Scientific Evolution’. Social Studies of Science, 3(3):319–321, July 1, 1973. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277300300306.

Weimer:1974:HPRa

[Wei74a] Walter B. Weimer. The history of psychology and its retrieval from historiography: I. The problematic nature of history. Social Studies of Science, 4(3):235–258, July 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277400400302.

Weimer:1974:HPRb

[Wei74b] Walter B. Weimer. The history of psychology and its retrieval from historiography II: Some lessons for the methodology of scientific research. Social Studies of Science, 4(4):367–396, October 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277400400403.

Weindling:1990:BRL

[Wei90] Paul Weindling. Book review: Linda Bryder, Below the Magic Mountain: a Social History of Tuberculosis in Twentieth-Century Britain. Social Studies of Science, 20(1):172–174, February 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631290020001009.

Weinberg:1994:RSF

[Wei94a] Steven Weinberg. Response to steve Fuller. Social Studies of Science, 24(4):748–751, November 1, 1994. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631279402400407.

Weiss:1994:BRS

[Wei94b] Sheila Faith Weiss. Book review: Styles of Scientific Thought: The German Genetics Community 1900–1933. Social Studies of Science, 24(2):409–415, May 1, 1994. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631279402400212.

Weisz:2010:DFF

[Wei10] George M. Weisz. Dr Fleck fighting Fleck typhus. Social Studies of Science, 40(1):145–153, February 1, 2010. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312709348569.

Welsh:2003:R

[Wel03] Ian Welsh. Review. Social Studies of Science, 33(2):311–314, April 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/03063127030332018.

Welsh:2004:JWY

[Wel04] Ian Welsh. Just when you thought you could relax and forget about the bomb. Social Studies of Science, 34(1):139–141, February 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/0306312704039216.

Werskey:1971:BSO

[Wer71] Paul Gary Werskey. British scientists and ‘outsider’ politics, 1931–1945. Social Studies of Science, 1(1):67–83, January 1, 1971. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277100100105.

Westrum:1977:SIA

[Wes77] Ron Westrum. Social intelligence about anomalies: The case of UFOs. Social Studies of Science, 7(3):271–302, August 1, 1977. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277700700302.

Westrum:1978:SSI

[Wes78] Ron Westrum. Science and social intelligence about anomalies: The case of meteorites. Social Studies of Science, 8(4): 461–493, November 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277800800403.

Westrum:1979:NIS

[Wes79] Ron Westrum. The notion of independent simultaneous invention or discovery to the Editors, social studies of science Edinburgh, 31 January 1979 dear sirs. Social Studies of Science, 9(4):509, November 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277900900408.

Westrum:1986:JBD

[Wes86] Ron Westrum. Joseph Ben-David (1920–86): Sociologist of science and of higher education. Social Studies of Science, 16 (3):565–567, August 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631286016003011.

Westrum:1988:BRE

[Wes88] Ron Westrum. Book review: Ernest Baümler, trans. Grant Edwards, Paul Ehrlich: Scientist for Life (New York and London: Holmes & Meier, 1984), xvi + 288 pp., $39.50. ISBN 0-8419-0837-0. Social Studies of Science, 18(1):189–190, February 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631288018001010.

Westrum:1989:SCT

[Wes89] Ron Westrum. The social construction of technological systems. Social Studies of Science, 19(1):189–191, February 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631289019001010.

West:2022:CIC

[Wes22] Sarah Myers West. Cryptography as information control. Social Studies of Science, 52(3):353–375, June 1, 2022. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/03063127221078314.

Westerlaken:2024:DTD

[Wes24] Michelle Westerlaken. Digital twins and the digital logics of biodiversity. Social Studies of Science, 54(4):575–597, August 1, 2024. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127241236809.

Wetmore:2007:DRR

[Wet07] Jameson M. Wetmore. Distributing risks and responsibilities: Flood hazard mitigation in New Orleans. Social Studies of Science, 37(1):119–126, February 1, 2007. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706069433.

Wolek:1974:PIC

[WG74] Francis W. Wolek and Belver C. Griffith. Policy and informal communications in applied science and technology. Social Studies of Science, 4(4):411–420, October 1, 1974. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277400400406.

Welsh:2006:CRU

[WG06] Rick Welsh and Leland Glenna. Considering the role of the University in conducting research on agri-biotechnologies. Social Studies of Science, 36(6):929–942, December 1, 2006. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706060062.

Wadmann:2021:SSR

[WH21] Sarah Wadmann and Amalie Martinus Hauge. Strategies of stratification: Regulating market access in the era of personalized medicine. Social Studies of Science, 51(4):628–653, August 1, 2021. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/03063127211005539.

Woodhouse:2002:SSA

[WHBM02] Edward Woodhouse, David Hess, Steve Breyman, and Brian Martin. Science studies and activism: Possibilities and problems for reconstructivist agendas. Social Studies of Science, 32(2):297–319, April 1, 2002. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312702032002004.

Whitley:1975:BSA

[Whi75a] R. D. Whitley. The British Sociological Association’s Sociology of Science Study Group. Social Studies of Science, 5(4): 485–488, November 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277500500407.

Whittemore:1975:WWP

[Whi75b] Gilbert F. Whittemore. World War I, poison gas research, and the ideals of American chemists. Social Studies of Science, 5 (2):135–163, May 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277500500202.

Whitley:1976:UPS

[Whi76] Richard Whitley. Umbrella and polytheistic scientific disciplines and their elites. Social Studies of Science, 6(3–4):471– 497, September 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277600600309.

Whittaker:1989:CCS

[Whi89] John Whittaker. Creativity and conformity in science: Titles, keywords and co-word analysis. Social Studies of Science, 19 (3):473–496, August 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631289019003004.

Wieser:2017:JAS

[Wie17] Bernhard Wieser. It’s not just about speed: Reviewing the recumbent bicycle once more. Social Studies of Science, 47 (2):300–304, April 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312716681521.

Williams:1973:SPA

[Wil73] Roger Williams. Some political aspects of the Rothschild Affair. Social Studies of Science, 3(1):31–46, January 1, 1973. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277300300103.

Williams:1988:AWH

[Wil88] Bill Williams. Accidents will happen. Social Studies of Science, 18(3):556–560, August 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631288018003009.

Willink:1991:OSG

[Wil91] Bastiaan Willink. Origins of the Second Golden Age of Dutch science after 1860: Intended and unintended consequences of educational reform. Social Studies of Science, 21(3):503–526, August 1, 1991. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631291021003004.

Winstanley:1976:ALC

[Win76] Monica Winstanley. Assimilation into the literature of a critical advance in molecular biology. Social Studies of Science, 6(3–4): 545–549, September 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277600600312.

Winskel:2002:AEN

[Win02a] Mark Winskel. Autonomy’s end: Nuclear power and the privatization of the British electricity supply industry. Social Studies of Science, 32(3):439–467, June 1, 2002. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312702032003004.

Winskel:2002:WSO

[Win02b] Mark Winskel. When systems are overthrown: The ‘dash for gas’ in the British electricity supply industry. Social Studies of Science, 32(4):563–598, August 1, 2002. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312702032004003.

Wirten:2015:PMC

[Wir15] Eva Hemmungs Wirtén. The pasteurization of marie Curie: a (meta)biographical experiment. Social Studies of Science, 45 (4):597–610, August 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312715589220.

Wise:2001:RRG

[Wis01] Steve Wise. Revolution in references: Give readers a chance by putting page numbers. Social Studies of Science, 31(2):309–310, April 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/0306312701031002009.

Woolgar:2013:WBT

[WL13] Steve Woolgar and Javier Lezaun. The wrong bin bag: a turn to ontology in science and technology studies? Social Studies of Science, 43(3):321–340, June 1, 2013. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312713488820.

Woolgar:2015:MQM

[WL15] Steve Woolgar and Javier Lezaun. Missing the (question) mark? What is a turn to ontology? Social Studies of Science, 45 (3):462–467, June 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312715584010.

Wade:2015:GRP

[WLBRS15] Peter Wade, Carlos López-Beltrán, Eduardo Restrepo, and Ricardo Ventura Santos. Genomic research, publics and experts in Latin America: Nation, race and body. Social Studies of Science, 45(6):775–796, December 1, 2015. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312715623108.

Winickoff:2017:PEJ

[WM17] David E. Winickoff and Matthieu Mondou. The problem of epistemic jurisdiction in global governance: The case of sustainability standards for biofuels. Social Studies of Science, 47 (1):7–32, February 1, 2017. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312716667855.

White-Nockleby:2022:GSB

[WN22] Caroline White-Nockleby. Grid-scale batteries and the politics of storage. Social Studies of Science, 52(5):689–709, October 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127221109605.

Wodak:1996:VCF

[WO96] Jo Wodak and David Oldroyd. ‘Vedic creationism’: a further twist to the evolution debate. Social Studies of Science, 26(1): 192–213, February 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631296026001012.

Wojick:1974:NRR

[Woj74] David Wojick. The norm of rationality or the rationality of norms. Social Studies of Science, 4(2):193–195, April 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277400400206.

Wolpert:1994:RSF

[Wol94] Lewis Wolpert. Response to Steve Fuller. Social Studies of Science, 24(4):745–747, November 1, 1994. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631279402400406. See [Ful94b, Ful94a].

Wolf:1995:ATO

[Wol95] Hans-Georg Wolf. An academy in transition: Organizational success and failure in the process of German unification. Social Studies of Science, 25(4):829–852, November 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631295025004011.

Woodward:1974:SGL

[Woo74] William R. Woodward. Scientific genius and loss of a parent. Social Studies of Science, 4(3):265–277, July 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277400400304.

Woolgar:1976:WIH

[Woo76] S. W. Woolgar. Writing an intellectual history of scientific development: The use of discovery accounts. Social Studies of Science, 6(3–4):395–422, September 1, 1976. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277600600306.

Woolgar:1981:CCT

[Woo81a] Steve Woolgar. Critique and criticism: Two readings of ethnomethodology. Social Studies of Science, 11(4):504–514, November 1, 1981. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631278101100406.

Woolgar:1981:IES

[Woo81b] Steve Woolgar. Interests and explanation in the social study of science. Social Studies of Science, 11(3):365–394, August 1, 1981. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631278101100304.

Woolgar:1982:LSC

[Woo82] Steve Woolgar. Laboratory studies: a comment on the state of the art. Social Studies of Science, 12(4):481–498, November 1, 1982. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631282012004001.

Woolgar:1986:ADB

[Woo86] Steve Woolgar. On the alleged distinction between discourse and praxis. Social Studies of Science, 16(2):309–317, May 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312786016002006.

Woolgar:1989:CCP

[Woo89] Steve Woolgar. A coffeehouse conversation on the possibility of mechanizing discovery and its sociological analysis. Social Studies of Science, 19(4):658–668, November 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019004008. See postscript [Mye89].

Woodgate:2006:TTA

[Woo06] Dawn Woodgate. Taking things apart: Ovario-hysterectomy — textbook knowledge and actual practice in veterinary surgery. Social Studies of Science, 36(3):367–397, June 1, 2006. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312706054518.

Weimer:1973:PNS

[WP73] Walter B. Weimer and David S. Palermo. Paradigms and normal science in psychology. Social Studies of Science, 3(3):211–244, July 1, 1973. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631277300300301.

Weimer:1974:SSD

[WP74] Walter Weimer and David Palermo. Standards, scholarship, and debate: a rejoinder to Warren. Social Studies of Science, 4(2):198–200, April 1, 1974. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277400400208.

Wray:2003:SRY

[Wra03] K. Brad Wray. Is science really a young man’s game? Social Studies of Science, 33(1):137–149, February 1, 2003. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312703033001961.

Wray:2005:RSS

[Wra05] K. Brad Wray. Rethinking scientific specialization. Social Studies of Science, 35(1):151–164, February 1, 2005. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312705045811.

Wright:1975:ASS

[Wri75] Peter Wright. Astrology and science in Seventeenth-Century England. Social Studies of Science, 5(4):399–422, November 1, 1975. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277500500402.

Wright:1980:ERR

[Wri80] Peter W. G. Wright. Essay review: The radical sociology of medicine: V. Navarro, Medicine Under Capitalism (New York: Prodist, and London: Croom Helm, 1976), 230 pp., $14.95 (£7.95). ISBN 0-88202-116-8. V. Navarro, Class Struggle, the State and Medicine: an Historical and Contemporary Analysis of the Medical Sector in Great Britain (London: Martin Robertson, 1978), 156 pp., £7.95. ISBN 0-85520-248-3. E. Krause, Power and Illness: The Political Sociology of Health and Medical Care (New York and London: Elsevier, 1977), 383 pp., $16.50 (£9.20), paperback (Amsterdam), $8.75 (£4.90). ISBN 0-444-99037-2 (paperback 99056-9). J. Ehrenreich (ed.), The Cultural Crisis of Modern Medicine (New York and London: Monthly Review Press, 1978), 300 pp., $15.00 (£8.85). ISBN 0-85345-438-8. Social Studies of Science, 10(1):103–120, February 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631278001000106.

Wright:1986:MBM

[Wri86] Susan Wright. Molecular biology or molecular politics? The production of scientific consensus on the hazards of recombinant DNA technology. Social Studies of Science, 16(4): 593–620, November 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631286016004003.

White:1979:ITE

[WSB79] D. Hywel White, Daniel Sullivan, and Edward J. Barboni. The interdependence of theory and experiment in revolutionary science: The case of parity violation. Social Studies of Science, 9 (3):303–327, August 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277900900302.

Wyatt:2007:RMT

[Wya07] Sally Wyatt. Review: Making time and taking time. Social Studies of Science, 37(5):821–824, October 1, 2007. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312707080044.

Wylie:2015:APA

[Wyl15] Caitlin Donahue Wylie. ‘The artist’s piece is already in the stone’: Constructing creativity in paleontology laboratories. Social Studies of Science, 45(1):31–55, February 1, 2015. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0306312714549794.

Wylie:2019:STS

[Wyl19] Caitlin Donahue Wylie. Socialization through stories of disaster in engineering laboratories. Social Studies of Science, 49(6): 817–838, December 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312719880266.

Wynne:1976:CGB

[Wyn76] Brian Wynne. C. G. Barkla and the J phenomenon: a case study in the treatment of deviance in physics. Social Studies of Science, 6(3–4):307–347, September 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277600600303.

Wynne:1988:UTP

[Wyn88] Brian Wynne. Unruly technology: Practical rules, impractical discourses and public understanding. Social Studies of Science, 18(1):147–167, February 1, 1988. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631288018001006.

Wynne:1992:CSP

[Wyn92a] Brian Wynne. Carving out science (and politics) in the regulatory jungle. Social Studies of Science, 22(4):745–758, November 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631292022004011.

Wynne:1992:RPC

[Wyn92b] Brian Wynne. Representing policy constructions and interests in SSK. Social Studies of Science, 22(3):575–580, August 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312792022003006.

Wynne:1996:SIP

[Wyn96] Brian Wynne. SSK’s identity parade: Signing-up, off-and-on. Social Studies of Science, 26(2):357–391, May 1, 1996. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631296026002007.

Wynne:1998:RR

[Wyn98] Brian Wynne. Reply to Radder. Social Studies of Science, 28(2):338–344, April 1, 1998. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631298028002007. See

[Rad98a, Rad98b].

Wynne:2003:STW

[Wyn03] Brian Wynne. Seasick on the third wave? Subverting the hegemony of propositionalism: Response to Collins & Evans (2002). Social Studies of Science, 33(3):401–417, June 1, 2003. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/03063127030333005. See [CE02].

Waters:2025:RDA

[WZH^{+}25] Colin N. Waters, Jan Zalasiewicz, Martin J. Head, Georg N. Schäfer, Francine M. G. McCarthy, and Simon D. Turner. Response to Damianos–Anthropocene angst: Authentic geology and stratigraphic sincerity. Social Studies of Science, 55(3):465–478, June 2025. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/03063127251343046.

Yair:2019:CCM

[Yai19] Gad Yair. Culture counts more than money: Israeli critiques of German science. Social Studies of Science, 49(6): 898–918, December 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312719878788.

Yaneva:2005:SET

[Yan05] Albena Yaneva. Scaling up and down: Extraction trials in architectural design. Social Studies of Science, 35(6):867– 894, December 1, 2005. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312705053053.

Yaneva:2022:AHL

[Yan22] Albena Yaneva. The architecture of the hybrid lab: Spacing graphene research. Social Studies of Science, 52(5):733–757, October 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
abs/10.1177/03063127221114721.

Yearley:1985:VFR

[Yea85] Steven Yearley. Vocabularies of freedom and resentment: a Strawsonian perspective on the nature of argumentation in science and the law. Social Studies of Science, 15(1):99–126, February 1, 1985. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631285015001004.

Yearley:1989:BSS

[Yea89a] Steven Yearley. Bog standards: Science and conservation at a public inquiry. Social Studies of Science, 19(3):421–438, August 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631289019003002.

Yearley:1989:ESS

[Yea89b] Steven Yearley. Environmentalism: Science and a social movement. Social Studies of Science, 19(2):343–355, May 1, 1989. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631289019002007.

Yearley:1990:RPB

[Yea90] Steven Yearley. Researching the Precambrian Biosphere: Constructing knowledge and shaping the organization of scientific work. Social Studies of Science, 20(2):313–332, May 1, 1990. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631290020002005.

Yearley:1992:SDI

[Yea92] Steven Yearley. Skills, deals and impartiality: The sale of environmental consultancy skills and public perceptions of scientific neutrality. Social Studies of Science, 22(3):435–453, August 1, 1992. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/0306312792022003001.

Yearley:1999:CMP

[Yea99] Steven Yearley. Computer models and the public’s understanding of science: a case-study analysis. Social Studies of Science, 29(6):845–866, December 1, 1999. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029006002.

Yearley:2001:MIS

[Yea01] Stevan Yearley. Mapping and interpreting societal responses to genetically modified crops and food. Social Studies of Science, 31(1):151–160, February 1, 2001. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631201031001008.

Lin:2014:CSL

[yLL14] Wen yuan Lin and John Law. A correlative STS: Lessons from a Chinese medical practice. Social Studies of Science, 44(6): 801–824, December 1, 2014. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312714531325.

Lin:2022:TDC

[yLL22] Wen yuan Lin and John Law. Thinking differently with Chinese medicine: ‘explanations’ and case studies for a postcolonial STS. Social Studies of Science, 52(4):491–511, August 1, 2022. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127221092180.

Yonay:1994:WBB

[Yon94] Yuval P. Yonay. When black boxes clash: Competing ideas of what science is in economics, 1924–39. Social Studies of Science, 24(1):39–80, February 1, 1994. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631279402400103.

Yonay:2004:WD

[Yon04] Yuval Yonay. Whose dreams? Social Studies of Science, 34 (4):621–629, August 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312704044906.

Young:1971:EBI

[You71] Robert M. Young. Evolutionary biology and ideology: Then and now. Social Studies of Science, 1(2):177–206, April 1, 1971. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631277100100204.

Yoxen:1984:SAC

[Yox84] E. J. Yoxen. Scepticism about the centrality of technology transfer in the Rockefeller Foundation Programme in Molecular Biology. Social Studies of Science, 14(2):248–252, May 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631284014002006.

Yoxen:1987:RRK

[Yox87] Edward Yoxen. Regulation, risk and knowledge. Social Studies of Science, 17(1):183–189, February 1, 1987. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631287017001011.

Zandvoort:1984:CNO

[Zan84] Henk Zandvoort. A comment on Nadel’s operationalization of incommensurability. Social Studies of Science, 14(3):443–446, August 1, 1984. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.
com/doi/pdf/10.1177/030631284014003005. See [Nad84].

Zhao:2023:MSD

[ZDI^{+}23] Helen Zhao, Marina DiMarco, Kelsey Ichikawa, Marion Boulicault, Meg Perret, Kai Jillson, Alexandra Fair, Kai DeJesus, and Sarah S. Richardson. Making a ‘sex-difference fact’: Ambien dosing at the interface of policy, regulation, women’s health, and biology. Social Studies of Science, 53(4):475–494, August 1, 2023. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/03063127231168371.

Ziegler:1988:WJD

[Zie88] Charles Ziegler. Waiting for Joe-1: Decisions leading to the detection of Russia’s first atomic bomb test. Social Studies of Science, 18(2):197–229, May 1, 1988. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631288018002002.

Ziewitz:2019:RGE

[Zie19] Malte Ziewitz. Rethinking gaming: The ethical work of optimization in web search engines. Social Studies of Science, 49 (5):707–731, October 1, 2019. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0306312719865607.

Zilsel:2000:SRS

[Zil00] Edgar Zilsel. The sociological roots of science. Social Studies of Science, 30(6):935–949, December 1, 2000. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631200030006006.

Zinberg:1971:WGA

[Zin71] Dorothy Zinberg. The widening gap: Attitudes of first-year students and staff towards chemistry, science, careers and commitment: An informal study of the Chemistry Department at an English university. Social Studies of Science, 1(3–4): 287–313, October 1, 1971. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277100100303.

Zinberg:1976:ETS

[Zin76] Dorothy S. Zinberg. Education through science: The early stages of career development in chemistry. Social Studies of Science, 6(2):215–246, May 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277600600203.

Zuiderent-Jerak:2009:CWR

[ZJ09] Teun Zuiderent-Jerak. Competition in the wild: Reconfiguring healthcare markets. Social Studies of Science, 39(5): 765–792, October 1, 2009. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/0306312709104433.